A Bibliography of Publications in Theoretical Computer Science: 2015–2019

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

12 March 2021 Version 1.107

Title word cross-reference

#SAT [1026, 1616].

(1 + 1) [2193]. (1 + ϵ) [1953]. (1 + λ) [10]. (1, λ) [2186]. (1.408 + ϵ) [529]. (3 + p) [2276]. (J, K) [360]. (μ, α) [212]. (n, 1) [1296]. (n, 3) [2064]. (n, k) [1484, 696, 2394, 1924, 1412, 1133]. (n^{2} − 1) [1702]. (r, l) [1829]. (s, t) [1365]. 1 [1641, 473, 2156, 1572, 2057, 288, 1160, 2312, 96]. 1.5 [387, 1153]. 1/3^{h} [725]. 13k [948]. 2 [873, 906, 1834, 2081, 2310, 500, 301, 542, 623, 1155, 724, 1657, 2108, 759, 1421, 609, 356, 804, 2058, 2360, 854, 1137, 1262, 1517, 165, 1717, 143, 39, 1403, 636, 2062, 2121]. {2, 3} [695]. 2 < p < ∞ [1922]. 2k [1764].

2n − 3 [1262]. 3 [2173, 351, 542, 869, 724, 545, ; 776, 287, 464, 1509, 2164, 1403, 1092, 799]. 4 ; [1447, 1990, 1210, 1080]. 7 [1762]. ^{R} [621]. ^{♯} ; [1772]. _{ew} [821]. _{n} [2305]. α ; [1466, 2146, 1890]. ∗ [1927]. b [1893, 534]. β ; [315, 1621]. C^{∞} [89]. C_{8} [208]. C_{9} [208]. Cb ; [1498]. D [1120, 707, 383]. D_{4} [2413]. Δ ; [2281]. DRH [1463]. ; F 2 → D|v = 1, c ≥ 1|C_{max} [1214]. F_{4} [149]. g ; [1925, 1913, 1238, 1482, 2237, 1924, 1413, ; 2051, 1412, 1133, 799, 361]. H ; [1830, 2174, 1406, 1496, 2342, 694]. H_{k} [59]. i ; [808]. ∞ [854]. K ; [93, 289, 2218, 2227, 1097, 1126, 752, 2154, ; 707, 851, 1665, 1170, 344, 623, 1663, 2313, ; 400, 797, 1892, 893, 1552, 1110, 1079, 325, ; 1575, 2174, 2320, 1157, 1532, 50, 2239, 470, ; 1689, 636, 1861, 1089, 1858, 1353, 2303]. K5

[346]. κ [1463]. L [378, 673, 1922]. L_{1} [2240, 337]. λ [506, 661]. λ_{2} [2170]. λβ [1163]. λβη [1839]. {⌊log_{b}(αn + β)⌋}_{n≥0} [1512]. log_{2} N [586]. M [1796, 1619, 2235]. F_{2} [1258]. R^{2} [693]. Z/pZ [2258]. Z^{3} [503]. μ [1748, 1670, 762, 1562]. N [1312, 1737, 623, 2033, 452, 861, 1349, 2305, 114, 1575, 2174, 2320, 1689, 1994, 799]. N^{k} [54]. O [1568]. O(n^{2} log m) [1538]. O(n^{3}) [1989]. P [1444, 563, 1652, 1651, 1650, 1455, 1456, 1231, 1649, 1460, 923, 1648, 926, 1647, 2038, 1859, 1922, 123, 625, 349]. P_{3} [521, 1486, 134]. P_{4} [195]. P_{k} [909]. π [672, 787]. Π_{1}^{1} [1017]. Π_{2} [1502]. P ||C_{max} [2281]. Q [179, 671]. Q_{8} [2413]. s [619, 638]. s_{2} [156]. σ [2186]. st [953]. T [129, 1574, 619, 25, 638]. t/k [2289, 1988]. Θ_{k}^{P} [1390]. X^{ω} [1287]. Z[x] [1914]. Z_{4}[u]/⟨u_{2} − 1⟩ [2046].

-abelian [356, 1110]. -agent [301]. -algebras [821]. -approval [383]. -approximate [1953, 2281]. -ary [623, 2174, 2320, 1689, 799]. -basis [1120]. -batch [123]. -binomial [143]. -Bonacci [2235]. -bound [59]. -Calculus [1748, 506, 672, 1670, 1839, 787]. -center [1859, 344, 2057, 470, 2062, 93]. -clique-colouring [724]. -clustering [797]. -Coloring [2360, 351, 1137, 2121]. -component [1990]. -composition [1552]. -Concave [1466]. -connected [869, 636]. -connectivity [1421, 1406, 1496]. -convex [1170, 2108]. -convexity [1486, 521]. -cube [2320, 799]. -cubes [623, 2174, 1689, 1994]. -cut-introduction [1502]. -cycles [1717, 1762]. -diagnosability [2289, 1988]. -dimensional [2033, 861, 1349, 2305]. -disjoint [623, 1262]. -distinct [752]. -domains [673]. -dominating [1155, 1892]. -domination [2313]. -edge [2342]. -equicontinuous [762]. -equivalence [2146]. -expansions [1621]. -extensions [25]. -extra [804, 1509, 1482, 2174, 2237, 1924, 2051, 361, 694]. -Fan-bundle-planar [1641]. -forest [2218]. -frame [1079]. -free [208, 1830, 909]. -gapped [1890]. -good-neighbor [1925, 1913, 1238, 1413, 2051, 1412, 1133, 799]. -graph [1829]. -graphs [953]. -group [349]. -invariant [149]. -isometric [39, 1403]. -kernel [948, 2303]. -kernelization [1764]. -length [1353]. -level [129]. -Mark [808]. -matching [1893, 534]. -maximal [2156]. -MaxSAT [2064]. -Means [289, 851]. -median [2038, 473, 2312, 1861]. -meet-semidistributive [1126]. -neighbor [1517]. -nested [1097]. -notation [1568]. -optimal [2170]. -out-of- [1575]. -packing [1657]. -partite [2303]. -partition [636]. -partitions [873, 906, 1834, 2081, 2310, 625]. -partners [195]. -path [2239, 2164, 1092]. -paths [1365]. -permutation [452]. -placement [325]. -planar [96]. -plane [1572, 1160]. -Player [114]. -protected [759]. -puzzle [1702]. -Quasicontinuous [156]. -Queens [1737]. -reduction [1163]. -regular [1447]. -relation [1796]. -resolution [179, 671]. -Restricted [695, 1157]. -reversible [50]. -SAT [2276, 1089]. -search [360]. -searcher [854]. -Self-Adaptation [2186]. -server [500]. -set [1665, 1532]. -sets [1498]. -sink [288]. -spaces [1922]. -spanners [212]. -spectrum [315]. -star [696, 2394, 1484, 1924, 1412, 1133]. -storage [400]. -sun-free [464]. -terms [661, 1562]. -tetrominoes [1574]. -time [1538]. -trees [2154, 165]. -unambiguous [1619]. -Uncut [1858]. -uniform [287]. -variables [1210]. -visibility [2227]. -word [1463]. -words [89].

1 [96]. 11th [2225]. 1st [234]. 1st-order [234].

2015 [1588, 1331]. 2019 [2345, 2386]. 21st [323]. 24 [2386]. 24th [732]. 2CNF [1679].

3 [2345, 2028]. 3-SAT [2028]. 34-36 [984].

411 [984]. 429 [68].

524 [474]. 552 [1329]. 557 [1004].

607 [1486]. 60th [879, 1534, 1416]. 646 [1023]. 657 [1354]. 696 [1496].

702 [2024]. 70th [2261]. 732 [2417]. 769 [2415]. 770 [2416]. 795 [2418]. 7th [505].

8th [505].

abelian [2272, 908, 2141, 1071, 356, 1110, 130, 455, 456, 1070, 858, 1333, 94]. Abelian-square-rich [1333]. absolute [404]. abstract [1226, 181, 1810]. abstraction [825, 1983, 1358]. abstractions [1586, 1731, 1746]. abstractor [1776]. AC [2146, 2273]. AC-automata [2273]. Accelerate [2220]. Accelerated [1301]. accelerating [969]. acceptors [1819]. access [1338, 254, 108, 2349, 2026, 1193, 85]. account [1443]. accounts [547]. Accumulating [2308]. accumulation [2238]. Accurate [1295]. Achieving [1952, 2169]. achromatic [1937]. across [2355]. action [843, 921]. activation [531, 206]. Active [1313, 1968, 1737, 2073, 2143, 1232, 103, 1738]. ACTL [1385]. Acyclic [1180]. ad [666, 254, 545, 656]. ad-hoc [545]. adaptation [937, 2186]. Adaptive [1381, 739, 938, 1590, 342, 334, 1018, 2193, 603, 2056, 840]. adaptively [947]. adaptively-secure [947]. adaptivity [56]. addition [1632, 515, 2343]. additive [2234, 212, 1781, 245]. additivity [780]. adjacency [1032]. adjacent [2321, 2311, 1148, 1762]. adjusting [249]. admit [1721]. Advancements [166]. Advances [1169, 1534, 774, 426, 934]. adversarial [2339]. adversaries [248]. adversary [2133]. adversary-dependent [2133]. advice [398, 575, 446]. advised [676]. affine [1734, 1902]. AG [2305]. against [1998, 634, 676]. agent [480, 301, 1475, 401, 2352, 1962]. agents [2331, 578, 1360, 1982, 1047, 710, 1489]. aggregate [201]. aggregating [1398]. aggregation [546, 994, 119]. Agreement [1929, 1665, 1811, 55, 2132, 1532, 2357, 613]. ahead [701]. airline [1860]. airspace [276]. AIXI [1591]. aka [1118]. al [914]. Alberto [1166]. algebra [2172, 763, 2106, 1523, 1176]. Algebraic [345, 2007, 1240, 904, 2065, 489, 491, 2101, 1678, 1805, 1961, 2111, 1772, 1838, 1752, 663, 2041]. Algebraic-coalgebraic [489]. algebras [1964, 2258, 73, 821]. Algorand [2098]. Algorithm [1003, 1657, 916, 2258, 1916, 500, 1872, 693, 149, 614, 1967, 473, 151, 526, 607, 1210, 1538, 334, 1792, 2156, 1609, 1155, 1892, 444, 10, 1996, 1788, 1707, 1289, 750, 1860, 1820, 200, 1733, 1297, 815, 64, 1893, 1914, 1781, 1365, 387, 1063, 461, 355, 779, 1769, 1764, 1419, 330, 1250, 1816, 1074, 1280, 2176, 2053, 102, 1568, 2013, 2415, 925, 2239, 2164, 599, 37, 636, 1861, 1653, 2161, 2067, 362, 638, 2055, 718, 1353, 1065, 494]. Algorithmic [2189, 1239, 77, 273, 1678, 1857, 1991, 732, 358, 1980, 268]. Algorithmics [525, 2195]. Algorithms [1061, 544, 1126, 1846, 936, 1534, 1897, 2231, 610, 1932, 2162, 1416, 2225, 2196, 1639, 600, 495, 1508, 1507, 2263, 1677, 1119, 2155, 1797, 548, 163, 1012, 1150, 34, 976, 440, 108, 110, 1873, 2326, 457, 1256, 344, 30, 55, 1026, 1725, 1669, 1088, 2313, 516, 2131, 1593, 240, 1218, 687, 104, 1077, 547, 417, 679, 2222, 1723, 2187, 239, 38, 1999, 963, 2223, 2221, 49, 224, 1612, 658, 359, 1364, 1660, 1066, 2408, 552, 47, 1246, 1547, 2042, 315, 499, 631, 2343, 1911, 1232, 124, 319, 912, 1084, 350, 1087, 1358, 532, 147, 88, 386, 1815, 446, 1410]. algorithms [2285, 1265, 2019, 534, 1522, 1039, 1435, 1086, 265, 1092, 819, 2064, 757, 1584, 535, 1030, 2279, 2199, 979]. Alignment [167, 2214, 223, 895, 1548, 1425]. alignment-free [1425]. aligns [1741]. all-pairs [1419]. All-to-all [604]. alliances [224]. allocation [802, 28, 306, 319]. allocations [1901, 803, 393]. almost [1182, 1514, 307, 13, 1518]. almost-confluent [307]. alphabet [1316, 2327, 1179, 1063, 870, 2270]. alphabet-independent [1063]. alphabetic [1649]. alphabets [1570]. ALT [732, 1588]. alternans [1985]. Alternating [2375, 1766, 1990, 1673, 2305, 493, 2060, 33]. Alternating-time [2375]. Alternation [1748, 2031]. alternative [854]. always [703, 1182]. Ambiguous [35]. Amihood [879]. Amir [879]. among [674]. amortized [2134, 1989]. amplifiers [1933]. analyses [823, 567]. analysing [763]. Analysis [495, 432, 522, 914, 251, 12, 2186, 817, 262, 2267, 1338, 281, 427, 282, 1404, 108, 876, 2175, 151, 1629, 253, 750, 23, 2079, 1857, 1515, 435, 939, 1985, 374, 931, 493, 832, 14, 748, 1185, 1842, 494, 521, 1486, 1518, 1740, 1568, 1688, 2241, 1942, 1653, 360, 1881, 2392, 1927]. analytic [481]. Analytical [2179]. Analyzing [2052, 616]. anarchy [2122]. ancestor [1065]. Ancestors [561]. Anchor [742]. anchored [1037]. Angelic [1917]. angle [900]. angular [1578]. annihilating [211]. annotated [1500]. Annual [2216]. annulus [1656, 1607]. anonymity [1818]. anonymization [515]. anonymized [2231]. Anonymous [216, 758, 654, 1181, 1089]. answering [193, 2118]. answers [2372, 1706]. ant [14]. antenna [250]. antennae [657, 314]. antennas [1479]. anti [1069, 1648]. anti-exponent [1069]. anti-spikes [1648]. antimatter [1456]. ants [580]. Antunes [2024]. Any [287]. Aperiodic [972, 1188, 1323]. Apollonian [1473]. Application [2206, 185, 723, 262, 2267, 270, 1345, 526, 1540, 1729, 567, 1951, 102, 1615, 2127, 969, 1962]. Applications [2292, 285, 1337, 427, 505, 1043, 2216, 2145, 413, 1872, 1062, 1382, 394, 2148, 345, 1760, 2349, 547, 306, 1806, 2133, 539, 1520, 251, 1551, 2264, 932, 1140, 380, 1854, 1065]. Applicative [263]. applied [672, 1866, 581, 1561, 943, 2186]. Applying [1370, 1111]. approach [2218, 302, 1269, 427, 1294, 1804, 1373, 1963, 1101, 71, 1713, 1925, 629, 253, 382, 1545, 1979, 336, 2322, 2166, 977, 1459, 1425, 874, 692, 1981, 169, 37, 663]. Approaches [1416, 1710]. approval [2168, 383]. Approximability [2181, 324, 1134, 3, 68, 2069, 2286, 1369, 1824]. Approximate [2280, 281, 1922, 1929, 546, 951, 894, 803, 1041, 792, 1436, 1012, 51, 998, 1953, 1364, 428, 2281, 1304, 1086]. Approximating [2218, 1156, 2057, 1236, 1089, 855]. Approximation [523, 628, 1725, 2131, 343, 47, 912, 350, 636, 819, 1858, 1697, 289, 1508, 1987, 440, 2326, 384, 55, 1155, 1920, 1759, 775, 545, 2221, 1396, 949, 2229, 708, 529, 1659, 697, 2125, 1911, 319, 1074, 532, 1301, 147, 386, 1715, 1039, 37, 1861, 2067, 1030, 2279, 979, 638]. approximations [465]. arbitrary [1607, 1235, 1108, 1327, 1640, 1384]. arborescenses [744]. Arc [526, 1347, 1235, 911, 1511, 330]. arc-consistency [1511]. arcs [1668, 1248]. argumentation [2149]. arguments [523, 1939]. arithmetic [2398, 1728, 2023, 1472]. arithmetization [443]. arrangements [2278]. array [1310]. arrays [437, 1539, 2408, 2409, 1271, 2412]. arrow [1384]. art [2409]. articulated [776]. ary [623, 2174, 2320, 1689, 799]. ASM [987, 1918]. ASP [431]. aspects [1797, 1678, 1991, 1802, 234, 358]. assembling [1630]. Assembly [1459, 614, 562, 1191, 1218, 1800, 1217]. assessment [1145]. asset [448]. assignable [755]. assignment [1946, 2281, 124, 718]. assignments [862, 849]. associated [1543].

Associative [127, 1143]. assumptions [2344, 2329]. Asymptotic [1418, 2177, 551, 10]. Asymptotically [1933]. asynchronous [16, 1444, 579, 1895, 1636, 1175, 1162, 697, 1630, 1532, 710, 1489, 2119]. asynchronously [1054, 2290]. at-most-one [1959]. ATL [2375]. atomic [1531, 1125, 1971]. attached [1530]. attack [1483, 2073, 1690, 1036]. attackers [2148]. attacks [1998, 634, 1952]. attractively [1219]. Attractor [429]. attribute [1642, 1197]. attribute-based [1642, 1197]. attributes [1345, 1193]. auctions [1821, 145, 131, 1686]. audio [1884]. auditing [2405]. augmentation [2230, 2218]. augmented [2033, 1209]. Automata [2206, 185, 1183, 1803, 1132, 1594, 2383, 752, 1169, 71, 407, 187, 1051, 457, 436, 1495, 2159, 838, 728, 2209, 193, 1175, 2308, 2104, 86, 1279, 1289, 565, 1230, 762, 1506, 593, 1278, 2384, 2273, 188, 1751, 62, 2381, 1203, 1731, 825, 1650, 1766, 190, 1321, 183, 765, 191, 329, 920, 1280, 1113, 205, 490, 1809, 106, 92, 2114, 2382, 1324, 2324, 650, 1022, 1894, 2211, 371, 1378, 612, 2188, 857, 1187, 1189, 1385]. automatic [2290, 2095, 1077, 25, 2315, 991, 795, 2231, 550, 2160]. automaticity [2315]. Automaton [1234]. Autonomous [597, 1810]. auxiliary [1827]. auxiliary-input [1827]. availability [61]. available [704]. Average [151, 590, 88, 627, 408, 402, 170, 538, 165]. Average-case [590, 408, 170]. Avoidability [1662, 1077, 1323, 455]. Avoidance [1754, 2257, 782, 2268]. Avoiding [143, 2093, 621]. aware [1150, 1409]. axiom [637]. axiomatization [2091, 685]. axioms [1225, 1375]. axis [2013, 2415]. axis-parallel [2013, 2415].

B [2024]. Back [559, 2209]. Backbone [208, 545]. backlog [496]. backtracking [718]. backward [2165]. bad [477, 39]. Balance [2397]. Balanced [385, 2288, 1509, 1950, 2117, 1752, 636, 2342]. Balancedness [2094]. balancing [520]. balls [855]. Bandit [995]. bandwidth [1524, 532]. barrier [198]. barriers [1932]. base [1554]. based [1120, 367, 1338, 302, 887, 1311, 1916, 972, 1804, 661, 270, 71, 1873, 1642, 850, 1088, 1609, 2168, 2046, 629, 775, 1422, 2238, 421, 1909, 1550, 2344, 1007, 939, 1673, 179, 931, 986, 343, 1998, 776, 493, 1840, 841, 1457, 1085, 1764, 1939, 1683, 2169, 1434, 1390, 2284, 1924, 366, 754, 1459, 568, 833, 1197, 932, 698, 692, 2371, 169, 588, 107, 2161, 2035]. bases [1352, 1797, 149, 814, 1754, 2343]. basic [1586, 1603, 317, 736, 58, 466]. basis [1120, 839]. batch [123, 1709, 1159, 207, 124, 237, 1769, 140, 714, 1435]. Batcher [874]. Batching [681]. battle [1629]. Baum [2178]. Bayesian [2302]. BC [209]. BC-subtrees [209]. BDI [1962]. be [727, 1800, 1717, 96]. Beachcombers’ [577]. Beating [349]. beaver [962]. beeping [2138]. behavior [1286, 1229]. behavioral [1637]. behaviors [417, 2112, 1113, 1340]. behaviour [1183, 762, 1746]. Behavioural [1776]. behaviours [1808]. being [129]. Belief [1756, 1886]. benefit [56]. benefits [2048]. Bertoni [1169, 1166]. Bertrand [1931]. best [1312, 1769]. Better [2079, 1079, 500, 703, 545, 1821, 1193]. between [74, 1636, 1728, 1556, 278, 1366, 87, 804, 1955, 2113, 650, 371, 1435]. betweenness [1429]. Beyond [2403, 1783, 1226, 728, 1259, 994, 2019]. BFS [270]. BFS-based [270]. BGP [445]. Bi [1207]. Bi-objective [1207]. biased [1074]. bichromatic [2066]. Biclique [1723]. bicliques [872]. bicriteria [1396, 229]. bidding [1272]. bidegree [1296]. bidirectional [894]. bifix [1168, 2208, 2210]. bifix-free [2208, 2210]. bifurcation [1985]. Big [1416, 965]. Bigraphs [180, 2334]. bijection [1104, 1907]. bilevel [386]. bilinear [30, 2236]. bimachine [2238]. Bin [465, 399, 913, 551, 469, 803, 446, 393]. Binary [1059, 1120, 2030, 1399, 291, 379, 908, 31, 2403, 1316, 924, 621, 1318, 1570, 2079, 1490, 725, 1768, 690, 1583, 1323, 1064, 1907, 1403]. binding [2185]. binning [1424]. Binomial [456, 143, 1956]. bins [2309]. biochemical [922, 283]. biological [2267, 281, 280, 109]. Biology [278, 426, 559]. bipancyclicity [471, 809, 807]. bipartite [97, 2070, 1400, 625, 1899, 40, 1391, 2020, 748, 2259, 2015, 88, 1615]. bipartition [1135]. birecurrent [1894]. bireversible [1506]. Birthday [1416, 879, 1534, 2261, 1036]. Bisimilar [1339]. bisimilarity [74, 1972]. Bisimulation [72, 428, 1979, 668, 372]. bistars [1828]. bit [1909, 1035]. bit-pair [1035]. Bivalent [481]. Bivariate [1518]. black [399, 502, 104, 595]. black-box [104]. blips [1325]. block [605, 607, 1059, 815, 112, 1140]. block-interchange [607]. block-interchanges [1140]. blockers [1830]. blocking [507, 784]. blood [567]. blowing [219]. Blue-split [1369]. Board [1687, 1695, 1704, 1711, 1726, 1745, 1758, 1765, 1654, 1667, 1671, 1675, 1682, 27, 7, 8, 264, 277, 284, 296, 308, 321, 333, 347, 365, 377, 391, 405, 410, 424, 433, 449, 459, 475, 484, 492, 504, 512, 524, 536, 555, 556, 21, 557, 69, 75, 82, 91, 98, 105, 113, 120, 128, 137, 144, 152, 158, 168, 177, 184, 194, 203, 214, 222, 232, 241, 255, 582, 583, 846, 856, 865, 875, 878, 897, 905, 918, 927, 930, 933, 946, 956, 964, 974, 985, 988]. Board [997, 1005, 1014, 1024, 1042, 1050, 1056, 1073, 584, 639, 640, 659, 665, 675, 684, 688, 699, 709, 715, 726, 731, 741, 751, 760, 773, 785, 790, 800, 811, 820, 828, 834, 1081, 1094, 1292, 1307, 1326, 1330, 1336, 1343, 1346, 1355, 1371, 1377, 1383, 1387, 1395, 1402, 1407, 1415, 1426, 1432, 1440, 1462, 1470, 1477, 1487, 1115, 1116, 1129, 1139, 1147, 1154, 1161, 1164, 1177, 1184, 1190, 1196, 1202, 1206, 1215, 1224, 1228, 1233, 1243, 1253, 1261, 1268, 1275, 1288, 1497, 1503, 1739, 1753, 1773]. Board [1782, 1790, 1801, 1813, 1825, 1837, 1844, 1852, 1862, 1870, 1877, 1504, 1516, 1525, 1533, 1549, 1560, 1564, 1573, 1580, 1587, 1597, 1605, 1613, 1620, 1627, 1634, 1638, 1644, 1661, 1889, 1896, 2120, 2128, 2135, 2144, 2152, 2163, 2171, 2182, 2194, 2205, 2215, 2224, 2233, 2245, 2260, 2274, 2291, 2301, 2346, 2368, 2376, 2387, 2400, 1906, 1915, 1919, 1926, 1934, 1941, 1947, 1954, 1960, 1965, 1976, 1986, 1992, 1993, 2001, 2008, 2014, 2025, 2034, 2044, 2054, 2068, 2077, 2088]. Boltzmann [419]. Bonacci [2235]. bondage [2359]. Boneh [1952]. Book [166, 1660]. boolean [992, 963, 146, 88, 1615, 1623, 826, 2172, 2403, 429, 127, 849, 1259, 1957, 2256, 2192, 1322, 233, 1752, 431]. Boosting [466]. bootstrap [19, 1517]. Borda [757]. Border [1061]. bordered [1335]. borders [622]. Born [947]. both [471]. bottleneck [161]. Bound [2287, 851, 613, 462, 225, 59, 404, 419, 2126, 447, 1959, 1524, 349, 1952]. Bound-decreasing [2287]. Boundary [646, 1380, 522, 1636, 189]. Bounded [1826, 76, 310, 1868, 1011, 887, 2029, 223, 1734, 2154, 2083, 311, 780, 1006, 876, 316, 1920, 43, 138, 1142, 816, 2375, 744, 1368, 2229, 1523, 520, 1855, 890, 521, 1486, 1135, 1376, 2378]. bounded-degree [521, 1486]. bounded-retrieval [1142]. Bounding [51, 220, 396, 477]. Bounds [2294, 2031, 1596, 1594, 2134, 703, 653, 1034, 1026, 869, 2366, 1045, 777, 1616, 1351, 872, 2353, 1072, 1710, 1079, 2059, 1902, 1493, 2042, 20, 765, 941, 2412, 1890, 1199]. box [104, 1680]. boxed [1538]. boxed-mesh [1538]. bracing [543]. Braids [2105]. brain [2106, 844]. Branch [600, 1820, 698]. branch-and-merge [1820]. Branch-and-reduce [600]. Branching [150, 644, 370, 668, 372]. branchings [385]. Brane [282, 1404]. breaches [408]. breakdown [960]. breaking [2049]. breakpoint [1717]. bribery [1255, 383, 1945]. broadcast [1642, 604, 2331, 160, 1998, 115, 656]. Broadcasting [254, 747, 53]. Brownian [1373]. Bruijn [1481, 1789, 1625]. Brzozowski [2403]. Bubble [1792, 952, 2130]. Bubble-Flip [1792]. bubble-sort [952, 2130]. bubblesort [1211]. Bucharest [1849]. Büchi [151]. Budget [1629, 802, 269, 2221]. Budget-constrained [1629]. budgets [145]. Buffer [298, 1513, 1602, 1246, 1643]. buffering [1246]. bunch [469]. bunched [686]. bundle [1641]. Burning [1835, 2294]. Burnside [1506]. burnt [589, 227]. Burrows [1540, 2162]. Busy [962, 57]. butterfly [1047]. button [1585]. buyback [2059]. BWT [1422, 1423]. BWT-based [1422]. Byzantine [613, 1531, 1929, 1532]. Byzantine-tolerant [1531].

C [2146, 1772]. cacti [899]. cactus [1859, 1657, 815]. cadences [1417]. Calculi [282, 74, 2147, 2007, 182, 1404]. Calculus [1748, 506, 686, 508, 77, 1341, 672, 1670, 2366, 1839, 787, 1342, 1376]. call [506]. call-by-value [506]. calls [1027]. Can [1717, 727, 2219]. Canadian [1697]. canalizing [1244, 2256]. Cancellation [311]. Cancellation-free [311]. candidate [1270]. candidate-sequential [1270]. canonical [629, 490]. Canonicity [2099]. Cantor [1287, 705]. Cantor-like [705]. capabilities [2148]. capability [1942, 2392, 1927]. capacitated [93, 681]. capacities [164, 2125]. capacity [1414, 844, 573]. capture [293]. Card [754]. Card-based [754]. cardiac [1985]. cardinalities [1110]. cardinality [624, 1511, 1430]. cards [1247]. cares [1542]. Carlo [938, 941, 942, 943]. carpenter [267]. cartesian [373, 1372, 604, 1351, 84, 689, 1204]. cascade [2203]. cascaded [2308]. cascades [353]. cascading [541]. case [600, 302, 394, 408, 151, 43, 138, 2100, 404, 590, 170, 2269, 306, 247, 1047, 102, 2027, 1303]. case-based [302]. cases [267, 1492]. Catalan [2093]. catalytic [764]. Catastrophic [541]. catch [810, 2293]. Categorial [197]. categorical [1980, 292]. categories [373]. Category [101, 1961, 1204]. catenation [2403]. caterpillar [1696]. Cauchy [79]. Causal [282, 155, 2003]. causalities [769]. causally [1501]. Cayley [130]. CCA [2344, 2307]. CCA2 [914]. CCA2-secure [914]. CDS [697]. Celebrity [980]. Cell [772, 2134, 564]. Cell-like [772]. cell-probe [2134]. cells [2181, 1741]. cellular [1132, 838, 1175, 1230, 762, 196, 498, 920, 205, 1187]. center [1507, 1859, 802, 93, 344, 903, 2057, 2087, 2394, 1988, 2226, 470, 2062, 1491]. centerpoint [693]. central [401, 2189]. centrality [1429]. century [323]. certain [2372, 2317, 25, 453, 2124]. certificate [756, 1550, 1955, 324]. certificate-based [1550]. certificateless [901]. certificates [1421]. ceteris [1125]. chain [871, 1614, 1574, 996]. chains [1632, 1431, 1841]. challenge [139]. change [740, 12]. changeover [744]. Changing [652]. channel [1381, 2148, 1642, 691, 1909, 428, 1401]. channels [254, 2349, 867, 931]. characterisation [257]. characteristic [2090, 1108, 1083, 1297]. characteristics [234]. Characterization [1260, 291, 1279, 2242, 16, 1100, 116, 2203, 248, 417, 1674, 921, 215, 670, 1390, 1568]. characterizations [1972, 1836, 1864, 2378]. characterized [394, 1650]. Characterizing [259, 370, 1806, 112]. characters [1101]. cheapest [679]. cheaters [1247]. Chebyshev [2312]. checkability [1530].

Checking [367, 768, 1916, 427, 1974, 876, 1728, 629, 1751, 986, 669, 169, 2060, 2061]. chemical [1980, 1979, 832, 2115, 1981]. Chern [1297]. cherry [1577]. cherry-picking [1577]. chess [939, 940]. Chinese [939, 413]. choice [1039]. cholesterol [567]. choosability [1148]. choosing [2168]. chop [1319]. Chord [1918]. chordal [439, 1967, 1817, 724, 2167, 2246, 1478, 175]. chordality [816]. chores [352]. chosen [603, 26]. chosen-key [26]. Christoffel [1793, 1106]. Chromar [1982]. chromatic [2306, 2341]. chromatin [2269]. Chronnectomics [1545]. chunks [160]. Church [307, 1465]. CIAA [2206, 185]. ciliates [562]. ciphertext [1642, 603, 1197, 588]. circle [2066, 1528, 2082]. circles [503]. Circuit [1472, 1616, 1710, 839]. circuits [722, 427, 311, 1026, 644, 2397, 253, 423]. Circular [900, 898, 526, 1149, 1107, 2271, 1662, 911, 2242, 330]. circular-arc [911, 330]. circulatory [567]. circumference [1136, 2041]. class [164, 2142, 294, 1102, 1200, 1297, 298, 1390, 955, 153, 1039, 1376, 1237, 1596, 96]. classes [1750, 2251, 1012, 503, 1611, 2, 65, 263, 678, 725, 2040, 991, 1650, 1110, 1009, 2124, 1380, 1465, 2200, 1783]. Classic [275, 481, 1040]. classic-like [481]. classical [789, 406, 151, 510, 1189, 1962]. classification [1757, 2220]. clauses [1679, 2108]. claw [1898]. claw-free [1898]. cleaners [574]. cleaning [498]. Clifford [2172, 1822]. clique [304, 2155, 1370, 1724, 724, 170, 519, 2246, 2016, 677, 346, 581]. clique-decomposition [1370]. clique-hypergraphs [346]. clique-width [304, 2155]. cliques [1724, 724, 2021, 602]. clones [1348, 2364]. close [295, 1494]. closed [2263, 373, 1204, 1372]. closest [2161, 2199, 435]. Closure [497, 2002, 805, 1123, 2100, 370, 1770, 276, 190, 2382, 1122]. closure-involution [2100]. closures [1797, 648]. cloud [1088, 2405]. cloud-based [1088]. clubs [2103, 1714]. clues [1848]. cluster [439, 1409]. clustered [129, 608, 2284]. clustered-level [129]. Clustering [1076, 797, 902, 1601, 1781, 447, 632, 1423]. clusters [440]. CNF [1959, 2064]. co [1478, 122, 102]. co-comparability [1478, 122]. co-optimization [102]. Coalgebraic [490, 491, 1809, 489]. coalgebraically [488]. coalitional [706]. coating [1222]. coboundaries [2094]. COCOA [2216]. code [43, 1849, 181, 566]. codes [1096, 1604, 341, 1174, 2046, 1201, 1940, 1579, 863, 2000, 1219]. coding [702, 336, 2161, 840]. coding-based [2161]. codomains [495]. CoFI [1777]. cognition [1887]. cognitive [843]. Cograph [1565]. cographs [2154, 1770]. coherence [1225]. coiling [1713]. Coin [2278]. coincide [2254]. coli [429]. collaboration [1461]. Collatz [725, 2]. collection [1434]. collections [1271, 2037]. Collective [2371]. colliding [1295]. colonies [1652]. colony [1650, 14]. color [1656, 971, 2198, 630, 855, 595]. color-spanning [2198, 630, 855]. colorable [2040, 2361]. colored [1101, 749, 468, 1541, 2395, 1676]. colorful [1759]. Coloring [1677, 1930, 2360, 346, 208, 351, 398, 2222, 1180, 1899, 1763, 1137, 2042, 1434, 1510, 1524, 6, 1004, 596, 1242, 917, 1762, 2121]. colorings [2154, 95, 1555, 1815]. colorless [1327]. colors [1930]. coloured [393]. colouring [2328, 2418, 898, 724, 658]. Colussi [725]. combination [1040]. combinations [1322]. Combinatori [626]. Combinatorial [305, 2136, 2216, 90, 131, 2351, 608, 341, 2269, 114, 1272, 340]. Combinatorics [622, 45, 1534, 1309, 1331, 1025, 422, 1670]. combinatory [1163]. combined [2403]. come [859]. Comer [2258]. comma [1096]. comma-free [1096]. Command [986]. Command-based [986]. Comment [725]. commitment [1909, 348]. committees [2168]. common [1411, 310, 1537, 889, 2365, 229, 124, 2285, 48, 1522, 1304, 1353, 1065]. communicating [1054, 1495, 1455]. Communication [651, 576, 242, 2318, 1029, 19, 867, 931, 1558, 1575, 2027, 2153]. communicationless [2308]. communications [1973]. communities [682, 1781]. community [2011, 244]. commutative [1143, 1476, 1735, 43, 138, 54, 127, 1766]. commutativity [806]. Compact [1527, 1064]. compacting [1642]. comparability [1478, 122]. comparative [2203]. compare [1425]. Comparing [1980, 2365, 1943]. comparison [1658, 717, 1390, 227, 1237]. comparisons [1480]. compartment [567]. compatibilities [681, 1435]. compatibility [4, 474, 237]. Competitive [516, 797, 23, 1246, 1388, 551, 679, 1045, 733, 360]. compiler [1854]. complement [368]. complementary [2201]. Complete [2203, 921, 1049, 116, 1450, 2070, 786, 1264, 1297, 1137, 2069, 1559, 683, 274, 637]. completely [1008, 2176]. Completeness [824, 1670, 667, 479, 1318, 593, 1581, 1242]. Completion [1581, 238, 755, 2335, 714, 1468]. Complex [1951, 842, 79, 382, 329, 1301]. Complex-demand [1951]. complexes [712, 334]. complexities [2073, 1783]. Complexity [722, 768, 439, 1903, 975, 2207, 1655, 256, 1666, 198, 2222, 2208, 1897, 95, 2075, 403, 576, 2000, 641, 2069, 2377, 1127, 1517, 521, 1486, 798, 172, 817, 1135, 549, 1411, 1239, 310, 1143, 1868, 1011, 1829, 1152, 186, 2310, 1684, 1987, 1734, 1010, 788, 1013, 2317, 303, 908, 712, 258, 935, 1623, 480, 515, 1600, 2096, 517, 1586, 2403, 822, 384, 2141, 445, 1257, 642, 1316, 327, 1698, 2417, 1850, 1175, 527, 706, 104, 1759, 263, 2190, 259, 83, 1279, 724, 590, 170, 441, 1595, 777, 849, 1601, 1368, 356, 1000, 1763, 1955, 1270, 435]. complexity [1002, 2385, 497, 2212, 791, 1755, 2080, 2381, 2409, 539, 2247, 648, 2111, 550, 247, 592, 1943, 119, 2020, 632, 805, 705, 1390, 2243, 1212, 852, 633, 571, 1157, 1284, 92, 494, 649, 2265, 794, 2359, 2390, 1518, 389, 456, 2268, 1146, 2153, 117, 2210, 2318, 383, 1824, 2378, 1945, 1705, 2010, 29, 2303, 1697, 2118, 651, 682, 242]. complexity-theoretic [2118]. Component [283, 915, 1563, 784, 1990]. components [1759, 2109, 776]. componentwise [1698, 2417, 1305]. composable [1041]. Composition [2178, 1078, 1552, 1237]. Compositional [1867, 2101, 1778]. compositionality [74, 481]. compound [1585]. Comprehensive [417]. Compressed [594, 1724, 188, 885, 1571, 1066, 895, 1958]. compressibility [1251]. compression [737, 343]. compressors [1536]. compromises [306]. Comput [2418, 1023, 474, 1329, 2417, 1354, 984, 68, 2416, 1496, 1004, 1486, 2415, 2024]. Computability [234, 1591]. computable [257, 2275, 1240]. Computation [9, 309, 476, 2119, 2225, 1069, 1872, 1294, 1062, 270, 2097, 429, 1149, 999, 2308, 2012, 1071, 1083, 32, 487, 2005, 1203, 669, 833, 2265, 1173, 2414, 218, 18, 2378, 637, 285]. Computational [445, 1175, 426, 592, 1459, 1217, 794, 1197, 103, 2073, 682, 1318, 2344, 2116, 2049, 923, 1647]. computationally [275]. Computations [728, 16, 2050, 443, 1300, 1304]. compute [1297, 1914, 779, 2162]. Computer [774, 278, 559, 934, 1879]. computers [999, 831, 1884]. Computing [2372, 891, 1607, 1141, 395, 908, 1481, 2185, 31, 770, 862, 624, 2250, 1798, 2341, 423, 2226, 1526, 15, 719, 412, 602, 1306, 2258, 99, 2227, 548, 693, 149, 1669, 1370, 867, 1707, 987, 1327, 382, 315, 841, 1520, 1816, 754, 1157, 2116, 1442, 1743, 698, 1802, 626, 1025]. concatenating [1789]. Concatenation [1282]. Concatenation-free [1282]. concave [1474, 2204, 1367, 1466]. Concentration [1191]. concepts [1093]. concerning [987, 2255]. Concurrent [36, 727, 1327, 1386, 511, 1340]. condition [1451, 2058, 2192, 1855, 1305]. Conditional [1672, 1078, 60, 1237, 807, 2373, 171, 1925, 1913, 804, 1007, 1238, 1948, 2289, 1988, 717, 482, 1482, 2237, 1924, 1225, 1412, 799, 618]. conditions [1869, 1345, 565, 73, 1582, 96]. Conference [2216, 732, 2225, 626, 1025]. confidence [941]. configuration [2045]. configurations [2185]. confirmer [716]. Conflict [573, 95, 2063, 1815]. conflict-free [95, 1815]. confluence [192]. confluent [307]. conformance [367]. Congest [1876]. CONGESTBC [1920]. Congested [581]. congestion [1902, 1971]. congruential [307]. conically [2186]. conjecture [2111, 2255, 1619, 39]. Conjectures [390, 2403]. Conjugacy [863, 1693, 1200]. conjunction [754]. conjunctions [1125]. conjunctive [721, 1179, 720]. Connected [252, 1859, 270, 517, 1967, 353, 1256, 1492, 1218, 2300, 869, 1194, 816, 352, 3, 68, 2286, 554, 1741, 1022, 134, 636, 1997]. connectedness [2394, 1691]. Connecting [2071]. connection [935, 1003, 611, 2058, 554]. Connectivity [314, 1221, 2230, 121, 2155, 1990, 1145, 1672, 1492, 327, 250, 657, 83, 1421, 1366, 952, 804, 695, 1755, 1020, 1509, 2289, 1988, 2358, 717, 853, 2174, 1157, 1406, 1496, 2389, 1928, 2041, 2378, 1569, 361, 694, 915]. connectivity-preserving [1755]. consecutive [890, 532]. consensual [1173]. consensus [1665, 1586, 248, 1546, 2009]. consequences [463]. conservation [414, 1446]. conservative [1125]. Considerations [2045]. considered [553]. consist [373]. consistency [1511, 818, 2371, 792]. consistent [912, 1067, 46, 968]. Conspiracy [944]. Constant [1953, 1074, 655, 1872, 814, 1920, 240, 2018, 746, 631, 907, 588]. constant-size [588]. constant-sized [631]. constant-time [1872, 2018]. Constrained [542, 367, 1381, 1445, 1834, 2081, 351, 316, 619, 1629, 1134, 1193, 530, 319, 2186, 1700]. constraint [627, 61, 1472, 1806, 1959, 2286, 916, 1561, 511, 2371, 169, 1401, 1093]. constraints [1841, 2325, 269, 1946, 448, 2140, 83, 246, 1798, 1601, 2221, 173, 1511, 990, 2197, 1266, 890, 1397, 141, 585]. construct [599]. Constructing [1789, 744, 1543, 219, 1008, 545]. Construction [2046, 2343, 547, 1142, 1433, 2238, 130, 888, 1615, 2072, 2035]. Constructions [818, 1752, 752, 1234, 32, 35, 490, 1064, 118, 345]. Constructive [1972]. constructs [1320]. consumption [1398, 1414]. contact [814, 1622]. containing [2393, 1248]. Contention [1265]. Contention-sensitive [1265]. contest [2231]. contests [1629]. Context [1735, 1260, 332, 43, 138, 1674, 786, 646, 192, 1324, 1285, 904, 736, 1378, 676]. Context-free [1735, 1260, 43, 138, 1674, 786, 646, 1285, 904, 736, 1378, 676]. contexts [1618, 261, 1055, 290]. Contextual [1310]. contextuality [1883]. continuation [1295]. continuity [1230]. Continuous [1269, 711, 1916, 1373, 376, 1114, 2035, 1076]. continuous-time [1916]. Continuously [1940]. continuum [442]. contour [1294, 568]. contour-based [568]. contours [1457]. contractibility [1254]. Contraction [1830, 954, 518]. control [1310, 1338, 305, 771, 1867, 1193, 2399, 383, 1499]. controllability [2370]. controlled [1318, 1818]. controller [1203]. Controlling [664, 1270, 2112]. Convergence [1931, 2, 419, 725, 697, 286]. convergent [955]. conversion [1709]. conversions [2243]. Convex [1685, 625, 544, 1466, 1608, 2252, 784, 1170, 2207, 2108, 1400, 2212, 2354]. convexities [606]. convexity [438, 521, 1486]. convolution [775]. convolution-based [775]. cooperating [664]. cooperation [1461]. cooperative [1267, 1738]. Coordinate [742]. Coordinating [125]. coordination [238]. cop [293]. coproducts [1964]. Cops [2298, 2297, 527, 1291, 950, 2143]. Core [228, 624, 803]. corners [2252]. corrected [1328]. Correction [2417, 743]. corrections [1496]. correctness [1362, 182]. Correlation [1026, 632]. Correspondence [452, 789, 77]. Corrigendum [2418, 1023, 474, 1329, 1354, 984, 68, 2416, 1004, 1486, 2415, 2024]. coset [2258]. coset-translation [2258]. Cost [959, 1208, 628, 1905, 28, 468, 744, 538, 2059, 2286, 1575, 530, 1485, 533, 206]. Cost-sharing [959]. Costas [1534]. costs [1563, 753, 966, 90, 2281, 531, 1787]. Counter [1173, 1728, 1730, 806, 2393]. counters [1735, 1836]. Counting [1098, 2138, 2129, 1508, 1404, 1021, 5, 1329, 1328, 849, 949, 146, 2110, 1456, 1721, 1390, 1212, 1615, 880, 1086, 466]. countings [546]. courteous [272]. Cover [1274, 600, 2280, 652, 1018, 83, 1241, 2250, 1278, 2385, 1138, 2125, 1764, 2063, 2013, 2415, 1615, 1622, 2164, 2228, 973, 134, 1092, 819, 1030, 2279]. coverability [1866]. coverage [198, 2221, 1659, 172]. covered [1829]. Covering [1608, 1420, 164, 1681, 2326, 2181, 2244, 2408, 2409, 592, 1932, 2412, 1787, 553, 535]. coverings [1724]. covers [1576, 623, 519, 2408, 1262, 848, 713, 2086, 599, 33]. covert [931]. crash [1532]. crash-prone [1532]. creation [148, 1603]. Crick [1495, 2159, 1453]. crisis [1545]. criterion [90]. Critical [2117, 1002, 870]. Crochemore [2235]. Cross [1168, 1653]. Cross-bifix-free [1168]. Cross-Entropy [1653]. crossed [777, 1913, 1406, 1496]. crossing [900, 1554, 117]. crossing-free [1554]. crossings [2173]. crossover [1185]. crown [1764]. CRT [2336]. CRT-exponent [2336]. cryptanalysis [2336]. cryptograms [85]. cryptographic [2402]. cryptography [1760, 595]. cryptoscheme [1035]. cryptosystems [1483, 2035]. crystallization [945]. CSP [1917]. CSPs [849, 1259, 1582]. cube [171, 2033, 2320, 1406, 1496, 2241, 799, 1997]. cube-connected [1997]. cubes [623, 1515, 1913, 695, 2174, 143, 1413, 1209, 1689, 1994]. cubic [702, 294, 1948, 2237, 618]. cumulative [711, 1185]. Curry [789]. curvature [1427]. curve [777]. curves [2407]. cut [163, 619, 552, 2020, 1502, 1621, 981, 610, 638]. cuts [978, 1400, 1002]. cutting [1700]. CVP [750]. cycle [2306, 2258, 1360, 1030, 2279, 2130]. Cycles [2311, 967, 468, 1789, 910, 1137, 300, 1717, 2211, 1242, 1689, 1643, 1762, 1997]. cyclic [1183, 2046, 2028, 612]. cyclically [2303]. cyclicity [1530]. cycling [652]. cylinder [390, 637]. cylindric [1096].

D [2173, 542, 545, 609, 776, 387, 1153]. D0L [1768, 2242]. D0L-systems [2242]. DAG [1049]. DAG-width [1049]. dagger [491]. dangerous [655]. dark [939]. Data [53, 1882, 903, 1409, 1416, 2372, 1803, 579, 318, 523, 965, 240, 2405, 221, 2018, 970, 1422, 921, 2087, 2394, 1772, 119, 587, 1988, 748, 1434, 1891, 1265, 2137, 470, 2049]. database [1998, 482, 633, 1376]. Databases [1124, 340]. dataflow [1088]. dataset [1088]. date [229]. DAWGs [2273]. DCell [1988, 599]. Deadline [2032]. Deadness [928]. Dealing [1210, 17, 1722]. decay [1, 984]. December [2345, 2386]. Decentralized [983, 17]. deceptive [2184]. Decidability [1693, 1010, 406, 1636, 672, 1751, 1864, 190].

Decidable [1865, 153]. deciding [1054]. decimation [1295]. Decision [1077, 313, 379, 151, 2380, 629, 1875, 994, 2006]. decisions [2373]. decomposability [1368]. decomposable [2079]. decomposed [1717]. decomposing [2131]. Decomposition [2115, 1167, 1872, 948, 1370, 1174, 1742, 1764, 150, 668, 2016, 1981]. decompositions [2300, 672, 2341, 951, 482, 698, 1305]. decontamination [1047]. decrease [1467, 1047, 2024, 2127]. decreasing [2287]. decryption [818]. dedicated [732]. deduction [789, 1227]. default [80]. defect [1011]. defensive [2297]. definability [2084]. defined [1103, 1344]. defining [1444]. definite [1123, 1908]. definiteness [2255]. definition [1109, 35]. definitional [1198]. degradation [1142]. degrading [1665]. Degree [1834, 2081, 310, 1617, 758, 515, 316, 2366, 1194, 1178, 2058, 1770, 1249, 2364, 521, 1486, 1302, 2361, 973, 1135, 2378, 620]. degree-anonymous [758]. degree-bounded [1135, 2378]. Degree-constrained [2081]. Degrees [328, 270]. Delay [1688, 2083, 1565, 1240, 690]. delays [418, 1905]. delete [2324]. deletes [63]. deletion [439, 1817, 1830, 2046, 1318, 150, 2362, 1855, 2217, 1518]. deliveries [124]. delivery [238, 755, 61, 140, 714]. demand [145, 1951, 1084]. demands [1156, 2070]. deniability [2169]. Denotational [1450]. dense [619, 1551, 1714]. density [1068, 2249, 2126]. dependencies [922]. Dependency [414, 197, 1589, 2180, 633, 671]. dependent [260, 2099, 2133, 348, 740, 1867, 1633, 1787, 489, 2352]. deployment [1145]. Depth [1748, 1467, 311, 1464, 2024, 2127]. derivated [1791]. derivative [79]. derivatives [1106, 775]. Deriving [1602, 1925]. Descendants [1700, 561]. description [2179]. Descriptional [641, 92, 805, 2377]. descriptive [1707]. Design [27, 1223, 493, 432, 279, 1593, 59, 2405, 1398, 28, 1718, 1220, 2047, 107, 969, 1041]. designated [716]. Designing [2199, 104]. Designs [1581]. destination [2027]. detect [1065]. detectability [1379]. detecting [1052]. detection [80, 522, 244, 654, 2021, 2056, 1358]. deteriorating [207]. determinantal [423]. determinants [2278, 25, 2315]. determined [1290, 378]. Determining [2402]. Deterministic [1018, 658, 190, 213, 473, 701, 1278, 949, 247, 2113, 2021, 805, 191, 2324, 2015, 1431, 2239]. development [1807]. Developments [256]. DFA [239, 968]. DFAs [186, 1013]. diagnosabilities [617]. diagnosability [1925, 1366, 2087, 1078, 1913, 804, 2305, 1238, 2289, 1988, 2356, 60, 717, 1482, 2237, 1413, 227, 2051, 1412, 1237, 1133, 2342, 799, 618]. Diagnosis [814, 567, 2284, 1942, 2392, 1927]. diagonal [1091]. diagrams [379]. dial [1824]. dial-a-ride [1824]. diameter [270, 2020, 1393, 1080]. DICE [256, 1610]. Dichotomy [146, 898, 2168, 403, 2020]. Dictionary [299, 2273, 188]. different [2038, 503, 10, 1360, 493]. differential [45, 2097, 430, 794, 234]. Differentially [738]. Differentially-private [738]. Diffusion [2175, 281, 246]. digital [775, 759, 776, 779, 2179]. digraph [42, 47]. digraphs [385, 873, 906, 1834, 1617, 2310, 1576, 1235, 1614, 84, 1136, 1347, 1048, 2303, 2217]. Dijkstra [487]. Dilemma [2184]. Dilworth [474, 4]. Dimension [1494, 1692, 1199, 1601, 815, 1719, 2124, 615, 1987]. dimensional [1096, 1132, 883, 1882, 344, 738, 2033, 861, 1349, 2305, 2229, 825, 447, 1262, 650, 470, 1306]. dimensionality [739]. dimensions [1168, 851]. Diophantine [1523, 2084]. Dipole [1219]. Direct [796, 1269, 1297]. Directed [1119, 700, 1136, 1128, 1665, 2175, 353, 2380, 400, 1045, 1421, 689, 1957, 1158]. directional [1479, 250, 314]. Disagreement [1593]. disambiguation [1280]. Discovery [1120, 1124, 1542, 247]. discrepancy [1431]. Discrete [774, 1453, 2340, 1286, 495, 2277, 780, 781, 448, 1241, 1983, 783, 2062]. discriminating [364]. disentangling [2053]. Disjoint [2086, 199, 2325, 1576, 623, 2321, 2070, 2103, 911, 359, 910, 130, 1262, 848, 713, 599, 33, 1714]. disjunction [1181]. disjunctive [1699]. disk [1241, 134]. dispersal [756, 324]. Displaying [2355]. disprove [928]. dissociation [1092]. Dissolution [1461]. distance [1757, 887, 148, 607, 2046, 998, 2165, 1953, 2214, 2363, 957, 1694, 1063, 1551, 1284, 2114, 1945]. distances [1535, 1669]. distinct [752, 982, 466]. distinguishability [1856]. Distinguishing [225, 1351, 1273, 596, 917, 1148]. Distributed [244, 1920, 499, 1526, 15, 218, 656, 1872, 2098, 1155, 1327, 196, 947, 251, 1425, 18, 388, 466, 2055]. Distributing [2404]. Distribution [2162, 837, 1429, 417, 2187, 1249, 1738, 973, 620]. distributional [736]. distributively [947]. diversity [12]. Dividing [352]. division [564, 808]. divisor [643, 1304]. DNA [2046, 2190, 1223, 2269, 204, 1457, 833, 566]. DNA-based [833]. Do [2288]. DOAMI [2055]. Document [864, 2037, 1833]. does [580, 2106, 1379, 859]. DOL [1779]. Dollar [1699]. DOM [1961]. domain [262, 1373]. domain-theoretic [1373]. domains [78, 1961, 373, 498, 794, 673, 1204, 1372]. Dominance [1923]. dominating [570, 1685, 65, 1155, 1892, 2140, 52, 816, 2167, 1000, 359, 2000, 1473, 2286, 2296, 554, 22, 2396, 1263, 2010]. Domination [1263, 2347, 1599, 975, 1685, 2314, 2313, 2297, 2052, 1199, 421, 1991, 2363, 1009, 330, 1437, 147, 2390, 2019, 813]. Dominator [1391]. Domineering [340, 945]. Don [1777]. don’t [1542]. Double [239, 657, 1615, 420]. doubly [628]. down [701]. downscaling [1513]. dozen [1743]. DP [2361, 1737]. DP-degree [2361]. DPDAs [886]. dragons [962]. Drags [2101]. Drawing [1578, 1923, 866, 2348, 1028, 528]. drawings [1641, 2332, 2334, 900]. dreadful [962]. drift [2350, 1740]. driven [190, 92, 720, 2114, 2382, 1105]. drop [714]. drop-line [714]. dual [2218, 253, 2108, 37, 819]. Duality [1567]. Dualization [1121, 1120, 723]. duals [132]. due [229, 124]. duo [1023, 958]. duples [1800]. duplication [2287]. Duration [1325, 413]. durations [253]. Durfee [1952]. during [938]. duty [652]. Dyck [860, 601, 2113]. Dynamic [2066, 1804, 2370, 970, 38, 1194, 549, 2149, 769, 753, 548, 2134, 1665, 1831, 2039, 1611, 2232, 381, 244, 1936, 801, 1058, 2273, 288, 514, 235, 884, 355, 912, 14, 12, 904, 818, 218, 588, 2137, 1843]. dynamical [1286, 1200, 146, 1453, 1957, 928, 489]. dynamicity [2169]. dynamics [1670, 983, 2330, 1545, 428, 101]. dynamos [392].

E. [429]. EA [2193]. Early [193]. ease [2169]. easier [501]. easy [266, 949]. eccentricity [1394]. ecological [563]. Eden [1389, 561]. Edge [2070, 2190, 1723, 519, 910, 809, 150, 1273, 1928, 1209, 1762, 2170, 2038, 2173, 1631, 2325, 1672, 578, 1572, 468, 83, 1180, 518, 1899, 1691, 300, 471, 2022, 2416, 1509, 2362, 2395, 2217, 1157, 534, 2342, 807, 694]. edge- [1672]. edge-bipancyclicity [809, 807]. Edge-clique [519]. edge-colored [468, 2395]. edge-connectivity [1509, 1157, 1928, 694]. Edge-disjoint [2070, 910, 2325]. Edge-independent [1209]. edge-isoperimetric [1631]. edge-length [2022, 2416]. Edge-outer [2190]. Edge-pancyclicity [809]. edge-recoloring [468]. edge-weighted [578, 534]. edges [1160, 1262, 1950, 2228, 807, 1245, 615]. Edit [2114, 1535, 1063]. Editing [328, 1178, 1266]. Editor [1705]. Editorial [1687, 1695, 1704, 1711, 1726, 1745, 1758, 1765, 1654, 1667, 1671, 1675, 1682, 1966, 1845, 1598, 411, 425, 558, 761, 829, 835, 919, 1216, 1308, 1441, 1645, 1736, 1878, 1977, 2183, 242, 1025, 576, 7, 8, 264, 277, 284, 296, 308, 321, 333, 347, 365, 377, 391, 405, 410, 424, 433, 449, 459, 475, 484, 492, 504, 512, 524, 536, 555, 556, 21, 557, 69, 75, 82, 91, 98, 105, 113, 120, 128, 137, 144, 152, 158, 168, 177, 184, 194, 203, 214, 222, 232, 241, 255, 582, 583, 846, 856]. Editorial [865, 875, 878, 897, 905, 918, 927, 930, 933, 946, 956, 964, 974, 985, 988, 997, 1005, 1014, 1024, 1042, 1050, 1056, 1073, 584, 639, 640, 659, 665, 675, 684, 688, 699, 709, 715, 726, 731, 741, 751, 760, 773, 785, 790, 800, 811, 820, 828, 834, 651, 1081, 1094, 1292, 1307, 1326, 1330, 1336, 1343, 1346, 1355, 1371, 1377, 1383, 1387, 1395, 1402, 1407, 1415, 1426, 1432, 1440, 1462, 1470, 1477, 1487, 1115, 1116, 1129, 1139, 1147, 1154, 1161, 1164, 1177, 1184, 1190, 1196, 1202]. Editorial [1206, 1215, 1224, 1228, 1233, 1243, 1253, 1261, 1268, 1275, 1288, 1497, 1503, 1739, 1753, 1773, 1782, 1790, 1801, 1813, 1825, 1837, 1844, 1852, 1862, 1870, 1877, 1504, 1516, 1525, 1533, 1549, 1560, 1564, 1573, 1580, 1587, 1597, 1605, 1613, 1620, 1627, 1634, 1638, 1644, 1661, 1889, 1896, 2120, 2128, 2135, 2144, 2152, 2163, 2171, 2182, 2194, 2205, 2215, 2224, 2233, 2245, 2260, 2274, 2291, 2301, 2346, 2368, 2376, 2387, 2400, 1906, 1915, 1919, 1926, 1934, 1941, 1947, 1954, 1960, 1965, 1976, 1986, 1992, 1993, 2001]. Editorial [2008, 2014, 2025, 2034, 2044, 2054, 2068, 2077, 2088]. Editors’ [989, 732, 1588, 1780, 159]. education [1443]. effect [258, 806, 1423, 1185]. Effective [837]. effectiveness [1151]. Effectivity [1708]. effects [1718]. efficiency [1193, 369, 2116, 103, 1461, 1738].

Efficient [2196, 2234, 1399, 1069, 318, 339, 2253, 976, 344, 30, 1083, 963, 2223, 2221, 1058, 2273, 1267, 1998, 1547, 1575, 783, 1358, 532, 588, 1522, 1912, 693, 163, 607, 965, 2098, 1060, 998, 1664, 2238, 1542, 1899, 1009, 1063, 1840, 115, 2047, 380, 599, 1353, 1065]. Efficiently [855, 1302, 458]. egos [276]. election [1132, 2262, 749, 20]. elections [305, 1690, 1270, 1686, 383]. electorates [117]. electromagnetism [925]. electromagnetism-like [925]. electronic [547]. element [226]. elementary [920]. elements [1352, 2322, 107]. elicitation [1690]. eligibility [141]. elimination [1616]. ellipsoid [771]. elliptic [2407]. Embedded [1020, 2158]. Embedding [1031, 1994, 608, 2190, 1660, 300, 166]. embeddings [1922, 653, 1701, 1408, 1409]. embodied [838]. emergent [2189]. Employing [1545]. emptiness [1285]. encode [2092, 1219]. encoded [1061]. Encoding [609, 1544, 1511]. encodings [1033, 1959, 832]. encrypted [867]. Encryption [2049, 2030, 1642, 1192, 603, 204, 1550, 1007, 1998, 2166, 914, 1193, 1939, 2307, 818, 692, 588, 2304, 139, 1912]. encryptions [1197]. end [806, 1949]. end-markers [806]. End-Nim [1949]. endedness [2189]. endgame [340]. ends [1201]. Energy [2331, 1664, 320, 1398, 1414]. Energy-optimal [2331]. engine [262]. enhanced [1271, 1393, 2389, 26]. enough [564]. enriched [486]. Entangling [2053]. entities [502]. Entropy [1653, 1034, 1188, 1788, 1794]. enumerable [734, 674, 906]. Enumerating [2093, 816, 210, 1996]. Enumeration [1898, 2167, 415, 209, 1849, 2016, 677, 349]. Enumerations [1439]. enumerators [1439]. environment [1652, 1052]. environments [2039, 509, 12]. envy [680]. envy-freeness [680]. enzymatic [1647]. Epigenomics [2269]. Epileptic [1545]. Epsilon [416]. Epsilon-net [416]. epsilons [491]. Equal [1366, 141]. Equality [1768, 1480, 1212, 719, 2304, 1912]. equality-free [719]. equalizer [2238]. equation [2264]. equational [1747]. equations [1399, 2380, 2308, 79, 430, 1091, 729, 1523, 2265, 794, 234]. equicontinuous [762]. equidistant [1628]. equilibria [531, 852, 317, 1688]. equilibrium [1474]. equilogical [1635]. Equivalence [1198, 1728, 2146, 43, 138, 54, 216, 261, 674, 1451, 854, 1110, 1067, 1746, 456]. equivalences [2097]. equivalent [1339, 1961, 1248, 1163]. Era [1416]. ergodic [1431]. Erratum [2417, 1496]. Erratum/Correction [2417]. Error [407, 743, 379]. Error-pruning [407]. errors [2339, 801]. essential [1749]. Establishing [867]. Estimating [1211, 1798, 1596]. estimation [837, 417, 775, 2187, 740, 380]. estimation-of-distribution [2187]. estimators [955, 1367]. Eternally [2296]. eukaryotic [2269]. Euler [1899, 1297]. Evacuating [1528]. evacuation [2340]. evaluate [965]. evaluation [523, 1019, 2356, 942]. evasion [2143, 1046]. evasive [287]. even [1247, 547, 1365, 874]. even-sized [1365]. event [522, 2056, 1808]. events [801]. eventual [215]. Every [300]. evidence [1880]. Evolution [2186, 838, 771, 1628]. Evolutionary [9, 1311, 924, 10]. Evolving [1317, 2184]. Exact [1012, 34, 1706, 1786, 1256, 1589, 1092, 757, 2155, 1820, 1364, 552, 147, 1410]. exactness [74]. examples [1234, 736]. except [2254]. Exception [1775]. exchange [301, 35, 1213, 874, 26]. exchanged [1406, 1496]. exclusion [1529, 1874]. Exclusive [1274]. Excuse [272]. execution [2370, 2180, 730]. executions [1327]. exhaustive [1170]. existence [1577, 672, 1957, 50]. Existential [2084, 1523]. exists [2140]. exits [1528]. expandable [1075]. expanded [1698, 2417]. expanding [498, 2343]. Expansion [179, 226]. Expansion-based [179]. expansions [1621]. Expected [827, 2189, 249, 1203, 996]. exploited [969]. Exploiting [240]. exploration [2331, 1045, 941]. Explore [502]. Exploring [655]. exponent [1069, 1098, 1159, 2336, 2265]. exponential [600, 1256, 2031, 1820, 336]. exponential/FPT [600]. exponents [2117]. exposure [1952]. expressed [727, 233]. expressibility [1348]. expressible [2265]. expression [1281]. expressions [2209, 2219, 805, 1956]. expressive [427, 793, 2374, 80]. Expressiveness [788]. expressivity [312, 1772]. extend [1689]. Extended [723, 595, 1473, 2069, 2150, 1266, 1376]. Extending [175, 615, 1943, 2343]. extension [2228]. extensions [479, 889, 25, 32, 2015, 1372]. extensive [670]. externalities [1821]. Extra [2305, 1366, 952, 804, 1509, 1988, 717, 1482, 2174, 2237, 1924, 2389, 2051, 361, 694]. extremal [422]. Extreme [994].

F [2024]. facets [1599]. facilitated [150]. Facility [1076, 1903, 1718, 868, 37, 2067]. factor [135, 529, 314, 1074]. factor- [529]. factorisations [719]. Factorization [817, 2263, 84, 689, 1066, 870, 2115]. factorizations [2074, 2235]. factors [1098, 1793, 1068, 2271, 2257, 1290, 1301]. failed [1146]. failures [320, 541]. Fair [2248, 544, 28, 35, 26]. fairly [352]. falling [810]. false [872]. false-twin-free [872]. families [2093, 2404, 2349, 1500, 2407, 2393, 1320, 1403]. family [1631, 86, 123, 2315, 1721]. Fan [1641, 303]. Fan-planarity [303]. Farey [2351]. Fast [2155, 270, 1360, 1071, 689, 1087, 575, 888, 18, 2158, 2258, 1872, 2370, 1876, 1593, 2297, 775, 1860, 1574, 2021, 1419, 1064]. Faster [1066, 1551, 907, 2239, 200, 2137]. Fault [1479, 212, 2087, 1509, 227, 2241, 93, 1672, 27, 1361, 300, 1484, 1948, 2356, 319, 1347, 1924, 2320, 1553, 2354, 1928, 1942, 807, 1994, 1401, 2170, 2392, 2130, 618, 1927]. fault-free [1994]. Fault-tolerant [1479, 1509, 93, 1672, 27, 1361, 319, 2354, 807, 1401]. Faults [1929, 562, 814, 1970, 471, 2289, 2284]. faulty [579, 623, 2321, 2266, 1900, 809, 2311, 1262, 58, 807, 1994]. Feasibility [169]. feature [885]. features [937]. feedback [948, 1880, 173, 633, 2028]. few [299, 2173, 291, 2404, 1510]. fewest [1848]. Fiat [1767]. Fibonacci [2074, 883, 1200, 1077, 2327, 357]. Fibonacci-automatic [1077]. Field [817, 2236, 841]. fields [837, 1091]. FIFO [2309]. filling [110, 176]. filtering [1808]. find [1256, 580, 1657]. Finding [1183, 1352, 42, 873, 906, 758, 2011, 199, 1312, 1679, 236, 210, 2300, 1480, 870, 1302, 1714, 310, 1537, 2310, 517, 2236, 2103, 1438, 338, 1400, 1365, 1182, 631, 2176, 1299, 2161]. finds [13]. finishers [1808]. Finite [727, 1446, 190, 1277, 2364, 2383, 1169, 1476, 814, 481, 1171, 1316, 2236, 1748, 1175, 78, 1279, 983, 1278, 200, 2385, 1091, 2381, 1957, 1321, 2213, 765, 329, 106, 649, 669, 1442, 331, 1189]. Finite-state [727, 1277]. finite-valued [481]. finitely [1693, 1151, 1498, 2375]. fire [363]. firefighter [1257, 678]. Firefighting [2299, 2157]. firing [2283]. First [404, 2350, 789, 479, 2156, 72, 1053]. First-hitting [2350]. first-order [789, 479, 1053]. fission [564]. Fit [404, 921]. fit-preserving [921]. fitness [2048]. five [2093, 1194]. Fives [1847]. Fixed [110, 173, 1791, 233, 1352, 2334, 1198, 1698, 2417, 1930, 154, 1601, 1703, 2249, 2126, 146, 1137, 387, 1832, 2020, 155, 1242]. fixed-density [2249]. fixed-mobile [2334]. Fixed-parameter [110, 173, 1930, 387, 1832]. fixed-point [1198, 155]. Fixing [1555]. FL [821]. Flat [766, 1744, 1728, 1649]. Flexible [1975, 1884]. Flip [1792]. flipping [589]. flood [176]. flood-filling [176]. floorplans [542]. flow [2091, 627, 1150, 822, 229, 1715]. flow-shops [1715]. flow-time [627]. flowchart [662]. flows [202, 243, 196, 2065]. Flowshop [1214, 1354, 1090]. flowshops [2202, 2085]. fly [728]. fly-automata [728]. FM [895, 1548]. FM-index [895, 1548]. focusing [1386, 1340]. folded [171, 1672, 300, 471, 809, 2311, 1482, 1553, 2241, 361, 694]. folding [1921]. Follow [1784]. fonts [273]. food [580]. Forbidden [1158, 2271, 916, 233]. forcing [451, 1724, 1146, 2200]. Forest [1518, 2218, 55, 2069]. forests [1677, 2294, 1994]. Foreword [1117, 1771, 1057, 1416, 1705, 989, 732, 159, 1780]. Forewords [1043]. form [727, 629, 290]. Formal [1341, 78, 1864, 641, 2377, 332, 1963, 2100, 2160]. formalisation [2146]. formalism [1918]. Formalization [1499, 478, 1460]. Formalized [2147]. formation [1831, 1600, 1603, 317]. formations [2413]. forming [1322]. forms [1227, 1130, 1914]. formula [969]. formulae [2090]. formulas [1123, 2236, 1754, 2219, 1662, 1144, 1589, 1523, 1519, 1615, 1128, 1118, 663, 2064]. formulations [1448]. Forward [876]. Foundations [826, 1344, 932, 2007]. four [1921, 1578, 1063]. Four-Russians [1063]. Fourier [1034]. FPT [600, 1677, 1085, 2019]. FPTAS [1354, 1090]. FPTRAS [1721]. fractals [294]. fractional [658, 1499]. fragment [406, 645]. fragmentable [465]. fragments [1974, 482, 2374]. Fraïssé [1750]. frame [1079]. Framework [1274, 1338, 707, 1471, 1822, 384, 2101, 823, 1422, 1805, 348, 2335, 102, 412, 1637, 1881]. Frameworks [1337, 505, 2145, 543]. Fréchet [2165]. Free [442, 1096, 1352, 1260, 1069, 1168, 311, 208, 1735, 43, 138, 1830, 909, 1674, 2208, 1530, 786, 95, 1898, 872, 67, 454, 2023, 1327, 646, 647, 1554, 1282, 464, 1249, 2217, 1592, 1425, 1285, 1815, 904, 719, 692, 1263, 1624, 736, 458, 2210, 1378, 1350, 676, 1994, 126, 2392]. freeness [680]. frequencies [12]. frequency [1334, 981]. frequent [1, 984, 23]. friend [1166]. friendly [2407, 1197]. Frobenius [1701]. frontier [1823, 2113, 1738]. Frontiers [525, 482, 2195]. Frugal [1255]. fuel [553]. fuel-considered [553]. Full [2092, 1192, 2393, 2169]. full-hiding [1192]. Fully [2030, 1936, 1142, 1335, 947, 1060, 1218, 949, 251, 117, 388]. fully-functional [1060]. Fun [273, 1845, 265]. functigraphs [2396]. function [2146, 2402, 523, 2391, 2, 1019, 2405, 1699, 1490, 2166, 1036, 1027, 725]. function-private [2166]. Functional [1348, 1060, 78, 633]. functions [1286, 495, 257, 1623, 1348, 1476, 1102, 345, 1760, 10, 154, 2017, 127, 200, 2204, 1812, 2110, 1219, 2005, 1244, 1134, 550, 1827, 2256, 942, 369, 155, 1367, 1114, 874, 1485, 2153, 1752, 1376, 118]. Functorial [1053]. functors [1964]. fundamental [1119]. Fundamentals [309, 662]. Further [1257, 1455, 2382, 2398]. fusion [1882]. Fuzzy [1460, 1669].

G [1772]. Gabriel [397]. Galaxies [268]. Game [1658, 2306, 2328, 2418, 41, 2277, 1828, 148, 66, 1044, 71, 975, 1655, 1500, 28, 252, 1731, 531, 1718, 1761, 2399, 803, 808, 340, 293, 2121]. game-based [71]. Games [591, 1195, 934, 2333, 275, 980, 959, 1916, 1828, 753, 1831, 935, 270, 71, 59, 687, 1585, 2107, 624, 1699, 176, 983, 827, 1729, 1778, 231, 1732, 211, 1267, 1603, 1291, 1902, 114, 467, 670, 1131, 1949, 1252, 389, 1971, 172, 206]. Gap [1814, 1259, 1556, 1640, 890, 803]. Gap-planar [1814]. gapped [1640, 1890]. gaps [299, 223, 2031, 867, 1548]. Garden [1389]. gardens [561]. gasket [1624]. Gate [1616, 1446]. gates [1822]. Gathering [654, 1970, 53, 710, 1489]. Gaussian [996]. gene [562, 374]. General [467, 107, 817, 1757, 1895, 707, 148, 1373, 394, 1471, 1446, 1925, 2156, 1400, 1396, 617, 2069, 1127, 205, 2127, 2352]. generalised [1635, 79, 1455]. Generalization [1594, 1466, 211, 456]. generalizations [66, 1283, 1725, 455]. Generalized [2148, 1652, 1887, 195, 1201, 1957, 1840, 1067, 329, 683, 2336, 1481, 1483, 481, 2141, 1892, 2102, 2139, 2363, 1914, 336, 1650, 1511, 1772, 2259, 2003, 973, 1187, 1093]. generated [1151, 1315, 2275, 2413]. Generating [375, 458, 1446, 1102, 1405, 1808, 1649, 874]. generation [1170, 1191, 1792, 341, 1565, 2264]. generator [349]. generator-enumeration [349]. generators [972, 676, 118]. Generic [486, 364, 823, 178, 2072, 2035]. generics [1779]. genes [1065]. Genetic [9, 494, 427, 1609, 104]. genetically [1984]. genome [614]. genomes [2269, 529]. genomics [1458]. genus [1920]. geodesic [606, 503, 2247]. geodetic [909, 1816]. Geographic [742]. Geometric [2333, 1944, 782, 2123, 1584, 1713, 1218, 221, 1996, 1659, 2010]. Geometry [774, 1880, 1176]. Giancarlo [1416]. Girard [661]. girth [1194, 403, 2043]. Give [2253]. Given [328, 1121, 2402, 2140, 777, 1178, 382, 1248, 1064, 1303, 2041]. glasses [2217]. glasses-free [2217]. Global [223, 1404, 1996, 1047]. glue [1219]. gluing [1579]. Gong [914]. Good [528, 873, 906, 1925, 1913, 2305, 1238, 1413, 2051, 1412, 1133, 799]. good-neighbor [2305]. goods [220, 306]. Graceful [1142]. Gracefully [1665]. graded [1750]. grammar [2275, 343, 708, 844]. grammar-based [343]. grammars [1310, 332, 1618, 1313, 1345, 1735, 197, 1674, 2379, 1179, 235, 2112, 366, 290, 1442, 650, 904, 736]. Grammatical [1836]. Graph [2292, 687, 1043, 328, 1178, 1266, 1397, 1338, 1829, 213, 655, 45, 1869, 1101, 913, 1873, 1448, 1345, 438, 5, 1329, 1328, 65, 2101, 2190, 591, 1318, 1045, 678, 1567, 1083, 1451, 2026, 363, 421, 2394, 658, 1763, 1543, 528, 1673, 2305, 2341, 2180, 776, 1009, 235, 1520, 366, 1032, 2302, 1835, 1157, 1380, 681, 833, 1717, 1263, 2153, 2357, 1038, 1342, 1435, 1427, 2378, 1720, 620, 2130, 1697, 2363]. graph-based [1338, 421, 1673]. graph-bin [913]. graph-controlled [1318]. Graphical [205, 1774]. graphs [1254, 2306, 2230, 1507, 2337, 1641, 2173, 97, 2029, 132, 414, 1814, 1859, 2251, 1895, 41, 2347, 2081, 1828, 2154, 1631, 163, 148, 1012, 397, 1044, 439, 270, 703, 108, 866, 1685, 1447, 1944, 208, 4, 474, 1611, 1967, 1817, 312, 353, 604, 316, 696, 1492, 2313, 437, 244, 1655, 2156, 2070, 135, 1936, 1155, 1920, 1748, 400, 502, 2388, 2348, 1003, 2297, 869, 1578, 1572, 682, 1830, 2057, 195, 862, 909, 136, 978, 953, 1724, 724, 1199, 1422, 95, 1555, 52, 1194, 1421, 162, 1400, 816, 911, 1898, 2167, 744, 1368, 1614, 625, 872, 611]. graphs [1366, 1953, 403, 518, 952, 617, 804, 40, 1991, 815, 292, 1660, 791, 472, 2040, 1391, 1473, 2246, 951, 1478, 1957, 1505, 745, 1365, 2069, 175, 1770, 2143, 2021, 499, 1127, 130, 1937, 2022, 2416, 112, 464, 977, 2020, 2043, 1046, 1488, 1911, 1160, 1238, 1716, 2289, 330, 717, 1250, 122, 2016, 1566, 1510, 1696, 2259, 1816, 2395, 2063, 847, 1369, 2132, 2124, 1551, 677, 532, 147, 848, 713, 2086, 521, 1486, 596, 2359, 2390, 1815, 798, 389, 1615, 346, 917, 1622, 2019, 573, 973, 134, 1148, 2351, 1910, 1938, 2200, 1001, 1942, 1135, 1133, 2041, 271, 1676, 813, 1015, 33, 615, 96, 1762]. graphs [2170, 2392, 2010]. Gray [702, 341, 1201, 1579, 1849]. greatest [729, 1304]. Greedy [2332, 1826, 653, 463, 37, 2079]. Green [202]. grid [249, 543, 745, 1365, 1951, 1015]. grid-like [1015]. grids [53, 654, 2296, 634, 2293, 85]. Gröbner [149]. Gromov [1370]. Group [2266, 1529, 1274, 337, 436, 1990, 334, 2221, 1673, 2305, 2322, 1939, 1683, 2169, 349, 33]. groups [1352, 1151, 199, 2290, 1785, 1864, 130, 1514]. Grover [2053]. Growth [621, 429]. Grundy [2391, 1699]. guaranteed [2037]. guarantees [381, 48]. guard [1153]. guarded [1635, 72]. guarding [387, 1153]. guards [2071, 652]. Guest [651, 576, 242, 989, 1588, 732, 1780, 159, 1705]. guide [268]. guided [1449].

Hadamard [1172, 2209, 2123, 2411]. half [1134]. half-product [1134]. halfspaces [992]. Halpern [2374]. halting [1141]. Hamiltonian [472, 745, 1365, 1950, 1689, 1994]. Hamiltonicity [58]. Hamming [1535, 887]. Hanabi [1247]. Hankel [25, 2315]. Hanoi [1473, 1488, 620]. hard [275, 1247, 758, 381, 1702, 2125, 2248, 1182, 1847, 1721, 1851]. harder [396]. Hardness [266, 1921, 1213, 147, 698, 1492, 176, 2309, 3, 68, 1832, 434, 2286, 230, 632, 1858, 1397]. Harmonious [2042]. Hash [2405, 2404, 1036, 2304]. having [4, 474, 724, 2250, 2043]. head [1321]. healing [1527]. heap [823]. height [1021, 1907]. heights [635]. help [2153]. Herbrand [789, 2004]. Hereditary [873]. Hermite [1914]. Heterogeneous [1162, 868, 578, 1603, 1630]. heuristic [962]. heuristics [1903, 2052]. hidden [1590, 2193, 1193, 1637, 2319]. Hider [2295]. hiding [1192]. Hierarchical [1593, 724, 1010, 841, 1948, 482, 2237, 520, 692, 2241, 618, 1468]. hierarchies [954, 963]. Hierarchy [1748, 392, 1283]. High [1361, 567, 1752]. Higher [1320, 397, 1198, 2005, 825, 787, 1561, 366]. higher-dimensional [825]. Higher-level [1320]. higher-order [397, 1198, 2005, 787, 1561, 366]. highly [2039, 740, 1767]. hijacking [445]. Hilbert [1882]. hip [1800]. HiperTM [1361]. history [692, 489]. history-dependent [489]. history-free [692]. hit [2013, 2415]. hitting [2350, 109]. Hoare [1984, 487, 667]. HOAS [509]. hoc [666, 254, 545, 656]. HOL4 [1499]. Holant [1822]. hold [448]. hole [1365]. holes [502, 543, 219]. holonomic [1399, 1102]. Holub [461]. homogeneous [1840]. homomorphic [2030, 850, 1701, 603, 914]. homomorphism [403]. homomorphisms [316]. homophyly [358]. homotopy [1295]. honest [1706]. Honor [1416, 879, 2261]. Honoring [1534]. hop [2138, 691, 2262]. Horn [1123, 2108, 1128, 169, 1118]. host [579, 148]. Howard [789]. HS [1974]. HTN [1962]. HTN-like [1962]. hub [122]. hull [1466, 909, 2247]. human [567, 1443]. humans [2118]. Hybrid [2374, 1942, 2392, 1927, 281, 1010, 763, 1344, 831, 2356, 1842, 928, 1809, 153, 375]. Hydras [1127, 1128]. Hyper [191]. Hyper-optimization [191]. hyperbolic [903]. hyperbolicity [1370]. Hypercube [1581, 390, 2284, 1008, 848, 713]. hypercube-based [2284]. hypercube-like [848, 713]. hypercube-variant [1008]. hypercubes [1672, 2321, 300, 471, 809, 2311, 1509, 2358, 853, 1482, 1950, 1393, 6, 1004, 1553, 2389, 2342, 807, 361, 694, 915, 1927]. hyperexcitable [842]. hypergraph [1590, 1120, 1312, 1883, 2175]. hypergraphs [1119, 2391, 402, 1996, 689, 211, 287, 631, 2361, 346, 1128]. Hyperopic [2298]. Hypersequent [1055]. hypervolume [11]. hypothesis [2378].

I/O [163]. IC [866]. IC-planar [866]. ICDCN [1526]. Ideal [1022, 1575]. ideas [1169]. Identical [1633, 440, 238, 755, 530, 1249, 2132, 1715]. Identification [109, 1199, 614, 283]. identifiers [658]. identifying [791, 2000, 1510]. identity [1007, 1998, 692, 588, 2035]. identity-based [1007, 1998, 692, 588, 2035]. II [906, 1077, 1163]. III [2363]. Iliopoulos [1534]. image [1457, 595]. Imagery [774]. images [971, 568, 1557]. IMDPs [1733]. immobile [2295]. immunity [1752]. immutability [1975]. impact [1900, 801]. impartial [114]. implement [1586]. Implementation [185, 2206]. implementations [722, 1979, 1981]. implementing [2199]. implication [1797]. implications [1118]. Implicit [256]. imply [1379]. importance [129, 986]. imprimitivity [461]. improper [1555]. Improved [579, 548, 440, 2326, 1088, 2236, 1155, 547, 2229, 552, 1493, 124, 1085, 319, 290, 494, 386, 1410, 1890, 2067, 2064, 1030, 301, 1210, 447, 2335, 2201, 941, 925, 118]. improvement [1145, 939]. Improvements [750, 1457]. Improving [59, 718]. In-place [1364]. Inapproximability [438, 52, 1581]. incentive [1824]. Incidence [6, 1004]. inclined [2202]. including [1439]. Inclusion [2113, 392, 1730]. incomplete [1639, 2372, 1171, 2365, 1690]. increase [916]. increasing [1907]. Incremental [1235, 2002, 262]. Independence [1624, 1693, 482]. Independent [22, 1191, 2203, 1756, 1149, 292, 1063, 1437, 1008, 2176, 1624, 1209, 1596]. Indeterminate [437, 1420, 1543]. Index [2219, 1351, 1542, 894, 895, 1548, 1273, 895]. indexes [364, 1610]. Indexing [971, 1571, 892, 230, 1958, 1061]. indicator [11]. individual [538]. indivisible [306]. Induced [501, 911, 1411, 517, 869, 295, 782, 1626, 674, 73, 1490, 463, 1135, 40]. Inducing [1271]. induction [1591]. inductive [735, 990]. Inefficiency [531, 1902]. Inequality [2002, 1207, 1610, 1074]. infection [606]. inference [302, 735, 990]. Inferring [1286, 430]. infinitary [2150]. Infinite [1208, 1403, 2272, 972, 1631, 1312, 2398, 1748, 2327, 2023, 2235, 1046, 2117, 456, 378]. Influence [246, 269, 1968, 1969]. influences [1831]. Information [651, 576, 242, 476, 2091, 409, 1404, 822, 1876, 1666, 2308, 1961, 673, 840]. informational [510, 1458]. inhibitor [1668]. inhibitors [1444]. inhomogeneous [1944]. inner [1192, 2166]. Input [720, 236, 1827, 190, 92, 2114, 2382, 1105]. Input-driven [720, 190, 92, 2114, 2382, 1105]. inputs [667]. Insecurity [1379]. insertion [1449, 2380, 1318]. insertion-deletion [1318]. insertions [1149]. insider [1998]. insights [1846, 924]. inspired [1881]. instance [348]. instance-dependent [348]. instances [851, 10, 3, 68, 1410]. instantiable [1767]. Institutions [1774]. Integer [1011, 817, 778, 236, 1735, 2397, 200, 1865, 949]. integer-manipulating [1865]. integers [2030, 1701, 729]. integral [1294, 533]. integrality [803]. Integrated [61, 140, 74]. integrated-time [74]. intellectual [2402]. Intensional [2005, 406]. Intensity [526]. Intensity-Modulated [526]. interaction [660]. Interactions [278, 1963, 1900]. Interactive [106, 936, 348, 947]. interchange [607]. interchanges [1140]. interconnected [429]. interconnection [1925, 903, 1928]. Interface [1385, 407]. Interfaces [929]. interference [1501, 53, 1362, 119, 2055]. interiors [648]. Interleaving [1746]. intermediate [1088, 219]. Internal [1632, 1716]. International [2216, 732, 2225]. interpolation [72]. interpretation [258, 181]. interpretations [259]. Intersection [880, 1749, 1297, 2010]. intersections [965, 210, 2179]. Interval [1668, 1469, 1733, 976, 108, 1974, 351, 2313, 1936, 2297, 81, 591, 1199, 38, 40, 1716, 681, 2086, 1435, 1841]. intervals [2222, 630, 1524]. intransitive [1501].

Intrinsic [2080]. introducing [2093].

Introduction [1119, 1588, 1502]. Intruder [1227]. intuitionistic [77]. invariant [845, 1294, 149, 1287, 1840]. invariants [375]. Inverse [2038, 2312, 1507, 916]. inverses [1227, 1806, 2178]. inversion [2093, 642, 335, 969, 1065]. inversions [167]. Investigating [769]. invisible [2228]. involution [2100]. involving [2403, 1380]. Invulnerability [1995]. ion [428]. IoT [1973]. Irreducible [1186, 1498, 1072]. irreducibles [1121]. irredundant [1898]. irregular [1614]. irreversible [392]. isometric [1685, 39, 1403]. isomorphic [2191]. isomorphism [1411, 313, 1448, 501, 1451, 349, 40]. isoperimetric [1631]. isotactics [1746]. Issue [1879, 1588, 505, 1897, 2216, 1309, 1331, 1526, 1217, 15, 1416, 2225, 2292, 285, 651, 879, 626, 2206, 2261, 1043, 732, 1863, 242, 1727, 2369, 1705]. issues [1759, 1134]. Item [1272]. items [1, 984, 23, 465, 142]. iterated [135, 1778, 2184]. iterating [1167]. iteration [1733, 1520, 46]. iterations [1299]. iterative [1036]. iterativity [1964].

jamming [1381]. jigsaw [266]. job [1006, 704, 237, 48, 1363]. jobs [1904, 440, 1392, 1664, 123, 229, 207, 2158, 1435, 2202, 2085]. joining [2321, 2086]. joy [1118]. Jumbled [1061]. jump [281]. Jumping [1048, 2383, 1279]. juntas [1786]. justifications [483].

Kannan [750]. Kautz [1347]. Kayles [34]. keeping [1856]. Kelly [1158]. Kelly-width [1158]. Kernel [1651, 948, 1085, 2201, 2303]. Kernelization [706, 535, 1684, 1764, 1720, 1084]. Kernels [2244, 45, 1135]. key [2307, 818, 1035, 1952, 26, 2304, 139, 1912]. keys [201]. keyword [1998]. KHL [460]. kidney [301]. Killing [491]. Kinetic [902, 1459]. Kleene [1708, 2406]. Knapsack [2076, 627, 87, 949, 49, 598, 2204, 1430, 1414, 2399, 386]. Know [1602]. knowledge [1939]. known [2295]. Kobayashi [1251]. Kochen [2411]. Kolmogorov [550]. König [2362]. Kripke [157, 1561]. Kuhn [718].

L [2024, 953, 745]. L-shaped [745]. L-shapes [953]. label [491, 981]. Labeled [978, 632, 44, 271]. labeling [2232, 912]. labelled [1869, 1943]. labels [1875]. laconic [1439]. Lagrangian [1609]. lambda [1376, 1341, 192]. Lambda-confluence [192]. landscape [2310]. Language [1879, 1124, 2380, 453, 476, 1649, 752, 1495, 781, 1982, 1819, 1772, 2113, 1277, 307, 1700, 1376]. Language-theoretic [453]. Languages [1290, 1309, 1809, 2413, 1310, 993, 1260, 1100, 1021, 2207, 2096, 1283, 1102, 605, 1315, 1316, 43, 138, 1029, 2100, 1107, 1701, 2208, 823, 786, 734, 963, 2214, 2133, 2385, 497, 646, 1319, 1785, 2212, 1819, 2110, 2393, 991, 189, 647, 1864, 1650, 648, 335, 2410, 1282, 1561, 1112, 1212, 1322, 1559, 2160, 720, 1285, 1105, 1022, 1621, 2210, 1378, 2318, 1465, 676, 662, 566]. Laplacian [2175, 1080]. Laplacians [1567]. Large [1757, 758, 1811, 2073, 1609, 210, 1626, 403, 2021, 2296, 2043]. larger [1544, 1033]. larger/smaller [1033]. largest [310, 1204]. latency [119]. lateness [123]. Latin [1581]. Lattice [850, 2169, 1428, 778, 2004, 1490, 1939, 1683, 423]. Lattice-based [850, 2169, 1939, 1683]. lattice-valued [2004]. Lattices [1797, 1121, 1126, 883, 669]. law [52, 973]. laws [1446]. lazy [976]. LBAs [886]. LCP [1072]. LCS [1891]. LCSk [881]. Leader [1132, 2262, 749, 1900, 20, 1784]. Leaf [1696, 798]. Leakage [1935, 1827, 409, 1760, 1142, 1550, 2307, 139, 2035]. Leakage-resilient [1935, 1760, 1142, 2307, 2035]. learnability [736]. Learning [992, 993, 1123, 1785, 732, 431, 1590, 845, 1594, 1706, 1786, 738, 734, 938, 1589, 991, 2080, 1710, 571, 1592, 1787, 1596, 1705]. learning-theoretic [1710]. Least [729, 843, 624, 724, 1194, 1087]. least-core [624]. ledger [2098]. Left [647]. leftmost [1839, 870]. lemma [2278, 2412, 1070, 1180]. Lemmings [274]. Lempel [1571, 2235]. length [477, 2038, 2142, 1061, 249, 395, 2325, 777, 2022, 2416, 350, 1583, 2084, 1031, 955, 1367, 1242, 981, 1353]. length-weighted [350, 1583]. Leontief [1474]. Lessons [581]. letter [1290]. level [129, 1320, 37]. levels [1379]. leverage [1087]. Lexicalization [1674]. lexicographical [341]. liar [147]. libraries [1563]. lies [300]. light [1909, 2049]. light-weight [1909]. lights [597, 211]. lights-out [211]. like [1536, 481, 2126, 467, 705, 848, 713, 923, 1931, 925, 772, 1015, 1962]. Limit [762, 205]. limited [1019, 195, 2384, 1144, 1324]. line [2230, 2263, 500, 1513, 116, 1471, 2181, 902, 1558, 1769, 2197, 1718, 1524, 2013, 2415, 714, 2055, 500]. Linear [304, 1618, 1062, 614, 1149, 896, 19, 444, 2384, 224, 461, 2316, 1740, 2013, 2415, 1208, 406, 260, 1967, 2147, 1026, 138, 54, 900, 261, 10, 2018, 2193, 1595, 2026, 911, 1529, 815, 64, 795, 207, 1659, 1565, 951, 1365, 2112, 1298, 2201, 330, 1250, 720, 1340, 153, 796, 375, 1700, 1843, 819, 2067, 2064, 2378, 1994, 585]. Linear-size [896, 1026]. Linear-space [1618, 2018, 64]. Linear-time [1062, 614, 444, 2384, 461, 1365, 330]. linear/submodular [819]. Linearizing [477]. linearly [850]. linguistics [1887]. link [199, 371, 602]. links [1969]. LinUCB [943]. lion [1761]. Lipschitz [79]. Lipschitzian [1167]. Lipschitzian-type [1167]. list [2154, 1763]. list-colorings [2154]. listing [2037]. lists [1639, 313, 1140]. liveness [928]. Load [753, 520]. loading [420]. Local [522, 1530, 1180, 1109, 1286, 223, 1155, 643, 1875, 658, 498, 325, 1362, 1683, 955, 1367, 2412, 1380, 1741, 13, 1089, 1245, 1783]. localised [1719]. localities [1454]. Locality [748, 393, 1409]. Locality-preserving [393]. Localization [1783]. Locally [316, 1626, 1510, 1227, 2154, 1967, 2349, 849, 695, 1105, 1413]. Locating [1436, 2396, 2000, 2390]. Locating-dominating [2396, 2000]. locating-total [2390]. Location [1076, 1507, 2038, 1903, 1199, 288, 514, 1236, 1911, 1718, 868, 37, 2067, 540]. location-domination [1199]. Löf [286, 1017]. log [1140]. log-Lists [1140]. Logarithmic [1513, 263]. logic [824, 686, 789, 1803, 1984, 406, 480, 1974, 1198, 728, 877, 72, 510, 81, 629, 2219, 1344, 645, 2375, 1806, 374, 1561, 1807, 1340, 669, 178, 932, 483, 1163, 412, 107, 2060, 1374, 2003, 431, 2374, 1843, 667, 637, 2061, 1384, 1804, 476]. Logical [2149, 1337, 505, 2145, 2127, 1467, 1972, 1100, 1464, 670, 2024]. logically [1339]. logics [1868, 481, 2147, 2004, 370, 1344, 1055, 1774, 371]. long [869]. Longest [889, 1537, 2365, 1522, 1353]. Look [701, 1247, 5, 1329, 1328, 2366]. Look-ahead [701]. lookahead [2083]. lookup [1063]. loop [1530]. Loopless [1849]. loosely [664]. lossy [1760]. Lost [1741, 1334]. Lovász [1180, 2412]. Low [221, 653, 738, 1770, 748, 12]. low-stretch [653]. Lower [869, 1045, 765, 851, 2134, 653, 225, 2366, 1616, 872, 1710, 447, 1493, 1959, 1524]. Lowness [1464]. LSFA [505]. Łukasiewicz [1562]. Lunch [442]. LWE [2329]. Lyndon [1068, 1539, 2102, 2249, 1066, 2270]. LZ78 [1536, 1066]. LZ78-like [1536].

M [2024]. machine [125, 1229, 680, 2012, 61, 1820, 123, 2393, 229, 207, 704, 960, 2281, 124, 530, 237, 1475, 531, 140, 1633, 2357, 2158, 141, 1363]. machines [432, 478, 959, 788, 1006, 238, 755, 1392, 419, 441, 962, 806, 229, 704, 2213, 1769, 57, 2132, 1173, 412, 764, 107, 714, 2127, 1435, 206, 1468, 2118]. MacPherson [1297]. macro [932]. macro-step-based [932]. made [949]. magic [547]. Magnifying [867]. Mainly [1118]. maintaining [23]. maintenance [1904]. majority [522, 970, 593]. make [396, 63]. makespan [123, 140, 1715]. malicious [634]. malleable [1940]. mammalian [844]. man [1761]. management [381, 298, 1246, 1643]. Manhattan [1546]. manifolds [738]. manipulating [1865]. manipulation [706, 1666, 1884, 757]. Manipulative [1690]. many [1599, 2033, 580, 1519, 848, 713, 2293, 718]. many-to-many [848, 713]. Map [2043, 990, 912, 685]. mapping [1023, 958]. mappings [1167, 73]. MapReduce [581]. maps [702, 2236, 79]. Marcus [1442]. Mark [808]. marked [1869]. markers [806]. Market [1474]. Markov [1841, 837, 151, 1289, 1614, 994, 1574, 1431, 380, 663]. Markovian [74]. MARLEDA [837]. Marriage [2075]. Martin [286, 1017]. Martin-Löf [1017]. martingales [2160]. mass [414, 921]. mass-action [921]. Mastermind [396]. match [63]. matches [1438]. Matching [171, 1349, 880, 1910, 1938, 1639, 299, 339, 594, 2198, 2096, 882, 1663, 696, 1538, 2156, 1905, 624, 893, 1078, 2273, 861, 1552, 188, 1893, 1267, 885, 894, 1182, 2020, 1250, 1721, 1566, 1067, 2259, 2015, 1397, 534, 1237, 1584, 610]. Matchings [397, 202, 1428, 1016, 295, 1626, 1999, 1207, 1473, 1249, 748, 463, 1689, 126]. Mathematical [567, 1185]. matrices [1099, 782, 1914, 2123, 2411, 1908, 1300, 228, 1796, 1619]. matrix [1310, 1021, 526, 590, 453, 2341, 1939, 1032, 1299, 1305, 969]. matrix-merging [526]. matrix-vector [1939]. matroid [916]. matroids [2391, 306, 1901, 423]. matter [1222, 1874]. matters [11]. Max [1930, 148, 682, 2015, 1485, 1030, 681, 1410, 1089, 1858]. Max-Batching [681]. max-distance [148]. max-sat [2015]. Maximal [2256, 2285, 1436, 891, 1069, 1098, 1967, 2156, 1149, 1003, 1438, 777, 1999, 849, 1898, 2201, 2016, 677, 723, 766, 2265, 1890, 602]. maximal-exponent [1098]. maximally [1928]. maximin [1901]. maximization [812, 1156, 384, 1969, 1079, 1134, 1388]. maximizing [11]. Maximum [295, 1473, 982, 1249, 748, 1183, 606, 1023, 958, 55, 238, 755, 1756, 1657, 123, 609, 2167, 2221, 2367, 292, 1659, 957, 1770, 1716, 1517, 463, 798, 1624, 1680, 714, 534, 1092]. maximum-box [1680]. maximum-duo [1023, 958]. maximum-weight [1756]. MaxSAT [1210, 2064]. may [1047]. MCTS [939]. MCTS-based [939]. MDPs [1733]. MDS [1920]. me [2253, 272]. Mealy [1506]. Mean [1732, 841, 1252, 573]. Mean-payoff [1732]. Means [289, 851, 1299]. measure [881, 2073, 336]. measurements [2253]. measures [2091, 887, 297, 1017, 799, 694]. Measuring [1973, 1423]. Mechanism [1398, 1718, 301, 401, 1041, 1027]. mechanisms [408, 51, 12]. Mechanizing [509]. median [2038, 802, 236, 473, 2312, 1861]. medians [1418]. meet [1126]. MELL [2006]. membership [992, 1706, 1786, 963, 1290]. Membrane [564, 418, 1744, 1450, 1232, 2116, 1743, 1461]. membranes [1737, 103, 1738]. memory [236, 240, 746, 1361, 396, 1505, 1362, 1442]. Memoryless [32]. Menger [1672, 2394, 1691, 1509, 2289]. Menger-edge-connectedness [1691]. Menger-vertex-connectedness [1691]. merge [1820]. merged [1522]. merging [526, 907]. mesh [1538]. Message [115, 248, 247, 656]. messages [1527, 2269]. meta [2147, 771]. meta-theory [2147]. Metagenomic [1424]. metaheuristic [1609]. metamorphism [181]. metatheory [1341]. method [508, 965, 2236, 1396, 1523, 1520, 568, 1302, 416, 2343]. Methods [426, 1722, 2340, 1304]. metric [1903, 473, 1450, 81, 1199, 739, 815, 2027, 615]. metrics [1513, 1943]. micro [932]. micro- [932]. microarray [1457, 568]. might [396]. migrating [926]. migration [753]. mildly [1410]. million [1302]. Mim [2363]. Mim-width [2363]. Min [682, 2028, 590, 1485, 1030, 638]. Min-max [682, 1030]. min-sum [590]. MINCCA [1368]. Minimal [563, 2271, 1248, 1568, 2179, 1325, 968, 891, 976, 65, 1892, 402, 1996, 816, 2167, 2367, 359, 229, 2302, 1080]. minimal-interval [976]. minimality [1823]. Minimax [288, 2062, 1596]. minimaxity [1784]. minimization [457, 239, 1396, 598, 1468]. minimize [249, 238, 755, 123, 1505, 2335, 1493, 140, 714, 2055]. Minimizing [627, 2281, 2252, 1150, 538, 533, 1715]. Minimum [1656, 2230, 1492, 1905, 2058, 973, 2071, 1183, 2066, 1607, 1608, 496, 163, 66, 395, 883, 2326, 1256, 628, 619, 1756, 1155, 400, 1394, 2018, 2222, 2165, 744, 1805, 1058, 3, 68, 324, 1473, 2246, 697, 1832, 325, 2286, 119, 1816, 1628, 165, 1263, 134, 1092, 1245, 1030, 2279, 362, 1152, 1593]. minimum-width [1607]. Mining [1038, 723]. minmax [548]. minors [1158]. mirror [2228]. mirror-edges [2228]. Misère [389]. mismatch [880]. mismatches [1663, 893]. missing [615]. Mitigating [931]. Mixed [1305, 1296, 2312, 2062]. mixes [1574]. mixture [286]. MKTP [2319]. MM [1927]. MM* [1913, 1238]. Mobile [1929, 2334, 613, 1531, 577, 2331, 502, 597, 1970, 578, 1755, 1630, 111, 710, 1489]. Modal [929, 480, 77, 1670, 1806, 1055, 2003, 1385]. modalities [877, 1386, 1342]. mode [221, 1647]. Model [428, 669, 2060, 2061, 768, 1916, 427, 2198, 157, 1635, 1974, 1469, 1876, 876, 850, 2203, 1728, 629, 1142, 1, 984, 1913, 771, 567, 2305, 986, 119, 1238, 60, 717, 901, 2015, 1810, 227, 747, 26, 388, 1086, 667, 1237, 656, 637, 1912, 1927, 1459]. model-checking [1728]. Model-order [428]. modeled [562]. modeling [852]. Modelling [767, 1883, 763, 563, 2148, 1651, 1459]. Models [285, 666, 1183, 1389, 279, 260, 913, 1882, 429, 109, 1821, 1865, 1809, 489, 431, 663]. modes [1480]. Modification [941, 1032]. modifications [2038]. modified [1984, 420]. Modular [832, 211, 150]. Modulated [526]. modulo [1225]. mold [2050]. Molecular [831, 432]. molecules [945]. monadic [728, 487, 1700]. monads [486, 487]. money [2122, 868, 1824]. Monitoring [1205, 1604]. monochromatic [468, 1137, 1242]. Monoid [1289]. monoids [1693, 453]. monophonic [606]. monopolies [1194]. monotone [1028, 287, 2197, 2028]. monotonic [154, 38]. Monotonically [2112]. Monte [938, 941, 942, 943]. Monte-Carlo [941, 942]. Moran [1933]. morphic [908, 461]. Morphism [1375]. morphisms [2142, 690, 1791, 1557]. morphogenesis [839]. Morse [2141, 356]. most [1959, 1907]. Motif [1542]. motion [1373, 1273]. Move [1489, 937]. Move-optimal [1489]. Movement [1195, 1012]. moves [293]. moving [950]. MSO [370]. MSOL [1254]. MSVL [1854, 1027]. MTL [479]. mu [77]. mu-calculus [77]. Multi [2143, 1131, 1949, 840, 262, 478, 1882, 1642, 1051, 1629, 691, 398, 2262, 2349, 2308, 2012, 547, 1612, 1985, 298, 196, 1246, 1321, 115, 1262, 586, 1740, 142, 1715, 1491, 1401, 2072, 1697]. multi- [586]. multi-access [2349]. multi-battle [1629]. multi-broadcast [115]. multi-channel [1642, 691, 1401]. multi-coloring [398]. multi-dimensional [1882, 1262]. multi-domain [262]. multi-head [1321]. multi-hop [2262]. multi-objective [1612, 1740]. multi-parameter [1985]. multi-party [2308, 2012]. multi-path [196]. Multi-periodic [840]. Multi-player [2143, 1131, 1949]. multi-queue [298, 1246]. multi-service [1491]. multi-stack [1051]. multi-stage [1715]. multi-tape [478]. multi-target [1697]. multi-type [142]. multi-user [547, 2072]. MultiAspect [1001]. multicolorings [532]. multidimensional [1392, 1719]. multifurcating [55]. Multigames [488]. multigraphs [592]. multigrid [955]. Multilevel [1501]. multimatroids [562]. multimodal [2048, 1881]. multioperator [371]. multipartite [135]. multiparty [1205, 1963]. Multipass [436]. Multiphase [663]. Multiple [1674, 514, 254, 1833, 1149, 1528, 220, 740, 1932, 754, 2359, 904, 1303, 1039, 2202, 2085, 656, 585]. multiple-choice [1039]. multiple-message [656]. Multiplicative [1623]. multiplicity [654]. multiplier [2047]. multiply [1943]. multiply-labelled [1943]. multiprocessor [812, 2284]. multiscale [2179]. multiset [2278]. multisets [1320]. Multistate [1244]. multithreaded [1063]. multivalued [1589]. Munkr [718]. Mutation [771, 2193, 1185]. Mutual [1692, 1874, 1529].

NAE [1089]. NAE- [1089]. naming [251]. Naor [1382]. Natural [99, 789, 258, 1277, 1442, 297, 1017]. Nature [1309]. navigational [1774]. NCL [1850]. Near [1659, 534, 1295]. near-colliding [1295]. Near-linear [1659]. nearest [1757, 1922, 1541, 1544, 382, 46, 1303, 572, 337]. nearest-neighbor [337]. Nearly [1028, 1300, 1699]. necessary [1519]. necklaces [2249, 2126]. needed [2293]. neededness [1226]. needs [1443]. Negation [1144]. Negation-limited [1144]. negative [1345, 644, 540]. Neighbor [1148, 1922, 1925, 1913, 2305, 1238, 1517, 596, 1413, 572, 337, 917, 2051, 1412, 1133, 799]. neighborhood [1187]. Neighbourhood [1719, 571]. Neighbourhood-preserving [1719]. neighbourhoods [2114]. nested [1097, 193, 1244, 2256]. net [1866, 416]. nets [506, 1668, 418, 1804, 1730, 1173]. Network [1145, 1195, 666, 2295, 1872, 2253, 148, 1181, 17, 59, 2103, 2394, 1979, 1603, 1079, 1236, 832, 1988, 1924, 1221, 317, 1981, 969, 1997, 1714]. network-decomposition [1872]. Networking [15]. networks [1479, 1904, 2189, 164, 1311, 249, 2011, 199, 548, 1665, 1205, 1600, 2185, 814, 2370, 1980, 2138, 1990, 2073, 429, 1925, 691, 269, 763, 1968, 903, 2262, 27, 225, 578, 326, 749, 2355, 657, 1162, 109, 565, 545, 1860, 246, 801, 593, 921, 2087, 1078, 546, 1691, 243, 288, 514, 2007, 861, 1349, 1552, 1673, 374, 1473, 1244, 1511, 1781, 1246, 1983, 314, 1211, 1020, 1388, 1484, 1249, 1262, 1948, 2356, 586, 2237, 1434, 2320, 571, 2302, 1008, 907, 233, 1176, 218, 842, 541, 572, 2371, 227, 1928, 431, 599, 2051, 1412, 1995, 1237, 1569, 656, 799, 2373, 358, 126, 1401, 2055, 2130]. networks [618, 60]. Neural [2106, 767, 1315, 763, 100, 1460, 772, 1648, 2119, 969, 840]. neurons [842]. neuropercolation [841]. Neutral [1108, 924]. Newton [2340, 1302]. Next [372, 1033]. Next-preserving [372]. next/previous [1033]. NFAs [649]. NIC [348]. Nim [1699, 961, 1949, 1826]. Nintendo [275]. Nivat [2111]. NLC [700]. NLC-width [700]. No [442, 266, 1530, 346]. No-Free-Lunch [442]. Node [1875, 235, 1145, 1541, 1002, 172]. nodes [759, 58]. noise [1909, 1258]. noisy [495, 711, 775, 1142]. Nominal [1749, 2146]. Non [1590, 213, 1075, 1099, 342, 1495, 335, 1362, 245, 955, 1367, 650, 1971, 1444, 320, 2030, 1501, 318, 1234, 167, 966, 2070, 240, 154, 1940, 1570, 87, 348, 2113, 286, 1298, 947, 1380, 375, 2211, 1956, 818, 1557, 1036, 1372, 1065]. non- [286]. Non-adaptive [1590, 342]. Non-additive [245]. Non-atomic [1971]. non-automaton [1234]. non-binary [2030, 1570]. non-binomial [1956]. non-bipartite [2070]. non-constant [240]. Non-deterministic [213, 2113]. non-dynamic [818]. Non-expandable [1075]. non-interactive [348, 947]. Non-interference [1362, 1501]. non-iterative [1036]. non-linear [375]. Non-local [955, 1367, 1380]. non-malleable [1940]. non-monotonic [154]. Non-overlapping [1099, 335, 1075, 167, 1065]. non-pointed [1372]. Non-recursive [650]. Non-regular [1495]. non-renewable [87]. non-semilinear [1444]. non-strict [318]. non-termination [1298]. non-trivial [2211]. non-uniform [320, 966]. nonadaptive [473]. Nonconvex [267]. nondeterminism [649]. Nondeterministic [929, 2212, 2153, 443, 1230, 1278, 490]. none [1717]. nonequivalent [982]. nonlinear [1567, 1840]. nonlinearity [1752]. nonmonotone [253]. nonnesting [341]. Nonparametric [740]. nonsimultaneous [704]. nonstandard [982]. normal [727, 1227, 2177, 1130, 1792, 1914, 779, 1696, 290]. normalisation [1226]. Normalization [2139, 789, 2099]. normalized [1628]. norms [2038, 2312]. notation [1568]. note [2254, 67, 1138, 119, 2259, 2265, 2390, 917, 460]. notion [1151, 661, 1746]. Notions [216, 822, 716]. Novel [1713, 2353, 1390, 1185]. novelty [2184]. NP [1851, 1247, 219, 1137, 1832, 434, 1581, 1397, 1242]. NP-complete [1137]. NP-completeness [1581, 1242]. NP-Hard [1851, 1247]. NP-hardness [1397, 1832, 434]. NP-intermediate [219]. NPDAs [1836]. nucleolus [624]. nucleosome [2269]. Number [817, 1239, 2306, 97, 655, 1037, 2177, 972, 223, 2074, 258, 4, 474, 1191, 65, 195, 909, 777, 801, 2167, 872, 611, 2126, 2058, 2360, 2341, 2247, 1554, 982, 1640, 1937, 554, 122, 1520, 1566, 1835, 944, 1263, 1624, 1910, 1938, 1015, 620]. numbers [1011, 1229, 2093, 2294, 329, 1305, 2200]. numeration [1931]. numeric [1582]. Numerical [2192, 1231, 926, 925, 1647].

O [163]. OBDD [1873]. OBDD-based [1873]. object [823, 1772, 1688]. object-oriented [823, 1772]. objective [1207, 1612, 1740]. objectives [151]. objects [776]. oblivious [248, 1575, 111, 533]. obnoxious [1507, 2038, 1718]. observant [962]. observation [1732]. Observational [182]. observations [1286, 1085]. obstacles [1761]. Obtaining [518, 1751]. Occam [1866]. occupancy [2269]. occurrence [1404]. occurrences [890]. octagon [1428]. octagonal [415]. Octal [1828]. octant [778]. octilinear [2131]. odd [421, 874, 1752]. odd-even [874]. odd-variable [1752]. Off [2263, 2227]. Off-line [2263]. offensive [224]. Offline [399]. offs [1571, 1438, 314, 650]. oligopolistic [125]. Ollivier [1427]. omega [993]. on-board [27]. On-line [2197, 2263, 1769, 1524, 2055]. once [1679, 814, 1519]. One [1576, 1321, 470, 33, 1923, 2383, 743, 480, 2254, 1283, 2138, 345, 344, 607, 516, 1730, 1179, 447, 1827, 1959, 2323, 864, 1718, 290, 1302, 1971, 1907, 293, 118]. one-agent [480]. one-cop-moves [293]. one-counter [1730]. One-dimensional [470, 344, 447]. one-facility [1718]. one-hop [2138]. one-round [1971]. one-rule [2323]. One-Sided [1923, 607, 290]. one-symbol [1179]. One-to-one [1576, 33]. one-unambiguous [1283]. One-way [1321, 2383, 743, 345, 516, 1827, 118]. ones [2028]. Online [320, 2083, 381, 398, 253, 469, 2339, 598, 2204, 2082, 1414, 446, 714, 142, 1435, 141, 1643, 360, 995, 1208, 1150, 1006, 551, 2222, 1045, 938, 49, 1612, 960, 447, 1079, 2335, 1493, 355, 520, 1592, 2015, 534]. only [236]. onset [1545]. open [2189, 1963, 1778, 1775, 297, 1017, 2357, 2060]. open-endedness [2189]. opening [2344]. operation [1405, 1172]. Operational [649]. operations [778, 2403, 642, 2100, 687, 2381, 350, 783]. operator [1339, 2175, 135, 1185, 1122]. operators [1585]. Opinion [1831, 983, 2330]. optima [13]. Optimal [1507, 743, 1531, 160, 2026, 538, 1612, 1434, 634, 1113, 572, 1536, 116, 2236, 2331, 1578, 2140, 338, 220, 1933, 73, 1028, 2000, 64, 1203, 247, 1300, 1489, 1752, 534, 293, 2170]. Optimality [2353, 2232]. Optimally [543, 976, 2013, 2415]. optimisation [12, 2048]. Optimistic [26, 35]. Optimization [2216, 995, 495, 711, 1811, 924, 123, 2191, 90, 14, 191, 1369, 102, 13, 925, 420, 1401, 2045]. optimizations [416]. optimize [1414]. Optimizing [802, 1681, 10, 276, 2056, 57]. optimum [2027]. option [245]. oracle [342, 603, 1953]. oracles [998, 201, 36]. orbit [2100]. orbits [78]. Order [892, 737, 893, 1796, 2372, 789, 257, 479, 397, 1198, 728, 72, 1059, 1053, 547, 486, 428, 2005, 787, 1561, 366, 234, 380, 1499]. order-incomplete [2372]. Order-preserving [892, 893]. order-theoretic [257]. Ordered [1051, 1121, 379, 721, 560, 229, 2324]. ordering [1632, 607, 1698, 2417]. orderings [1540]. orders [477, 871, 795, 585]. ordinal [2275]. ordinary [430]. Ore [359]. organization [2269]. organizing [243]. orientation [97, 899, 1607]. orientations [2029, 126]. oriented [823, 403, 1772]. origami [1223]. oritatami [2191, 1220]. orthogonal [2328, 2418]. other [1141, 1717]. otter [962]. Out-degree [1617]. outer [2190]. outerplanar [1003, 2022, 2416, 1816]. Outfix [1449]. Outfix-guided [1449]. outlier [80]. outliers [2168]. output [2112, 779]. output-sensitive [779]. outsourcing [30]. overabundant [2267].

Overdetermined [46]. overlapping [1075, 1099, 394, 167, 1781, 335, 1454, 1065]. overlay [3, 68]. OWA [90].

P [2024, 1310, 2142, 1315, 2192, 100, 766, 103, 764, 772, 2119]. P4s [1510]. packet [1381, 2083, 2339, 1246, 634]. packets [1643]. packing [399, 913, 551, 2244, 1657, 910, 465, 56, 2082, 1721, 803, 446, 393, 979, 1152, 2229]. packings [967, 1037, 195]. page [1660]. PageRank [18]. Pages [2345, 2386]. paging [108]. pair [1685, 1035, 2161]. pairability [2070]. Paired [623, 1262, 848, 330]. paired-domination [330]. pairing [2407]. pairing-friendly [2407]. pairings [30]. pairs [1294, 1419]. pairwise [727, 4, 474]. palindrome [1069]. palindrome-free [1069]. palindromes [67, 1890]. Palindromic [2142, 781, 2235, 1537]. palletizing [2309]. pancake [227]. pancakes [589]. Pancyclicity [471, 809]. Papers [2225, 505]. paradigm [1382]. Parallel [1433, 991, 2180, 1363, 2196, 320, 1339, 440, 1495, 1354, 1090, 1996, 1392, 987, 419, 704, 1505, 315, 2343, 1488, 530, 1475, 1769, 668, 57, 2124, 574, 1633, 2013, 2415, 1715]. parallel-batch [1769]. parallel-machine [1475]. parallelism [1865, 766]. parallelizable [1664]. Parameter [1731, 1983, 1811, 110, 1930, 2045, 1985, 173, 1478, 387, 1832]. parameter-rich [1811]. parameterised [1982]. parameterizations [1411, 2157]. Parameterized [666, 1639, 1152, 1987, 752, 1023, 958, 1600, 517, 55, 2338, 2168, 1759, 1723, 756, 1368, 1763, 435, 1002, 539, 2042, 1911, 1721, 1125, 1815, 2164, 383, 1824, 979, 549, 2196, 1143, 1829, 2310, 121, 594, 2317, 2155, 404, 1722, 83, 170, 2076, 552, 2247, 1153, 1902, 632, 2239, 2378, 535]. parameters [438, 1763, 219, 47, 1136, 698, 2179, 118]. Parametric [81, 1841]. parentage [204].

Pareto [123]. paribus [1125]. Parikh [228, 1796, 1619, 1465]. Parikh-reducing [1465]. parity [687, 173]. Parry [1931, 94]. parsable [1105]. parsimony [957]. parsing [1536, 1281]. Part [1486, 1163]. Partial [734, 1779, 368, 710, 1952, 622, 186, 2154, 1747, 1420, 1666, 1570, 1732, 2080, 2125, 175, 1547, 1581, 916, 1489, 234, 1987]. partially [721, 1113]. participants [2026, 2231]. Particle [2045, 13, 420]. partite [2303]. Partition [1084, 2124, 1348, 1899, 636]. Partitioned [1793, 1660, 166]. partitioning [1239, 544, 1609, 2246, 1160]. partitions [1389, 873, 906, 1834, 1617, 2081, 2310, 341, 1572, 625]. partners [195]. party [2308, 2012]. Pascal [2126, 1931]. Pascal-like [2126, 1931]. passing [1950]. passive [1652, 1291]. Path [2017, 2130, 2306, 249, 2151, 2240, 548, 2294, 1576, 623, 2300, 1394, 998, 1860, 288, 514, 196, 1262, 14, 423, 1040, 2116, 848, 713, 2086, 2239, 2164, 599, 1092, 33, 981, 1401]. path-forests [2294]. path-reporting [998]. Paths [2395, 199, 503, 2325, 2321, 2070, 869, 1235, 468, 911, 1490, 745, 1365, 130, 1950, 783, 1551, 161, 2065, 1994, 535, 540]. pathway [1981]. pathways [283]. Patricia [635]. Pattern [841, 339, 594, 2093, 2096, 1538, 782, 893, 1542, 1785, 2111, 2021, 1067, 723]. pattern-avoiding [2093]. Pattern-based [841]. patterns [1833, 1795, 1707, 501, 1182, 1323, 133, 842]. payment [51]. payoff [1732, 1252]. payoffs [48]. PCF [260]. PCGs [2251, 136]. PD [1499]. PDAs [886]. PDL [1339]. peak [1505]. penalties [1138, 1861, 819, 2067]. Penalty [530]. perception [1902]. perception-parameterized [1902]. percolation [1517]. Perennial [2012]. Perfect [2347, 1428, 1101, 1830, 1905, 724, 2040, 1250, 1625, 1689, 126]. Performance [2356, 48, 1361, 2037, 102]. Period [1535, 2288, 907]. periodic [1132, 579, 1621, 840]. Periodicity [451, 215, 1067, 2265, 1070]. periods [1064]. permutahedron [995]. permutation [1352, 2402, 1316, 1538, 591, 452, 1199, 40, 705, 585]. permutations [860, 462, 1989, 1795, 1678, 1418, 1583]. permute [2335]. Permuted [882]. persist [1602]. persistent [1101]. personal [1857, 2223]. personalized [53]. perspective [608, 1980, 1629, 1809]. perturbation [748]. pessimistic [1366, 617]. Pete [683]. Petersen [171]. Petri [1668, 418, 1804, 1866, 1173]. Pfaffian [126]. phase [841]. phylogenetic [2355, 957, 1694, 1943, 1427, 585]. phylogenetic-tree [1943]. phylogeny [1101]. physical [119]. picking [1577]. picture [1285, 2318]. pictures [1075, 2092]. piecewise [1734, 1474, 1812, 1322, 1559, 796]. pigeonhole [174]. Pike [683]. pile [1389, 961]. pivot [677]. pixels [777]. PKE [2344]. place [1364]. placement [325]. plaintexts [2030]. plan [2049]. Planar [2334, 363, 1680, 1254, 2230, 1508, 1641, 132, 702, 1814, 1037, 948, 866, 208, 1967, 1155, 1748, 869, 1578, 1953, 403, 951, 1554, 2217, 521, 1486, 153, 596, 798, 1615, 917, 1148, 1995, 813, 96, 1762]. Planarity [2388, 129, 121, 303, 608]. plane [2266, 1572, 198, 2406, 953, 1400, 779, 1160, 337, 1907]. planes [781]. planning [1962]. plausible [483]. play [231, 1658]. played [389]. player [2143, 1131, 1949, 114]. players [125, 1603, 114]. playing [939]. Playout [937]. playouts [938]. PMC [1913, 2305, 1238, 60, 1927]. point [1352, 2227, 1037, 1198, 1698, 2417, 154, 162, 2076, 146, 1554, 740, 155, 233]. pointed [1372]. points [2071, 1608, 802, 1681, 693, 902, 1703, 1791, 1306]. polarization [1461]. polarizations [2119]. polarized [1311]. policies [1338, 938]. Policy [1818, 937]. politician [276]. poly [1378]. poly-slender [1378]. Polygon [2252, 2227, 2277, 1245]. polygonal [165]. polygons [2071, 544, 2240, 2131, 1452]. polyhedral [1761]. polylogarithms [2406]. Polymerization [945]. polymorphism [2151]. Polynomial [1476, 1669, 464, 1816, 1956, 508, 2156, 2300, 687, 1, 984, 259, 2366, 1083, 200, 949, 382, 1258, 1091, 2229, 1914, 1716, 423, 1301, 1299, 794, 2316, 1303, 1306]. Polynomial-time [1669, 687, 1914]. polynomials [1296, 1840, 2414, 1302, 1303]. polyomino [1250]. polyominoes [784, 1170, 2108]. polytime [257]. polytopes [2354]. polytrees [338]. Population [1900, 11]. posets [156]. position [2254, 1700, 2352]. position-dependent [2352]. positions [1334]. Positive [644, 2200, 1235, 2107, 1724, 1908, 2255, 736, 1306, 540]. positivity [616]. possible [2372, 706, 1769]. Post [139, 452]. Post-challenge [139]. postal [747]. Posterior [1811]. Postfix [62]. potential [2011, 931]. potentialities [508]. Power [2314, 336, 80, 2326, 381, 793, 52, 1800, 119, 1456, 455, 2049, 541, 2268, 412, 533, 923, 573, 973, 2374, 2119, 1647]. powered [2093]. powerful [224]. powers [858, 2363]. Practical [388, 2166]. practice [600, 1302, 2371]. Precise [1351]. precision [1734]. preclusion [171, 696, 861, 1349, 1552, 1566, 2259, 1910, 1938]. precondition [486]. predecessor [1957, 50]. predecessor-existence [50]. predicate [1358]. predict [1545]. Predictors [1744]. Preemptive [238]. Preface [2292, 285, 1853, 2089, 485, 1165, 1276, 2401, 1978, 569, 1897, 774, 2216, 1095, 513, 732, 1356, 830, 1646, 1606, 1871, 836, 1293, 1082, 537, 1712, 1705]. preferences [2234, 1690, 1125]. Prefix [1314, 1334, 647, 2177, 1062, 2207, 1130, 462, 437, 1792, 1174, 863, 1419, 350, 1696, 1284, 1700]. prefix-convex [2207]. Prefix-free [647]. Prefix-suffix [1314]. prefixal [2323]. prefixes [2074]. Preparing [1886]. Presburger [1728]. prescribed [1201, 1950, 2086, 1994]. presence [2168]. Presentation [465]. presented [1693]. preservation [1023, 958]. preserve [73]. preserves [1511]. preserving [1964, 892, 893, 921, 204, 1755, 2307, 1719, 1221, 393, 1038, 372]. pressure [2330]. prevention [567]. previous [1033]. price [1513, 680, 1492, 59]. pricing [142]. prima [778]. Primal [37, 2218, 253, 819]. Primal-dual [37, 2218, 253, 819]. prime [2090]. primitive [778, 369]. principle [843, 174, 2238]. principles [1458]. prints [1619]. prior [1596]. priorities [565]. priority [1874]. Prisoner [2184]. Privacy [204, 408, 2073]. Privacy-preserving [204]. private [738, 2166]. Probabilistic [2337, 250, 1985, 2284, 736, 1359, 302, 563, 1972, 1044, 1883, 2148, 1289, 1203, 1731, 482, 251, 1093]. probabilities [1141, 1229]. probability [1185]. probe [2134, 112]. probing [56, 2056]. Problem [2075, 1508, 313, 2030, 1227, 1859, 570, 2402, 496, 1799, 1023, 958, 163, 1156, 608, 439, 1463, 975, 1737, 351, 1967, 1448, 1921, 2283, 384, 93, 344, 628, 607, 1725, 1257, 1538, 327, 966, 1609, 400, 1698, 2417, 1892, 1701, 1848, 326, 1354, 1090, 1394, 448, 2193, 2190, 1, 984, 678, 590, 1241, 1860, 1820, 38, 1866, 272, 1506, 1368, 598, 2204, 1612, 1763, 435, 288, 229, 2246, 1138, 2340, 2059, 1153, 1951, 1430, 1832, 1546, 2281, 610, 2009, 1214, 1502, 530, 916, 1716, 1419, 22, 632, 1630, 520, 1696, 1628, 50, 147, 2312, 1285, 386, 2027, 798, 48, 1485]. problem [1680, 161, 2186, 2006, 2357, 420, 2158, 134, 1039, 1824, 37, 636, 1861, 2067, 1491, 2279, 362, 2199, 360, 981, 638, 1858, 718, 29, 2303, 1697, 24, 2172, 577, 452, 2319]. Problems [2408, 1727, 549, 2196, 164, 1507, 2038, 166, 1693, 1684, 121, 1010, 2198, 2155, 2134, 1012, 1903, 1811, 2326, 1448, 2232, 1817, 436, 604, 2073, 619, 2244, 171, 2313, 2380, 1756, 2338, 1420, 706, 1759, 2397, 679, 1722, 1723, 2076, 1805, 87, 1396, 49, 514, 1364, 453, 791, 2363, 1864, 219, 173, 2409, 552, 539, 90, 1957, 2069, 1236, 230, 1863, 1911, 56, 1085, 1514, 14, 150, 1721, 2362, 1266, 1213, 847, 1369, 1380, 1040, 1220, 521, 1486, 446, 1397, 2211, 553, 393, 1653, 1092, 819, 2062, 757, 1720, 2352, 1030, 979, 1189, 1962, 1043, 1076]. procedure [629, 786]. procedures [824, 251]. process [1357, 74, 763, 1933, 1741, 109]. processes [686, 1373, 151, 672, 994, 996, 664, 50, 1917]. processing [237, 1457, 520, 1277, 1633, 1435, 141, 1643, 2352]. processors [1311]. product [604, 1192, 590, 1351, 84, 519, 292, 1134, 2166, 1910]. production [61, 140, 1231]. products [2257, 689, 1835]. Prof. [1534, 2261]. Professor [1416]. Profinite [1289, 1114]. profit [1388]. Program [261, 1500, 939, 940, 2113, 182]. programmability [342]. programmable [1222]. programmer [369]. programming [1208, 823, 1344, 1561, 511, 904, 1093]. programs [727, 1339, 1883, 1198, 644, 259, 1865, 183, 1298, 88, 730]. progressions [2398, 2023]. Progressive [940]. projection [629, 2186, 2060, 1304, 637, 2061]. Projections [2150, 2316]. projective [1297]. projects [1272]. proliferation [564]. Promise [1189]. promoters [766, 923]. prone [1532]. Prony [1295]. Proof [506, 2149, 1702, 1839, 854, 922, 1242, 2304]. proofs [479, 174, 644, 2353, 106]. propagation [1756]. Proper [899, 611, 554, 2121, 129, 97, 2029, 2207, 2313, 40, 2058]. properly [468]. Properties [2383, 303, 873, 906, 279, 2185, 422, 682, 782, 2327, 497, 472, 1751, 219, 884, 190, 805, 1918, 1380, 2382, 181, 736, 94, 2351, 566, 1997]. property [2402, 884, 287, 1908, 2170]. Proportional [2059, 1194, 117, 1521]. propositional [174, 629, 1055, 482, 637, 1804]. protected [759]. protects [725]. proteomes [1425]. protocol [1909, 1918]. protocols [666, 409, 216, 1900, 1767, 754]. prototype [1757]. Provably [1550, 901, 1036, 388]. providing [1404]. Proving [1918, 1345, 1810]. proxy [1007, 901]. prune [1427]. pruning [407]. pseudo [1861]. pseudo-solution [1861]. pseudofractal [1263, 1624]. Pseudoknot [1405]. Pseudoknot-generating [1405]. Pseudorandom [676, 118, 972]. pseudostandard [2139]. PSPACE [1049, 593, 1264, 683, 2264, 274]. PSPACE-complete [1049, 1264, 683, 274]. PSPACE-completeness [593]. PTAS [24, 2027, 134]. PTime [1559]. PTime-complete [1559]. Ptolemaic [464]. Public [2304, 201, 2307, 1272, 818, 1035, 139, 1912]. public-key [2307, 818, 139]. pullback [368]. pumpable [605]. pure [1118]. purely [764]. Pursuit [1046, 2143]. pursuits [2337]. push [1585]. push-the-button [1585]. pushdown [187, 1051, 2104, 92, 2114, 2382, 857, 2061]. Pushing [1823]. pushout [1779]. puzzle [1702]. puzzles [266]. PV [730]. PV-programs [730]. pyramid [58].

QANIZK [2133]. QBF [179]. QoS [298, 1246, 1401]. QPTAS [1554]. Quad [707]. Quad- [707]. quadraginta [778]. quadratic [961, 88]. quadrilaterals [2033]. quadtree [777]. Qualitative [374, 1291, 2007, 2371]. quality [1398]. Quantified [721]. quantifiers [72, 1212]. Quantifying [409, 1029, 2282]. quantiles [546]. Quantum [1879, 1169, 1742, 631, 1885, 722, 1888, 1886, 421, 765, 1087, 106, 412, 1884, 1881, 1189]. quantum-inspired [1881]. Quartet [2009]. Quasi [2290, 2095, 1621, 477, 37]. Quasi-automatic [2290, 2095]. quasi-greedy [37]. quasi-orders [477]. Quasi-periodic [1621]. Quasicontinuous [156, 373]. Quasipolynomial [174]. quaternion [2414]. QUBO [1448]. qudit [1885]. Queens [1737]. Quell [354]. queries [992, 1123, 2240, 721, 883, 1706, 1786, 2017, 609, 64, 1033, 587, 572, 337, 2137]. Query [1124, 679, 1742, 1799, 193, 221, 2018, 1953, 1058, 633]. Query-competitive [679]. querying [971]. Question [2118]. questions [1708]. queue [298, 1246, 190]. queues [2309]. quotient [647].

R [500]. R-LINE [500]. R1Q [1799]. RaceTrack [1846]. Rademacher [1783]. radio [53, 691, 2262, 571]. radius [270, 1080]. radix [2243]. Raffaele [1416]. rainbow [1003, 362]. raised [947]. Random [1229, 297, 837, 1399, 1013, 342, 1944, 1692, 1191, 1429, 1722, 759, 603, 611, 211, 1418, 2180, 114, 499, 201, 218, 1431, 85, 1017, 36, 973, 2276, 466]. randomization [408]. Randomized [1508, 1873, 49, 1430, 1532, 533, 500, 409, 1018, 2052, 20, 362]. Randomness [2160, 1382, 859]. randoms [1494]. Range [587, 1799, 883, 221, 2018, 970, 609, 546, 1058, 1033, 1153]. Rank [1999, 304, 2341, 748, 1032, 1186]. Rank-maximal [1999]. rank-width [304]. Ranked [1833]. Ranking [871, 2249, 2231]. rapid [2056]. Rate [1019, 1467, 429, 621, 419, 2024, 2127, 1596]. Rate-limited [1019]. rates [2193]. ratio [1610, 2022, 2416]. rational [1172, 1703, 497, 1112]. ratios [292]. Rauzy [294]. Raw [742]. Razor [1866]. re [1007]. re-encryption [1007]. Reachability [1841, 1860, 211, 1863, 1727, 1734, 71, 1530, 565, 827, 1985, 1203, 1842]. reachability-time [827]. Reaching [1461]. Reaction [422, 414, 768, 1445, 769, 1980, 561, 1317, 430, 1451, 921, 1979, 832, 2115, 413, 1325, 1981, 1186]. reactive [1972, 1807]. read [1679, 814, 236, 1519]. read-once [1679, 814, 1519]. read-only [236]. reads [1424]. Real [258, 1299, 270, 2370, 381, 1729, 329, 932, 1303, 1306]. real-time [2370, 381, 1729, 932]. real-world [270]. realisability [1869]. Realistic [1245, 1584]. Realizability [260, 1622]. realization [1696]. reallocation [2234]. really} [708]. reals [297, 1017]. rearrangements [1583]. Reasoning [2369, 1804, 482, 1460]. rebalancing [753]. receiver [2344]. reciprocal [529]. Recognition [2348, 1378, 1618, 749, 464]. recognizable [2318, 1700]. recognize [977]. Recognizing [866, 1829, 112, 1894]. recoloring [468]. Recombination [562]. recommendations [733]. Recommending [1969]. recompression [343]. reconciliation [743]. Reconfiguration [2040, 898, 1000, 1763]. Reconstructing [1539]. Reconstruction [784, 2108]. recovery [1535, 1975, 1781, 2201, 1252]. rectangle [1037]. rectangular [1365]. rectilinear [2332, 1031, 2312, 1245]. recurrence [211]. recursion [1635, 1708, 489]. Recursive [2107, 824, 788, 1865, 2113, 1020, 369, 650, 58, 1569]. recursively [734, 674]. Red [1369]. Red/Blue [1369]. Red/Blue-split [1369]. redistricting [2248]. reduce [600]. Reducibilities [674]. reducibility [661]. Reducing [1032, 1617, 1465]. Reduction [957, 743, 1314, 739, 1839, 428, 1856, 1719, 1163, 169, 716]. Reductions [87, 1490, 2232, 1113]. Reeb [776]. Rees [1021]. Refined [677, 881]. refinement [480, 70, 786, 921]. refresh [1940]. refutations [1679]. region [2227, 948, 2226]. regions [1111]. register [1803, 1531, 1490, 2243, 764]. regraft [1427]. regression [1087]. regret [711, 548, 288]. Regular [1760, 2109, 1310, 993, 1281, 1100, 1631, 1447, 2207, 2096, 1283, 1495, 43, 138, 2208, 786, 963, 1366, 2214, 617, 1691, 497, 646, 189, 648, 2410, 2356, 717, 805, 798, 1022, 2210, 1465, 1942, 118, 566]. regularity [1266, 1512]. Regulated [1324]. regulating [941]. regulatory [109, 374, 1983]. rejection [704, 1475, 1363]. Rekeying [1274]. relabeling [235]. Related [1077, 1389, 2391, 1903, 1611, 1448, 2073, 1723, 1134, 1369]. Relating [717]. relation [2258, 1366, 1955, 1796, 694]. relational [1750]. relations [661, 1172, 1887, 1230, 674, 863, 1939, 1067, 2265, 857]. Relationship [804, 1021]. Relative [108, 944]. relaxation [1609]. Relaxed [1610, 1601, 1907]. relay [1145]. relevance [1880]. Reliability [1515, 1924, 694, 1673, 1211]. relocatable [198]. remainder [413]. removal [701]. Rendezvous [578, 746, 1895, 1360, 575]. renewable [87]. rental [245]. Repackable [586]. repair [502, 633, 2186]. repeats [1640, 1890]. repetends [725]. Repetition [2257]. repetitions [858]. repetitive [2037, 719]. replacement [235]. replica [1409]. replicas [325]. replication [1738]. reporter [2190, 1223]. Reporting [890, 364, 998]. represent [2219]. representability [200]. Representation [2369, 1239, 712, 89, 2179, 117, 1521, 44, 380, 673, 1884]. representations [2173, 845, 395, 924, 1936, 591, 953, 127, 175, 1122, 1064, 612, 1465]. Representing [1011]. research [1743]. reset [2104]. resets [1735]. resilience [379, 1550]. resiliency [2338]. resilient [1760, 240, 1142, 1935, 1827, 2307, 2035]. resistant [2407]. resisting [1036]. Resolution [2064, 1679, 1210, 1578, 179, 671]. Resolvability [2354]. resolving [628]. resource [262, 1868, 2218, 1946, 247, 319]. resource-bounded [1868]. resources [686, 87]. respect [276, 1928, 981, 2170]. response [1787]. response-dependent [1787]. responsive [2234]. restart [2324]. restart-delete [2324]. restarting [1324, 857]. restrained [813]. Restricted [80, 1946, 695, 1254, 959, 1010, 2057, 419, 1055, 633, 1157, 848, 713, 1363]. restrictions [467, 1031, 1945]. results [2050, 266, 1234, 1492, 438, 2398, 1572, 775, 1930, 1279, 32, 162, 146, 791, 2000, 1455, 1085, 1437, 1517, 1213, 2132, 11, 147, 340, 1858]. retrieval [1879, 579, 1833, 1142, 864]. return [357]. revenue [1156, 220]. reversal [1754, 239, 1418, 647]. reversals [1447, 1556, 1323, 1140]. Reversibility [787]. reversible [722, 1446, 1321, 50, 2047, 107, 2127, 662]. reversibly [2213]. revisited [2283, 1142, 1776, 145, 1730, 2360, 1214, 22, 1460, 874, 1680, 968, 1258]. Revisiting [1707, 362]. Revocable [692]. revocation [1683]. reward [663]. rewards [2107]. rewrite [661, 2323, 1637]. rewriting [2101, 2105, 2150, 366, 1225, 368, 192, 1700]. RFID [2220]. RGB [2217]. RGB-digraphs [2217]. Ricci [1427]. Rich [1350, 1811, 1333]. ride [1824]. ridesharing [1857, 2223]. Riemann [79]. right [2383, 900, 1806, 647, 2112, 1907]. right-angle [900]. Rigorous [11]. ring [508, 1970, 2046, 28, 1935]. rings [579, 1476, 200, 111, 710]. risk [199, 852]. risk-modeling [852]. risks [2048]. RLE [1423]. RNA [339, 1921]. road [326]. robber [527, 2297, 1291, 810, 2143]. Robbers [2298, 950, 1048]. robots [2039, 577, 1528, 2266, 654, 597, 1970, 831, 1755, 498, 1630, 111]. Robust [71, 2075, 1876, 1259, 1842]. robustness [279, 1430]. role [121, 862]. Roman [2347]. roofs [1245]. root [1299]. root-finding [1299]. rooted [1427]. Roots [2414, 1068, 1301, 1302]. Rosser [307, 1465]. rotating [2209]. rotation [1752]. Rouen [1059]. round [1078, 1971]. rounding [88]. rounds [2136]. routing [1527, 253, 2286, 742, 533, 553, 2055]. rows [2404]. RQL [1124]. RSA [1483, 2336, 1952]. Rule [1124, 366, 1220, 2323, 996, 1459]. Rule-based [366, 1459]. rulers [267]. rules [2370, 957, 1055, 366, 589, 1648]. Ruleset [2191]. rulesets [1585]. Ruling [1677]. run [1061]. run-length [1061]. runs [1436, 1071]. Runtime [14, 495, 826, 11, 1653]. runtimes [10]. Russians [1063]. RVCCC [1997]. Rytter [2261].

S [1534, 445]. S-BGP [445]. Safe

[196, 1842, 240, 697]. safety [1729, 1810]. SAGBI [149]. salesman [847, 1653]. sampled [1418]. Sampling [1428, 183, 1399, 318, 419, 1614, 986, 633, 466]. sand [1389, 86]. Sannella [1777]. sat [2015, 2028, 1410, 2276, 1089]. satisfaction [1472]. SATisfiability [2172, 2148, 1472, 1259, 821, 539, 1093]. Satisfying [585, 849, 1074, 1615]. saturated [2272]. Saving [1930]. scaffold [110, 1223]. scaffolding [384, 1725]. Scalable [691]. Scale [1592, 1249, 1263, 1624, 126]. Scale-free [1592, 1249, 1263, 1624, 126]. scaled [882]. scaling [812, 1269, 1398, 434, 355, 1719]. scan [634]. Scattered [967, 795]. scattering [1755]. scenario [367]. scenario-based [367]. scheduled [2193]. Schedulers [1808]. schedules [2180, 2056]. Scheduling [1904, 1392, 704, 1505, 2085, 206, 125, 320, 1381, 959, 202, 680, 2083, 440, 1006, 238, 755, 1946, 1398, 61, 1664, 2339, 38, 1729, 123, 2122, 87, 207, 960, 1951, 2335, 1493, 2281, 119, 124, 530, 237, 1475, 1769, 531, 140, 2197, 2132, 401, 1040, 1633, 446, 48, 714, 2158, 1435, 2202, 141, 2352, 1363, 1468]. schema [1185]. scheme [522, 551, 603, 204, 949, 2113, 1575, 1683, 901, 1715, 573, 460]. schemes [737, 19, 2026, 1007, 1935, 1659, 348, 115, 818, 671, 2329]. schools [1016]. Schwartz [1297]. Sci [2418, 1023, 474, 1329, 2417, 1354, 984, 68, 2416, 1496, 1004, 1486, 2415, 2024]. Science [278, 559, 1879]. scientist [1166]. scores [1087, 944]. SDNs [1530]. Search [2295, 954, 1922, 966, 2266, 2136, 2052, 1821, 2079, 1612, 1998, 493, 2123, 941, 943, 2053, 944, 1163, 1089, 360, 1697, 938, 942]. search-based [493]. Search-space [954]. searcher [854]. searches [493]. Searching [2292, 116, 1043, 133, 213, 1208, 1471, 577, 1626, 2176]. searchlight [326]. second [5, 1329, 1328, 728]. second-order [728]. secondary [339]. Secret [1686, 2026, 421, 595]. section [2377]. Secure [999, 30, 2098, 1019, 2308, 2012, 1550, 2344, 2322, 914, 947, 901, 1036, 1035, 388, 2072, 2329]. securely [754]. Security [1379, 1973, 822, 603, 1007, 2307, 460, 716, 1015]. seeds [1424, 1547]. SEF [1855]. SEFE [166]. segmentation [776, 1457]. Segmenting [568]. segments [2129, 2181, 2013, 2415, 2228]. segregation [298]. Select [2335]. Selected [505, 2225]. selection [1469, 70, 473, 1511, 677, 401]. selective [2344]. selectors [2349]. Self [249, 2039, 1181, 243, 325, 1193, 1531, 1527, 1191, 2156, 797, 1892, 1218, 359, 1800, 1893, 697, 111, 1741, 1217, 2186]. Self-adjusting [249]. self-assembly [1191, 1218, 1800, 1217]. self-healing [1527]. Self-organizing [243]. self-stabilization [111, 1741]. Self-stabilizing [2039, 1181, 325, 1531, 2156, 797, 1892, 359, 1893, 697]. Self-updatable [1193]. Selfish [1291, 2143, 1392]. sell [448]. selling [220]. Semantic [1337, 505, 2145]. semantically [1035]. semantics [1879, 1668, 1281, 976, 1883, 481, 1198, 1450, 1053, 1500, 2375, 486, 1983, 1561, 668, 1779, 1125, 833, 76]. Semi [1006, 960, 520, 138, 54, 765, 1908, 2255, 1625]. semi-definite [1908]. semi-definiteness [2255]. semi-linear [138, 54]. Semi-online [1006, 960, 520]. semi-perfect [1625]. semi-quantum [765]. semicomplete [1834]. semidefinite [2200]. semidistributive [1126]. semigroups [2095, 1021, 1234]. semilinear [1444]. semiregular [1239]. semiring [73]. semiring-induced [73]. semirings [1766]. semitotal [1991]. sensitive [666, 779, 1265]. sensitivity [2031, 1052, 1955, 2256]. sensor [657, 546, 314, 1434, 572, 2055]. sensors [522, 198]. sentiment [1881]. Separability [1559, 786]. separable [416]. Separating [1604, 1670, 1182, 2066, 2317, 877, 1556]. separation [824, 270]. separations [735]. separator [2367]. Sequence [887, 880, 2267, 223, 523, 1149, 1577, 1178, 357, 1546, 2178, 1038, 217]. Sequences [1103, 477, 1097, 2093, 1481, 1692, 2365, 2141, 1077, 25, 1789, 2079, 1494, 1768, 2315, 1391, 705, 1522]. sequent [2147]. Sequential [201, 1585, 1270, 1709, 315, 841, 107, 2282, 1647]. Serbanuta [1619]. serial [160, 140]. serializability [832]. series [1612, 1505, 740, 2124]. series-parallel [1505]. server [500]. service [1398, 1491]. session [1205, 1636]. Set [880, 553, 1037, 570, 743, 1681, 693, 1665, 948, 628, 341, 1756, 1155, 1892, 2219, 679, 2140, 1657, 52, 1000, 292, 173, 2408, 1554, 1454, 2286, 22, 1816, 1532, 1220, 2264, 1248, 1064, 2013, 2415, 698, 1092, 535, 2010]. sets [1444, 1075, 2092, 1121, 1608, 1168, 971, 1332, 157, 395, 721, 1446, 1171, 1256, 965, 65, 54, 1107, 1108, 1498, 1595, 816, 1898, 2167, 359, 226, 497, 646, 1768, 2040, 729, 1473, 1182, 554, 2411, 205, 2396, 2086, 2028, 1894, 1263, 1624, 297, 1017]. setting [1827, 1784, 2072]. several [1722, 230, 1040]. shadows [1035]. Shamir [1767]. shaped [745]. shapes [953]. Shapley [59]. share [1901]. Shared [1688, 199, 1382, 547]. Shared-object [1688]. shares [947]. sharing [959, 1344, 2026, 421, 2048, 180, 1824, 961]. sharp [1156]. Shift [1287, 1794]. Shift-invariant [1287]. Shifts [1097, 601]. Shinohara [1707]. Shoham [2374]. shop [229, 1040, 2357]. shops [1715]. Short [2322, 2288, 67, 201, 947, 670, 1197]. Shorter [443]. Shortest [630, 2240, 1235, 1394, 1438, 1860, 1364, 64, 1547, 130, 14, 1551, 1040, 2065]. should [1443]. shuffles [454]. sibling [944]. side [2148]. side-channel [2148]. Sided [1923, 332, 607, 290, 534]. Sierpiński [1263, 1624]. Sieve [817]. sight [670]. Signal [1454]. signaling [431]. signals [775, 1884]. signature [1935, 1683, 901, 1112, 36, 2329]. signatures [850, 1142, 2322, 201, 947, 2169, 388, 716, 2072]. signed [1937]. signing [2322]. silent [1181, 1892]. Sillke [1104]. Silver [1699]. similar [895, 1522]. Similarity [2162, 887, 881, 1908]. Simple [711, 1060, 1498, 16, 2071, 1269, 2240, 2277, 2370, 1967, 1655, 1702, 2104, 2018, 64, 708, 498, 1085, 2373, 494]. Simpler [1660, 1232, 2201]. simplicial [157, 712]. simplicity [920]. simplification [2252]. Simplified [1159, 987]. simulating [764]. Simulation [2344]. Simulation-based [2344]. Simultaneous [953, 1033, 2349]. Simultaneously [950]. Single [207, 237, 1903, 691, 1235, 61, 1820, 1290, 960, 124, 117, 2158]. single-crossing [117]. single-hop [691]. Single-machine [207, 237, 124, 2158]. single-source [1235]. single-swap [1903]. singleton [2371]. singleton-based [2371]. sink [288, 514, 2086]. sinks [548]. SINR [573, 656]. SINR-capacity [573]. Site [2380]. Site-directed [2380]. six [270, 2026]. size [1608, 954, 174, 1256, 1026, 619, 896, 724, 1601, 1616, 1137, 765, 1856, 2270, 11, 588]. size-constrained [619]. sized [1365, 631]. sizes [886, 1006, 27, 1318, 2412]. skeleton [84]. ski [245]. Skolem [1472]. Skolemization [2004]. skyline [2222]. slack [2253]. sleep [434, 355]. slender [1378]. slenderness [2110]. slicing [1052]. sliding [444]. slime [2050]. slopes [1578]. SLP [1066]. Small [1444, 764, 992, 1352, 1026, 1315, 2300, 2250, 1159, 2021, 100, 2336, 907]. small-width [2300]. smaller [45, 1789, 1033, 2201]. smallest [708, 2270]. smart [1951, 634]. smartphones [1414]. Smoothed [635]. SMT [1798]. snake [1850]. snap [1359, 793]. snap-stabilization [1359, 793]. SNC [1293]. Snowflake [945]. Snowman [1264]. Social [2330, 1831, 1600, 1968, 246, 243, 1388, 1714]. Sofic [601]. software [493]. solo [1327]. Solomon [1442]. Solomonoff [286, 1591]. Solution [39, 1295, 116, 2232, 1798, 247, 2264, 796, 2158, 1861, 1093]. solutions [73, 949, 2000, 729, 2281, 88, 1615, 2276, 1962]. solvable [248, 616]. solved [1189]. Solving [1737, 2232, 1258, 1091, 1562, 1716, 718, 179, 610, 794, 420]. Some [2251, 860, 2398, 1930, 162, 96, 291, 2253, 2198, 2155, 65, 195, 1699, 501, 1366, 617, 1911, 1085, 1114, 455, 1465, 1030, 2121]. Sørensen [1610]. sort [952, 2130]. sortedness [1397]. Sorting [1447, 1989, 1583, 462, 529, 350, 1140]. sound [1767]. Soundness [671]. source [1235, 1236, 2086, 2027]. source-destination [2027]. sources [2317, 1880, 1905]. Souto [2024]. Space [693, 653, 998, 1438, 2238, 1899, 380, 817, 2227, 1618, 954, 1571, 1882, 2232, 19, 221, 2018, 1798, 1529, 1287, 64, 1063, 1983, 2378, 1353]. Space-efficient [693, 998, 2238, 1899, 380, 1063]. Space-time [1438]. spaced [1424]. spaces [1757, 1922, 1635, 1450, 1885, 730]. spanner [2333, 1832]. spanners [1479, 212]. spanning [1656, 2198, 66, 1967, 1756, 1805, 528, 630, 855, 1716, 1520, 165, 1008, 2176, 2027, 798, 1485, 1209, 1569, 620, 2069]. Sparse [1167, 775, 1421, 2030, 341, 2016, 1410, 2019]. sparsity [2031]. spatial [1806, 922]. Special [1879, 651, 879, 626, 2206, 1588, 2261, 505, 1897, 2216, 2377, 1863, 1309, 1331, 242, 1727, 1526, 1217, 15, 1416, 2369, 2225, 2292, 285, 1012, 3, 68, 732, 1705, 1043]. Specializations [66]. specialized [523]. specifications [367, 332, 1776]. specified [1113]. Specker [2411]. spectral [2175, 1781, 1080]. spectrum [315]. Specular [1332]. speed [812, 1269, 1398, 434, 355]. speed-scaling [812]. speeds [1360]. sphere [778, 2082]. spherical [503]. spiders [2294]. spikes [1648]. Spiking [1648, 767, 1315, 100, 1460, 772, 2119]. Spiral [268]. splay [2078]. splicing [1107, 1649]. split [2154, 862, 1940, 239, 518, 464, 977, 1369, 1720, 60]. Split-Star [60]. split-state [1940]. splitting [2306, 251]. sponsored [1821]. Sprague [2391, 1699]. Spread [269]. spreading [1876, 1968]. Sprouts [936]. Spy [1655]. Spy-game [1655]. squad [2283]. Square [454, 1607, 1428, 1314, 547, 1333, 67, 2023, 1800, 458, 1350, 24, 1061]. Square-free [454, 67, 2023, 458, 1350]. squarefree [2398]. squares [291, 1681, 394, 1009, 982, 464, 977, 1087, 143, 2013, 2415]. stability [289, 59]. stabilization [1359, 793, 1267, 111, 1741]. stabilizing [1531, 2039, 2156, 797, 1181, 1892, 359, 1893, 196, 697, 325]. Stable [1016, 2075, 1639, 1227, 2185, 1213, 1204]. Stack [2243, 1051, 1751]. Stackelberg [66, 2399]. stage [755, 1354, 1090, 1715, 2202, 2085]. staged [1218]. stages [2176]. Standard [457, 560, 479, 850, 1106, 2344, 1511, 901, 1557, 388, 667, 1912, 2049]. Star [2069, 199, 1021, 696, 2394, 952, 1236, 1484, 1238, 2289, 1924, 1412, 1133, 60]. star-height [1021]. stars [1828, 2406, 910, 535]. State [2096, 2403, 642, 1316, 2379, 2381, 1284, 2378, 727, 186, 788, 1013, 1940, 1742, 994, 2409, 648, 434, 355, 2213, 1867, 1277, 1113, 649, 1442, 1176]. state-dependent [1867]. states [1445, 416]. static [282, 1404, 2056]. statistic [1087]. statistical [427, 986]. statistics [2220]. Steady [1445]. steganography [1379]. Steiner [2071, 1031]. step [932]. steps [1257]. stepwise [1807, 48]. Stern [217]. stick [2348]. Stochastic [1653, 1916, 279, 1373, 2052, 2107, 1731, 1784, 56, 1252]. storage [1088, 400, 2405]. store [1819]. stores [2379]. story [322]. Stoughton [1341]. straight [2230]. straight-line [2230]. strand [2190, 833]. strands [1223]. Strategies [574, 1729, 771, 1430, 488, 2184]. Strategy [1252, 961, 498, 1784, 1700, 2186]. stream [826, 259, 1602, 23]. streamed [2388]. streaming [1469, 1, 984]. streams [318, 5, 1329, 1328, 193, 602, 466]. strength [771, 939, 2412]. stretch [653, 1493, 314, 165]. stretching [469]. Strict [1225, 318]. strictly [155]. String [1417, 435, 2263, 1436, 1684, 1023, 958, 1663, 1539, 1540, 1543, 2112, 1546, 894, 230, 2323, 2084, 1271, 719, 2199]. Strings [1534, 1416, 1061, 1535, 2280, 1069, 1062, 31, 167, 437, 1420, 1530, 1545, 335, 1182, 1583, 63, 895]. Strip [652, 1241, 2201, 1152, 2229]. Strong [696, 657, 861, 1552, 1770, 1566, 789, 1672, 135, 84, 1509, 2289, 60, 1015]. strong-factor [135]. Strongly [1192, 2394, 1691, 353, 1022, 1619]. Structural [576, 1797, 2157, 651, 242]. Structure [1009, 2358, 853, 2307, 2320, 1553, 2258, 1632, 491, 2018, 1999, 550, 1808, 317, 1891, 612, 2137]. Structure-preserving [2307]. Structured [1521, 1300]. structures [1879, 1750, 339, 523, 240, 221, 560, 970, 1422, 674, 2026, 324, 1240, 1554, 2410, 1885, 1774, 1956, 85, 2354, 1265, 1038, 2003]. Studies [331]. Study [2075, 600, 491, 402, 78, 1040, 2078, 1093]. Sturmian [1793, 457, 858, 1791, 331, 1557, 217]. sub [631, 2378]. sub-hypergraphs [631]. sub-linear [2378]. subclass [1127]. subclasses [439, 2212, 175, 813]. subcubic [1578]. subdirectly [1498]. subdivided [1828]. subdivision [42, 346]. Subexponential [511, 539]. subexponential-time [539]. subexponentials [1386]. Subgraph [40, 1411, 619, 501, 362]. subgraphs [758, 517, 1248]. Subgroup [2319].

Sublinear [20, 473]. sublinear-time [473]. Submodular [1787, 1138, 819]. submodularity [2203]. subnetworks [233]. subsequence [1537, 2365, 2197, 2285, 1522, 1353]. Subset [1274, 2295, 2030, 2193, 173, 2329]. Subset-Cover [1274]. subshifts [1151, 2111]. substantially [501]. substitution [1341]. substring [891, 1364, 64, 1067, 890, 1891]. substring-substring [1891]. substrings [1438, 1353]. substructure [2358, 853]. subsumption [2282]. subsystem [1211]. subtraction [226, 1131, 808]. subtree [310, 1427]. subtrees [209]. subtyping [1636]. Subword [2268, 1021]. subwords [648]. successive [160]. Successor [589]. Succinct [364, 44, 2137, 1060, 1433, 1729]. Succinctness [1122]. Sufficient [1869, 565, 96]. suffix [1314, 896, 884, 1419, 350, 1271, 380]. suffixes [891]. suitable [1109]. sum [2030, 871, 590, 1072, 596, 917, 1148, 2329]. summands [1519]. Sums [1519, 1149, 1699, 336]. sun [464]. Super [2276, 1569]. superbubble [614]. supercomputing [17]. supergrid [472]. superwords [648]. supply [1084]. supported [1498]. supporting [2349]. suprema [996]. surely [13]. surfaces [1761]. surveillance [252]. Survey [1277, 1119, 1797, 1540, 643, 2409, 1118]. survivable [1236]. survival [1185]. survivor [2293]. Survivors [1044]. SVM [2353]. swap [1903]. Swapping [271, 1676]. swaps [63]. Swarm [2045, 13, 420]. Sweet [2178]. switch [27, 1585, 1614]. switches [298, 1079, 1246]. switching [281, 2109, 1867]. symbol [1179]. Symbolic [2097, 1203, 2094, 1500, 1545, 101]. symbols [2146, 1313, 1921]. Symmetric [507, 1479, 225, 804, 1752]. symmetries [2406]. synapses [1648]. synchronised [987]. synchronizability [1054].

Synchronization [2211, 1357].

synchronized [1029]. Synchronizing [690, 1171, 1514, 1022]. synchronous [1531, 822]. Syntactic [2410, 2210]. Synthesis [1563, 376, 1203, 1731, 1840, 1867]. Synthesising [1729]. synthetic [685]. System [279, 1286, 2287, 1886, 1200, 2284, 46, 1688, 1186, 1818, 489, 2304, 637]. Systems [426, 2002, 641, 1286, 1310, 1444, 418, 767, 281, 2149, 414, 768, 1445, 563, 769, 1744, 1054, 1295, 1734, 753, 1010, 1972, 1797, 2094, 1869, 491, 1679, 1737, 280, 1495, 1315, 1669, 1450, 1728, 561, 422, 1317, 430, 1318, 1162, 1451, 1651, 1806, 831, 2191, 1961, 146, 1453, 1258, 2220, 1091, 1998, 336, 1957, 1454, 2242, 1455, 2323, 742, 1456, 2192, 1807, 1838, 1842, 1532, 928, 100, 1386, 922, 1220, 192, 766, 1231, 1649, 1460, 153, 796, 375, 932, 413, 1325, 378, 103, 923, 764, 1931, 169, 1186, 2060, 1461, 1738, 772, 673, 1306, 1648, 2119, 616, 926, 1647, 1499]. systems [2061, 1962, 2377].

table [1062, 1063]. tableaux [481]. tables [1120]. tailor [2188]. take [580, 1443]. tame [962]. tape [478]. tardiness [1820]. tardy [229]. target [2317, 2112, 1697]. Tarski [2417, 1698, 1523]. task [1505]. tasks [80, 320, 2231, 1596]. Taxi [1824]. Taxi-sharing [1824]. TCS [323, 1309, 1331, 322, 1217, 1293]. teacher [1706]. teachers [1016]. teaching [1595]. Team [1630, 1600]. technique [1642, 643, 2016, 819]. techniques [1289, 469]. technology [1277]. Temperature [1150]. tempering [419]. template [1840]. template-based [1840]. temporal [279, 1974, 2370, 81, 629, 2219, 2375, 374, 847, 922, 932, 2060, 1374, 2373, 637, 2061, 2369]. Ten [1443]. tensor [519]. Term [821, 178, 2105, 1700]. Term-generic [178]. Terminal [2070, 1408].

Terminal-pairability [2070]. Termination [661, 135, 1183, 1298]. terms [661, 1562, 233]. ternary [2139, 1185, 1994, 585]. terrain [1153]. terrains [387, 1584]. Test [70, 1018, 204, 1159, 2304, 1912]. testability [1109]. testable [1322, 1559]. Testing [1812, 566, 2129, 334, 1651, 200, 1660, 493, 461, 1093, 429]. tetrominoes [1574]. text [1542, 1066, 1958]. theatregoer [272]. their [722, 2267, 132, 1247, 258, 1021, 1760, 1106, 402, 1540, 547, 1556, 2133, 1244, 1791, 1140, 2374]. Theor [2418, 2417, 2415, 2024]. Theorem [1345, 789, 898, 1708, 154, 777, 359, 1839, 413]. theorems [442, 2004, 1259, 233]. Theoret [1023, 474, 1329, 1354, 984, 68, 2416, 1496, 1004, 1486]. theoretic [257, 1373, 453, 336, 1710, 922, 2118]. Theoretical [1416, 1802]. theoretically [2308]. theories [80, 1227, 1747, 263, 1637]. Theory [2292, 1043, 660, 309, 732, 1309, 2414, 2371, 2188, 1447, 2147, 2099, 1561, 1520, 155, 1838, 1105, 101, 2065, 489, 340, 1705, 285, 685]. Therapy [526]. there [2033]. thermodynamic [2185]. thesis [987]. theta [703]. theta-graphs [703]. thin [2349, 781]. think [2106]. Three [394, 961, 100, 814, 2073, 226, 592, 2176, 389]. three-element [226]. Three-pile [961]. three-vertex [592]. Threes [1847]. threshold [1611, 250, 624, 1083, 947, 818, 85, 588]. thresholds [1194, 1231]. throttling [2314]. Throughput [812, 1079, 1688]. Thue [2141, 356, 231]. ties [1639]. Tight [851, 613, 777, 872, 2125, 1902, 703, 225, 404]. Tighter [462, 1072, 2353, 1955]. Tightly [2329, 2322, 2072]. tile [1191, 1459]. Tiling [24, 1574]. tilings [1188, 1452, 415]. TIME [2369, 1571, 1193, 1374, 666, 2189, 367, 1183, 1916, 2227, 1872, 1062, 1734, 606, 627, 74, 1150, 614, 2370, 1967, 312, 473, 381, 238, 755, 1669, 1989, 1538, 269, 1149, 900, 444, 2300, 687, 2018, 259, 827, 376, 441, 1438, 1860, 246, 1729, 911, 2375, 1953, 2384, 949, 224, 1612, 1529, 961, 815, 64, 1755, 704, 1659, 1914, 1203, 951, 539, 1365, 740, 2335, 461, 464, 115, 237, 1716, 330, 1517, 57, 1816, 494, 1633, 932, 2013, 2415, 714, 969, 293, 2352, 1468]. time-constrained [367]. time-dependent [1633]. time-optimal [64]. time-sensitive [666]. Time-space [1571, 2227]. time-varying [312, 969]. Timed [788, 1183, 1100, 71, 1162, 1203, 1731, 103]. times [755, 2350, 520, 1633, 141]. tissue [1310, 1454, 923]. tissue-like [923]. TNFS [2407]. tokens [655, 444, 271, 1676]. tolerable [2342]. tolerance [1484, 1948, 1347, 2320, 1553, 227, 1928, 2241, 2170, 2130, 618]. tolerant [1479, 1531, 212, 93, 1672, 27, 1361, 300, 1509, 319, 2354, 807, 1401]. Tomographic [2108]. tomography [780]. toolbox [1232]. tools [1223, 2318]. top [701]. top-down [701]. topic [3, 68]. topics [1743]. Topological [735, 838, 825, 503, 1297, 1794, 1997]. topologies [1655, 1287, 1114]. Topology [749, 1558]. toroidal [2293]. torus [390, 861, 1349, 1262]. Total [1431, 1468, 2313, 701, 1820, 1493, 1520, 596, 2390, 917, 1148]. tournaments [2184]. Tower [1849, 1473, 620]. Trace [1730, 876, 1315]. traces [70, 1109]. Track [1526]. tracking [1775, 546]. Tractability [176, 1709, 1582, 2362, 1023, 958, 2103, 1930, 173, 1832, 1122]. tractable [1974, 1492, 131, 1242]. trade [2227, 1571, 1438, 314, 650]. trade-off [2227]. trade-offs [1571, 1438, 314, 650]. tradeoff [74]. Trading [320, 516]. traffic [1604, 445, 1810]. trails [2395]. Trainyard [1851]. transactional [1361, 1362]. Transducers [1311, 1103, 701, 2115]. Transducing [2213]. transductions [1703]. transfer [1575, 1523, 840]. transform [1540, 1059]. transformation [366]. transformations [792]. transformers [1221]. transforming [1637]. transition [1286, 1869, 1669]. transitions [841]. transitive [1501, 370]. translation [2258]. translations [182]. translocations [167, 529]. transmissions [2349]. transmitting [1643]. transport [567]. Transposing [1772]. transpositions [462, 1989, 1140]. transversals [1492, 402, 1996]. trapdoor [345]. traveler [1697]. Traveling [847, 1653, 29]. treatment [1778, 567]. Tree [938, 795, 1856, 942, 1794, 548, 66, 1967, 966, 1756, 2331, 1108, 701, 1674, 2250, 1541, 127, 2079, 1805, 2214, 324, 951, 884, 1943, 1716, 941, 1434, 191, 943, 2176, 669, 2027, 798, 1485, 371, 1700, 1995, 535, 540]. Tree-automatic [795]. tree-based [1434]. tree-searching [2176]. trees [304, 2038, 213, 2029, 313, 1399, 2154, 707, 889, 712, 1527, 55, 1060, 2299, 654, 444, 2355, 1429, 2036, 176, 759, 1433, 370, 1194, 538, 1490, 224, 1028, 528, 552, 957, 1694, 1127, 1943, 1084, 1520, 1031, 888, 2270, 165, 1008, 2176, 2312, 2078, 1489, 747, 44, 2137, 380, 1907, 1209, 209, 1080, 620, 2121, 792, 585]. Treewidth [1694, 2029, 121, 316, 744]. trellis [1378]. triadic [1770]. Triangle [2217, 1207, 1610, 1074, 1622, 2392]. triangle-free [2392]. triangles [5, 1329, 1328, 2033, 1931, 1148]. triangulations [2302]. trie [1542]. tries [896, 635]. trigger [2189]. Triggering [353]. Trimming [187, 1579]. trios [2393]. trivial [2211]. tropical [1256]. true [322]. truncated [380]. Truthful [145, 51, 1041]. TSP [2032, 1074]. Turing [478, 2012, 441, 412, 2127]. turn [1208]. turns [190]. turnstile [318]. Tutte [423]. Tversky [1610]. twin [872]. twisted [1515, 695, 2358, 1413]. Two [1357, 1096, 332, 2050, 883, 2325, 755, 1747, 2002, 1475, 401, 463, 1064, 2357, 1086, 2352, 540, 1132, 1168, 851, 4, 474, 1006, 1198, 167, 1103, 1528, 2136, 2355, 1354, 1090, 1905, 1007, 1660, 567, 229, 2229, 2311, 245, 1514, 2132, 1008, 2176, 650, 2013, 2415, 1242, 534, 2200, 37, 2202, 2085, 1995, 1584, 1468]. Two-agent [1475, 2352]. two-coloring [1242]. two-compartment [567]. Two-dimensional [1096, 1132, 650]. two-level [37]. Two-machine [2357, 229]. two-option [245]. two-page [1660]. Two-process [1357]. Two-sided [332, 534]. Two-stage [755, 1354, 1090, 2202, 2085]. two-stages [2176]. two-tree [1995]. two-way [1103]. type [1167, 1483, 509, 2099, 1561, 142, 1689]. Typed [2151]. types [1749, 1205, 814, 1772]. typography [273].

UDGs [697]. ultimate [292, 616]. Ultrafilters [645]. umpire [29]. Unambiguous [1179, 1283, 2381, 92, 1619]. Unary [2159, 1795, 1399, 1495, 649]. unary-binary [1399]. Unavoidable [1107]. unaware [574]. unbalanced [638]. Unbiasedness [2187]. unbounded [311, 516, 794, 714]. uncertain [2129, 90, 587, 470]. uncertainty [2370, 679, 2373]. unconstrained [1134]. Uncut [1858]. Undecidability [1285, 1636, 1384]. undecidable [2113, 2364]. undependable [17]. understanding [1257]. undirected [163, 437, 1850, 59]. unexpected [1751, 960]. unfolding [1983]. unidirectional [867, 710]. unification [1143]. unified [382, 977, 1637]. Uniform [479, 1230, 320, 395, 908, 1006, 966, 72, 690, 287, 119, 1769, 111, 1861, 206, 1468]. uniformization [1916]. uniformization-based [1916]. uniformly [1114]. union [466]. Unique [2000, 668, 672, 1438, 658, 1364, 64]. uniqueness [1975]. unit [1241, 145, 447, 134]. unit-demand [145]. univariate [1301]. Universal [1222, 1570, 1958, 2012, 1789, 286, 100, 2067]. universality [103]. unknown [655, 1528, 253, 2112, 1414]. unordered [870]. Unpaired [713]. unpredictable [1909]. unranking [2249]. Unraveling [920]. unrelated [125, 959, 704]. unreliable [522]. unsharp [412]. unshuffling [1678]. unsigned [529, 2047]. Unsupervised [845]. until [663]. Unveiling [181]. unweighted [163, 757]. Unwinding [280]. updatable [1193]. update [990, 1384]. updates [1530]. updating [538]. UPGMA [1628]. Upgrading [1485]. Upper [1034, 2412, 1599, 462, 869, 1045, 1616, 2042, 941, 1890]. Upward [132]. urban [1810]. useful [1478]. user [547, 1344, 2072]. user-defined [1344]. Using [942, 63, 1841, 279, 562, 1737, 1882, 2232, 429, 1492, 1450, 763, 1341, 2331, 502, 1500, 953, 259, 2108, 1052, 204, 2309, 428, 2341, 1731, 1511, 1246, 894, 1918, 1459, 423, 2412, 2048, 568, 1035, 420, 819]. utilities [1474]. Utilizing [1880]. UTP [1917].

Vaccinate [2036]. valid [2074]. valuation [73, 371]. valuations [1313, 852]. value [506, 823, 1602, 1033, 88]. valued [1623, 481, 2004]. values [448, 1544, 1072]. Vandermonde [2316]. variability [1269]. Variable [1304, 164, 27, 229, 493, 2412, 1752]. Variables [2002, 1210, 1323, 754, 926]. variance [325]. variant [1008]. variants [1850, 756, 2226, 1324]. variation [1431]. variational [1997]. Variations [1751, 161, 2178]. varieties [1297]. various [1411, 219, 1485]. varying [312, 969]. VC [1987, 1596, 1783]. vector [1623, 327, 2322, 779, 1939]. vectors [885, 2161]. Vegas [743]. vehicle [553]. vehicles [1857, 2223]. verifiable [1686]. Verification [418, 666, 767, 826, 1651, 1159, 131, 832]. verifier [1683]. verifier-local [1683].

Verifying [1500, 441, 1979, 1981]. version [1044, 2278, 461, 2150, 389]. versus [1609, 1018, 239, 179, 564]. Vertex [1817, 2321, 1137, 2063, 600, 1507, 948, 515, 83, 2250, 1691, 1138, 2125, 592, 300, 471, 130, 1764, 2028, 698, 534, 2164, 973, 134, 1092, 819, 1876]. Vertex-Congest [1876]. Vertex-disjoint [2321, 130]. vertex-fault-tolerant [300]. vertex-weighted [534]. Vertical [89]. vertices [1312, 2321, 2311, 807]. via [418, 1399, 1922, 1295, 523, 1635, 948, 2365, 2052, 2105, 771, 2341, 343, 2408, 841, 832, 2343, 1719, 1423, 2160, 1917, 218, 692, 234, 2228, 37, 1861, 2304]. view [510, 2076, 632, 369, 1855, 922, 2118]. views [225, 1053]. violations [633]. virtual [1713, 253, 1409, 545, 742]. viruses [770]. Visibility [2129, 2228, 1508, 2173, 2227, 527, 953, 162]. visible [1637]. Visibly [1843, 187, 1051]. Visual [85, 595]. Vitányi [2024]. Volume [1152, 1798]. Volumetric [526]. Voronoi [41, 2277]. Vortex [2045]. vote [522]. voter [305]. votes [2168]. voting [706, 1255, 1666]. VPDAs [1836]. vulnerability [1393].

wait [1327]. wait-free [1327]. wake [570, 691]. wake-up [691]. WALCOM [2225]. Walking [577]. Walks [1452, 211, 1418, 218, 1431, 1971]. Wallace [2047]. Watson [1495, 2159, 1453]. Watson-Crick [1495]. wavelet [888]. way [2383, 743, 345, 516, 1103, 1827, 1321, 118]. Weak [1961, 1923, 74, 1685, 509, 421, 668, 1796]. weakest [486]. weakly [1239, 270, 724, 287]. web [1263, 1624, 1888]. Weierstrass [46]. weight [2230, 1756, 1909, 2367, 2059, 916, 2399, 134]. Weighted [1677, 2029, 1803, 2154, 2104, 2165, 857, 813, 767, 1594, 1062, 1428, 1312, 2313, 269, 2209, 578, 370, 224, 2246, 2335, 1766, 350, 1583, 1437, 2395, 1280, 2027, 747, 534, 757, 2009]. weights [992, 1235, 540]. Well [1632, 477, 1829]. well-covered [1829]. Well-ordering [1632]. Wheeler [1540, 1422, 2162]. whether [2402]. Which [1974, 2254, 248, 307]. white [399, 595]. who [1247]. whole [1425]. Wide [1888]. width [1656, 304, 1049, 1757, 1607, 2155, 2300, 700, 2363, 2229, 1478, 47, 1136, 1158, 698]. widths [250]. wild [1281]. wildcard [864]. Wildcards [435]. Window [2343, 246, 124, 1374]. windows [1860]. winner [706]. winning [961]. wirelength [390]. wireless [164, 1381, 545, 546, 1558, 314, 1434, 656, 1401, 2055]. within [1241]. within-strip [1241]. without [2030, 654, 999, 2122, 603, 1240, 1137, 2021, 201, 2211, 868, 1242, 36, 58, 1148, 1245, 1762]. Witness [1298, 1306]. witnesses [2043]. WK [58, 1569]. WK-recursive [58, 1569]. Wojciech [2261]. WoLLIC [476]. Word [1864, 622, 186, 1463, 436, 193, 1788, 2327, 356, 982, 2265, 2264]. Words [450, 1331, 2267, 2142, 2272, 1803, 291, 1098, 2177, 972, 860, 364, 395, 908, 2288, 2254, 1130, 1793, 457, 1103, 1747, 1792, 1420, 621, 2398, 1106, 451, 2102, 1077, 1556, 89, 858, 1333, 2271, 2139, 645, 1570, 1109, 67, 454, 2023, 2249, 1335, 357, 2235, 1791, 1547, 1696, 2117, 1625, 331, 456, 1064, 1557, 458, 378, 228, 94, 1350, 39, 1403, 566, 217, 792]. work [1047]. workers [17]. Working [725, 1647]. Workshops [2225, 505]. world [270, 1775, 1888]. Worst [408, 306, 247, 1488, 102]. Worst- [408]. worst-case [247, 102]. WSTS [876]. Wythoff [467]. Wythoff-like [467].

XML [193]. XPath [193]. xxx [621].

Yung [1382].

Zero [1939, 1021, 527, 1724, 382, 1303, 1146, 2200]. Zero-knowledge [1939]. zeromatrix [1021]. zero-visibility [527]. Zhang [969]. Zimin [133]. Ziv [1571, 2235]. zombies [2293, 1044]. zoom [560].

References

Feigenblat:2010:FIP

[1] Guy Feigenblat, Ofra Itzhaki, and Ely Porat. The frequent items problem, under polynomial decay, in the streaming model. Theoretical Computer Science, 411(34–36):3048–3054, July 17, 2010. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See corrigendum [984].

Colussi:2011:CCC

[2] Livio Colussi. The convergence classes of Collatz function. Theoretical Computer Science, 412(39):5409–5419, September 9, 2011. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). See comment [725].

Hosoda:2012:AHM

[3] Jun Hosoda, Juraj Hromkovic, Taisuke Izumi, Hirotaka Ono, Monika Steinová, and Koichi Wada. On the approximability and hardness of minimum topic connected overlay and its special instances. Theoretical Computer Science, 429(1):144– 154, April 20, 2012. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                              //www.sciencedirect.com/science/
                                              article/pii/S0304397511009996. See

corrigendum [68].

Calamoneri:2014:PCG

[4] T. Calamoneri and R. Petreschi. On pairwise compatibility graphs having Dilworth number two. Theoretical Computer Science, 524(??): 34–40, March 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397513009274. See

corrigendum [474].

Cormode:2014:SLC

[5] Graham Cormode and Hossein Jowhari. A second look at counting triangles in graph streams. Theoretical Computer Science, 552(??):44– 51, October 2, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005945. See

corrigendum [1329].

Pai:2014:ICH

[6] Kung-Jui Pai, Jou-Ming Chang, Jinn-Shyong Yang, and Ro-Yu Wu. Incidence coloring on hypercubes. Theoretical Computer Science, 557(??): 59–65, November 6, 2014. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006380. See

corrigendum [1004].

Anonymous:2015:EBa

[7] Anonymous. Editorial Board. Theoretical Computer Science, 561 (part A) (??):iii–ix, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009049.

Anonymous:2015:EBb

[8] Anonymous. Editorial Board. Theoretical Computer Science, 561 (part B) (??):iii–ix, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514009189.

Friedrich:2015:GEC

[9] Tobias Friedrich, Jun He, Thomas Jansen, and Alberto Moraglio. Genetic and evolutionary computation. Theoretical Computer Science, 561 (part A)(??):1–2, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514009116.

Doerr:2015:OLF

[10] Benjamin Doerr and Marvin Künnemann. Optimizing linear functions with the (1 + λ) evolutionary algorithm — different asymptotic runtimes for different instances. Theoretical Computer Science, 561 (part A)(??):3– 23, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514002060.

Nguyen:2015:PSM

[11] Anh Quang Nguyen, Andrew M. Sutton, and Frank Neumann. Population size matters: Rigorous runtime results for maximizing the hypervolume indicator. Theoretical Computer Science, 561 (part A)(??): 24–36, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004599.

Oliveto:2015:ADM

[12] Pietro S. Oliveto and Christine Zarges. Analysis of diversity mechanisms for optimisation in dynamic environments with low frequencies of change. Theoretical Computer Science, 561 (part A) (??):37–56, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008202.

Schmitt:2015:PSO

[13] Manuel Schmitt and Rolf Wanka. Particle swarm optimization almost surely finds local optima. Theoretical Computer Science, 561 (part A) (??):57–72, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004150.

Lissovoi:2015:RAA

[14] Andrei Lissovoi and Carsten Witt. Runtime analysis of ant colony optimization on dynamic shortest path problems. Theoretical Computer Science, 561 (part A)(??):73–85, January 4, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004952.

Raynal:2015:SID

[15] Michel Raynal and Franck Petit. Special issue on distributed computing and networking. Theoretical Computer Science, 561 (part B) (??):87, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009256.

Afek:2015:SCA

[16] Yehuda Afek and Eli Gafni. A simple characterization of asynchronous computations. Theoretical Computer Science, 561 (part B)(??):88– 95, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005659.

Davtyan:2015:DUW

[17] Seda Davtyan, Kishori M. Konwar, Alexander Russell, and Alexander A. Shvartsman. Dealing with undependable workers in decentralized network supercomputing. Theoretical Computer Science, 561 (part B)(??): 96–112, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007774.

Sarma:2015:FDP

[18] Atish Das Sarma, Anisur Rahaman Molla, Gopal Pandurangan, and Eli Upfal. Fast distributed PageRank computation. Theoretical Computer Science, 561 (part B)(??):113– 121, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002709.

Delporte-Gallet:2015:LSB

[19] Carole Delporte-Gallet, Hugues Fauconnier, Eli Gafni, and Sergio Rajsbaum. Linear space bootstrap communication schemes. Theoretical Computer Science, 561 (part B)(??):122– 133, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007750.

Kutten:2015:SBR

[20] Shay Kutten, Gopal Pandurangan, David Peleg, Peter Robinson, and Amitabh Trehan. Sublinear bounds for randomized leader election. Theoretical Computer Science, 561 (part B)(??): 134–143, January 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001029.

Anonymous:2015:EBc

[21] Anonymous. Editorial Board. Theoretical Computer Science, 562(??): iii–ix, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400930X.

Liu:2015:IDS

[22] Ching-Hao Liu, Sheung-Hung Poon, and Jin-Yong Lin. Independent dominating set problem revisited. Theoretical Computer Science, 562(??): 1–22, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006537.

Giannakopoulos:2015:CAM

[23] Yiannis Giannakopoulos and Elias Koutsoupias. Competitive analysis of maintaining frequent items of a stream. Theoretical Computer Science, 562(??): 23–32, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006768.

Amir:2015:PST

[24] Amihood Amir, Alberto Apostolico, Gad M. Landau, Ely Porat, and Oren Sar Shalom. A PTAS for the Square Tiling Problem. Theoretical Computer Science, 562(??): 33–45, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400677X.

Fu:2015:EHD

[25] Hao Fu and Guo-Niu Han. On textensions of the Hankel determinants of certain automatic sequences. Theoretical Computer Science, 562(??): 46–56, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007105.

Wang:2015:OFE

[26] Yang Wang, Man Ho Au, and Willy Susilo. Optimistic fair exchange in the enhanced chosen-key model. Theoretical Computer Science, 562(??): 57–74, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007117.

Delmas:2015:DFT

[27] O. Delmas, F. Havet, M. Montassier, and S. Pérennes. Design of fault-tolerant on-board networks with variable switch sizes. Theoretical Computer Science, 562(??): 75–89, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007129.

Fanelli:2015:RDG

[28] Angelo Fanelli, Dariusz Leniowski, Gianpiero Monaco, and Piotr Sankowski. The ring design game with fair cost allocation. Theoretical Computer Science, 562(??):90–100, January 11, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007130.

deOliveira:2015:CTU

[29] Lucas de Oliveira, Cid C. de Souza, and Tallys Yunes. On the complexity of the traveling umpire problem. Theoretical Computer Science, 562(??): 101–111, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007270.

Chen:2015:EAS

[30] Xiaofeng Chen, Willy Susilo, Jin Li, Duncan S. Wong, Jianfeng Ma, Shaohua Tang, and Qiang Tang. Efficient algorithms for secure outsourcing of bilinear pairings. Theoretical Computer Science, 562(??):112– 121, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007282.

Bu:2015:CBS

[31] Tian-Ming Bu, Chen Yuan, and Peng Zhang. Computing on binary strings. Theoretical Computer Science, 562(??): 122–128, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007294.

Gadouleau:2015:MCN

[32] Maximilien Gadouleau and Søren Riis. Memoryless computation: New results, constructions, and extensions. Theoretical Computer Science, 562(??): 129–145, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007300.

You:2015:OOD

[33] Lantao You, Jianxi Fan, Yuejuan Han, and Xiaohua Jia. One-to-one disjoint path covers on alternating group graphs. Theoretical Computer Science, 562(??):146– 164, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007312.

Bodlaender:2015:EAK

[34] Hans L. Bodlaender, Dieter Kratsch, and Sjoerd T. Timmer. Exact algorithms for Kayles. Theoretical Computer Science, 562(??):165– 176, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007324.

Huang:2015:AOF

[35] Qiong Huang, Guomin Yang, Duncan S. Wong, and Willy Susilo. Ambiguous optimistic fair exchange: Definition and constructions. Theoretical Computer Science, 562(??):177– 193, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007336.

Tan:2015:CSR

[36] Xiao Tan, Qiong Huang, and Duncan S. Wong. Concurrent signature without random oracles. Theoretical Computer Science, 562(??):194– 212, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007348.

Wu:2015:PDA

[37] Chenchen Wu, Donglei Du, and Dachuan Xu. Primal-dual approximation algorithm for the two-level facility location problem via a dual quasi-greedy approach. Theoretical Computer Science, 562(??):213– 226, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400735X.

Gavruskin:2015:DAM

[38] Alexander Gavruskin, Bakhadyr Khoussainov, Mikhail Kokho, and Jiamou Liu. Dynamic algorithms for monotonic interval scheduling problem. Theoretical Computer Science, 562(??): 227–242, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007361.

Wei:2015:SCW

[39] Jianxin Wei and Heping Zhang. Solution to a conjecture on words that are bad and 2-isometric. Theoretical Computer Science, 562(??):243– 251, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007634.

Heggernes:2015:ISI

[40] Pinar Heggernes, Pim van ’t Hof, Daniel Meister, and Yngve Villanger. Induced Subgraph Isomorphism on proper interval and bipartite permutation graphs. Theoretical Computer Science, 562(??):252– 269, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007646.

Bandyapadhyay:2015:VGG

[41] Sayan Bandyapadhyay, Aritra Banik, Sandip Das, and Hirak Sarkar. Voronoi game on graphs. Theoretical Computer Science, 562(??):270– 282, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007658.

Bang-Jensen:2015:FSD

[42] Jørgen Bang-Jensen, Frédéric Havet, and A. Karolinna Maia. Finding a subdivision of a digraph. Theoretical Computer Science, 562(??):283– 303, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400766X.

DAlessandro:2015:CEBa

[43] Flavio D’Alessandro and Benedetto Intrigila. On the commutative equivalence of bounded context-free and regular languages: the code case. Theoretical Computer Science, 562(??): 304–319, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007671.

Tsur:2015:SRL

[44] Dekel Tsur. Succinct representation of labeled trees. Theoretical Computer Science, 562(??):320– 329, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007683.

Bermudo:2015:CSK

[45] Sergio Bermudo and Henning Fernau. Combinatorics for smaller kernels: the differential of a graph. Theoretical Computer Science, 562(??): 330–345, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007695.

Ruatta:2015:OWI

[46] Olivier Ruatta, Mark Sciabica, and Agnes Szanto. Overdetermined Weierstrass iteration and the nearest consistent system. Theoretical Computer Science, 562(??):346– 364, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007701.

Kintali:2015:AAD

[47] Shiva Kintali, Nishad Kothari, and Akash Kumar. Approximation algorithms for digraph width parameters. Theoretical Computer Science, 562(??): 365–376, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007713.

Seddik:2015:PGS

[48] Yasmina Seddik, Christophe Gonzales, and Safia Kedad-Sidhoum. Performance guarantees for a scheduling problem with common stepwise job payoffs. Theoretical Computer Science, 562(??):377–394, January 11, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007725.

Han:2015:RAO

[49] Xin Han, Yasushi Kawase, and Kazuhisa Makino. Randomized algorithms for online knapsack problems. Theoretical Computer Science, 562(??): 395–405, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007798.

Oliveira:2015:PEP

[50] Leonardo I. L. Oliveira, Valmir C. Barbosa, and Fábio Protti. The predecessor-existence problem for kreversible processes. Theoretical Computer Science, 562(??):406–418, January 11, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007804.

Calinescu:2015:BPA

[51] Gruia Calinescu. Bounding the payment of approximate truthful mechanisms. Theoretical Computer Science, 562(??):419–435, January 11, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007816.

Gast:2015:IDS

[52] Mikael Gast, Mathias Hauptmann, and Marek Karpinski. Inapproximability of dominating set on power law graphs. Theoretical Computer Science, 562(??):

436–452, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400783X.

Bermond:2015:DGP

[53] Jean-Claude Bermond, Bi Li, Nicolas Nisse, Hervé Rivano, and Min-Li Yu. Data gathering and personalized broadcasting in radio grids with interference. Theoretical Computer Science, 562(??):453–475, January 11, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008214.

DAlessandro:2015:CES

[54] Flavio D’Alessandro and Benedetto Intrigila. On the commutative equivalence of semi-linear sets of N^{k}. Theoretical Computer Science, 562(??): 476–495, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008226.

Chen:2015:PAA

[55] Jianer Chen, Jia-Hao Fan, and Sing-Hoi Sze. Parameterized and approximation algorithms for maximum agreement forest in multifurcating trees. Theoretical Computer Science, 562(??): 496–512, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008238.

Levin:2015:BAS

[56] Asaf Levin and Aleksander Vainer. The benefit of adaptivity in stochastic packing problems with probing. Theoretical Computer Science, 562(??): 513–523, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400824X.

Mertzios:2015:OBT

[57] George B. Mertzios, Mordechai Shalom, Ariella Voloshin, Prudence W. H. Wong, and Shmuel Zaks. Optimizing busy time on parallel machines. Theoretical Computer Science, 562(??): 524–541, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008251.

Wang:2015:HBW

[58] Yi-Chun Wang and Justie Su-Tzu Juan. Hamiltonicity of the basic WK-recursive pyramid with and without faulty nodes. Theoretical Computer Science, 562(??):542– 556, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008263.

Disser:2015:IBP

[59] Yann Disser, Andreas Emil Feldmann, Max Klimm, and Matús Mihalák. Improving the H_{k}-bound on the price of stability in undirected Shapley network design games. Theoretical Computer Science, 562(??):

557–564, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008299.

Lin:2015:CDS

[60] Limei Lin, Li Xu, and Shuming Zhou. Conditional diagnosability and strong diagnosability of Split-Star Networks under the PMC model. Theoretical Computer Science, 562(??):565– 580, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400838X.

Fan:2015:ISP

[61] Jing Fan, Xiwen Lu, and Peihai Liu. Integrated scheduling of production and delivery on a single machine with availability constraint. Theoretical Computer Science, 562(??): 581–589, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008391.

Jing:2015:PA

[62] Maohua Jing, Yixian Yang, Ning Lu, Wenbo Shi, and Changyong Yu. Postfix automata. Theoretical Computer Science, 562(??):590– 605, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008421.

Meister:2015:USD

[63] Daniel Meister. Using swaps and deletes to make strings match. Theoretical Computer Science, 562(??): 606–620, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008457.

Ileri:2015:SYT

[64] Atalay Mert Ileri, M. Oguzhan Külekci, and Bojian Xu. A simple yet timeoptimal and linear-space algorithm for shortest unique substring queries. Theoretical Computer Science, 562(??): 621–633, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008470.

Couturier:2015:NMD

[65] Jean-François Couturier, Romain Letourneur, and Mathieu Liedloff. On the number of minimal dominating sets on some graph classes. Theoretical Computer Science, 562(??): 634–642, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008494.

Bilo:2015:SGS

[66] Davide Bilò, Luciano Gualà, Stefano Leucci, and Guido Proietti. Specializations and generalizations of the Stackelberg minimum spanning tree game. Theoretical Computer Science, 562(??): 643–657, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008524.

Harju:2015:NSP

[67] Tero Harju and Mike Müller. A note on short palindromes in square-free words. Theoretical Computer Science, 562(??): 658–659, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008329.

Hosoda:2015:CAH

[68] Jun Hosoda, Juraj Hromkovic, Taisuke Izumi, Hirotaka Ono, Monika Steinová, and Koichi Wada. Corrigendum to “On the approximability and hardness of minimum topic connected overlay and its special instances” [Theoret. Comput. Sci. 429 (2012) 144–154]. Theoretical Computer Science, 562(??): 660–661, January 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008445. See

[3].

Anonymous:2015:EBd

[69] Anonymous. Editorial Board. Theoretical Computer Science, 563(??): iii–ix, January 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514009621.

Cavalcanti:2015:TST

[70] Ana Cavalcanti and Marie-Claude Gaudel. Test selection for traces refinement. Theoretical Computer Science, 563(??):1–42, January 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514006331.

Bouyer:2015:RRT

[71] Patricia Bouyer, Nicolas Markey, and Ocan Sankur. Robust reachability in timed automata and games: a game-based approach. Theoretical Computer Science, 563(??):43– 74, January 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006355.

DAgostino:2015:BQU

[72] Giovanna D’Agostino and Giacomo Lenzi. Bisimulation quantifiers and uniform interpolation for guarded first order logic. Theoretical Computer Science, 563(??):75–85, January 19, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006367.

Guan:2015:CMP

[73] Xuechong Guan and Yongming Li. On conditions for mappings to preserve optimal solutions of semiringinduced valuation algebras. Theoretical Computer Science, 563(??): 86–98, January 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007786.

Bernardo:2015:TBC

[74] Marco Bernardo. On the tradeoff between compositionality and exactness in weak bisimilarity for integratedtime Markovian process calculi. Theoretical Computer Science, 563(??): 99–143, January 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008172.

Anonymous:2015:EBe

[75] Anonymous. Editorial Board. Theoretical Computer Science, 564(??): iii–ix, January 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514009748.

Zhang:2015:BS

[76] Wenhui Zhang. Bounded semantics. Theoretical Computer Science, 564(??): 1–29, January 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008184.

Conradie:2015:ACI

[77] Willem Conradie, Yves Fomatati, Alessandra Palmigiano, and Sumit Sourabh. Algorithmic correspondence for intuitionistic modal mu-calculus. Theoretical Computer Science, 564(??): 30–62, January 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008196.

Dufourd:2015:FSF

[78] Jean-François Dufourd. Formal study of functional orbits in finite domains. Theoretical Computer Science, 564(??): 63–88, January 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008330.

Edalat:2015:DCL

[79] Abbas Edalat. A derivative for complex Lipschitz maps with generalised Cauchy–Riemann equations. Theoretical Computer Science, 564(??):89– 106, January 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008469.

Angiulli:2015:RDT

[80] Fabrizio Angiulli, Rachel Ben-Eliyahu-Zohary, and Luigi Palopoli. Restricted default theories: Expressive power and outlier detection tasks. Theoretical Computer Science, 564(??): 107–130, January 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008536.

DiGiampaolo:2015:PMI

[81] Barbara Di Giampaolo, Salvatore La Torre, and Margherita Napoli. Parametric metric interval temporal logic. Theoretical Computer Science, 564(??): 131–148, January 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008627.

Anonymous:2015:EBf

[82] Anonymous. Editorial Board. Theoretical Computer Science, 565(??): iii–ix, February 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400989X.

Fernau:2015:PCV

[83] Henning Fernau, Fedor V. Fomin, Geevarghese Philip, and Saket Saurabh. On the parameterized complexity of vertex cover and edge cover with connectivity constraints. Theoretical Computer Science, 565(??):1– 15, February 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008275.

Hellmuth:2015:CSF

[84] Marc Hellmuth and Tilen Marc. On the Cartesian skeleton and the factorization of the strong product of digraphs. Theoretical Computer Science, 565(??): 16–29, February 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008378.

Shyu:2015:VCR

[85] Shyong Jian Shyu. Visual cryptograms of random grids for threshold access structures. Theoretical Computer Science, 565(??):30– 49, February 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008408.

Faulkner:2015:FSA

[86] Nicholas Faulkner and Reem Yassawi. A family of sand automata. Theoretical Computer Science, 565(??): 50–62, February 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008482.

Gyorgyi:2015:RBS

[87] Péter Györgyi and Tamás Kis. Reductions between scheduling problems with non-renewable resources and knapsack problems. Theoretical Computer Science, 565(??):63– 76, February 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008500.

Punnen:2015:AVS

[88] Abraham P. Punnen, Piyashat Sripratak, and Daniel Karapetyan. Average value of solutions for the bipartite boolean quadratic programs and rounding algorithms. Theoretical Computer Science, 565(??):77– 89, February 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008512.

Fedou:2015:VRW

[89] Jean-Marc Fédou and Gabriele Fici. Vertical representation of C^{∞}-words. Theoretical Computer Science, 565(??): 90–101, February 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008548.

Kasperski:2015:COP

[90] Adam Kasperski and Paweł Zieliński. Combinatorial optimization problems with uncertain costs and the OWA criterion. Theoretical Computer Science, 565(??):102–112, February 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397514008561.

Anonymous:2015:EBg

[91] Anonymous. Editorial Board. Theoretical Computer Science, 566(??): iii–ix, February 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010020.

Okhotin:2015:DCU

[92] Alexander Okhotin and Kai Salomaa. Descriptional complexity of unambiguous input-driven pushdown automata. Theoretical Computer Science, 566(??): 1–11, February 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008585.

Chechik:2015:FTC

[93] Shiri Chechik and David Peleg. The fault-tolerant capacitated K-center problem. Theoretical Computer Science, 566(??):12–25, February 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397514008603.

Turek:2015:APP

[94] Ondrej Turek. Abelian properties of Parry words. Theoretical Computer Science, 566(??):26– 38, February 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009384.

Gargano:2015:CCF

[95] Luisa Gargano and Adele A. Rescigno. Complexity of conflict-free colorings of graphs. Theoretical Computer Science, 566(??):39–49, February 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397514009463.

Zhang:2015:SSC

[96] Wenwen Zhang and Jian-Liang Wu. Some sufficient conditions for 1-planar graphs to be Class 1. Theoretical Computer Science, 566(??):50– 58, February 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009499.

Araujo:2015:PON

[97] Julio Araujo, Nathann Cohen, Susanna F. de Rezende, Frédéric Havet, and Phablo F. S. Moura. On the proper orientation number of bipartite graphs. Theoretical Computer Science, 566(??): 59–75, February 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009554.

Anonymous:2015:EBh

[98] Anonymous. Editorial Board. Theoretical Computer Science, 567(??): iii–ix, ???? 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010202.

Anonymous:2015:NC

[99] Anonymous. Natural computing. Theoretical Computer Science, 567 (??):1, ???? 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010275.

Neary:2015:TSU

[100] Turlough Neary. Three small universal spiking neural P systems. Theoretical Computer Science, 567(??): 2–20, ???? 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006719.

Salo:2015:CTS

[101] Ville Salo and Ilkka Törmä. Category theory of symbolic dynamics. Theoretical Computer Science, 567 (??):21–45, ???? 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008020.

Popovici:2015:FCO

[102] Elena Popovici and Ezra Winston. A framework for co-optimization algorithm performance and its application to worst-case optimization.

Theoretical Computer Science, 567 (??):46–73, ???? 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008305.

Song:2015:CEU

[103] Bosheng Song and Linqiang Pan. Computational efficiency and universality of timed P systems with active membranes. Theoretical Computer Science, 567(??):74–86, ???? 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008433.

Doerr:2015:BBC

[104] Benjamin Doerr, Carola Doerr, and Franziska Ebel. From black-box complexity to designing new genetic algorithms. Theoretical Computer Science, 567(??):87–104, ???? 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009451.

Anonymous:2015:EBi

[105] Anonymous. Editorial Board. Theoretical Computer Science, 568(??): iii–ix, February 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000067.

Nishimura:2015:IPQ

[106] Harumichi Nishimura and Tomoyuki Yamakami. Interactive proofs with quantum finite automata. Theoretical Computer Science, 568(??):1– 18, February 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009487.

Tang:2015:GDR

[107] Ming-Xiao Tang, Jia Lee, and Kenichi Morita. General design of reversible sequential machines based on reversible logic elements. Theoretical Computer Science, 568(??):19– 27, February 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009505.

Boyar:2015:RIA

[108] Joan Boyar, Sushmita Gupta, and Kim S. Larsen. Relative interval analysis of paging algorithms on access graphs. Theoretical Computer Science, 568(??):28–48, February 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009530.

Folschette:2015:IBR

[109] Maxime Folschette, Loïc Paulevé, Katsumi Inoue, Morgan Magnin, and Olivier Roux. Identification of biological regulatory networks from Process Hitting models. Theoretical Computer Science, 568(??):49– 71, February 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009578.

Bulteau:2015:FPA

[110] Laurent Bulteau, Anna Paola Carrieri, and Riccardo Dondi. Fixed-parameter algorithms for scaffold filling. Theoretical Computer Science, 568(??): 72–83, February 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009827.

Ooshita:2015:SSM

[111] Fukuhito Ooshita and Sébastien Tixeuil. On the self-stabilization of mobile oblivious robots in uniform rings. Theoretical Computer Science, 568(??): 84–96, February 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009967.

Le:2015:CRP

[112] Van Bang Le and Sheng-Lung Peng. Characterizing and recognizing probe block graphs. Theoretical Computer Science, 568(??):97– 102, February 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514010135.

Anonymous:2015:EBj

[113] Anonymous. Editorial Board. Theoretical Computer Science, 569(??): iii–ix, March 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000225.

Krawec:2015:PIC

[114] Walter O. Krawec. n-Player impartial combinatorial games with random players. Theoretical Computer Science, 569(??):1–12, March 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009694.

Levin:2015:MTE

[115] Liron Levin, Dariusz R. Kowalski, and Michael Segal. Message and time efficient multi-broadcast schemes. Theoretical Computer Science, 569(??): 13–23, March 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009839.

Bose:2015:SLC

[116] Prosenjit Bose, Jean-Lou De Carufel, and Stephane Durocher. Searching on a line: a complete characterization of the optimal solution. Theoretical Computer Science, 569(??): 24–42, March 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009840.

Skowron:2015:CFP

[117] Piotr Skowron, Lan Yu, Piotr Faliszewski, and Edith Elkind. The complexity of fully proportional representation for single-crossing electorates. Theoretical Computer Science, 569 (??):43–57, March 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514010111.

Yu:2015:PGR

[118] Yu Yu, Xiangxue Li, and Jian Weng. Pseudorandom generators from regular one-way functions: New constructions with improved parameters. Theoretical Computer Science, 569(??): 58–69, March 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514010123.

Lam:2015:NCM

[119] Nhat X. Lam, Tien Tran, Min Kyung An, and Dung T. Huynh. A note on the complexity of minimum latency data aggregation scheduling with uniform power in physical interference model. Theoretical Computer Science, 569(??):70–73, March 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009529.

Anonymous:2015:EBk

[120] Anonymous. Editorial Board. Theoretical Computer Science, 570(??): iii–ix, March 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000559.

Baste:2015:RPC

[121] Julien Baste and Ignasi Sau. The role of planarity in connectivity problems parameterized by treewidth. Theoretical Computer Science, 570(??): 1–14, March 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010093.

Liu:2015:HNC

[122] Jia-Jie Liu, Cindy Tzu-Hsin Wang, Yue-Li Wang, and William Chung-Kung Yen. The hub number of co-comparability graphs. Theoretical Computer Science, 570(??):15– 21, March 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751401010X.

Geng:2015:POS

[123] Zhichao Geng and Jinjiang Yuan. Pareto optimization scheduling of family jobs on a p-batch machine to minimize makespan and maximum lateness. Theoretical Computer Science, 570(??):22–29, March 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010329.

Li:2015:IAS

[124] Chung-Lun Li. Improved algorithms for single-machine common due window assignment and scheduling with batch deliveries. Theoretical Computer Science, 570(??): 30–39, March 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010330.

Abed:2015:COP

[125] Fidaa Abed and Chien-Chung Huang. Coordinating oligopolistic players in unrelated machine scheduling. Theoretical Computer Science, 570(??): 40–54, March 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000134.

Zhang:2015:POP

[126] Zhongzhi Zhang and Bin Wu. Pfaffian orientations and perfect matchings of scale-free networks. Theoretical Computer Science, 570(??): 55–69, March 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000158.

Genitrini:2015:ACT

[127] Antoine Genitrini, Bernhard Gittenberger, Veronika Kraus, and Cécile Mailler. Associative and commutative tree representations for Boolean functions. Theoretical Computer Science, 570(??):70–101, March 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500016X.

Anonymous:2015:EBl

[128] Anonymous. Editorial Board. Theoretical Computer Science, 571(??): iii–ix, March 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000821.

Angelini:2015:IBP

[129] Patrizio Angelini, Giordano Da Lozzo, Giuseppe Di Battista, Fabrizio Frati, and Vincenzo Roselli. The importance of being proper: (in clusteredlevel planarity and T -level planarity). Theoretical Computer Science, 571 (??):1–9, March 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010317.

Lai:2015:CAS

[130] Cheng-Nan Lai. On the construction of all shortest vertex-disjoint paths in Cayley graphs of abelian groups. Theoretical Computer Science, 571(??): 10–20, March 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000146.

Krysta:2015:CAV

[131] Piotr Krysta and Carmine Ventre. Combinatorial auctions with verification are tractable. Theoretical Computer Science, 571(??): 21–35, March 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000171.

Auer:2015:UPG

[132] Christopher Auer, Christian Bachmaier, Franz J. Brandenburg, Andreas Gleißner, and Kathrin Hanauer. Upward planar graphs and their duals. Theoretical Computer Science, 571 (??):36–49, March 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000304.

Rytter:2015:SZP

[133] Wojciech Rytter and Arseny M. Shur. Searching for Zimin patterns. Theoretical Computer Science, 571(??): 50–57, March 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000316.

Wang:2015:PMW

[134] Limin Wang, Xiaoyan Zhang, Zhao Zhang, and Hajo Broersma. A PTAS for the minimum weight connected vertex cover P_{3} problem on unit disk graphs. Theoretical Computer Science, 571(??):58– 66, March 16, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000328.

Crespelle:2015:TIS

[135] Christophe Crespelle, Thi Ha Duong Phan, and The Hung Tran. Termination of the iterated strong-factor operator on multipartite graphs. Theoretical Computer Science, 571(??): 67–77, March 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000365.

Durocher:2015:GP

[136] Stephane Durocher, Debajyoti Mondal, and Md. Saidur Rahman. On graphs that are not PCGs. Theoretical Computer Science, 571(??): 78–87, March 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000389.

Anonymous:2015:EBm

[137] Anonymous. Editorial Board. Theoretical Computer Science, 572(??): iii–ix, March 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001231.

DAlessandro:2015:CEBb

[138] Flavio D’Alessandro and Benedetto Intrigila. On the commutative equivalence of bounded context-free and regular languages: the semi-linear case. Theoretical Computer Science, 572 (??):1–24, March 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000353.

Zhang:2015:PCL

[139] Zongyang Zhang, Sherman S. M. Chow, and Zhenfu Cao. Post-challenge leakage in public-key encryption. Theoretical Computer Science, 572(??): 25–49, March 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000377.

Lu:2015:IPD

[140] Lingfa Lu, Liqi Zhang, and Long Wan. Integrated production and delivery scheduling on a serial batch machine to minimize the makespan. Theoretical Computer Science, 572(??): 50–57, March 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000419.

Xu:2015:OSE

[141] Jia Xu and Zhaohui Liu. Online scheduling with equal processing times and machine eligibility constraints. Theoretical Computer Science, 572(??): 58–65, March 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000420.

Ting:2015:OPM

[142] H. F. Ting and Xiangzhong Xiang. Online pricing for multi-type of items. Theoretical Computer Science, 572(??): 66–82, March 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000432.

Rao:2015:ABS

[143] Michaël Rao, Michel Rigo, and Pavel Salimov. Avoiding 2-binomial squares and cubes. Theoretical Computer Science, 572(??):83–91, March 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000651.

Anonymous:2015:EBn

[144] Anonymous. Editorial Board. Theoretical Computer Science, 573(??): iii–ix, March 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001346.

Henzinger:2015:TUD

[145] Monika Henzinger and Veronika Loitzenbauer. Truthful unit-demand auctions with budgets revisited. Theoretical Computer Science, 573(??): 1–15, March 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000699.

Homan:2015:DRF

[146] Christopher M. Homan and Sven Kosub. Dichotomy results for fixed point counting in boolean dynamical systems. Theoretical Computer Science, 573(??):16–25, March 30, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000766.

Panda:2015:HRA

[147] B. S. Panda, S. Paul, and D. Pradhan. Hardness results, approximation and exact algorithms for liar’s domination problem in graphs. Theoretical Computer Science, 573(??): 26–42, March 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000778.

Bilo:2015:MDN

[148] Davide Bilò, Luciano Gualà, Stefano Leucci, and Guido Proietti. The max-distance network creation game on general host graphs. Theoretical Computer Science, 573(??): 43–53, March 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000912.

Boroujeni:2015:IAC

[149] Marziyeh Boroujeni, Abdolali Basiri, Sajjad Rahmany, and Annick Valibouze. F_{4}-invariant algorithm for computing SAGBI–Gröbner bases. Theoretical Computer Science, 573(??): 54–62, March 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000924.

Liu:2015:EDP

[150] Yunlong Liu, Jianxin Wang, Jie You, Jianer Chen, and Yixin Cao. Edge deletion problems: Branching facilitated by modular decomposition. Theoretical Computer Science, 573(??): 63–70, March 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000961.

Chatterjee:2015:ACA

[151] Krishnendu Chatterjee, Manas Joglekar, and Nisarg Shah. Average case analysis of the classical algorithm for Markov decision processes with Büchi objectives. Theoretical Computer Science, 573(??):71–89, March 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397515000973.

Anonymous:2015:EBo

[152] Anonymous. Editorial Board. Theoretical Computer Science, 574(??): iii–ix, April 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001437.

Prabhakar:2015:DCP

[153] Pavithra Prabhakar, Vladimeros Vladimerou, Mahesh Viswanathan, and Geir Dullerud. A decidable class of planar linear hybrid systems. Theoretical Computer Science, 574(??):1– 17, April 6, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008615.

Esik:2015:FPT

[154] Zoltán Ésik and Panos Rondogiannis. A fixed point theorem for non-monotonic functions. Theoretical Computer Science, 574(??):18– 38, April 6, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000687.

Matsikoudis:2015:FPT

[155] Eleftherios Matsikoudis and Edward A. Lee. The fixed-point theory of strictly causal functions. Theoretical Computer Science, 574(??):39– 77, April 6, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000729.

Zhang:2015:QP

[156] Wenfeng Zhang and Xiaoquan Xu. s_{2}-quasicontinuous posets. Theoretical Computer Science, 574(??):78– 85, April 6, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000730.

Bezem:2015:KMS

[157] Marc Bezem and Thierry Coquand. A Kripke model for simplicial sets. Theoretical Computer Science, 574 (??):86–91, April 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000717.

Anonymous:2015:EBp

[158] Anonymous. Editorial Board. Theoretical Computer Science, 575(??): iii–ix, April 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001814.

Pal:2015:GEF

[159] Sudebkumar Prasant Pal and Kunihiko Sadakane. Guest Editors’ foreword. Theoretical Computer Science, 575(??):1–2, April 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001887.

Fujita:2015:OSB

[160] Satoshi Fujita. Optimal serial broadcast of successive chunks. Theoretical Computer Science, 575(??):3– 9, April 13, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009517.

Shinn:2015:VBP

[161] Tong-Wook Shinn and Tadao Takaoka. Variations on the bottleneck paths problem. Theoretical Computer Science, 575(??):10–16, April 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S030439751400841X.

Ghosh:2015:SRP

[162] Subir Kumar Ghosh and Bodhayan Roy. Some results on point visibility graphs. Theoretical Computer Science, 575(??):17–32, April 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008342.

Bhushan:2015:EAM

[163] Alka Bhushan and G. Sajith. I/O efficient algorithms for the minimum cut problem on unweighted undirected graphs. Theoretical Computer Science, 575(??):33–41, April 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008354.

Akl:2015:CCP

[164] Selim Akl, Robert Benkoczi, Daya Ram Gaur, Hossam Hassanein, Shahadat Hossain, and Mark Thom. On a class of covering problems with variable capacities in wireless networks. Theoretical Computer Science, 575 (??):42–55, April 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008366.

Narayanaswamy:2015:MAS

[165] N. S. Narayanaswamy and G. Ramakrishna. On minimum average stretch spanning trees in polygonal 2trees. Theoretical Computer Science, 575(??):56–70, April 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009566.

Angelini:2015:ASP

[166] Patrizio Angelini, Giordano Da Lozzo, and Daniel Neuwirth. Advancements on SEFE and Partitioned Book Embedding problems. Theoretical Computer Science, 575(??): 71–89, April 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008597.

Cho:2015:ANO

[167] Da-Jung Cho, Yo-Sub Han, and Hwee Kim. Alignment with nonoverlapping inversions and translocations on two strings. Theoretical Computer Science, 575(??):90– 101, April 13, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514008287.

Anonymous:2015:EBq

[168] Anonymous. Editorial Board. Theoretical Computer Science, 576(??): iii–ix, April 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002066.

Subramani:2015:FCH

[169] K. Subramani and James Worthington. Feasibility checking in Horn constraint systems through a reduction based approach. Theoretical Computer Science, 576(??):1–17, April 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010159.

Fountoulakis:2015:ACC

[170] Nikolaos Fountoulakis, Tobias Friedrich, and Danny Hermelin. On the averagecase complexity of parameterized clique. Theoretical Computer Science, 576(??):18–29, April 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000894.

Cheng:2015:MPC

[171] Eddie Cheng, Robert Connolly, and Christoper Melekian. Matching preclusion and conditional matching preclusion problems for the folded Petersen cube. Theoretical Computer Science, 576(??):30–44, April 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000936.

Wang:2015:CNC

[172] Farn Wang, Sven Schewe, and Jung-Hsuan Wu. Complexity of node coverage games. Theoretical Computer Science, 576(??):45– 60, April 20, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000997.

Kakimura:2015:FPT

[173] Naonori Kakimura and Ken ichi Kawarabayashi. Fixed-parameter tractability for subset feedback set problems with parity constraints. Theoretical Computer Science, 576(??): 61–76, April 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001012.

Buss:2015:QSP

[174] Sam Buss. Quasipolynomial size proofs of the propositional pigeonhole principle. Theoretical Computer Science, 576 (??):77–84, April 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001024.

Klavik:2015:EPR

[175] Pavel Klavík, Jan Kratochvíl, Yota Otachi, and Toshiki Saitoh. Extending partial representations of subclasses of chordal graphs. Theoretical Computer Science, 576(??): 85–101, April 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001048.

Fellows:2015:THF

[176] Michael R. Fellows, Uéverton dos Santos Souza, Fábio Protti, and Maise Dantas da Silva. Tractability and hardness of flood-filling games on trees. Theoretical Computer Science, 576(??): 102–116, April 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500105X.

Anonymous:2015:EBr

[177] Anonymous. Editorial Board. Theoretical Computer Science, 577(??): iii–ix, April 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002303.

Popescu:2015:TGL

[178] Andrei Popescu and Grigore Rosu. Term-generic logic. Theoretical Computer Science, 577(??):1–24, April 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397515000948.

Janota:2015:EBQ

[179] Mikolás Janota and Joao Marques-Silva. Expansion-based QBF solving versus Q-resolution. Theoretical Computer Science, 577(??):25– 42, April 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500095X.

Sevegnani:2015:BS

[180] Michele Sevegnani and Muffy Calder. Bigraphs with sharing. Theoretical Computer Science, 577(??):43– 73, April 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001085.

Preda:2015:UMA

[181] Mila Dalla Preda, Roberto Giacobazzi, and Saumya Debray. Unveiling metamorphism by abstract interpretation of code properties. Theoretical Computer Science, 577(??): 74–97, April 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001541.

Schmidt-Schauss:2015:OPC

[182] Manfred Schmidt-Schauß, David Sabel, Joachim Niehren, and Jan Schwinghammer. Observational program calculi and the correctness of translations. Theoretical Computer Science, 577(??): 98–124, April 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001577.

Li:2015:SAP

[183] Qin Li and Zhe Dang. Sampling automata and programs. Theoretical Computer Science, 577(??): 125–140, April 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002169.

Anonymous:2015:EBs

[184] Anonymous. Editorial Board. Theoretical Computer Science, 578(??): iii–ix, May 3, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002522.

Konstantinidis:2015:IAA

[185] Stavros Konstantinidis. Implementation and application of automata (CIAA 2013). Theoretical Computer Science, 578(??): 1, May 3, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002595.

Balkanski:2015:SCP

[186] Eric Balkanski, F. Blanchet-Sadri, Matthew Kilgore, and B. J. Wyatt. On the state complexity of partial word DFAs. Theoretical Computer Science, 578(??):2–12, May 3, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000481.

Caralp:2015:TVP

[187] Mathieu Caralp, Pierre-Alain Reynier, and Jean-Marc Talbot. Trimming visibly pushdown automata. Theoretical Computer Science, 578(??):

13–29, May 3, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000456.

I:2015:CAD

[188] Tomohiro I., Takaaki Nishimoto, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Compressed automata for dictionary matching. Theoretical Computer Science, 578(??): 30–41, May 3, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000468.

Jirasek:2015:BRL

[189] Jozef Jirásek and Galina Jirásková. On the boundary of regular languages. Theoretical Computer Science, 578 (??):42–57, May 3, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000493.

Kutrib:2015:DID

[190] Martin Kutrib, Andreas Malcher, Carlo Mereghetti, Beatrice Palano, and Matthias Wendlandt. Deterministic input-driven queue automata: Finite turns, decidability, and closure properties. Theoretical Computer Science, 578(??):58–71, May 3, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000390.

Maletti:2015:HOD

[191] Andreas Maletti. Hyper-optimization for deterministic tree automata. Theoretical Computer Science, 578(??): 72–87, May 3, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500047X.

Otto:2015:LCC

[192] Friedrich Otto and Frantisek Mráz. Lambda-confluence for context rewriting systems. Theoretical Computer Science, 578(??):88–99, May 3, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397515000407.

Debarbieux:2015:ENW

[193] Denis Debarbieux, Olivier Gauwin, Joachim Niehren, Tom Sebastian, and Mohamed Zergaoui. Early nested word automata for XPath query answering on XML streams. Theoretical Computer Science, 578(??): 100–125, May 3, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000444.

Anonymous:2015:EBt

[194] Anonymous. Editorial Board. Theoretical Computer Science, 579(??): iii–ix, May 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500273X.

Dobson:2015:GLP

[195] M. P. Dobson, E. Hinrichsen, and V. Leoni. Generalized limited packings of some graphs with a limited number of P_{4}-partners. Theoretical Computer Science, 579(??): 1–8, May 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514008573.

Johnson:2015:SSD

[196] Taylor T. Johnson and Sayan Mitra. Safe and stabilizing distributed multi-path cellular flows. Theoretical Computer Science, 579(??):9– 32, May 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500050X.

Dekhtyar:2015:CDG

[197] Michael Dekhtyar, Alexander Dikovsky, and Boris Karlov. Categorial dependency grammars. Theoretical Computer Science, 579(??):33– 63, May 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000900.

Dobrev:2015:CBC

[198] Stefan Dobrev, Stephane Durocher, Mohsen Eftekhari, Konstantinos Georgiou, Evangelos Kranakis, Danny Krizanc, Lata Narayanan, Jaroslav Opatrny, Sunil Shende, and Jorge Urrutia. Complexity of barrier coverage with relocatable sensors in the plane. Theoretical Computer Science, 579(??):64–73, May 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001036.

Bermond:2015:FDP

[199] Jean-Claude Bermond, David Coudert, Gianlorenzo D’Angelo, and Fatima Zahra Moataz. Finding disjoint paths in networks with star shared risk link groups. Theoretical Computer Science, 579(??):74– 87, May 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001097.

Guha:2015:FAT

[200] Ashwin Guha and Ambedkar Dukkipati. A faster algorithm for testing polynomial representability of functions over finite integer rings. Theoretical Computer Science, 579(??):88– 99, May 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001103.

Lee:2015:SAS

[201] Kwangsu Lee, Dong Hoon Lee, and Moti Yung. Sequential aggregate signatures with short public keys without random oracles. Theoretical Computer Science, 579(??):100– 125, May 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001164.

Bampis:2015:GSF

[202] Evripidis Bampis, Dimitrios Letsios, and Giorgio Lucarelli. Green scheduling, flows and matchings. Theoretical Computer Science, 579(??): 126–136, May 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001176.

Anonymous:2015:EBu

[203] Anonymous. Editorial Board. Theoretical Computer Science, 580(??): iii–ix, May 17, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002844.

Gritti:2015:PPE

[204] Clémentine Gritti, Willy Susilo, Thomas Plantard, and Khin Than Win. Privacy-preserving encryption scheme using DNA parentage test. Theoretical Computer Science, 580 (??):1–13, May 17, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001188.

Muller:2015:GLS

[205] Johannes Müller and Hengrui Jiang. Graphical limit sets for general cellular automata. Theoretical Computer Science, 580(??):14–27, May 17, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001528.

Xie:2015:SGU

[206] Fang Xie, Zhe Xu, Yuzhong Zhang, and Qingguo Bai. Scheduling games on uniform machines with activation cost. Theoretical Computer Science, 580(??):28–35, May 17, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500153X.

Ji:2015:SMB

[207] Min Ji, Qinyun Yang, Danli Yao, and T. C. E. Cheng. Single-machine batch scheduling of linear deteriorating jobs. Theoretical Computer Science, 580(??):36–49, May 17, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001553.

Bu:2015:BCP

[208] Yuehua Bu and Xudong Bao. Backbone coloring of planar graphs for C_{8}-free or C_{9}-free. Theoretical Computer Science, 580(??):50–58, May 17, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001565.

Yang:2015:EBS

[209] Yu Yang, Hongbo Liu, Hua Wang, and Scott Makeig. Enumeration of BC-subtrees of trees. Theoretical Computer Science, 580(??):59– 74, May 17, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001589.

Damaschke:2015:FEL

[210] Peter Damaschke. Finding and enumerating large intersections. Theoretical Computer Science, 580(??): 75–82, May 17, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500170X.

Istrate:2015:RRM

[211] Gabriel Istrate. Reachability and recurrence in a modular generalization of annihilating random walks (and lights-out games) to hypergraphs. Theoretical Computer Science, 580 (??):83–93, May 17, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001711.

Braunschvig:2015:FTA

[212] Gilad Braunschvig, Shiri Chechik, David Peleg, and Adam Sealfon. Fault tolerant additive and (μ, α)-spanners. Theoretical Computer Science, 580 (??):94–100, May 17, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001723.

Amini:2015:NDG

[213] Omid Amini, David Coudert, and Nicolas Nisse. Non-deterministic graph searching in trees. Theoretical Computer Science, 580(??): 101–121, May 17, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001747.

Anonymous:2015:EBv

[214] Anonymous. Editorial Board. Theoretical Computer Science, 581(??): iii–ix, May 24, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002996.

Kamae:2015:CEP

[215] Teturo Kamae and Dong Han Kim. A characterization of eventual periodicity. Theoretical Computer Science, 581(??):1–8, May 24, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001759.

DArco:2015:APN

[216] Paolo D’Arco and Alfredo De Santis. Anonymous protocols: Notions and equivalence. Theoretical Computer Science, 581(??):9– 25, May 24, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001917.

deLuca:2015:SWS

[217] Aldo de Luca and Alessandro De Luca. Sturmian words and the Stern sequence. Theoretical Computer Science, 581(??):26–44, May 24, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001929.

Sarma:2015:DCD

[218] Atish Das Sarma, Anisur Rahaman Molla, and Gopal Pandurangan. Distributed computation in dynamic networks via random walks. Theoretical Computer Science, 581(??): 45–66, May 24, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001930.

Jonsson:2015:CNI

[219] Peter Jonsson, Victor Lagerkvist, and Gustav Nordh. Constructing NPintermediate problems by blowing holes with parameters of various properties. Theoretical Computer Science, 581 (??):67–82, May 24, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002133.

Giannakopoulos:2015:BOR

[220] Yiannis Giannakopoulos. Bounding the optimal revenue of selling multiple goods. Theoretical Computer Science, 581(??):83–96, May 24, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002145.

Durocher:2015:LSD

[221] Stephane Durocher, Hicham El-Zein, J. Ian Munro, and Sharma V. Thankachan. Low space data structures for geometric range mode query. Theoretical Computer Science, 581 (??):97–101, May 24, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002157.

Anonymous:2015:EBw

[222] Anonymous. Editorial Board. Theoretical Computer Science, 582(??): iii–ix, May 31, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003321.

Barton:2015:GLS

[223] Carl Barton, Tomás Flouri, Costas S. Iliopoulos, and Solon P. Pissis. Global and local sequence alignment with a bounded number of gaps. Theoretical Computer Science, 582(??): 1–16, May 31, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002200.

Harutyunyan:2015:LTA

[224] Ararat Harutyunyan and Sylvain Legay. Linear time algorithms for weighted offensive and powerful alliances in trees. Theoretical Computer Science, 582(??):17– 26, May 31, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002236.

Dereniowski:2015:DVS

[225] Dariusz Dereniowski, Adrian Kosowski, and Dominik Pajak. Distinguishing views in symmetric networks: a tight lower bound. Theoretical Computer Science, 582(??):27– 34, May 31, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002248.

Ho:2015:ETE

[226] Nhan Bao Ho. On the expansion of three-element subtraction sets. Theoretical Computer Science, 582(??): 35–47, May 31, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002418.

Song:2015:FTD

[227] Sulin Song, Xiaoyan Li, Shuming Zhou, and Mi Chen. Fault tolerance and diagnosability of burnt pancake networks under the comparison model. Theoretical Computer Science, 582 (??):48–59, May 31, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002431.

Teh:2015:CWP

[228] Wen Chean Teh and Kiam Heong Kwa. Core words and Parikh matrices. Theoretical Computer Science, 582 (??):60–69, May 31, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002649.

Ilic:2015:VCD

[229] Aleksandar Ilić. On the variable common due date, minimal tardy jobs bicriteria two-machine flow shop problem with ordered machines. Theoretical Computer Science, 582(??): 70–73, May 31, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500239X.

Larsen:2015:HSS

[230] Kasper Green Larsen, J. Ian Munro, Jesper Sindahl Nielsen, and Sharma V. Thankachan. On hardness of several string indexing problems. Theoretical Computer Science, 582(??): 74–82, May 31, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500242X.

Grytczuk:2015:HPT

[231] Jaroslaw Grytczuk, Karol Kosiński, and Michal Zmarz. How to play Thue games. Theoretical Computer Science, 582(??):83–88, May 31, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002637.

Anonymous:2015:EBx

[232] Anonymous. Editorial Board. Theoretical Computer Science, ??(??): iii–ix, ???? 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500345X.

Richard:2015:FPT

[233] Adrien Richard. Fixed point theorems for Boolean networks expressed in terms of forbidden subnetworks. Theoretical Computer Science, 583 (??):1–26, ???? 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002650.

Sun:2015:CAO

[234] Shu-Ming Sun and Ning Zhong. Computability aspects for 1st-order partial differential equations via characteristics. Theoretical Computer Science, 583(??):27–39, ???? 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002662.

Kim:2015:NRG

[235] Changwook Kim and Mahiro Ando. Node replacement graph grammars with dynamic node relabeling. Theoretical Computer Science, 583(??): 40–50, ???? 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002674.

Chan:2015:FMR

[236] Timothy M. Chan, J. Ian Munro, and Venkatesh Raman. Finding median in read-only memory on integer input. Theoretical Computer Science, 583(??):51–56, ???? 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002893.

Li:2015:SMB

[237] Shisheng Li, T. C. E. Cheng, C. T. Ng, and Jinjiang Yuan. Singlemachine batch scheduling with job processing time compatibility. Theoretical Computer Science, 583(??):

57–66, ???? 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500290X.

Chen:2015:PSI

[238] Youjun Chen, Lingfa Lu, and Jinjiang Yuan. Preemptive scheduling on identical machines with delivery coordination to minimize the maximum delivery completion time. Theoretical Computer Science, 583 (??):67–77, ???? 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002935.

Garcia:2015:DMD

[239] Pedro García, Damián López, and Manuel Vázquez de Parga. DFA minimization: Double reversal versus split minimization algorithms. Theoretical Computer Science, 583 (??):78–85, ???? 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003084.

DeStefani:2015:ENC

[240] Lorenzo De Stefani and Francesco Silvestri. Exploiting non-constant safe memory in resilient algorithms and data structures. Theoretical Computer Science, 583(??):86– 97, ???? 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003096.

Anonymous:2015:EBy

[241] Anonymous. Editorial Board. Theoretical Computer Science, 584(??): iii–ix, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003606.

Moscibroda:2015:GES

[242] Thomas Moscibroda and Adele A. Rescigno. Guest editorial: Special issue on Structural Information and Communication Complexity. Theoretical Computer Science, 584(??):1–2, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003667.

Hegde:2015:SOF

[243] Nidhi Hegde, Laurent Massoulié, and Laurent Viennot. Self-organizing flows in social networks. Theoretical Computer Science, 584(??): 3–18, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001152.

Clementi:2015:DCD

[244] Andrea Clementi, Miriam Di Ianni, Giorgio Gambosi, Emanuele Natale, and Riccardo Silvestri. Distributed community detection in dynamic graphs. Theoretical Computer Science, 584(??):19–41, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009402.

Levi:2015:NAT

[245] Amir Levi and Boaz Patt-Shamir. Nonadditive two-option ski rental. Theoretical Computer Science, 584(??): 42–52, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000742.

Gargano:2015:IDS

[246] Luisa Gargano, Pavol Hell, Joseph G. Peters, and Ugo Vaccaro. Influence diffusion in social networks under time window constraints. Theoretical Computer Science, 584(??): 53–66, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001127.

Kniesburges:2015:DWC

[247] Sebastian Kniesburges, Andreas Koutsopoulos, and Christian Scheideler. A deterministic worst-case message complexity optimal solution for resource discovery. Theoretical Computer Science, 584(??):67– 79, June 13, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514009414.

Coulouma:2015:COM

[248] Étienne Coulouma, Emmanuel Godard, and Joseph Peters. A characterization of oblivious message adversaries for which consensus is solvable. Theoretical Computer Science, 584(??):80–90, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000602.

Avin:2015:SAG

[249] Chen Avin, Michael Borokhovich, Bernhard Haeupler, and Zvi Lotker. Self-adjusting grid networks to minimize expected path length. Theoretical Computer Science, 584(??): 91–102, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009542.

Daltrophe:2015:PCT

[250] Hadassa Daltrophe, Shlomi Dolev, and Zvi Lotker. Probabilistic connectivity threshold for directional antenna widths. Theoretical Computer Science, 584(??):103– 114, June 13, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002443.

Metivier:2015:AFD

[251] Y. Métivier, J. M. Robson, and A. Zemmari. Analysis of fully distributed splitting and naming probabilistic procedures and applications. Theoretical Computer Science, 584(??): 115–130, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001139.

Giroire:2015:CSG

[252] F. Giroire, I. Lamprou, D. Mazauric, N. Nisse, S. Pérennes, and R. Soares. Connected surveillance game. Theoretical Computer Science, 584(??): 131–143, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514009396.

Even:2015:NAP

[253] Guy Even and Moti Medina. A nonmonotone analysis with the primaldual approach: Online routing of virtual circuits with unknown durations. Theoretical Computer Science, 584(??): 144–154, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514007841.

Anantharamu:2015:BAH

[254] Lakshmi Anantharamu and Bogdan S. Chlebus. Broadcasting in ad hoc multiple access channels. Theoretical Computer Science, 584(??): 155–176, June 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002613.

Anonymous:2015:EBz

[255] Anonymous. Editorial Board. Theoretical Computer Science, 585(??): iii–ix, June 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003825.

DalLago:2015:DIC

[256] Ugo Dal Lago and Guillaume Bonfante. Developments in Implicit Complexity (DICE 2012). Theoretical Computer Science, 585(??):1– 2, June 20, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001942.

Avanzini:2015:NOT

[257] Martin Avanzini, Naohi Eguchi, and Georg Moser. A new ordertheoretic characterisation of the polytime computable functions. Theoretical Computer Science, 585(??): 3–24, June 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001966.

Bonfante:2015:RNN

[258] Guillaume Bonfante, Florian Deloup, and Antoine Henrot. Real or natural number interpretation and their effect on complexity. Theoretical Computer Science, 585(??): 25–40, June 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001978.

Feree:2015:CPT

[259] Hugo Férée, Emmanuel Hainry, Mathieu Hoyrup, and Romain Péchoux. Characterizing polynomial time complexity of stream programs using interpretations. Theoretical Computer Science, 585(??):41–54, June 20, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002017.

Brunel:2015:RML

[260] Aloïs Brunel and Marco Gaboardi. Realizability models for a linear dependent PCF. Theoretical Computer Science, 585(??):55–70, June 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439751500198X.

Deng:2015:PEL

[261] Yuxin Deng and Yu Zhang. Program equivalence in linear contexts. Theoretical Computer Science, 585 (??):71–90, June 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001991.

Albert:2015:MDI

[262] Elvira Albert, Jesús Correas, Germán Puebla, and Guillermo Román-Díez. A multi-domain incremental analysis engine and its application to incremental resource analysis. Theoretical Computer Science, 585(??): 91–114, June 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001954.

Eberhard:2015:ATL

[263] Sebastian Eberhard. Applicative theories for logarithmic complexity classes. Theoretical Computer Science, 585(??):

115–135, June 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002005.

Anonymous:2015:EBba

[264] Anonymous. Editorial Board. Theoretical Computer Science, 586(??): iii–ix, June 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004041.

Widmayer:2015:FA

[265] Peter Widmayer. FUN with algorithms. Theoretical Computer Science, 586(??):1, June 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004119.

Brand:2015:NEP

[266] Michael Brand. No easy puzzles: Hardness results for jigsaw puzzles. Theoretical Computer Science, 586 (??):2–11, June 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001607.

Chen:2015:NCC

[267] Ke Chen and Adrian Dumitrescu. Nonconvex cases for carpenter’s rulers. Theoretical Computer Science, 586 (??):12–25, June 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001619.

Fertin:2015:TAG

[268] Guillaume Fertin, Shahrad Jamshidi, and Christian Komusiewicz. Towards an algorithmic guide to spiral galaxies. Theoretical Computer Science, 586 (??):26–39, June 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001620.

Cicalese:2015:SIW

[269] Ferdinando Cicalese, Gennaro Cordasco, Luisa Gargano, Martin Milanic, Joseph Peters, and Ugo Vaccaro. Spread of influence in weighted networks under time and budget constraints. Theoretical Computer Science, 586(??):40–58, June 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001632.

Borassi:2015:FDR

[270] Michele Borassi, Pierluigi Crescenzi, Michel Habib, Walter A. Kosters, Andrea Marino, and Frank W. Takes. Fast diameter and radius BFS-based computation in (weakly connected) real-world graphs: With an application to the six degrees of separation games. Theoretical Computer Science, 586(??):59–80, June 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001644.

Yamanaka:2015:SLT

[271] Katsuhisa Yamanaka, Erik D. Demaine, Takehiro Ito, Jun Kawahara, Masashi Kiyomi, Yoshio Okamoto, Toshiki Saitoh, Akira Suzuki, Kei Uchizawa, and Takeaki Uno. Swapping labeled tokens on graphs. Theoretical Computer Science, 586(??): 81–94, June 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001656.

Georgiou:2015:EMC

[272] Konstantinos Georgiou, Evangelos Kranakis, and Danny Krizanc. Excuse me! or the courteous theatregoers’ problem. Theoretical Computer Science, 586(??):95–110, June 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001668.

Demaine:2015:FFA

[273] Erik D. Demaine and Martin L. Demaine. Fun with fonts: Algorithmic typography. Theoretical Computer Science, 586(??):111– 119, June 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500167X.

Viglietta:2015:LPC

[274] Giovanni Viglietta. Lemmings is PSPACE-complete. Theoretical Computer Science, 586(??):120– 134, June 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001681.

Aloupis:2015:CNG

[275] Greg Aloupis, Erik D. Demaine, Alan Guo, and Giovanni Viglietta. Classic Nintendo games are (computationally) hard. Theoretical Computer Science, 586(??):135– 160, June 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001735.

Kostitsyna:2015:OAC

[276] Irina Kostitsyna, Maarten Löffler, and Valentin Polishchuk. Optimizing airspace closure with respect to politicians’ egos. Theoretical Computer Science, 586(??): 161–175, June 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003217.

Anonymous:2015:EBbb

[277] Anonymous. Editorial Board. Theoretical Computer Science, 587(??): iii–ix, July 4, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004326.

Giannini:2015:IBC

[278] Paola Giannini, Emanuela Merelli, and Angelo Troina. Interactions between computer science and biology. Theoretical Computer Science, 587 (??):1–2, July 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004399.

Bartocci:2015:SDS

[279] Ezio Bartocci, Luca Bortolussi, Laura Nenzi, and Guido Sanguinetti. System design of stochastic models using robustness of temporal properties. Theoretical Computer Science, 587(??):3–25, July 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002224.

Casagrande:2015:UBS

[280] Alberto Casagrande and Carla Piazza. Unwinding biological systems. Theoretical Computer Science, 587 (??):26–48, July 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002212.

Angius:2015:AAB

[281] Alessio Angius, Gianfranco Balbo, Marco Beccuti, Enrico Bibbona, Andras Horvath, and Roberta Sirovich. Approximate analysis of biological systems by hybrid switching jump diffusion. Theoretical Computer Science, 587(??):49–72, July 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002194.

Bodei:2015:CSA

[282] Chiara Bodei, Roberta Gori, and Francesca Levi. Causal static analysis for brane calculi. Theoretical Computer Science, 587(??):73– 103, July 4, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002182.

Pardini:2015:CIB

[283] Giovanni Pardini, Paolo Milazzo, and Andrea Maggiolo-Schettini. Component identification in biochemical pathways. Theoretical Computer Science, 587(??):104–124, July 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002170.

Anonymous:2015:EBbc

[284] Anonymous. Editorial Board. Theoretical Computer Science, 588(??): iii–ix, July 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004466.

Anonymous:2015:PSI

[285] Anonymous. Preface for the special issue on Theory and Applications of Models of Computation. Theoretical Computer Science, 588 (??):1, July 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500448X.

Lattimore:2015:MLN

[286] Tor Lattimore and Marcus Hutter. On Martin–Löf (non-)convergence of Solomonoff’s universal mixture. Theoretical Computer Science, 588(??): 2–15, July 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009815.

Kulkarni:2015:MPU

[287] Raghav Kulkarni, Youming Qiao, and Xiaoming Sun. Any monotone property of 3-uniform hypergraphs is weakly evasive. Theoretical Computer Science, 588(??):16–23, July 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400855X.

Higashikawa:2015:MRS

[288] Yuya Higashikawa, John Augustine, Siu-Wing Cheng, Mordecai J. Golin, Naoki Katoh, Guanqun Ni, Bing Su, and Yinfeng Xu. Minimax regret 1-sink location problem in dynamic path networks. Theoretical Computer Science, 588(??):24–36, July 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514001030.

Agarwal:2015:MUA

[289] Manu Agarwal, Ragesh Jaiswal, and Arindam Pal. k-means++ under approximation stability. Theoretical Computer Science, 588(??):37– 51, July 11, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003734.

Okhotin:2015:INF

[290] Alexander Okhotin. Improved normal form for grammars with one-sided contexts. Theoretical Computer Science, 588(??):52–72, July 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002686.

Badkobeh:2015:CSB

[291] Golnaz Badkobeh and Pascal Ochem. Characterization of some binary words with few squares. Theoretical Computer Science, 588(??):73– 80, July 11, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002911.

Hon:2015:MIS

[292] Wing-Kai Hon, Ton Kloks, Ching-Hao Liu, Hsiang-Hsuan Liu, Sheung-Hung Poon, and Yue-Li Wang. On maximum independent set of categorical product and ultimate categorical ratios of graphs. Theoretical Computer Science, 588(??):81–95, July 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002923.

Yang:2015:OCT

[293] Boting Yang and William Hamilton. The optimal capture time of the one-cop-moves game. Theoretical Computer Science, 588(??): 96–113, July 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003126.

Bastos:2015:CCR

[294] J. Bastos, A. Messaoudi, T. Rodrigues, and D. Smania. A class of cubic Rauzy fractals. Theoretical Computer Science, 588(??): 114–130, July 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500314X.

Duarte:2015:MIM

[295] Márcio A. Duarte, Felix Joos, Lucia D. Penso, Dieter Rautenbach, and Uéverton S. Souza. Maximum induced matchings close to maximum matchings. Theoretical Computer Science, 588(??):131– 137, July 11, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003072.

Anonymous:2015:EBbd

[296] Anonymous. Editorial Board. Theoretical Computer Science, 589(??): iii–ix, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004600.

Sureson:2015:RRM

[297] Claude Sureson. Random reals as measures of natural open sets. Theoretical Computer Science, 589(??): 1–23, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003163.

Itoh:2015:BMM

[298] Toshiya Itoh and Seiji Yoshimoto. Buffer management of multi-queue QoS switches with class segregation. Theoretical Computer Science, 589 (??):24–33, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003175.

Amir:2015:DMF

[299] Amihood Amir, Avivit Levy, Ely Porat, and B. Riva Shalom. Dictionary matching with a few gaps. Theoretical Computer Science, 589 (??):34–46, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003187.

Kuo:2015:EEL

[300] Che-Nan Kuo. Every edge lies on cycles embedding in folded hypercubes with vertex-fault-tolerant. Theoretical Computer Science, 589(??): 47–52, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003199.

Caragiannis:2015:IAK

[301] Ioannis Caragiannis, Aris Filos-Ratsikas, and Ariel D. Procaccia. An improved 2-agent kidney exchange mechanism. Theoretical Computer Science, 589(??): 53–60, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003205.

Anthony:2015:PAC

[302] Martin Anthony and Joel Ratsaby. A probabilistic approach to case-based inference. Theoretical Computer Science, 589(??):61–75, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003230.

Binucci:2015:FPP

[303] Carla Binucci, Emilio Di Giacomo, Walter Didimo, Fabrizio Montecchiani, Maurizio Patrignani, Antonios Symvonis, and Ioannis G. Tollis. Fanplanarity: Properties and complexity. Theoretical Computer Science, 589 (??):76–86, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003515.

Adler:2015:LRW

[304] Isolde Adler and Mamadou Moustapha Kanté. Linear rank-width and linear clique-width of trees. Theoretical Computer Science, 589(??): 87–98, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003527.

Bulteau:2015:CVC

[305] Laurent Bulteau, Jiehua Chen, Piotr Faliszewski, Rolf Niedermeier, and Nimrod Talmon. Combinatorial voter control in elections. Theoretical Computer Science, 589(??):

99–120, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003540.

Gourves:2015:WCC

[306] Laurent Gourvès, Jérôme Monnot, and Lydia Tlilane. Worst case compromises in matroids with applications to the allocation of indivisible goods. Theoretical Computer Science, 589(??): 121–140, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003722.

ODunlaing:2015:ACC

[307] Colm Ó Dúnlaing. An almost-confluent congruential language which is not Church–Rosser congruential. Theoretical Computer Science, 589(??): 141–146, July 19, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003254.

Anonymous:2015:EBbe

[308] Anonymous. Editorial Board. Theoretical Computer Science, 590(??): iii–ix, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004867.

Gasieniec:2015:FCT

[309] Leszek Gasieniec, Russell Martin, Frank Wolter, and Prudence W. H. Wong. Fundamentals of computation theory. Theoretical Computer Science, 590(??):1, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004934.

Akutsu:2015:CFL

[310] Tatsuya Akutsu, Takeyuki Tamura, Avraham A. Melkman, and Atsuhiro Takasu. On the complexity of finding a largest common subtree of bounded degree. Theoretical Computer Science, 590(??):2–16, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007749.

Boyar:2015:CFC

[311] Joan Boyar and Magnus Gausdal Find. Cancellation-free circuits in unbounded and bounded depth. Theoretical Computer Science, 590(??): 17–26, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514007762.

Casteigts:2015:ETV

[312] Arnaud Casteigts, Paola Flocchini, Emmanuel Godard, Nicola Santoro, and Masafumi Yamashita. On the expressivity of time-varying graphs. Theoretical Computer Science, 590 (??):27–37, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003102.

Arvind:2015:IPD

[313] V. Arvind, Johannes Köbler, Sebastian Kuhnert, Gaurav Rattan, and Yadu Vasudev. On the isomorphism problem for decision trees and decision lists. Theoretical Computer Science, 590(??):38–54, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000614.

Kranakis:2015:CSF

[314] Evangelos Kranakis, Fraser MacQuarrie, and Oscar Morales Ponce. Connectivity and stretch factor tradeoffs in wireless sensor networks with directional antennae. Theoretical Computer Science, 590(??): 55–72, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003916.

Kowaluk:2015:NSP

[315] Mirosław Kowaluk and Gabriela Majewska. New sequential and parallel algorithms for computing the βspectrum. Theoretical Computer Science, 590(??):73–85, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515002601.

Chaplick:2015:LCH

[316] Steven Chaplick, Jirí Fiala, Pim van ’t Hof, Daniël Paulusma, and Marek Tesar. Locally constrained homomorphisms on graphs of bounded treewidth and bounded degree. Theoretical Computer Science, 590(??):

86–95, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500064X.

Nikoletseas:2015:SEB

[317] S. Nikoletseas, P. Panagopoulou, C. Raptopoulos, and P. G. Spirakis. On the structure of equilibria in basic network formation. Theoretical Computer Science, 590(??): 96–105, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002455.

Barkay:2015:ESN

[318] Neta Barkay, Ely Porat, and Bar Shalem. Efficient sampling of nonstrict turnstile data streams. Theoretical Computer Science, 590(??): 106–117, July 26, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000626.

Liao:2015:IAA

[319] Kewen Liao, Hong Shen, and Longkun Guo. Improved approximation algorithms for constrained fault-tolerant resource allocation. Theoretical Computer Science, 590(??):118–128, July 26, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001590.

Anta:2015:OPS

[320] Antonio Fernández Anta, Chryssis Georgiou, Dariusz R. Kowalski, and Elli Zavou. Online parallel scheduling of non-uniform tasks: Trading failures for energy. Theoretical Computer Science, 590(??):129– 146, July 26, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000638.

Anonymous:2015:EBbf

[321] Anonymous. Editorial Board. Theoretical Computer Science, 591(??): iii–ix, August 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005733.

Nivat:2015:TST

[322] Maurice Nivat. The true story of TCS. Theoretical Computer Science, 591(??):1–2, August 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004260.

Ausiello:2015:TC

[323] Giorgio Ausiello, Lila Kari, Grzegorz Rozenberg, and Donald Sannella. TCS in the 21st century. Theoretical Computer Science, 591 (??):3–4, August 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004272.

Izumi:2015:AMC

[324] Taisuke Izumi, Tomoko Izumi, Hirotaka Ono, and Koichi Wada. Approximability of minimum certificate dispersal with tree structures. Theoretical Computer Science, 591(??): 5–14, August 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000341.

Kohler:2015:SSL

[325] S. Köhler and V. Turau. Selfstabilizing local k-placement of replicas with local minimum variance. Theoretical Computer Science, 591(??): 15–27, August 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003503.

Dereniowski:2015:SPR

[326] Dariusz Dereniowski, Hirotaka Ono, Ichiro Suzuki, Łukasz Wrona, Masafumi Yamashita, and Paweł Zyliński. The searchlight problem for road networks. Theoretical Computer Science, 591(??):28–59, August 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515003692.

Cicalese:2015:CVC

[327] Ferdinando Cicalese, Martin Milanic, and Romeo Rizzi. On the complexity of the vector connectivity problem. Theoretical Computer Science, 591(??): 60–71, August 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003758.

Golovach:2015:EGG

[328] Petr A. Golovach. Editing to a graph of given degrees. Theoretical Computer Science, 591(??):72– 84, August 2, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003771.

Meer:2015:GFA

[329] Klaus Meer and Ameen Naif. Generalized finite automata over real and complex numbers. Theoretical Computer Science, 591(??):85– 98, August 2, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003928.

Lin:2015:LTA

[330] Ching-Chi Lin and Hai-Lun Tu. A linear-time algorithm for paireddomination on circular-arc graphs. Theoretical Computer Science, 591(??): 99–105, August 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500393X.

Reutenauer:2015:SFS

[331] Christophe Reutenauer. Studies on finite Sturmian words. Theoretical Computer Science, 591(??):106– 133, August 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003941.

Barash:2015:TSC

[332] Mikhail Barash and Alexander Okhotin. Two-sided context specifications in formal grammars. Theoretical Computer Science, 591(??):134–153, August 2, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003953.

Anonymous:2015:EBbg

[333] Anonymous. Editorial Board. Theoretical Computer Science, 592(??): iii–ix, August 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005952.

Chodoriwsky:2015:AAG

[334] Jacob Chodoriwsky and Lucia Moura. An adaptive algorithm for group testing for complexes. Theoretical Computer Science, 592(??):1– 8, August 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003965.

Kim:2015:NOI

[335] Hwee Kim and Yo-Sub Han. Nonoverlapping inversion on strings and languages. Theoretical Computer Science, 592(??):9–22, August 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003990.

Kamiya:2015:PES

[336] Yuichi Kamiya, Tatsuya Okada, Takeshi Sekiguchi, and Yasunobu Shiota. Power and exponential sums for generalized coding systems by a measure theoretic approach. Theoretical Computer Science, 592(??): 23–38, August 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004144.

Son:2015:GNN

[337] Wanbin Son, Sang Won Bae, and Hee-Kap Ahn. Group nearest-neighbor queries in the L_{1} plane. Theoretical Computer Science, 592(??): 39–48, August 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004156.

Gaspers:2015:FOP

[338] Serge Gaspers, Mikko Koivisto, Mathieu Liedloff, Sebastian Ordyniak, and Stefan Szeider. On finding optimal polytrees. Theoretical Computer Science, 592(??):49– 58, August 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004168.

Beal:2015:EPM

[339] Richard Beal and Donald Adjeroh. Efficient pattern matching for RNA secondary structures. Theoretical Computer Science, 592(??):59– 71, August 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500420X.

Uiterwijk:2015:NRD

[340] Jos W. H. M. Uiterwijk and Michael Barton. New results for domineering from combinatorial game theory endgame databases. Theoretical Computer Science, 592(??):72– 86, August 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004211.

Conflitti:2015:GCL

[341] Alessandro Conflitti and Ricardo Mamede. Gray codes and lexicographical combinatorial generation for nonnesting and sparse nonnesting set partitions. Theoretical Computer Science, 592(??):87–96, August 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004259.

Bhattacharyya:2015:NAP

[342] Rishiraj Bhattacharyya and Pratyay Mukherjee. Non-adaptive programmability of random oracle. Theoretical Computer Science, 592(??):97– 114, August 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004673.

Jez:2015:AGB

[343] Artur Jez. Approximation of grammarbased compression via recompression. Theoretical Computer Science, 592(??): 115–134, August 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004685.

Chen:2015:EAO

[344] Danny Z. Chen, Jian Li, and Haitao Wang. Efficient algorithms for the onedimensional k-center problem. Theoretical Computer Science, 592(??): 135–142, August 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004697.

Catalano:2015:ATO

[345] Dario Catalano, Dario Fiore, Rosario Gennaro, and Konstantinos Vamvourellis. Algebraic (trapdoor) one-way functions: Constructions and applications. Theoretical Computer Science, 592(??): 143–165, August 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004703.

Shan:2015:CCH

[346] Erfang Shan and Liying Kang. Coloring clique-hypergraphs of graphs with no subdivision of K5. Theoretical Computer Science, 592(??): 166–175, August 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004715.

Anonymous:2015:EBbh

[347] Anonymous. Editorial Board. Theoretical Computer Science, 593(??): iii–ix, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006064.

Kapron:2015:FNI

[348] Bruce Kapron, Lior Malka, and Venkatesh Srinivasan. A framework for non-interactive instance-dependent commitment schemes (NIC). Theoretical Computer Science, 593(??): 1–15, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004727.

Rosenbaum:2015:BGE

[349] David J. Rosenbaum and Fabian Wagner. Beating the generatorenumeration bound for p-group isomorphism. Theoretical Computer Science, 593(??):16–25, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515004776.

Lintzmayer:2015:AAS

[350] Carla Negri Lintzmayer, Guillaume Fertin, and Zanoni Dias. Approximation algorithms for sorting by lengthweighted prefix and suffix operations. Theoretical Computer Science, 593(??): 26–41, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004818.

Byrka:2015:ICC

[351] Jaroslaw Byrka, Andreas Karrenbauer, and Laura Sanità. The interval constrained 3-coloring problem. Theoretical Computer Science, 593(??): 42–50, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004946.

Heydrich:2015:DCC

[352] Sandy Heydrich and Rob van Stee. Dividing connected chores fairly. Theoretical Computer Science, 593(??): 51–61, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004958.

Chang:2015:TCS

[353] Ching-Lueh Chang and Yuh-Dauh Lyuu. Triggering cascades on strongly connected directed graphs. Theoretical Computer Science, 593(??): 62–69, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004971.

Jiang:2015:Q

[354] Minghui Jiang, Pedro J. Tejada, and Haitao Wang. Quell. Theoretical Computer Science, 593(??): 70–78, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004983.

Kumar:2015:NOA

[355] Gunjan Kumar and Saswata Shannigrahi. New online algorithm for dynamic speed scaling with sleep state. Theoretical Computer Science, 593(??): 79–87, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004995.

Greinecker:2015:ACT

[356] Florian Greinecker. On the 2-abelian complexity of the Thue–Morse word. Theoretical Computer Science, 593(??): 88–105, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005010.

Huang:2015:SRW

[357] Yuke Huang and Zhiying Wen. The sequence of return words of the Fibonacci sequence. Theoretical Computer Science, 593(??):106–116, August 16, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005022.

Zhang:2015:AAH

[358] Peng Zhang and Angsheng Li. Algorithmic aspects of homophyly of networks. Theoretical Computer Science, 593(??):117–131, August 16, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005058.

Hedetniemi:2015:TOS

[359] Stephen T. Hedetniemi, David P. Jacobs, and K. E. Kennedy. A theorem of Ore and self-stabilizing algorithms for disjoint minimal dominating sets. Theoretical Computer Science, 593(??): 132–138, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500506X.

Zhang:2015:OSP

[360] Wenming Zhang, E. Zhang, and Feifeng Zheng. Online (J, K)-search problem and its competitive analysis. Theoretical Computer Science, 593(??): 139–145, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005071.

Zhang:2015:ECF

[361] Mi-Mi Zhang and Jin-Xin Zhou. On g-extra connectivity of folded hypercubes. Theoretical Computer Science, 593(??):146–153, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515005101.

Yuan:2015:RRA

[362] Chen Yuan and Haibin Kan. Revisiting a randomized algorithm for the minimum rainbow subgraph problem. Theoretical Computer Science, 593(??): 154–159, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500496X.

Gordinowicz:2015:PGF

[363] Przemysław Gordinowicz. Planar graph is on fire. Theoretical Computer Science, 593(??):160–164, August 16, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005046.

Biswas:2015:SIR

[364] Sudip Biswas, Manish Patil, Rahul Shah, and Sharma V. Thankachan. Succinct indexes for reporting discriminating and generic words. Theoretical Computer Science, 593(??): 165–173, August 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005095.

Anonymous:2015:EBbi

[365] Anonymous. Editorial Board. Theoretical Computer Science, 594(??): iii–ix, August 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006180.

Machado:2015:RBT

[366] Rodrigo Machado, Leila Ribeiro, and Reiko Heckel. Rule-based transformation of graph rewriting rules: Towards higher-order graph grammars. Theoretical Computer Science, 594(??): 1–23, August 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000705.

Akshay:2015:CCT

[367] S. Akshay, Paul Gastin, Madhavan Mukund, and K. Narayan Kumar. Checking conformance for time-constrained scenario-based specifications. Theoretical Computer Science, 594(??):24–43, August 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515002467.

Nasab:2015:PPC

[368] A. R. Shir Ali Nasab and S. N. Hosseini. Partial pullback complement rewriting. Theoretical Computer Science, 594(??):44–64, August 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003138.

Matos:2015:EPR

[369] Armando B. Matos. The efficiency of primitive recursive functions: a programmer’s view. Theoretical Computer Science, 594(??): 65–81, August 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003539.

Fulop:2015:CWM

[370] Zoltán Fülöp and Heiko Vogler. Characterizing weighted MSO for trees by branching transitive closure logics. Theoretical Computer Science, 594(??): 82–105, August 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003552.

Teichmann:2015:LBM

[371] Markus Teichmann and Johannes Osterholzer. A link between multioperator and tree valuation automata and logics. Theoretical Computer Science, 594(??):106–119, August 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500376X.

Yatapanage:2015:NPB

[372] Nisansala Yatapanage and Kirsten Winter. Next-preserving branching bisimulation. Theoretical Computer Science, 594(??):120–142, August 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500417X.

Jia:2015:ACC

[373] Xiaodong Jia, Achim Jung, Hui Kou, Qingguo Li, and Haoran Zhao. All cartesian closed categories of quasicontinuous domains consist of domains. Theoretical Computer Science, 594(??): 143–150, August 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004181.

Ito:2015:QAG

[374] Sohei Ito, Takuma Ichinose, Masaya Shimakawa, Naoko Izumi, Shigeki Hagihara, and Naoki Yonezaki. Qualitative analysis of gene regulatory networks by temporal logic. Theoretical Computer Science, 594(??): 151–179, August 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005277.

Rebiha:2015:GIN

[375] Rachid Rebiha, Arnaldo V. Moura, and Nadir Matringe. Generating invariants for non-linear hybrid systems. Theoretical Computer Science, 594(??): 180–200, August 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005289.

French:2015:SCT

[376] Tim French, John McCabe-Dansted, and Mark Reynolds. Synthesis for continuous time. Theoretical Computer Science, 594(??):201–222, August 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005423.

Anonymous:2015:EBbj

[377] Anonymous. Editorial Board. Theoretical Computer Science, 595(??): iii–ix, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006519.

Smith:2015:IWD

[378] Tim Smith. On infinite words determined by L systems. Theoretical Computer Science, 595(??): 1–10, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005137.

Bernasconi:2015:ERO

[379] Anna Bernasconi, Valentina Ciriani, and Lorenzo Lago. On the error resilience of ordered binary decision diagrams. Theoretical Computer Science, 595(??):11–33, August 30, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005149.

Vitale:2015:SER

[380] Luciana Vitale, Álvaro Martín, and Gadiel Seroussi. Space-efficient representation of truncated suffix trees, with applications to Markov order estimation. Theoretical Computer Science, 595(??):34–45, August 30, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500523X.

Chen:2015:ODP

[381] Jian-Jia Chen, Mong-Jen Kao, D. T. Lee, Ignaz Rutter, and Dorothea Wagner. Online dynamic power management with hard real-time guarantees. Theoretical Computer Science, 595(??): 46–64, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005241.

Hu:2015:UAC

[382] Wenyu Hu, Xingjun Luo, and Zhongxuan Luo. A unified approach to computing the nearest complex polynomial with a given zero. Theoretical Computer Science, 595(??): 65–81, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005253.

Wang:2015:PCC

[383] Jianxin Wang, Weimin Su, Min Yang, Jiong Guo, Qilong Feng, Feng Shi, and Jianer Chen. Parameterized complexity of control and bribery for d-approval elections. Theoretical Computer Science, 595(??):82– 91, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005265.

Chateau:2015:CAF

[384] A. Chateau and R. Giroudeau. A complexity and approximation framework for the maximization scaffolding problem. Theoretical Computer Science, 595(??):92–106, August 30, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005332.

Bang-Jensen:2015:BBD

[385] J. Bang-Jensen and A. Yeo. Balanced branchings in digraphs. Theoretical Computer Science, 595(??): 107–119, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005368.

Qiu:2015:IAA

[386] Xian Qiu and Walter Kern. Improved approximation algorithms for a bilevel knapsack problem. Theoretical Computer Science, 595(??): 120–129, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500537X.

Khodakarami:2015:FPA

[387] Farnoosh Khodakarami, Farzad Didehvar, and Ali Mohades. A fixedparameter algorithm for guarding 1.5D terrains. Theoretical Computer Science, 595(??):130–142, August 30, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005381.

Wang:2015:PFD

[388] Yujue Wang, Duncan S. Wong, Qianhong Wu, Sherman S. M. Chow, Bo Qin, Jianwei Liu, and Yong Ding. Practical (fully) distributed signatures provably secure in the standard model. Theoretical Computer Science, 595(??): 143–158, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005393.

Renault:2015:CMV

[389] Gabriel Renault and Simon Schmidt. On the complexity of the Misère version of three games played on graphs. Theoretical Computer Science, 595(??): 159–167, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005447.

Arockiaraj:2015:CWH

[390] Micheal Arockiaraj and Arul Jeya Shalini. Conjectures on wirelength of hypercube into cylinder and torus. Theoretical Computer Science, 595(??): 168–171, August 30, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005216.

Anonymous:2015:EBbk

[391] Anonymous. Editorial Board. Theoretical Computer Science, 596(??): iii–ix, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006738.

Brunetti:2015:IHI

[392] S. Brunetti, E. Lodi, and W. Quattrociocchi. An inclusion hierarchy of irreversible dynamos. Theoretical Computer Science, 596(??):1– 11, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005459.

Twigg:2015:LPA

[393] Andrew Twigg and Eduardo C. Xavier. Locality-preserving allocations problems and coloured bin packing. Theoretical Computer Science, 596(??): 12–22, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005460.

Bland:2015:TOS

[394] Widmer Bland and W. F. Smyth. Three overlapping squares: the general case characterized & applications. Theoretical Computer Science, 596(??): 23–40, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005472.

Blanchet-Sadri:2015:CML

[395] F. Blanchet-Sadri and Andrew Lohr. Computing minimum length representations of sets of words of uniform length. Theoretical Computer Science, 596(??):41–54, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515005484.

Jager:2015:BMM

[396] Gerold Jäger and Marcin Peczarski. Bounding memory for Mastermind might not make it harder. Theoretical Computer Science, 596(??):55– 66, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005496.

Biniaz:2015:MHO

[397] Ahmad Biniaz, Anil Maheshwari, and Michiel Smid. Matchings in higher-order Gabriel graphs. Theoretical Computer Science, 596(??): 67–78, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005502.

Christ:2015:OMC

[398] Marie G. Christ, Lene M. Favrholdt, and Kim S. Larsen. Online multicoloring with advice. Theoretical Computer Science, 596(??):79– 91, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500554X.

Balogh:2015:OBW

[399] János Balogh, József Békési, György Dósa, Leah Epstein, Hans Kellerer, Asaf Levin, and Zsolt Tuza. Offline black and white bin packing. Theoretical Computer Science, 596(??):92– 101, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005551.

DAngelo:2015:MSP

[400] Gianlorenzo D’Angelo, Daniele Diodati, Alfredo Navarra, and Cristina M. Pinotti. The minimum k-storage problem on directed graphs. Theoretical Computer Science, 596(??):102– 108, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005599.

Nicosia:2015:TAS

[401] Gaia Nicosia, Andrea Pacifici, and Ulrich Pferschy. Two agent scheduling with a central selection mechanism. Theoretical Computer Science, 596(??): 109–123, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005617.

David:2015:ASH

[402] Julien David, Loïck Lhote, Arnaud Mary, and François Rioult. An average study of hypergraphs and their minimal transversals. Theoretical Computer Science, 596(??):124– 141, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005629.

Guegan:2015:CDO

[403] Guillaume Guégan and Pascal Ochem. Complexity dichotomy for oriented homomorphism of planar graphs with large girth. Theoretical Computer Science, 596(??):142–148, September 6, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515005514.

Dosa:2015:TAB

[404] Gyorgy Dosa. The tight absolute bound of first fit in the parameterized case. Theoretical Computer Science, 596(??):149–154, September 6, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005563.

Anonymous:2015:EBbl

[405] Anonymous. Editorial Board. Theoretical Computer Science, 597(??): iii–ix, September 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006994.

Bimbo:2015:DIF

[406] Katalin Bimbó. The decidability of the intensional fragment of classical linear logic. Theoretical Computer Science, 597(??):1– 17, September 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005290.

Bujtor:2015:EPI

[407] Ferenc Bujtor and Walter Vogler. Error-pruning in interface automata. Theoretical Computer Science, 597(??): 18–39, September 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005575.

Boreale:2015:WAC

[408] Michele Boreale and Michela Paolini. Worst- and average-case privacy breaches in randomization mechanisms. Theoretical Computer Science, 597(??):40–61, September 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515006313.

Biondi:2015:QIL

[409] Fabrizio Biondi, Axel Legay, Pasquale Malacaria, and Andrzej Wasowski. Quantifying information leakage of randomized protocols. Theoretical Computer Science, 597(??):62– 87, September 13, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006684.

Anonymous:2015:EBbm

[410] Anonymous. Editorial Board. Theoretical Computer Science, 598(??): iii–ix, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007355.

Kari:2015:Ea

[411] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 598(??):1, September 20, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007422.

Shang:2015:CPT

[412] Yun Shang, Xian Lu, and Ruqian Lu. Computing power of Turing machines in the framework of unsharp quantum logic. Theoretical Computer Science, 598(??):2– 14, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010147.

Salomaa:2015:ACR

[413] Arto Salomaa. Applications of the Chinese remainder theorem to reaction systems with duration. Theoretical Computer Science, 598(??):15– 22, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514010305.

Azimi:2015:DGM

[414] Sepinoud Azimi, Cristian Gratie, Sergiu Ivanov, and Ion Petre. Dependency graphs and mass conservation in reaction systems. Theoretical Computer Science, 598(??):23– 39, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001115.

Hutchinson:2015:EOT

[415] M. Hutchinson and M. Widom. Enumeration of octagonal tilings. Theoretical Computer Science, 598(??):40– 50, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500225X.

Shi:2015:ENM

[416] Yaoyun Shi and Xiaodi Wu. Epsilonnet method for optimizations over separable states. Theoretical Computer Science, 598(??):51–63, September 20, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002479.

Echegoyen:2015:CCB

[417] Carlos Echegoyen, Roberto Santana, Alexander Mendiburu, and Jose A. Lozano. Comprehensive characterization of the behaviors of estimation of distribution algorithms. Theoretical Computer Science, 598(??):64– 86, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003229.

Aman:2015:VMS

[418] Bogdan Aman and Gabriel Ciobanu. Verification of membrane systems with delays via Petri nets with delays. Theoretical Computer Science, 598(??):87– 101, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003278.

Fischer:2015:BCR

[419] Asja Fischer and Christian Igel. A bound for the convergence rate of parallel tempering for sampling restricted Boltzmann machines. Theoretical Computer Science, 598(??):102– 117, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004235.

Tlili:2015:SDL

[420] Takwa Tlili and Saoussen Krichen. On solving the double loading problem using a modified particle swarm optimization. Theoretical Computer Science, 598(??):118–128, September 20, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004788.

Gravier:2015:WOD

[421] Sylvain Gravier, Jérôme Javelle, Mehdi Mhalla, and Simon Perdrix. On weak odd domination and graph-based quantum secret sharing. Theoretical Computer Science, 598(??):129– 137, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004806.

Dennunzio:2015:RSE

[422] Alberto Dennunzio, Enrico Formenti, and Luca Manzoni. Reaction systems and extremal combinatorics properties. Theoretical Computer Science, 598(??): 138–149, September 20, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005034.

Morton:2015:CTP

[423] Jason Morton and Jacob Turner. Computing the Tutte polynomial of lattice path matroids using determinantal circuits. Theoretical Computer Science, 598(??):150–156, September 20, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006878.

Anonymous:2015:EBbn

[424] Anonymous. Editorial Board. Theoretical Computer Science, 599(??): iii–ix, September 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007501.

Kari:2015:Eb

[425] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 599(??):1, September 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007434.

Gilbert:2015:ACM

[426] David Gilbert and Monika Heiner. Advances in computational methods in systems biology. Theoretical Computer Science, 599(??):2– 3, September 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007446.

Ballarini:2015:AES

[427] Paolo Ballarini and Marie Duflot. Applications of an expressive statistical model checking approach to the analysis of genetic circuits. Theoretical Computer Science, 599(??):4– 33, September 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004223.

Islam:2015:MOR

[428] Md. Ariful Islam, Abhishek Murthy, Ezio Bartocci, Elizabeth M. Cherry, Flavio H. Fenton, James Glimm, Scott A. Smolka, and Radu Grosu. Model-order reduction of ion channel dynamics using approximate bisimulation. Theoretical Computer Science, 599(??):34–46, September 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514002205.

Chaves:2015:ACU

[429] Madalena Chaves and Alfonso Carta. Attractor computation using interconnected Boolean networks: Testing growth rate models in E. coli. Theoretical Computer Science, 599(??):47– 63, September 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004575.

Fages:2015:IRS

[430] François Fages, Steven Gay, and Sylvain Soliman. Inferring reaction systems from ordinary differential equations. Theoretical Computer Science, 599(??):64–78, September 27, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514006197.

Videla:2015:LBL

[431] Santiago Videla, Carito Guziolowski, Federica Eduati, Sven Thiele, Martin Gebser, Jacques Nicolas, Julio Saez-Rodriguez, Torsten Schaub, and Anne Siegel. Learning Boolean logic models of signaling networks with ASP. Theoretical Computer Science, 599(??): 79–101, September 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514004587.

Angione:2015:ADM

[432] C. Angione, J. Costanza, G. Carapezza, P. Lió, and G. Nicosia. Analysis and design of molecular machines. Theoretical Computer Science, 599(??): 102–117, September 27, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000663.

Anonymous:2015:EBbo

[433] Anonymous. Editorial Board. Theoretical Computer Science, 600(??): iii–ix, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007999.

Kumar:2015:NHS

[434] Gunjan Kumar and Saswata Shannigrahi. On the NP-hardness of speed scaling with sleep state. Theoretical Computer Science, 600(??): 1–10, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005228.

Hermelin:2015:PCA

[435] Danny Hermelin and Liat Rozenberg. Parameterized complexity analysis for the Closest String with Wildcards problem. Theoretical Computer Science, 600(??):11–18, October 4, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005538.

Ceccherini-Silberstein:2015:MAG

[436] Tullio Ceccherini-Silberstein, Michel Coornaert, Francesca Fiorenzi, Paul E. Schupp, and Nicholas W. M. Touikan. Multipass automata and group word problems. Theoretical Computer Science, 600(??):19–33, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397515005642.

Christodoulakis:2015:ISP

[437] Manolis Christodoulakis, P. J. Ryan, W. F. Smyth, and Shu Wang. Indeterminate strings, prefix arrays & undirected graphs. Theoretical Computer Science, 600(??):34– 48, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005666.

Coelho:2015:IRG

[438] Erika M. M. Coelho, Mitre C. Dourado, and Rudini M. Sampaio. Inapproximability results for graph convexity parameters. Theoretical Computer Science, 600(??):49– 58, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005794.

Bonomo:2015:CCD

[439] Flavia Bonomo, Guillermo Durán, and Mario Valencia-Pabon. Complexity of the cluster deletion problem on subclasses of chordal graphs. Theoretical Computer Science, 600(??): 59–69, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005800.

Bougeret:2015:IAA

[440] Marin Bougeret, Pierre-François Dutot, Denis Trystram, Klaus Jansen, and Christina Robenek. Improved approximation algorithms for scheduling parallel jobs on identical clusters. Theoretical Computer Science, 600(??): 70–85, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006246.

Gajser:2015:VTC

[441] David Gajser. Verifying time complexity of Turing machines. Theoretical Computer Science, 600(??): 86–97, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006623.

Alabert:2015:NFL

[442] Aureli Alabert, Alessandro Berti, Ricard Caballero, and Marco Ferrante. No-free-lunch theorems in the continuum. Theoretical Computer Science, 600(??):98–106, October 4, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006635.

Chiesa:2015:SAN

[443] Alessandro Chiesa and Zeyuan Allen Zhu. Shorter arithmetization of nondeterministic computations. Theoretical Computer Science, 600(??): 107–131, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006647.

Demaine:2015:LTA

[444] Erik D. Demaine, Martin L. Demaine, Eli Fox-Epstein, Duc A. Hoang, Takehiro Ito, Hirotaka Ono, Yota Otachi, Ryuhei Uehara, and Takeshi Yamada. Linear-time algorithm for sliding tokens on trees. Theoretical Computer Science, 600(??):132– 142, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006829.

Chiesa:2015:CCT

[445] Marco Chiesa, Giuseppe Di Battista, Thomas Erlebach, and Maurizio Patrignani. Computational complexity of traffic hijacking under BGP and S-BGP. Theoretical Computer Science, 600(??):143– 154, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006830.

Renault:2015:OAA

[446] Marc P. Renault, Adi Rosén, and Rob van Stee. Online algorithms with advice for bin packing and scheduling problems. Theoretical Computer Science, 600(??):155–170, October 4, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007069.

Kawahara:2015:ILB

[447] Jun Kawahara and Koji M. Kobayashi. An improved lower bound for one-dimensional online unit clustering. Theoretical Computer Science, 600(??): 171–173, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005654.

Du:2015:DSH

[448] Ye Du. The discrete sell or hold problem with constraints on asset values. Theoretical Computer Science, 600(??): 174–179, October 4, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006659.

Anonymous:2015:EBbp

[449] Anonymous. Editorial Board. Theoretical Computer Science, 601(??): iii–ix, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008166.

Karhumaki:2015:W

[450] J. Karhumäki and L. Zamboni. Words 2013. Theoretical Computer Science, 601(??):1, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008233.

Day:2015:PFW

[451] Joel D. Day, Daniel Reidenbach, and Johannes C. Schneider. Periodicity forcing words. Theoretical Computer Science, 601(??):2– 14, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500780X.

Ernvall:2015:PPC

[452] Mari Ernvall, Vesa Halava, and Tero Harju. On the n-permutation Post Correspondence Problem. Theoretical Computer Science, 601(??):15– 20, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006428.

Honkala:2015:LTP

[453] Juha Honkala. Language-theoretic problems in certain matrix monoids.

Theoretical Computer Science, 601(??): 21–28, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500643X.

Harju:2015:SFS

[454] Tero Harju and Mike Müller. Squarefree shuffles of words. Theoretical Computer Science, 601(??):29– 38, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006453.

Rao:2015:SGA

[455] Michaël Rao. On some generalizations of abelian power avoidability. Theoretical Computer Science, 601(??): 39–46, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006477.

Rigo:2015:AGA

[456] M. Rigo and P. Salimov. Another generalization of abelian equivalence: Binomial complexity of infinite words. Theoretical Computer Science, 601(??): 47–57, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006465.

Castiglione:2015:SSW

[457] G. Castiglione and M. Sciortino. Standard Sturmian words and automata minimization algorithms. Theoretical Computer Science, 601(??):58– 66, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006441.

Shur:2015:GSF

[458] Arseny M. Shur. Generating squarefree words efficiently. Theoretical Computer Science, 601(??):67– 72, October 11, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006489.

Anonymous:2015:EBbq

[459] Anonymous. Editorial Board. Theoretical Computer Science, 602(??): iii–ix, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008348.

Weng:2015:NSK

[460] Jian Weng, Yunlei Zhao, Robert H. Deng, Shengli Liu, Yanjiang Yang, and Kouichi Sakurai. A note on the security of KHL scheme. Theoretical Computer Science, 602(??): 1–6, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007070.

Kociumaka:2015:LTV

[461] T. Kociumaka, J. Radoszewski, W. Rytter, and T. Waleń. Lineartime version of Holub’s algorithm for morphic imprimitivity testing. Theoretical Computer Science, 602(??):

7–21, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007112.

Chitturi:2015:TUB

[462] Bhadrachalam Chitturi. Tighter upper bound for sorting permutations with prefix transpositions. Theoretical Computer Science, 602(??):22– 31, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500715X.

Rautenbach:2015:TGC

[463] Dieter Rautenbach. Two greedy consequences for maximum induced matchings. Theoretical Computer Science, 602(??):32–38, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515007173.

Le:2015:PTR

[464] Van Bang Le, Andrea Oversberg, and Oliver Schaudt. Polynomial time recognition of squares of Ptolemaic graphs and 3-sun-free split graphs. Theoretical Computer Science, 602(??): 39–49, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007185.

LeCun:2015:PFI

[465] Bertrand LeCun, Thierry Mautor, Franck Quessette, and Marc-Antoine Weisser. Bin packing with fragmentable items: Presentation and approximations. Theoretical Computer Science, 602(??):50–59, October 18, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007240.

Xu:2015:BDR

[466] Bojian Xu. Boosting distinct random sampling for basic counting on the union of distributed streams. Theoretical Computer Science, 602(??): 60–79, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007586.

Liu:2015:GRW

[467] Wen An Liu and Meile Zhao. General restrictions of Wythoff-like games. Theoretical Computer Science, 602(??): 80–88, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007616.

Faria:2015:ERC

[468] Luerbio Faria, Laurent Gourvès, Carlos A. Martinhon, and Jérôme Monnot. The edge-recoloring cost of monochromatic and properly edgecolored paths and cycles. Theoretical Computer Science, 602(??):89– 102, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007628.

Gabay:2015:OSB

[469] Michaël Gabay, Vladimir Kotov, and Nadia Brauner. Online bin stretching with bunch techniques. Theoretical Computer Science, 602(??):103– 113, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500763X.

Wang:2015:ODC

[470] Haitao Wang and Jingru Zhang. One-dimensional k-center on uncertain data. Theoretical Computer Science, 602(??):114–124, October 18, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007641.

Kuo:2015:PBF

[471] Che-Nan Kuo. Pancyclicity and bipancyclicity of folded hypercubes with both vertex and edge faults. Theoretical Computer Science, 602(??): 125–131, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007653.

Hung:2015:HPS

[472] Ruo-Wei Hung, Chih-Chia Yao, and Shang-Ju Chan. The Hamiltonian properties of supergrid graphs. Theoretical Computer Science, 602(??): 132–148, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007719.

Chang:2015:DST

[473] Ching-Lueh Chang. A deterministic sublinear-time nonadaptive algorithm for metric 1-median selection. Theoretical Computer Science, 602(??): 149–157, October 18, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007148.

Calamoneri:2015:CPC

[474] T. Calamoneri and R. Petreschi. Corrigendum to “On pairwise compatibility graphs having Dilworth number two” [Theoret. Comput. Sci. 524 (2014) 34–40]. Theoretical Computer Science, 602(??):158–159, October 18, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007677. See

[4].

Anonymous:2015:EBbr

[475] Anonymous. Editorial Board. Theoretical Computer Science, 603(??): iii–ix, October 25, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008610.

Ong:2015:LLI

[476] Luke Ong and Ruy J. G. B. de Queiroz. Logic, Language, Information and Computation (WoLLIC 2012). Theoretical Computer Science, 603(??): 1–2, October 25, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006325.

Abriola:2015:LWQ

[477] Sergio Abriola, Santiago Figueira, and Gabriel Senno. Linearizing well quasi-orders and bounding the length of bad sequences. Theoretical Computer Science, 603(??): 3–22, October 25, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006337.

Asperti:2015:FMT

[478] Andrea Asperti and Wilmer Ricciotti. A formalization of multitape Turing machines. Theoretical Computer Science, 603(??):23– 42, October 25, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006349.

Baldi:2015:UPS

[479] Paolo Baldi and Agata Ciabattoni. Uniform proofs of standard completeness for extensions of first-order MTL. Theoretical Computer Science, 603(??): 43–57, October 25, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006350.

Bozzelli:2015:COA

[480] Laura Bozzelli, Hans van Ditmarsch, and Sophie Pinchinat. The complexity of one-agent refinement modal logic. Theoretical Computer Science, 603(??): 58–83, October 25, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006362.

Caleiro:2015:BSG

[481] Carlos Caleiro, João Marcos, and Marco Volpe. Bivalent semantics, generalized compositionality and analytic classic-like tableaux for finitevalued logics. Theoretical Computer Science, 603(??):84–110, October 25, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006374.

Link:2015:FPR

[482] Sebastian Link. Frontiers for propositional reasoning about fragments of probabilistic conditional independence and hierarchical database decompositions. Theoretical Computer Science, 603(??):111–131, October 25, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006386.

Schechter:2015:LPJ

[483] L. Menasché Schechter. A logic of plausible justifications. Theoretical Computer Science, 603(??):132– 145, October 25, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006398.

Anonymous:2015:EBbs

[484] Anonymous. Editorial Board. Theoretical Computer Science, 604(??): iii–ix, November 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008865.

Bonsangue:2015:P

[485] Marcello Bonsangue. Preface. Theoretical Computer Science, 604(??): 1, November 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006842.

Hasuo:2015:GWP

[486] Ichiro Hasuo. Generic weakest precondition semantics from monads enriched with order. Theoretical Computer Science, 604(??):2– 29, November 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002947.

Jacobs:2015:DHM

[487] Bart Jacobs. Dijkstra and Hoare monads in monadic computation. Theoretical Computer Science, 604(??): 30–45, November 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002364.

Lenisa:2015:MSC

[488] Marina Lenisa. Multigames and strategies, coalgebraically. Theoretical Computer Science, 604(??):46– 62, November 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006416.

TranconyWidemann:2015:ACR

[489] B. Trancón y Widemann and M. Hauhs. Algebraic-coalgebraic recursion theory of history-dependent dynamical system models. Theoretical Computer Science, 604(??):63–80, November 2, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500568X.

Myers:2015:CCC

[490] Robert S. R. Myers, Jirí Adámek, Stefan Milius, and Henning Urbat. Coalgebraic constructions of canonical nondeterministic automata. Theoretical Computer Science, 604(??): 81–101, November 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002625.

Bonchi:2015:KED

[491] Filippo Bonchi, Stefan Milius, Alexandra Silva, and Fabio Zanasi. Killing epsilons with a dagger: a coalgebraic study of systems with algebraic label structure. Theoretical Computer Science, 604(??):102– 126, November 2, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002406.

Anonymous:2015:EBbt

[492] Anonymous. Editorial Board. Theoretical Computer Science, 605(??): iii–ix, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009056.

Kempka:2015:DAD

[493] Joseph Kempka, Phil McMinn, and Dirk Sudholt. Design and analysis of different alternating variable searches for search-based software testing. Theoretical Computer Science, 605(??): 1–20, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514009979.

Oliveto:2015:ITC

[494] Pietro S. Oliveto and Carsten Witt. Improved time complexity analysis of the Simple Genetic Algorithm. Theoretical Computer Science, 605(??): 21–41, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000298.

Akimoto:2015:ARO

[495] Youhei Akimoto, Sandra Astete-Morales, and Olivier Teytaud. Analysis of runtime of optimization algorithms for noisy functions over discrete codomains. Theoretical Computer Science, 605(??):42–50, November 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003151.

Bender:2015:MBP

[496] Michael A. Bender, Sándor P. Fekete, Alexander Kröller, Vincenzo Liberatore, Joseph S. B. Mitchell, Valentin Polishchuk, and Jukka Suomela. The minimum backlog problem. Theoretical Computer Science, 605(??): 51–61, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007744.

Holzer:2015:CPC

[497] Andreas Holzer, Christian Schallhart, Michael Tautschnig, and Helmut Veith. Closure properties and complexity of rational sets of regular languages. Theoretical Computer Science, 605(??): 62–79, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007823.

Klein:2015:LSC

[498] Rolf Klein, David Kriesel, and Elmar Langetepe. A local strategy for cleaning expanding cellular domains by simple robots. Theoretical Computer Science, 605(??):80– 94, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007835.

Krzywdzinski:2015:DAR

[499] K. Krzywdziński and K. Rybarczyk. Distributed algorithms for random graphs. Theoretical Computer Science, 605(??):95–105, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515007847.

Bang:2015:RLB

[500] Lucas Bang, Wolfgang Bein, and Lawrence L. Larmore. R-LINE: a better randomized 2-server algorithm on the line. Theoretical Computer Science, 605(??):106– 118, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007902.

Floderus:2015:ISI

[501] Peter Floderus, Mirosław Kowaluk, Andrzej Lingas, and Eva-Marta Lundell. Induced subgraph isomorphism: Are some patterns substantially easier than others? Theoretical Computer Science, 605(??):119– 128, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007914.

DEmidio:2015:ERG

[502] Mattia D’Emidio, Daniele Frigioni, and Alfredo Navarra. Explore and repair graphs with black holes using mobile entities. Theoretical Computer Science, 605(??):129–145, November 9, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007926.

Biswas:2015:DTC

[503] Ranita Biswas and Partha Bhowmick. On different topological classes of spherical geodesic paths and circles in Z^{3}. Theoretical Computer Science, 605(??):146–163, November 9, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397515007938.

Anonymous:2015:EBbu

[504] Anonymous. Editorial Board. Theoretical Computer Science, 606(??): iii–ix, November 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009251.

Finger:2015:SIS

[505] Marcelo Finger and Delia Kesner. Special issue: Selected papers of the 7th and 8th workshops on Logical and Semantic Frameworks with Applications (LSFA). Theoretical Computer Science, 606(??):1, November 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009329.

Accattoli:2015:PNC

[506] Beniamino Accattoli. Proof nets and the call-by-value λ-calculus. Theoretical Computer Science, 606(??):2– 24, November 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007264.

Areces:2015:SB

[507] Carlos Areces and Ezequiel Orbe. Symmetric blocking. Theoretical Computer Science, 606(??):25–41, November 16, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005307.

Carnielli:2015:MPR

[508] Walter Carnielli and Mariana Matulovic. The method of polynomial ring calculus and its potentialities. Theoretical Computer Science, 606(??): 42–56, November 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004193.

Ciaffaglione:2015:MTE

[509] Alberto Ciaffaglione and Ivan Scagnetto. Mechanizing type environments in weak HOAS. Theoretical Computer Science, 606(??):57–78, November 16, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006404.

DAgostino:2015:IVC

[510] Marcello D’Agostino. An informational view of classical logic. Theoretical Computer Science, 606(??): 79–97, November 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005678.

Olarte:2015:SCC

[511] Carlos Olarte, Elaine Pimentel, and Vivek Nigam. Subexponential concurrent constraint programming. Theoretical Computer Science, 606(??):98– 120, November 16, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005411.

Anonymous:2015:EBbv

[512] Anonymous. Editorial Board. Theoretical Computer Science, 607 (part 1)(??): iii–ix, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009548.

Gu:2015:P

[513] Qianping Gu, Pavol Hell, and Boting Yang. Preface. Theoretical Computer Science, 607 (part 1)(??): 1, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009615.

Higashikawa:2015:MSL

[514] Yuya Higashikawa, Mordecai J. Golin, and Naoki Katoh. Multiple sink location problems in dynamic path networks. Theoretical Computer Science, 607 (part 1)(??): 2–15, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005186.

Bredereck:2015:CDA

[515] Robert Bredereck, Vincent Froese, Sepp Hartung, André Nichterlein, Rolf Niedermeier, and Nimrod Talmon. The complexity of degree anonymization by vertex addition. Theoretical Computer Science, 607 (part 1)(??):16– 34, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006258.

Chin:2015:CAU

[516] Francis Y. L. Chin, Bin Fu, Jiuling Guo, Shuguang Han, Jueliang Hu, Minghui Jiang, Guohui Lin, Hing-Fung Ting, Luping Zhang, Yong Zhang, and Diwei Zhou. Competitive algorithms for unbounded one-way trading. Theoretical Computer Science, 607 (part 1)(??):35– 48, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004752.

Cai:2015:PCF

[517] Leizhen Cai and Junjie Ye. Parameterized complexity of finding connected induced subgraphs. Theoretical Computer Science, 607 (part 1)(??): 49–59, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004247.

Guo:2015:OSG

[518] Chengwei Guo and Leizhen Cai. Obtaining split graphs by edge contraction. Theoretical Computer Science, 607 (part 1)(??):60–67, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001693.

Hon:2015:ECC

[519] Wing-Kai Hon, Ton Kloks, Hsiang-Hsuan Liu, and Yue-Li Wang. Edgeclique covers of the tensor product. Theoretical Computer Science, 607 (part 1)(??):68–74, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005320.

Luo:2015:SOH

[520] Taibo Luo and Yinfeng Xu. Semionline hierarchical load balancing problem with bounded processing times. Theoretical Computer Science, 607 (part 1)(??):75–82, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003266.

Penso:2015:CAC

[521] Lucia Draque Penso, Fábio Protti, Dieter Rautenbach, and Uéverton dos Santos Souza. Complexity analysis of P_{3}-convexity problems on boundeddegree and planar graphs. Theoretical Computer Science, 607 (part 1)(??): 83–95, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/

article/pii/S0304397515008749. See

corrigendum [1486].

Brass:2015:LEB

[522] Peter Brass, Hyeon-Suk Na, and Chan-Su Shin. Local event boundary detection with unreliable sensors: Analysis of the majority vote scheme. Theoretical Computer Science, 607 (part 1)(??): 96–112, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500626X.

Biswas:2015:AFE

[523] Tamal T. Biswas and Kenneth W. Regan. Approximation of function evaluation over sequence arguments via specialized data structures. Theoretical Computer Science, 607 (part 1)(??): 113–123, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003114.

Anonymous:2015:EBbw

[524] Anonymous. Editorial Board. Theoretical Computer Science, 607 (part 2)(??): iii–ix, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009688.

Chen:2015:FA

[525] Jianer Chen and John E. Hopcroft. Frontiers of algorithmics. Theoretical Computer Science, 607 (part 2)(??): 125, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009755.

Chen:2015:CMM

[526] Danny Z. Chen, David Craft, and Lin Yang. A circular matrixmerging algorithm with application in Volumetric Intensity-Modulated Arc Therapy. Theoretical Computer Science, 607 (part 2)(??):126– 134, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003710.

Dereniowski:2015:CZV

[527] Dariusz Dereniowski, Danny Dyer, Ryan M. Tifenbach, and Boting Yang. The complexity of zero-visibility cops and robber. Theoretical Computer Science, 607 (part 2)(??):135– 148, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002388.

Hossain:2015:GST

[528] Md. Iqbal Hossain and Md. Saidur Rahman. Good spanning trees in graph drawing. Theoretical Computer Science, 607 (part 2)(??):149– 165, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500794X.

Jiang:2015:FAS

[529] Haitao Jiang, Lusheng Wang, Binhai Zhu, and Daming Zhu. A factor- (1.408 + ϵ) approximation for sorting unsigned genomes by reciprocal translocations. Theoretical Computer Science, 607 (part 2)(??):166– 180, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500479X.

Li:2015:PCC

[530] Weidong Li, Jianping Li, Xuejie Zhang, and Zhibin Chen. Penalty cost constrained identical parallel machine scheduling problem. Theoretical Computer Science, 607 (part 2)(??):181– 192, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008786.

Lin:2015:IES

[531] Ling Lin, Xiaochen Xian, Yujie Yan, Xing He, and Zhiyi Tan. Inefficiency of equilibria for scheduling game with machine activation costs. Theoretical Computer Science, 607 (part 2)(??): 193–207, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515004764.

Obata:2015:EAA

[532] Yuji Obata and Takao Nishizeki. Efficient approximation algorithms for bandwidth consecutive multicolorings of graphs. Theoretical Computer Science, 607 (part 2)(??):208– 220, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007082.

Shi:2015:ROI

[533] Yangguang Shi, Fa Zhang, Jie Wu, and Zhiyong Liu. Randomized oblivious integral routing for minimizing power cost. Theoretical Computer Science, 607 (part 2)(??):221– 246, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006283.

Ting:2015:NOA

[534] H. F. Ting and Xiangzhong Xiang. Near optimal algorithms for online maximum edge-weighted b-matching and two-sided vertex-weighted bmatching. Theoretical Computer Science, 607 (part 2)(??):247–256, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004739.

You:2015:KPA

[535] Jie You, Jianxin Wang, Qilong Feng, and Feng Shi. Kernelization and parameterized algorithms for covering a tree by a set of stars or paths. Theoretical Computer Science, 607 (part 2)(??): 257–270, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500540X.

Anonymous:2015:EBbx

[536] Anonymous. Editorial Board. Theoretical Computer Science, 607 (part 3)(??): iii–ix, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500986X.

Zhang:2015:P

[537] Zhao Zhang, Lidong Wu, and Ding-Zhu Du. Preface. Theoretical Computer Science, 607 (part 3)(??): 271, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009937.

Guo:2015:OTM

[538] Sicen Guo, Minming Li, and Yingchao Zhao. Optimal trees for minimizing average individual updating cost. Theoretical Computer Science, 607 (part 3)(??):272–281, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500777X.

Kanj:2015:PST

[539] Iyad Kanj and Stefan Szeider. Parameterized and subexponential-time complexity of satisfiability problems and applications. Theoretical Computer Science, 607 (part 3)(??):282– 295, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007768.

Zhou:2015:TPL

[540] Jianjie Zhou, Liying Kang, and Erfang Shan. Two paths location of a tree with positive or negative weights. Theoretical Computer Science, 607 (part 3)(??): 296–305, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007720.

Seo:2015:CCF

[541] Jungtaek Seo, Subhankar Mishra, Xiang Li, and My T. Thai. Catastrophic cascading failures in power networks. Theoretical Computer Science, 607 (part 3)(??):306–319, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007689.

Chang:2015:CF

[542] Yi-Jun Chang and Hsu-Chun Yen. Constrained floorplans in 2D and 3D. Theoretical Computer Science, 607 (part 3)(??):320–336, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007227.

Ito:2015:OBG

[543] Yoshihiko Ito, Yuki Kobayashi, Yuya Higashikawa, Naoki Katoh, Sheung-Hung Poon, and Maria Saumell. Optimally bracing grid frameworks with holes. Theoretical Computer Science, 607 (part 3)(??):337–350, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007215.

Armaselu:2015:AFP

[544] Bogdan Armaselu and Ovidiu Daescu. Algorithms for fair partitioning of convex polygons. Theoretical Computer Science, 607 (part 3)(??):351– 362, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007203.

Gao:2015:BAC

[545] Xiaofeng Gao, Jun Li, and Guihai Chen. A better approximation for constructing virtual backbone in 3D wireless ad-hoc networks. Theoretical Computer Science, 607 (part 3)(??):363– 380, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007197.

He:2015:AAT

[546] Zaobo He, Zhipeng Cai, Siyao Cheng, and Xiaoming Wang. Approximate aggregation for tracking quantiles and range countings in wireless sensor networks. Theoretical Computer Science, 607 (part 3)(??):381– 390, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007124.

Duan:2015:IEO

[547] Zhenhua Duan, Jin Liu, Jie Li, and Cong Tian. Improved even order magic square construction algorithms and their applications in multi-user shared electronic accounts. Theoretical Computer Science, 607 (part 3)(??):391– 410, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007094.

Bhattacharya:2015:IAC

[548] Binay Bhattacharya and Tsunehiko Kameda. Improved algorithms for computing minmax regret sinks on dynamic path and tree networks. Theoretical Computer Science, 607 (part 3)(??): 411–425, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006301.

Abu-Khzam:2015:PCD

[549] Faisal N. Abu-Khzam, Judith Egan, Michael R. Fellows, Frances A. Rosamond, and Peter Shaw. On the parameterized complexity of dynamic problems. Theoretical Computer Science, 607 (part 3)(??):426–434, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005630.

Kjos-Hanssen:2015:KSF

[550] Bjørn Kjos-Hanssen. Kolmogorov structure functions for automatic complexity. Theoretical Computer Science, 607 (part 3)(??):435–445, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005174.

Chen:2015:ACS

[551] Lin Chen, Deshi Ye, and Guochuan Zhang. An asymptotic competitive scheme for online bin packing. Theoretical Computer Science, 607 (part 3)(??):446–454, November 23, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005150.

Kanj:2015:IPE

[552] Iyad Kanj, Guohui Lin, Tian Liu, Weitian Tong, Ge Xia, Jinhui Xu, Boting Yang, Fenghui Zhang, Peng Zhang, and Binhai Zhu. Improved parameterized and exact algorithms for cut problems on trees. Theoretical Computer Science, 607 (part 3)(??):455– 470, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005125.

Song:2015:SCF

[553] Liang Song, Haibin Chen, Hao Gu, Hejiao Huang, and Hongwei Du. Set covering in fuel-considered vehicle routing problems. Theoretical Computer Science, 607 (part 3)(??):471– 479, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005113.

Li:2015:PCN

[554] Xueliang Li, Meiqin Wei, and Jun Yue. Proper connection number and connected dominating sets. Theoretical Computer Science, 607 (part 3)(??): 480–487, November 23, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005083.

Anonymous:2015:EBby

[555] Anonymous. Editorial Board. Theoretical Computer Science, 608 (part 1)(??): iii–ix, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009986.

Anonymous:2015:EBbz

[556] Anonymous. Editorial Board. Theoretical Computer Science, 608 (part 2)(??): iii–ix, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515010221.

Anonymous:2015:EBca

[557] Anonymous. Editorial Board. Theoretical Computer Science, 608 (part 3)(??): iii–ix, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751501035X.

Kari:2015:Ec

[558] Lila Kari. Editorial. Theoretical Computer Science, 608 (part 1)(??): 1, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515010063.

Merelli:2015:CSB

[559] Emanuela Merelli and Ion Petre. From computer science to biology and back. Theoretical Computer Science, 608 (part 1)(??):2–3, December 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010051.

Ehrenfeucht:2015:SOZ

[560] Andrzej Ehrenfeucht and Grzegorz Rozenberg. Standard and ordered zoom structures. Theoretical Computer Science, 608 (part 1)(??):4– 15, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006854.

Dennunzio:2015:ADG

[561] Alberto Dennunzio, Enrico Formenti, Luca Manzoni, and Antonio E. Porreca. Ancestors, descendants, and gardens of Eden in reaction systems. Theoretical Computer Science, 608 (part 1)(??):16–26, December 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005009.

Brijder:2015:RFG

[562] Robert Brijder. Recombination faults in gene assembly in ciliates modeled using multimatroids. Theoretical Computer Science, 608 (part 1)(??):27– 35, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007859.

Barbuti:2015:MPS

[563] Roberto Barbuti, Pasquale Bove, Paolo Milazzo, and Giovanni Pardini. Minimal probabilistic P systems for modelling ecological systems. Theoretical Computer Science, 608 (part 1)(??): 36–56, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006805.

Macias-Ramos:2015:MFV

[564] Luis F. Macías-Ramos, Mario J. Pérez-Jiménez, Agustín Riscos-Núñez, and Luis Valencia-Cabrera. Membrane fission versus cell division: When membrane proliferation is not enough. Theoretical Computer Science, 608 (part 1)(??):57–65, December 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005356.

Folschette:2015:SCR

[565] Maxime Folschette, Loïc Paulevé, Morgan Magnin, and Olivier Roux. Sufficient conditions for reachability in automata networks with priorities. Theoretical Computer Science, 608 (part 1)(??):66–83, December 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007872.

Zaccagnino:2015:TDC

[566] Rocco Zaccagnino, Rosalba Zizza, and Carlo Zottoli. Testing DNA code words properties of regular languages. Theoretical Computer Science, 608 (part 1)(??):84–97, December 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007811.

Hrydziuszko:2015:MAT

[567] Olga Hrydziuszko, Joanna Balbus, Magdalena Zulpo, Artur Wrona, and Krystian Kubica. Mathematical analyses of two-compartment model of human cholesterol circulatory transport in application to high blood cholesterol prevention, diagnosis and treatment. Theoretical Computer Science, 608 (part 1)(??):98–107, December 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007136.

Paun:2015:SMI

[568] Mihaela Paun, Yang Li, Yuan Cheng, Iris Tusa, and Andrei Paun. Segmenting microarray images using a contour-based method. Theoretical Computer Science, 608 (part 1)(??): 108–118, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006817.

Flocchini:2015:P

[569] Paola Flocchini and Jie Gao. Preface. Theoretical Computer Science, 608 (part 2)(??):119, December 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008531.

Bannoura:2015:WDS

[570] Amir Bannoura, Christian Ortolf, Leonhard Reindl, and Christian Schindelhauer. The wake up dominating set problem. Theoretical Computer Science, 608 (part 2)(??):120– 134, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500033X.

Miller:2015:CNL

[571] Avery Miller. On the complexity of neighbourhood learning in radio networks. Theoretical Computer Science, 608 (part 2)(??):135– 145, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003709.

Sharma:2015:ONN

[572] Gokarna Sharma and Costas Busch. Optimal nearest neighbor queries in sensor networks. Theoretical Computer Science, 608 (part 2)(??):146– 165, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003977.

Tonoyan:2015:CGS

[573] Tigran Tonoyan. Conflict graphs and the SINR-capacity of the mean power scheme. Theoretical Computer Science, 608 (part 2)(??):166– 177, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515000675.

Ortolf:2015:SPU

[574] Christian Ortolf and Christian Schindelhauer. Strategies for parallel unaware cleaners. Theoretical Computer Science, 608 (part 2)(??):178– 189, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008555.

Miller:2015:FRA

[575] Avery Miller and Andrzej Pelc. Fast rendezvous with advice. Theoretical Computer Science, 608 (part 2)(??): 190–198, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008543.

Halldorsson:2015:GES

[576] Magnús M. Halldórsson. Guest editorial: Structural information and communication complexity. Theoretical Computer Science, 608 (part 3)(??): 199–200, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010373.

Czyzowicz:2015:BPW

[577] Jurek Czyzowicz, Leszek Gasieniec, Konstantinos Georgiou, Evangelos Kranakis, and Fraser MacQuarrie. The Beachcombers’ Problem: Walking and searching with mobile robots. Theoretical Computer Science, 608 (part 3)(??): 201–218, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008269.

Dereniowski:2015:RHM

[578] Dariusz Dereniowski, Ralf Klasing, Adrian Kosowski, and Łukasz Kuszner. Rendezvous of heterogeneous mobile agents in edge-weighted networks. Theoretical Computer Science, 608 (part 3)(??):219–230, December 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005204.

Bampas:2015:IPD

[579] Evangelos Bampas, Nikos Leonardos, Euripides Markou, Aris Pagourtzis, and Matoula Petrolia. Improved periodic data retrieval in asynchronous rings with a faulty host. Theoretical Computer Science, 608 (part 3)(??): 231–254, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008488.

Emek:2015:HMA

[580] Yuval Emek, Tobias Langner, David Stolz, Jara Uitto, and Roger Wattenhofer. How many ants does it take to find the food? Theoretical Computer Science, 608 (part 3)(??):255– 267, December 10, 2015. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005198.

Hegeman:2015:LCC

[581] James W. Hegeman and Sriram V. Pemmaraju. Lessons from the Congested Clique applied to MapReduce. Theoretical Computer Science, 608 (part 3)(??):268–281, December 10, 2015. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008695.

Anonymous:2016:EBa

[582] Anonymous. Editorial Board. Theoretical Computer Science, 609 (part 1) (??):iii–ix, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010488.

Anonymous:2016:EBb

[583] Anonymous. Editorial Board. Theoretical Computer Science, 609 (part 2) (??):iii–ix, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010592.

Anonymous:2016:EBc

[584] Anonymous. Editorial Board. Theoretical Computer Science, 609 (part 3) (??):iii–ix, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515010701.

vanIersel:2016:STP

[585] Leo van Iersel, Steven Kelk, Nela Lekić, and Simone Linz. Satisfying ternary permutation constraints by multiple linear orders or phylogenetic trees. Theoretical Computer Science, 609 (part 1)(??):1– 21, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005319.

Lin:2016:RMN

[586] Bey-Chi Lin and Chin-Tau Lea. Repackable multi-log_{2} N networks. Theoretical Computer Science, 609 (part 1)(??):22–31, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397515007732.

Li:2016:RQU

[587] Jian Li and Haitao Wang. Range queries on uncertain data. Theoretical Computer Science, 609 (part 1) (??):32–48, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008075.

Susilo:2016:EDT

[588] Willy Susilo, Fuchun Guo, and Yi Mu. Efficient dynamic threshold identity-based encryption with constant-size ciphertext. Theoretical Computer Science, 609 (part 1)(??):

49–59, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008087.

Sawada:2016:SRF

[589] J. Sawada and A. Williams. Successor rules for flipping pancakes and burnt pancakes. Theoretical Computer Science, 609 (part 1)(??):60– 75, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008099.

Fong:2016:ACC

[590] Ken C. K. Fong, Minming Li, Hongyu Liang, Linji Yang, and Hao Yuan. Average-case complexity of the minsum matrix product problem. Theoretical Computer Science, 609 (part 1) (??):76–86, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008105.

Enright:2016:GIP

[591] Jessica Enright and Lorna Stewart. Games on interval and permutation graph representations. Theoretical Computer Science, 609 (part 1)(??): 87–103, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008117.

Kratochvil:2016:CCC

[592] Jan Kratochvíl, Jan Arne Telle, and Marek Tesar. Computational complexity of covering three-vertex multigraphs. Theoretical Computer Science, 609 (part 1)(??):104–117, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008282.

Goles:2016:PCM

[593] Eric Goles, Pedro Montealegre, Ville Salo, and Ilkka Törmä. PSPACEcompleteness of majority automata networks. Theoretical Computer Science, 609 (part 1)(??):118–128, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008294.

Beal:2016:CPP

[594] Richard Beal and Donald Adjeroh. Compressed parameterized pattern matching. Theoretical Computer Science, 609 (part 1)(??):129–142, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008440.

Yang:2016:ECV

[595] Ching-Nung Yang, Li-Zhe Sun, and Song-Ruei Cai. Extended color visual cryptography for black and white secret image. Theoretical Computer Science, 609 (part 1)(??):143– 161, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008452.

Qu:2016:NSD

[596] Cunquan Qu, Guanghui Wang, Jianliang Wu, and Xiaowei Yu. On the neighbor sum distinguishing total coloring of planar graphs. Theoretical Computer Science, 609 (part 1)(??): 162–170, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008464.

Das:2016:AMR

[597] Shantanu Das, Paola Flocchini, Giuseppe Prencipe, Nicola Santoro, and Masafumi Yamashita. Autonomous mobile robots with lights. Theoretical Computer Science, 609 (part 1)(??):171–184, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008476.

Han:2016:OMK

[598] Xin Han and Kazuhisa Makino. Online minimization knapsack problem. Theoretical Computer Science, 609 (part 1)(??):185–196, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008506.

Wang:2016:EAC

[599] Xi Wang, Jianxi Fan, Xiaohua Jia, and Cheng-Kuan Lin. An efficient algorithm to construct disjoint path covers of DCell networks. Theoretical Computer Science, 609 (part 1)(??):

197–210, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008518.

Akiba:2016:BRE

[600] Takuya Akiba and Yoichi Iwata. Branch-and-reduce exponential/FPT algorithms in practice: a case study of vertex cover. Theoretical Computer Science, 609 (part 1)(??):211– 225, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500852X.

Beal:2016:SDS

[601] Marie-Pierre Béal, Michel Blockelet, and Cătălin Dima. Sofic– Dyck shifts. Theoretical Computer Science, 609 (part 1)(??):226–244, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008567.

Viard:2016:CMC

[602] Jordan Viard, Matthieu Latapy, and Clémence Magnien. Computing maximal cliques in link streams. Theoretical Computer Science, 609 (part 1)(??): 245–252, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008701.

Gong:2016:HES

[603] Linming Gong, Shundong Li, Qing Mao, Daoshun Wang, and Jiawei Dou. A homomorphic encryption scheme with adaptive chosen ciphertext security but without random oracle. Theoretical Computer Science, 609 (part 1)(??):253–261, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008725.

Chang:2016:AAB

[604] Fei-Huang Chang, Ma-Lian Chia, David Kuo, Sheng-Chyang Liaw, and Jen-Chun Ling. All-to-all broadcast problems on Cartesian product graphs. Theoretical Computer Science, 609 (part 1)(??):262–271, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008737.

Chak:2016:BPL

[605] Christopher Hanrui Chak, Rusins Freivalds, Frank Stephan, and Henrietta Tan Wan Yik. On block pumpable languages. Theoretical Computer Science, 609 (part 1)(??):272– 285, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008798.

Benevides:2016:MIT

[606] Fabrício Benevides, Victor Campos, Mitre C. Dourado, Rudini M. Sampaio, and Ana Silva. The maximum infection time in the geodesic and monophonic convexities. Theoretical Computer Science, 609 (part 2)(??):

287–295, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008816.

Chen:2016:EAO

[607] Kun-Tze Chen, Chi-Long Li, Hsien-Tai Chiu, and Chin Lung Lu. An efficient algorithm for one-sided block ordering problem under blockinterchange distance. Theoretical Computer Science, 609 (part 2)(??):296– 305, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008932.

Blasius:2016:NPC

[608] Thomas Bläsius and Ignaz Rutter. A new perspective on clustered planarity as a combinatorial embedding problem. Theoretical Computer Science, 609 (part 2)(??):306– 315, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008944.

Golin:2016:ERM

[609] Mordecai Golin, John Iacono, Danny Krizanc, Rajeev Raman, Srinivasa Rao Satti, and Sunil Shende. Encoding 2D range maximum queries. Theoretical Computer Science, 609 (part 2)(??): 316–327, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008956.

Kratsch:2016:ASM

[610] Dieter Kratsch and Van Bang Le. Algorithms solving the Matching Cut problem. Theoretical Computer Science, 609 (part 2)(??):328– 335, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008993.

Gu:2016:PCN

[611] Ran Gu, Xueliang Li, and Zhongmei Qin. Proper connection number of random graphs. Theoretical Computer Science, 609 (part 2)(??): 336–343, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009007.

Tian:2016:SRC

[612] Jing Tian, Xianzhong Zhao, and Yong Shao. On structure and representations of cyclic automata. Theoretical Computer Science, 609 (part 2)(??): 344–360, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009111.

Bonnet:2016:TBM

[613] François Bonnet, Xavier Défago, Thanh Dang Nguyen, and Maria Potop-Butucaru. Tight bound on mobile Byzantine Agreement. Theoretical Computer Science, 609 (part 2)(??): 361–373, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009123.

Brankovic:2016:LTS

[614] Ljiljana Brankovic, Costas S. Iliopoulos, Ritu Kundu, Manal Mohamed, Solon P. Pissis, and Fatima Vayani. Linear-time superbubble identification algorithm for genome assembly. Theoretical Computer Science, 609 (part 2)(??):374– 383, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009147.

Zejnilovic:2016:EMD

[615] Sabina Zejnilović, Dieter Mitsche, João Gomes, and Bruno Sinopoli. Extending the metric dimension to graphs with missing edges. Theoretical Computer Science, 609 (part 2)(??):384– 394, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009159.

Xu:2016:AUP

[616] Ming Xu, Cheng-Chao Huang, Zhi-Bin Li, and Zhenbing Zeng. Analyzing ultimate positivity for solvable systems. Theoretical Computer Science, 609 (part 2)(??):395– 412, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009160.

Hao:2016:PDS

[617] Rong-Xia Hao, Mei-Mei Gu, and Yan-Quan Feng. The pessimistic diagnosabilities of some general regular graphs. Theoretical Computer Science, 609 (part 2)(??):413–420, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009184.

Zhou:2016:CFT

[618] Shuming Zhou, Sulin Song, Xiaoxue Yang, and Lanxiang Chen. On conditional fault tolerance and diagnosability of hierarchical cubic networks. Theoretical Computer Science, 609 (part 2)(??):421–433, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009366.

Chen:2016:SCM

[619] Wenbin Chen, Nagiza F. Samatova, Matthias F. Stallmann, William Hendrix, and Weiqin Ying. On sizeconstrained minimum s–t cut problems and size-constrained dense subgraph problems. Theoretical Computer Science, 609 (part 2)(??):434– 442, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009378.

Zhang:2016:NDD

[620] Zhongzhi Zhang, Shunqi Wu, Mingyun Li, and Francesc Comellas. The number and degree distribution of spanning trees in the Tower of Hanoi graph. Theoretical Computer Science, 609 (part 2)(??):443– 455, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500938X.

Currie:2016:GRB

[621] James Currie and Narad Rampersad. Growth rate of binary words avoiding xxx^{R}. Theoretical Computer Science, 609 (part 2)(??):456– 468, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009780.

Allen:2016:CPW

[622] Emily Allen, F. Blanchet-Sadri, Michelle Bodnar, Brian Bowers, Joe Hidakatsu, and John Lensmire. Combinatorics on partial word borders. Theoretical Computer Science, 609 (part 2)(??):469– 493, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009809.

Chen:2016:PDP

[623] Xie-Bin Chen. Paired 2-disjoint path covers of faulty k-ary ncubes. Theoretical Computer Science, 609 (part 2)(??):494–499, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009810.

Fang:2016:CLC

[624] Qizhi Fang, Bo Li, Xiaoming Sun, Jia Zhang, and Jialin Zhang. Computing the least-core and nucleolus for threshold cardinality matching games. Theoretical Computer Science, 609 (part 2)(??):500–510, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010129.

Grippo:2016:CPB

[625] Luciano N. Grippo, Martín Matamala, Martín D. Safe, and Maya J. Stein. Convex p-partitions of bipartite graphs. Theoretical Computer Science, 609 (part 2)(??):511– 514, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010117.

Cai:2016:SIC

[626] Zhipeng Cai and Alexander Zelikovsky. Special issue on Computing and Combinatori’s Conference. Theoretical Computer Science, 609 (part 3)(??): 515, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010798.

Bera:2016:MAF

[627] Suman K. Bera, Syamantak Das, and Amit Kumar. Minimizing average flow-time under knapsack constraint. Theoretical Computer Science, 609 (part 3)(??):516–525, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003746.

Chen:2016:AMC

[628] Xujin Chen, Xiaodong Hu, and Changjun Wang. Approximation for the minimum cost doubly resolving set problem. Theoretical Computer Science, 609 (part 3)(??):526– 543, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003060.

Duan:2016:CFB

[629] Zhenhua Duan, Cong Tian, and Nan Zhang. A canonical form based decision procedure and model checking approach for propositional projection temporal logic. Theoretical Computer Science, 609 (part 3)(??): 544–560, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007860.

Jiang:2016:SCS

[630] Minghui Jiang and Haitao Wang. Shortest color-spanning intervals. Theoretical Computer Science, 609 (part 3)(??):561–568, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515000754.

LeGall:2016:QAF

[631] François Le Gall, Harumichi Nishimura, and Seiichiro Tani. Quantum algorithms for finding constant-sized sub-hypergraphs. Theoretical Computer Science, 609 (part 3)(??):569– 582, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008774.

Liu:2016:HLC

[632] Xianmin Liu, Jianzhong Li, and Hong Gao. On the hardness of labeled correlation clustering problem: a parameterized complexity view. Theoretical Computer Science, 609 (part 3)(??): 583–593, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515002376.

Miao:2016:CSQ

[633] Dongjing Miao, Xianmin Liu, and Jianzhong Li. On the complexity of sampling query feedback restricted database repair of functional dependency violations. Theoretical Computer Science, 609 (part 3)(??):594– 605, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001073.

Mishra:2016:OPS

[634] Subhankar Mishra, Thang N. Dinh, My T. Thai, Jungtaek Seo, and Incheol Shin. Optimal packet scan against malicious attacks in smart grids. Theoretical Computer Science, 609 (part 3)(??):

606–619, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007100.

Tong:2016:SHT

[635] Weitian Tong, Randy Goebel, and Guohui Lin. Smoothed heights of tries and Patricia tries. Theoretical Computer Science, 609 (part 3)(??): 620–626, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001061.

Wu:2016:AAB

[636] Di Wu, Zhao Zhang, and Weili Wu. Approximation algorithm for the balanced 2-connected k-partition problem. Theoretical Computer Science, 609 (part 3)(??):627–638, January 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515000985.

Zhang:2016:CAS

[637] Nan Zhang, Zhenhua Duan, and Cong Tian. A complete axiom system for propositional projection temporal logic with cylinder computation model. Theoretical Computer Science, 609 (part 3)(??):639– 657, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515003989.

Zhang:2016:NAA

[638] Peng Zhang. A new approximation algorithm for the unbalanced Min s–t Cut problem. Theoretical Computer Science, 609 (part 3)(??): 658–665, January 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001760.

Anonymous:2016:EBd

[639] Anonymous. Editorial Board. Theoretical Computer Science, 610 (part B) (??):iii–ix, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011068.

Anonymous:2016:EBe

[640] Anonymous. Editorial Board. Theoretical Computer Science, 610 (part A) (??):iii–ix, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010920.

Jurgensen:2016:DCF

[641] Helmut Jürgensen, Juhani Karhumäki, and Alexander Okhotin. Descriptional complexity of formal systems. Theoretical Computer Science, 610 (part A)(??):1, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010956.

Cho:2016:SCI

[642] Da-Jung Cho, Yo-Sub Han, Sang-Ki Ko, and Kai Salomaa. State complexity of inversion operations. Theoretical Computer Science, 610 (part A) (??):2–12, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515003242.

Diekert:2016:SLD

[643] Volker Diekert and Manfred Kufleitner. A survey on the local divisor technique. Theoretical Computer Science, 610 (part A)(??):13– 23, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006295.

Dorzweiler:2016:PNP

[644] Olga Dorzweiler, Thomas Flamm, Andreas Krebs, and Michael Ludwig. Positive and negative proofs for circuits and branching programs. Theoretical Computer Science, 610 (part A)(??): 24–36, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007884.

Gehrke:2016:UWF

[645] Mai Gehrke, Andreas Krebs, and Jean-Éric Pin. Ultrafilters on words for a fragment of logic. Theoretical Computer Science, 610 (part A)(??): 37–58, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007276.

Holzer:2016:BSR

[646] Markus Holzer and Sebastian Jakobi. Boundary sets of regular and contextfree languages. Theoretical Computer Science, 610 (part A)(??):59– 77, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006660.

Jirasek:2016:PFL

[647] Jozef Jirásek, Galina Jirásková, Monika Krausová, Peter Mlynárcik, and Juraj Sebej. Prefix-free languages: Left and right quotient and reversal. Theoretical Computer Science, 610 (part A) (??):78–90, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007781.

Karandikar:2016:SCC

[648] P. Karandikar, M. Niewerth, and Ph. Schnoebelen. On the state complexity of closures and interiors of regular languages with subwords and superwords. Theoretical Computer Science, 610 (part A)(??):91– 107, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008683.

Palioudakis:2016:OSC

[649] Alexandros Palioudakis, Kai Salomaa, and Selim G. Akl. Operational state complexity of unary NFAs with finite nondeterminism. Theoretical Computer Science, 610 (part A)(??):108– 120, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006271.

Prusa:2016:NRT

[650] Daniel Prusa. Non-recursive tradeoffs between two-dimensional automata and grammars. Theoretical Computer Science, 610 (part A)(??):121– 132, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515004740.

Anonymous:2016:GES

[651] Anonymous. Guest Editorial: Special issue on Structural Information and Communication Complexity. Theoretical Computer Science, 610 (part B)(??):133–134, January 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011317.

Bar-Noy:2016:CGS

[652] Amotz Bar-Noy, Ben Baumer, and Dror Rawitz. Changing of the guards: Strip cover with duty cycling. Theoretical Computer Science, 610 (part B)(??):135–148, January 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514006549.

Caragiannis:2016:SLB

[653] Ioannis Caragiannis and Christos Kalaitzis. Space lower bounds for lowstretch greedy embeddings. Theoretical Computer Science, 610 (part B)(??): 149–157, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515001000.

DAngelo:2016:GRA

[654] Gianlorenzo D’Angelo, Gabriele Di Stefano, Ralf Klasing, and Alfredo Navarra. Gathering of robots on anonymous grids and trees without multiplicity detection. Theoretical Computer Science, 610 (part B)(??):158– 168, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397514005131.

Balamohan:2016:EUD

[655] B. Balamohan, S. Dobrev, P. Flocchini, and N. Santoro. Exploring an unknown dangerous graph with a constant number of tokens. Theoretical Computer Science, 610 (part B)(??): 169–181, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751400543X.

Yu:2016:DMM

[656] Dongxiao Yu, Qiang-Sheng Hua, Yuexuan Wang, Haisheng Tan, and Francis C. M. Lau. Distributed multiple-message broadcast in wireless ad hoc networks under the SINR model. Theoretical Computer Science, 610 (part B)(??):182– 191, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397514005106.

Eftekhari:2016:SCS

[657] Mohsen Eftekhari, Evangelos Kranakis, Fraser MacQuarrie, Oscar Morales-Ponce, and Lata Narayanan. Strong connectivity of sensor networks with double antennae. Theoretical Computer Science, 610 (part B)(??):192– 203, January 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515001140.

Hasemann:2016:DLA

[658] Henning Hasemann, Juho Hirvonen, Joel Rybicki, and Jukka Suomela. Deterministic local algorithms, unique identifiers, and fractional graph colouring. Theoretical Computer Science, 610 (part B)(??):204–217, January 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751400512X.

Anonymous:2016:EBf

[659] Anonymous. Editorial Board. Theoretical Computer Science, 611(??): iii–ix, January 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011512.

Fu:2016:TI

[660] Yuxi Fu. Theory of interaction. Theoretical Computer Science, 611(??): 1–49, January 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500688X.

Blanqui:2016:TRR

[661] Frédéric Blanqui. Termination of rewrite relations on λ-terms based on Girard’s notion of reducibility. Theoretical Computer Science, 611(??): 50–86, January 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006908.

Yokoyama:2016:FRF

[662] Tetsuo Yokoyama, Holger Bock Axelsen, and Robert Glück. Fundamentals of reversible flowchart languages. Theoretical Computer Science, 611(??): 87–115, January 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500691X.

Xu:2016:MUF

[663] Ming Xu, Lijun Zhang, David N. Jansen, Huibiao Zhu, and Zongyuan Yang. Multiphase until formulas over Markov reward models: an algebraic approach. Theoretical Computer Science, 611(??):116– 135, January 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006921.

Muscholl:2016:CLC

[664] Anca Muscholl and Sven Schewe. Controlling loosely cooperating processes.

Theoretical Computer Science, 611(??): 136–141, January 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006891.

Anonymous:2016:EBg

[665] Anonymous. Editorial Board. Theoretical Computer Science, 612(??): iii–ix, January 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011627.

Abdulla:2016:PVT

[666] Parosh Aziz Abdulla, Giorgio Delzanno, Othmane Rezine, Arnaud Sangnier, and Riccardo Traverso. Parameterized verification of time-sensitive models of ad hoc network protocols. Theoretical Computer Science, 612(??): 1–22, January 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006933.

Xu:2016:CHL

[667] Zhaowei Xu, Yuefei Sui, and Wenhui Zhang. Completeness of Hoare logic with inputs over the standard model. Theoretical Computer Science, 612(??): 23–28, January 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007239.

Luttik:2016:UPD

[668] Bas Luttik. Unique parallel decomposition in branching and weak bisimulation semantics. Theoretical Computer Science, 612(??):29– 44, January 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008968.

Pan:2016:MCC

[669] Haiyu Pan, Yongming Li, Yongzhi Cao, and Zhanyou Ma. Model checking computation tree logic over finite lattices. Theoretical Computer Science, 612(??): 45–62, January 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500897X.

Liu:2016:LCE

[670] Chanjuan Liu, Fenrong Liu, Kaile Su, and Enqiang Zhu. A logical characterization of extensive games with short sight. Theoretical Computer Science, 612(??):63– 82, January 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515008981.

Slivovsky:2016:SRD

[671] Friedrich Slivovsky and Stefan Szeider. Soundness of Q-resolution with dependency schemes. Theoretical Computer Science, 612(??):83– 101, January 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009135.

Dreier:2016:EDU

[672] Jannik Dreier, Cristian Ene, Pascal Lafourcade, and Yassine Lakhnech. On the existence and decidability of unique decompositions of processes in the applied π-calculus. Theoretical Computer Science, 612(??):102– 125, January 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011196.

Wu:2016:RDI

[673] Mingyuan Wu, Lankun Guo, and Qingguo Li. A representation of L-domains by information systems. Theoretical Computer Science, 612(??):126– 136, January 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011238.

Gavryushkin:2016:RAE

[674] Alex Gavryushkin, Bakhadyr Khoussainov, and Frank Stephan. Reducibilities among equivalence relations induced by recursively enumerable structures. Theoretical Computer Science, 612(??):137–152, January 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011287.

Anonymous:2016:EBh

[675] Anonymous. Editorial Board. Theoretical Computer Science, 613(??): iii–ix, February 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011925.

Yamakami:2016:PGA

[676] Tomoyuki Yamakami. Pseudorandom generators against advised context-free languages. Theoretical Computer Science, 613(??):1–27, February 1, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009196.

Naude:2016:RPS

[677] Kevin A. Naudé. Refined pivot selection for maximal clique enumeration in graphs. Theoretical Computer Science, 613(??):28– 37, February 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010130.

Fomin:2016:FPG

[678] Fedor V. Fomin, Pinar Heggernes, and Erik Jan van Leeuwen. The Firefighter problem on graph classes. Theoretical Computer Science, 613(??): 38–50, February 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010853.

Erlebach:2016:QCA

[679] Thomas Erlebach, Michael Hoffmann, and Frank Kammer. Querycompetitive algorithms for cheapest set problems under uncertainty. Theoretical Computer Science, 613(??):

51–64, February 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515010865.

Bilo:2016:PEF

[680] Vittorio Bilò, Angelo Fanelli, Michele Flammini, Gianpiero Monaco, and Luca Moscardelli. The price of envyfreeness in machine scheduling. Theoretical Computer Science, 613(??): 65–78, February 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011123.

Nonner:2016:CMB

[681] Tim Nonner. Capacitated Max-Batching with interval graph compatibilities. Theoretical Computer Science, 613(??):79–93, February 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397515011147.

DiIanni:2016:MMC

[682] Miriam Di Ianni, Giorgio Gambosi, Gianluca Rossi, and Paola Vocca. Minmax communities in graphs: Complexity and computational properties. Theoretical Computer Science, 613(??): 94–114, February 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011202.

Meyer:2016:GPP

[683] Bernd Meyer. Generalized Pete’s Pike is PSPACE-complete. Theoretical Computer Science, 613(??):115– 125, February 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011226.

Anonymous:2016:EBi

[684] Anonymous. Editorial Board. Theoretical Computer Science, 614(??): iii–ix, February 8, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012128.

Berline:2016:SAM

[685] Chantal Berline and Klaus Grue. A synthetic axiomatization of Map Theory. Theoretical Computer Science, 614 (??):1–62, February 8, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011135.

Anderson:2016:CLB

[686] Gabrielle Anderson and David Pym. A calculus and logic of bunched resources and processes. Theoretical Computer Science, 614(??):63– 96, February 8, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011214.

Dittmann:2016:GOP

[687] Christoph Dittmann, Stephan Kreutzer, and Alexandru I. Tomescu. Graph operations on parity games and polynomial-time algorithms. Theoretical Computer Science, 614(??):97– 108, February 8, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011305.

Anonymous:2016:EBj

[688] Anonymous. Editorial Board. Theoretical Computer Science, 615(??): iii–ix, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515012268.

Hellmuth:2016:FFC

[689] Marc Hellmuth and Florian Lehner. Fast factorization of Cartesian products of (directed) hypergraphs. Theoretical Computer Science, 615(??): 1–11, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751501124X.

Klouda:2016:SDB

[690] Karel Klouda and Katerina Medková. Synchronizing delay for binary uniform morphisms. Theoretical Computer Science, 615(??):12– 22, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011299.

Chlebus:2016:SWM

[691] Bogdan S. Chlebus, Gianluca De Marco, and Dariusz R. Kowalski. Scalable wake-up of multi-channel single-hop radio networks. Theoretical Computer Science, 615(??):23– 44, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011342.

Seo:2016:RHI

[692] Jae Hong Seo and Keita Emura. Revocable hierarchical identity-based encryption via history-free approach. Theoretical Computer Science, 615(??): 45–60, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011354.

Bhattacharya:2016:SEA

[693] Binay K. Bhattacharya, Subhas C. Nandy, and Sasanka Roy. Spaceefficient algorithm for computing a centerpoint of a set of points in R^{2}. Theoretical Computer Science, 615(??): 61–70, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011366.

Zhang:2016:RMR

[694] Mingzu Zhang, Lianzhu Zhang, and Xing Feng. Reliability measures in relation to the h-extra edge-connectivity of folded hypercubes. Theoretical Computer Science, 615(??):71– 77, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011378.

Hsieh:2016:RCL

[695] Sun-Yuan Hsieh, Hong-Wen Huang, and Chia-Wei Lee. {2, 3}-restricted connectivity of locally twisted cubes. Theoretical Computer Science, 615(??): 78–90, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751501138X.

Cheng:2016:SMP

[696] Eddie Cheng, Justin Kelm, and Joseph Renzi. Strong matching preclusion of (n, k)-star graphs. Theoretical Computer Science, 615(??):91– 101, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011391.

Kamei:2016:ASS

[697] Sayaka Kamei, Tomoko Izumi, and Yukiko Yamauchi. An asynchronous self-stabilizing approximation for the minimum CDS with safe convergence in UDGs. Theoretical Computer Science, 615(??):102–119, February 15, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011408.

Saether:2016:HCW

[698] Sigve Hortemo Sæther and Martin Vatshelle. Hardness of computing width parameters based on branch decompositions over the vertex set. Theoretical Computer Science, 615(??): 120–125, February 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011251.

Anonymous:2016:EBk

[699] Anonymous. Editorial Board. Theoretical Computer Science, 616(??): iii–ix, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000104.

Gurski:2016:DNW

[700] Frank Gurski, Egon Wanke, and Eda Yilmaz. Directed NLC-width. Theoretical Computer Science, 616 (??):1–17, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009779.

Engelfriet:2016:LAR

[701] Joost Engelfriet, Sebastian Maneth, and Helmut Seidl. Look-ahead removal for total deterministic topdown tree transducers. Theoretical Computer Science, 616(??):18– 58, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011445.

Avgustinovich:2016:GCC

[702] Sergey Avgustinovich, Sergey Kitaev, Vladimir N. Potapov, and Vincent Vajnovszki. Gray coding cubic planar maps. Theoretical Computer Science, 616(??):59–69, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011743.

Bose:2016:TTB

[703] Prosenjit Bose, Jean-Lou De Carufel, Pat Morin, André van Renssen, and Sander Verdonschot. Towards tight bounds on theta-graphs: More is not always better. Theoretical Computer Science, 616(??):70– 93, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011780.

Jiang:2016:SJR

[704] Dakui Jiang and Jiayin Tan. Scheduling with job rejection and nonsimultaneous machine available time on unrelated parallel machines. Theoretical Computer Science, 616(??): 94–99, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011810.

Lu:2016:PCC

[705] Xiao-Tao Lü, Jin Chen, Ying-Jun Guo, and Zhi-Xiong Wen. On the permutation complexity of the Cantor-like sequences. Theoretical Computer Science, 616(??):100– 110, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011834.

Dey:2016:KCP

[706] Palash Dey, Neeldhara Misra, and Y. Narahari. Kernelization complexity of possible winner and coalitional manipulation problems in voting. Theoretical Computer Science, 616 (??):111–125, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011846.

Bereczky:2016:QTG

[707] Nikolett Bereczky, Amalia Duch, Krisztián Németh, and Salvador Roura. Quad-k d trees: a general framework for k d trees and quad trees. Theoretical Computer Science, 616 (??):126–140, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012001.

Jez:2016:RSA

[708] Artur Jez. A really simple approximation of smallest grammar. Theoretical Computer Science, 616(??): 141–150, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012025.

Anonymous:2016:EBl

[709] Anonymous. Editorial Board. Theoretical Computer Science, 617(??): iii–ix, February 29, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000347.

Shibata:2016:PGM

[710] Masahiro Shibata, Shinji Kawai, Fukuhito Ooshita, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Partial gathering of mobile agents in asynchronous unidirectional rings. Theoretical Computer Science, 617(??): 1–11, February 29, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008270.

Astete-Morales:2016:SCR

[711] Sandra Astete-Morales, Marie-Liesse Cauwet, Jialin Liu, and Olivier Teytaud. Simple and cumulative regret for continuous noisy optimization. Theoretical Computer Science, 617(??): 12–27, February 29, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515008804.

Boissonnat:2016:CRS

[712] Jean-Daniel Boissonnat and Dorian Mazauric. On the complexity of the representation of simplicial complexes by trees. Theoretical Computer Science, 617(??):28– 44, February 29, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515012049.

Park:2016:UMM

[713] Jung-Heum Park. Unpaired manyto-many disjoint path covers in restricted hypercube-like graphs. Theoretical Computer Science, 617(??):

45–64, February 29, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012062.

Tian:2016:OSU

[714] Ji Tian, Qian Wang, Ruyan Fu, and Jinjiang Yuan. Online scheduling on the unbounded drop-line batch machines to minimize the maximum delivery completion time. Theoretical Computer Science, 617(??):65– 68, February 29, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000025.

Anonymous:2016:EBm

[715] Anonymous. Editorial Board. Theoretical Computer Science, 618(??): iii–ix, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000761.

Xia:2016:RSN

[716] Yingjie Xia, Xuejiao Liu, Fubiao Xia, and Guilin Wang. A reduction of security notions in designated confirmer signatures. Theoretical Computer Science, 618(??):1–20, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751501186X.

Lin:2016:REC

[717] Limei Lin, Li Xu, and Shuming Zhou. Relating the extra connectivity and the conditional diagnosability of regular graphs under the comparison model. Theoretical Computer Science, 618 (??):21–29, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012013.

Zhu:2016:SMM

[718] Haibin Zhu, Dongning Liu, Siqin Zhang, Yu Zhu, Luyao Teng, and Shaohua Teng. Solving the Many to Many assignment problem by improving the Kuhn–Munkr’s algorithm with backtracking. Theoretical Computer Science, 618(??): 30–41, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000037.

Schmid:2016:CEF

[719] Markus L. Schmid. Computing equality-free and repetitive string factorisations. Theoretical Computer Science, 618(??):42–51, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516000177.

Okhotin:2016:IDL

[720] Alexander Okhotin. Input-driven languages are linear conjunctive. Theoretical Computer Science, 618(??): 52–71, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000189.

Bova:2016:QCQ

[721] Simone Bova, Robert Ganian, and Stefan Szeider. Quantified conjunctive queries on partially ordered sets. Theoretical Computer Science, 618 (??):72–84, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000219.

Abdessaied:2016:CRC

[722] Nabila Abdessaied, Matthew Amy, Rolf Drechsler, and Mathias Soeken. Complexity of reversible circuits and their quantum implementations. Theoretical Computer Science, 618(??): 85–106, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000220.

Nourine:2016:EDA

[723] Lhouari Nourine and Jean-Marc Petit. Extended dualization: Application to maximal pattern mining. Theoretical Computer Science, 618(??): 107–121, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751600044X.

Filho:2016:HCC

[724] Hélio B. Macêdo Filho, Raphael C. S. Machado, and Celina M. H. de Figueiredo. Hierarchical complexity of 2-clique-colouring weakly chordal graphs and perfect graphs having cliques of size at least 3. Theoretical Computer Science, 618(??):

122–134, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000517.

Hew:2016:WBP

[725] Patrick Chisan Hew. Working in binary protects the repetends of 1/3^{h}: Comment on Colussi’s ‘The convergence classes of Collatz function’. Theoretical Computer Science, 618(??): 135–141, March 7, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515012037. See

[2].

Anonymous:2016:EBn

[726] Anonymous. Editorial Board. Theoretical Computer Science, 619(??): iii–ix, March 14, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000670.

Attie:2016:FSC

[727] Paul C. Attie. Finite-state concurrent programs can be expressed in pairwise normal form. Theoretical Computer Science, 619(??): 1–31, March 14, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011184.

Courcelle:2016:CFA

[728] Bruno Courcelle and Irène Durand. Computations by fly-automata beyond monadic second-order logic. Theoretical Computer Science, 619(??): 32–67, March 14, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011871.

Jez:2016:LGS

[729] Artur Jez and Alexander Okhotin. Least and greatest solutions of equations over sets of integers. Theoretical Computer Science, 619(??): 68–86, March 14, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000244.

Ziemianski:2016:ESP

[730] Krzysztof Ziemiański. On execution spaces of PV-programs. Theoretical Computer Science, 619(??): 87–98, March 14, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000293.

Anonymous:2016:EBo

[731] Anonymous. Editorial Board. Theoretical Computer Science, 620(??): iii–ix, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001067.

Jain:2016:GEF

[732] Sanjay Jain, Rémi Munos, Frank Stephan, and Thomas Zeugmann. Guest Editors’ foreword [preface to the special issue dedicated to the 24th International Conference on Algorithmic Learning Theory (ALT 2013)]. Theoretical Computer Science, 620 (??):1–3, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009391. Held

at the National University of Singapore, Singapore, October 6–9, 2013.

Uitto:2016:CR

[733] Jara Uitto and Roger Wattenhofer. On competitive recommendations. Theoretical Computer Science, 620(??): 4–14, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009408.

Gao:2016:PLR

[734] Ziyuan Gao, Frank Stephan, and Sandra Zilles. Partial learning of recursively enumerable languages. Theoretical Computer Science, 620(??): 15–32, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500941X.

Case:2016:TSI

[735] John Case and Timo Kötzing. Topological separations in inductive inference. Theoretical Computer Science, 620(??): 33–45, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009421.

Shibata:2016:PLC

[736] Chihiro Shibata and Ryo Yoshinaka. Probabilistic learnability of contextfree grammars with basic distributional properties from positive examples. Theoretical Computer Science, 620(??): 46–72, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009433.

Darnstadt:2016:OCS

[737] Malte Darnstädt, Thorsten Kiss, Hans Ulrich Simon, and Sandra Zilles. Order compression schemes. Theoretical Computer Science, 620(??): 73–90, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009445.

Choromanska:2016:DPL

[738] Anna Choromanska, Krzysztof Choromanski, Geetha Jagannathan, and Claire Monteleoni. Differentiallyprivate learning of low dimensional manifolds. Theoretical Computer Science, 620(??):91–104, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515009457.

Gottlieb:2016:AMD

[739] Lee-Ad Gottlieb, Aryeh Kontorovich, and Robert Krauthgamer. Adaptive metric dimensionality reduction. Theoretical Computer Science, 620(??): 105–118, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009469.

Khaleghi:2016:NMC

[740] Azadeh Khaleghi and Daniil Ryabko. Nonparametric multiple change point estimation in highly dependent time series. Theoretical Computer Science, 620 (??):119–133, March 21, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009470.

Anonymous:2016:EBp

[741] Anonymous. Editorial Board. Theoretical Computer Science, 621(??): iii–ix, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001237.

Leone:2016:GRV

[742] Pierre Leone and Kasun Samarasinghe. Geographic routing on Virtual Raw Anchor Coordinate systems. Theoretical Computer Science, 621(??):1–13, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011998.

Belazzougui:2016:OVR

[743] Djamal Belazzougui. Optimal Las Vegas reduction from one-way set reconciliation to error correction. Theoretical Computer Science, 621(??): 14–21, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000438.

Gozupek:2016:CMC

[744] Didem Gözüpek, Hadas Shachnai, Mordechai Shalom, and Shmuel Zaks. Constructing minimum changeover cost arborescenses in bounded treewidth graphs. Theoretical Computer Science, 621(??):22–36, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000463.

Keshavarz-Kohjerdi:2016:HPS

[745] Fatemeh Keshavarz-Kohjerdi and Alireza Bagheri. Hamiltonian paths in L-shaped grid graphs. Theoretical Computer Science, 621(??):37– 56, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000487.

Flocchini:2016:RCM

[746] P. Flocchini, N. Santoro, G. Viglietta, and M. Yamashita. Rendezvous with constant memory. Theoretical Computer Science, 621(??): 57–72, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000499.

Su:2016:BWT

[747] Yu-Hsuan Su, Ching-Chi Lin, and D. T. Lee. Broadcasting in weighted trees under the postal model. Theoretical Computer Science, 621(??):

73–81, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000554.

Liu:2016:MBM

[748] Xingwu Liu and Shang-Hua Teng. Maximum bipartite matchings with low rank data: Locality and perturbation analysis. Theoretical Computer Science, 621(??):82–91, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397516000578.

Dereniowski:2016:TRL

[749] Dariusz Dereniowski and Andrzej Pelc. Topology recognition and leader election in colored networks. Theoretical Computer Science, 621(??):92– 102, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751600061X.

Fu:2016:IAK

[750] Lianzhi Fu. Improvements in the analysis of Kannan’s CVP algorithm. Theoretical Computer Science, 621 (??):103–112, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000621.

Anonymous:2016:EBq

[751] Anonymous. Editorial Board. Theoretical Computer Science, 622(??): iii–ix, April 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001493.

Ben-Basat:2016:DLP

[752] Ran Ben-Basat, Ariel Gabizon, and Meirav Zehavi. The k-distinct language: Parameterized automata constructions. Theoretical Computer Science, 622(??):1–15, April 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516000530.

Belikovetsky:2016:LRG

[753] Sofia Belikovetsky and Tami Tamir. Load rebalancing games in dynamic systems with migration costs. Theoretical Computer Science, 622(??): 16–33, April 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000542.

Mizuki:2016:CBP

[754] Takaaki Mizuki. Card-based protocols for securely computing the conjunction of multiple variables. Theoretical Computer Science, 622(??): 34–44, April 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000839.

Chen:2016:TSS

[755] Youjun Chen, Lingfa Lu, and Jinjiang Yuan. Two-stage scheduling on identical machines with assignable delivery times to minimize the maximum delivery completion time. Theoretical Computer Science, 622(??): 45–65, April 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000888.

Garnero:2016:PCD

[756] Valentin Garnero and Mathias Weller. Parameterized certificate dispersal and its variants. Theoretical Computer Science, 622(??):66–78, April 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397516000906.

Yang:2016:EAW

[757] Yongjie Yang and Jiong Guo. Exact algorithms for weighted and unweighted Borda manipulation problems. Theoretical Computer Science, 622(??):79–89, April 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001134.

Bazgan:2016:FLD

[758] Cristina Bazgan, Robert Bredereck, Sepp Hartung, André Nichterlein, and Gerhard J. Woeginger. Finding large degree-anonymous subgraphs is hard. Theoretical Computer Science, 622 (??):90–110, April 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001146.

Fuchs:2016:PNR

[759] M. Fuchs, C.-K. Lee, and G.- R. Yu. On 2-protected nodes in random digital trees. Theoretical Computer Science, 622(??):111– 122, April 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001171.

Anonymous:2016:EBr

[760] Anonymous. Editorial Board. Theoretical Computer Science, 623(??): iii–ix, April 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001675.

Kari:2016:Ea

[761] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 623(??):1, April 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001742.

Garcia-Ramos:2016:LBE

[762] Felipe García-Ramos. Limit behaviour of μ-equicontinuous cellular automata. Theoretical Computer Science, 623 (??):2–14, April 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007306.

Colvin:2016:MAN

[763] Robert J. Colvin. Modelling and analysing neural networks using a hybrid process algebra. Theoretical Computer Science, 623(??):15– 64, April 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007665.

Sosik:2016:SPC

[764] Petr Sosík and Miroslav Langer. Small (purely) catalytic P systems simulating register machines. Theoretical Computer Science, 623(??): 65–74, April 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751500849X.

Li:2016:LBS

[765] Lvzhou Li and Daowen Qiu. Lower bounds on the size of semi-quantum finite automata. Theoretical Computer Science, 623(??):75–82, April 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397515008713.

Pan:2016:FMP

[766] Linqiang Pan, Gheorghe Paun, and Bosheng Song. Flat maximal parallelism in P systems with promoters. Theoretical Computer Science, 623 (??):83–91, April 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515009202.

Aman:2016:MVW

[767] Bogdan Aman and Gabriel Ciobanu. Modelling and verification of weighted spiking neural systems. Theoretical Computer Science, 623(??):92– 102, April 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009792.

Azimi:2016:CMC

[768] Sepinoud Azimi, Cristian Gratie, Sergiu Ivanov, Luca Manzoni, Ion Petre, and Antonio E. Porreca. Complexity of model checking for reaction systems. Theoretical Computer Science, 623(??):103–113, April 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515011263.

Barbuti:2016:IDC

[769] Roberto Barbuti, Roberta Gori, Francesca Levi, and Paolo Milazzo. Investigating dynamic causalities in reaction systems. Theoretical Computer Science, 623(??): 114–145, April 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011275.

Chen:2016:CV

[770] Xu Chen, Mario J. Pérez-Jiménez, Luis Valencia-Cabrera, Beizhan Wang, and Xiangxiang Zeng. Computing with viruses. Theoretical Computer Science, 623(??):146–159, April 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515011457.

Hellwig:2016:MSC

[771] Michael Hellwig and Hans-Georg Beyer. Mutation strength control via meta evolution strategies on the ellipsoid model. Theoretical Computer Science, 623(??):160– 179, April 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751501172X.

Wu:2016:CLS

[772] Tingfang Wu, Zhiqiang Zhang, Gheorghe Paun, and Linqiang Pan. Celllike spiking neural P systems. Theoretical Computer Science, 623(??): 180–189, April 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515012190.

Anonymous:2016:EBs

[773] Anonymous. Editorial Board. Theoretical Computer Science, 624(??): iii–ix, April 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001870.

Frosini:2016:ADG

[774] Andrea Frosini and Simone Rinaldi. Advances in discrete geometry for computer imagery: Preface. Theoretical Computer Science, 624(??): 1, April 18, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001948.

Esbelin:2016:SCB

[775] Henri-Alex Esbelin and Rémy Malgouyres. Sparse convolution-based digital derivatives, fast estimation for noisy signals and approximation results. Theoretical Computer Science, 624(??):2–24, April 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011792.

Karmakar:2016:RGB

[776] Nilanjana Karmakar, Arindam Biswas, and Partha Bhowmick. Reeb graph based segmentation of articulated components of 3D digital objects. Theoretical Computer Science, 624(??): 25–40, April 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010105.

Gerard:2016:TBQ

[777] Yan Gerard, Antoine Vacavant, and Jean-Marie Favreau. Tight bounds in the quadtree complexity theorem and the maximal number of pixels crossed by a curve of given length. Theoretical Computer Science, 624 (??):41–55, April 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011767.

Biswas:2016:PQO

[778] Ranita Biswas and Partha Bhowmick. From prima quadraginta octant to lattice sphere through primitive integer operations. Theoretical Computer Science, 624(??):56– 72, April 18, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515010178.

Lachaud:2016:OSA

[779] Jacques-Olivier Lachaud, Xavier Provençal, and Tristan Roussillon. An outputsensitive algorithm to compute the normal vector of a digital plane. Theoretical Computer Science, 624(??): 73–88, April 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515010828.

Brunetti:2016:BAD

[780] S. Brunetti, P. Dulio, and C. Peri. On bounded additivity in discrete tomography. Theoretical Computer Science, 624(??):89–100, April 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439751501083X.

Domenjoud:2016:PLT

[781] Eric Domenjoud, Xavier Provençal, and Laurent Vuillon. Palindromic language of thin discrete planes. Theoretical Computer Science, 624(??): 101–108, April 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515010841.

Frosini:2016:GPM

[782] Andrea Frosini, Veronica Guerrini, and Simone Rinaldi. Geometric properties of matrices induced by pattern avoidance. Theoretical Computer Science, 624(??):109– 120, April 18, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010099.

Masse:2016:EOD

[783] Alexandre Blondin Massé, Srecko Brlek, and Hugo Tremblay. Efficient operations on discrete paths. Theoretical Computer Science, 624(??): 121–135, April 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006672.

Brocchi:2016:RCP

[784] Stefano Brocchi, Paolo Dulio, and Silvia M. C. Pagani. Reconstruction of convex polyominoes with a blocking component. Theoretical Computer Science, 624(??):136– 146, April 18, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011779.

Anonymous:2016:EBt

[785] Anonymous. Editorial Board. Theoretical Computer Science, 625(??): iii–ix, April 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002140.

Gange:2016:CRP

[786] Graeme Gange, Jorge A. Navas, Peter Schachte, Harald Søndergaard, and Peter J. Stuckey. A complete refinement procedure for regular separability of context-free languages.

Theoretical Computer Science, 625 (??):1–24, April 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000505.

Lanese:2016:RHO

[787] Ivan Lanese, Claudio Antares Mezzina, and Jean-Bernard Stefani. Reversibility in the higher-order π-calculus. Theoretical Computer Science, 625(??): 25–84, April 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001419.

Benerecetti:2016:TRS

[788] Massimo Benerecetti and Adriano Peron. Timed recursive state machines: Expressiveness and complexity. Theoretical Computer Science, 625 (??):85–124, April 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001432.

Aschieri:2016:NDC

[789] Federico Aschieri and Margherita Zorzi. On natural deduction in classical first-order logic: Curry–Howard correspondence, strong normalization and Herbrand’s theorem. Theoretical Computer Science, 625(??): 125–146, April 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001596.

Anonymous:2016:EBu

[790] Anonymous. Editorial Board. Theoretical Computer Science, 626(??): iii–ix, May 2, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002280.

Hudry:2016:MRC

[791] Olivier Hudry and Antoine Lobstein. More results on the complexity of identifying problems in graphs. Theoretical Computer Science, 626 (??):1–12, May 2, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000451.

deRougemont:2016:ACT

[792] Michel de Rougemont and Adrien Vieilleribière. Approximate consistency for transformations on words and trees. Theoretical Computer Science, 626 (??):13–39, May 2, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000566.

Cournier:2016:EPS

[793] Alain Cournier, Ajoy K. Datta, Stéphane Devismes, Franck Petit, and Vincent Villain. The expressive power of snap-stabilization. Theoretical Computer Science, 626(??): 40–66, May 2, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000608.

Pouly:2016:CCS

[794] Amaury Pouly and Daniel S. Graça. Computational complexity of solving polynomial differential equations over unbounded domains. Theoretical Computer Science, 626(??):67– 82, May 2, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000918.

Jain:2016:TAS

[795] Sanjay Jain, Bakhadyr Khoussainov, Philipp Schlicht, and Frank Stephan. Tree-automatic scattered linear orders. Theoretical Computer Science, 626(??):83–96, May 2, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001183.

Radons:2016:DSP

[796] Manuel Radons. Direct solution of piecewise linear systems. Theoretical Computer Science, 626(??): 97–109, May 2, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001304.

Datta:2016:CSS

[797] Ajoy K. Datta, Stéphane Devismes, Karel Heurtefeux, Lawrence L. Larmore, and Yvan Rivierre. Competitive self-stabilizing k-clustering. Theoretical Computer Science, 626(??): 110–133, May 2, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001316.

Reich:2016:CML

[798] Alexander Reich. Complexity of the maximum leaf spanning tree problem on planar and regular graphs. Theoretical Computer Science, 626(??): 134–143, May 2, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001328.

Yuan:2016:GNC

[799] Jun Yuan, Aixia Liu, Xiao Qin, Jifu Zhang, and Jing Li. g-Good-neighbor conditional diagnosability measures for 3-ary n-cube networks. Theoretical Computer Science, 626(??): 144–162, May 2, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751600133X.

Anonymous:2016:EBv

[800] Anonymous. Editorial Board. Theoretical Computer Science, 627(??): iii–ix, May 9, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300056.

Glacet:2016:IDE

[801] Christian Glacet, Nicolas Hanusse, and David Ilcinkas. The impact of dynamic events on the number of errors in networks. Theoretical Computer Science, 627(??):1–12, May 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001341.

Ben-Moshe:2016:OBA

[802] Boaz Ben-Moshe, Michael Elkin, Lee-Ad Gottlieb, and Eran Omri. Optimizing budget allocation for center and median points. Theoretical Computer Science, 627(??):13– 25, May 9, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001353.

Qiu:2016:ACA

[803] Xian Qiu and Walter Kern. Approximate core allocations and integrality gap for the bin packing game. Theoretical Computer Science, 627 (??):26–35, May 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001407.

Hao:2016:RBC

[804] Rong-Xia Hao, Zeng-Xian Tian, and Jun-Ming Xu. Relationship between conditional diagnosability and 2-extra connectivity of symmetric graphs. Theoretical Computer Science, 627 (??):36–53, May 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001559.

Losemann:2016:CPD

[805] Katja Losemann, Wim Martens, and Matthias Niewerth. Closure properties and descriptional complexity of deterministic regular expressions. Theoretical Computer Science, 627 (??):54–70, May 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001584.

Ibarra:2016:EEM

[806] Oscar H. Ibarra and Ian McQuillan. The effect of end-markers on counter machines and commutativity. Theoretical Computer Science, 627 (??):71–81, May 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001766.

Yang:2016:CFT

[807] Da-Wei Yang and Mei-Mei Gu. Conditional fault-tolerant edge-bipancyclicity of hypercubes with faulty vertices and edges. Theoretical Computer Science, 627(??):82–89, May 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001791.

Sopena:2016:MNS

[808] Éric Sopena. i-Mark: a new subtraction division game. Theoretical Computer Science, 627(??):90– 101, May 9, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001808.

Kuo:2016:EPE

[809] Che-Nan Kuo and Iain A. Stewart. Edge-pancyclicity and edgebipancyclicity of faulty folded hypercubes. Theoretical Computer Science, 627(??):102–106, May 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001602.

Kinnersley:2016:CFR

[810] William B. Kinnersley, Paweł Pralat, and Douglas B. West. To catch a falling robber. Theoretical Computer Science, 627(??):107–111, May 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001614.

Anonymous:2016:EBw

[811] Anonymous. Editorial Board. Theoretical Computer Science, 630(??): iii–ix, May 30, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300767.

Angel:2016:TMM

[812] Eric Angel, Evripidis Bampis, Vincent Chau, and Nguyen Kim Thang. Throughput maximization in multiprocessor speed-scaling. Theoretical Computer Science, 630(??):1– 12, May 30, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751600222X.

Yen:2016:WRD

[813] William Chung-Kung Yen. Weighted restrained domination in subclasses of planar graphs. Theoretical Computer Science, 630(??):13–25, May 30, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002334.

Busbait:2016:DTT

[814] Monther Busbait and Mikhail Moshkov. Diagnosis of three types of constant faults in read-once contact networks over finite bases. Theoretical Computer Science, 630(??):26– 42, May 30, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002346.

Hoffmann:2016:LTA

[815] Stefan Hoffmann, Alina Elterman, and Egon Wanke. A linear time algorithm for metric dimension of cactus block graphs. Theoretical Computer Science, 630(??):43–62, May 30, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002358.

Golovach:2016:EMC

[816] Petr A. Golovach, Pinar Heggernes, and Dieter Kratsch. Enumerating minimal connected dominating sets in graphs of bounded chordality. Theoretical Computer Science, 630(??): 63–75, May 30, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300159.

Wang:2016:SCA

[817] Qi Wang, Xiubin Fan, Hongyan Zang, and Yu Wang. The space complexity analysis in the general number field sieve integer factorization. Theoretical Computer Science, 630 (??):76–94, May 30, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300172.

Sakai:2016:CDN

[818] Yusuke Sakai, Keita Emura, Jacob C. N. Schuldt, Goichiro Hanaoka, and Kazuo Ohta. Constructions of dynamic and non-dynamic threshold public-key encryption schemes with decryption consistency. Theoretical Computer Science, 630(??): 95–116, May 30, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300317.

Xu:2016:AAS

[819] Dachuan Xu, Fengmin Wang, Donglei Du, and Chenchen Wu. Approximation algorithms for submodular vertex cover problems with linear/submodular penalties using primal-dual technique. Theoretical Computer Science, 630(??): 117–125, May 30, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300330.

Anonymous:2016:EBx

[820] Anonymous. Editorial Board. Theoretical Computer Science, 631(??): iii–ix, June 6, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300858.

Hanikova:2016:TSF

[821] Zuzana Haniková and Petr Savický. Term satisfiability in FL_{ew}-algebras. Theoretical Computer Science, 631 (??):1–15, June 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001997.

Cassez:2016:CSN

[822] Franck Cassez, Ron van der Meyden, and Chenyi Zhang. The complexity of synchronous notions of information flow security. Theoretical Computer Science, 631(??):16– 42, June 6, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002048.

Ferrara:2016:GFH

[823] Pietro Ferrara. A generic framework for heap and value analyses of object-oriented programming languages. Theoretical Computer Science, 631(??):43–72, June 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300299.

AlAmeen:2016:CRP

[824] Mahmudul Faisal Al Ameen and Makoto Tatsuta. Completeness for recursive procedures in separation logic.

Theoretical Computer Science, 631 (??):73–96, June 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300329.

Kahl:2016:TAH

[825] Thomas Kahl. Topological abstraction of higher-dimensional automata. Theoretical Computer Science, 631 (??):97–117, June 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300573.

Bozzelli:2016:FBS

[826] Laura Bozzelli and César Sánchez. Foundations of Boolean stream runtime verification. Theoretical Computer Science, 631(??):118–138, June 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397516300585.

Forejt:2016:ERT

[827] Vojtech Forejt, Marta Kwiatkowska, Gethin Norman, and Ashutosh Trivedi. Expected reachability-time games. Theoretical Computer Science, 631(??): 139–160, June 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300603.

Anonymous:2016:EBy

[828] Anonymous. Editorial Board. Theoretical Computer Science, 632(??): iii–ix, June 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300986.

Kari:2016:Eb

[829] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 632(??):1, June 13, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301001.

Kwiatkowska:2016:P

[830] Marta Kwiatkowska, Andrew Phillips, and Chris Thachuk. Preface. Theoretical Computer Science, 632(??): 2–3, June 13, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301013.

Hagiya:2016:MCM

[831] Masami Hagiya, Nathanael Aubert-Kato, Shaoyu Wang, and Satoshi Kobayashi. Molecular computers for molecular robots as hybrid systems. Theoretical Computer Science, 632 (??):4–20, June 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009767.

Lakin:2016:MVC

[832] Matthew R. Lakin, Darko Stefanovic, and Andrew Phillips. Modular verification of chemical reaction network encodings via serializability analysis. Theoretical Computer Science, 632 (??):21–42, June 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005435.

Petersen:2016:SGS

[833] Rasmus L. Petersen, Matthew R. Lakin, and Andrew Phillips. A strand graph semantics for DNAbased computation. Theoretical Computer Science, 632(??):43–73, June 13, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515006866.

Anonymous:2016:EBz

[834] Anonymous. Editorial Board. Theoretical Computer Science, 633(??): iii–ix, June 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301293.

Kari:2016:Ec

[835] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 633(??):1, June 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301323.

Siegelmann:2016:P

[836] Hava T. Siegelmann. Preface. Theoretical Computer Science, 633(??): 2–3, June 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301311.

Alden:2016:MED

[837] Matthew Alden and Risto Miikkulainen. MARLEDA: Effective distribution estimation through Markov random fields. Theoretical Computer Science, 633(??):4– 18, June 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006945.

Cheney:2016:TEE

[838] Nick Cheney and Hod Lipson. Topological evolution for embodied cellular automata. Theoretical Computer Science, 633(??):19–27, June 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005344.

Harrington:2016:CBM

[839] Kyle I. S. Harrington. A circuit basis for morphogenesis. Theoretical Computer Science, 633(??):28– 36, June 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515006234.

Yoo:2016:MPN

[840] Yongseok Yoo, O. Ozan Koyluoglu, Sriram Vishwanath, and Ila Fiete. Multi-periodic neural coding for adaptive information transfer. Theoretical Computer Science, 633(??): 37–53, June 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001572.

Kozma:2016:PBC

[841] Robert Kozma and Marko Puljic. Pattern-based computing via sequential phase transitions in hierarchical mean field neuropercolation. Theoretical Computer Science, 633(??): 54–70, June 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007252.

Schindewolf:2016:CPN

[842] Craig Schindewolf, Dongwook Kim, Andrea Bel, and Horacio G. Rotstein. Complex patterns in networks of hyperexcitable neurons. Theoretical Computer Science, 633(??): 71–82, June 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005162.

Betti:2016:PLC

[843] Alessandro Betti and Marco Gori. The principle of least cognitive action. Theoretical Computer Science, 633(??):83–99, June 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515005526.

Rodriguez:2016:GMB

[844] A. Rodriguez and R. Granger. The grammar of mammalian brain capacity. Theoretical Computer Science, 633 (??):100–111, June 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516002231.

Anselmi:2016:ULI

[845] Fabio Anselmi, Joel Z. Leibo, Lorenzo Rosasco, Jim Mutch, Andrea Tacchetti, and Tomaso Poggio. Unsupervised learning of invariant representations. Theoretical Computer Science, 633(??):112–121, June 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515005587.

Anonymous:2016:EBba

[846] Anonymous. Editorial Board. Theoretical Computer Science, 634(??): iii–ix, June 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301505.

Michail:2016:TSP

[847] Othon Michail and Paul G. Spirakis. Traveling salesman problems in temporal graphs. Theoretical Computer Science, 634(??): 1–23, June 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300342.

Park:2016:PMM

[848] Jung-Heum Park. Paired manyto-many disjoint path covers in restricted hypercube-like graphs. Theoretical Computer Science, 634(??): 24–34, June 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300354.

Goldberg:2016:CCL

[849] Leslie Ann Goldberg and Mark Jerrum. The complexity of counting locally maximal satisfying assignments of Boolean CSPs. Theoretical Computer Science, 634(??): 35–46, June 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300366.

Chen:2016:LBL

[850] Wenbin Chen, Hao Lei, and Ke Qi. Lattice-based linearly homomorphic signatures in the standard model. Theoretical Computer Science, 634(??): 47–54, June 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300378.

Bhattacharya:2016:TLB

[851] Anup Bhattacharya, Ragesh Jaiswal, and Nir Ailon. Tight lower bound instances for k-means++ in two dimensions. Theoretical Computer Science, 634(??):55–66, June 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300512.

Mavronicolas:2016:CER

[852] Marios Mavronicolas and Burkhard Monien. The complexity of equilibria for risk-modeling valuations. Theoretical Computer Science, 634(??): 67–96, June 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300524.

Lin:2016:SCS

[853] Cheng-Kuan Lin, Lili Zhang, Jianxi Fan, and Dajin Wang. Structure connectivity and substructure connectivity of hypercubes. Theoretical Computer Science, 634(??):97– 107, June 27, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300536.

Kameda:2016:APE

[854] Tsunehiko Kameda, Ichiro Suzuki, and Masafumi Yamashita. An alternative proof for the equivalence of ∞-searcher and 2-searcher. Theoretical Computer Science, 634(??): 108–119, June 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630055X.

Khanteimouri:2016:EAC

[855] Payam Khanteimouri, Ali Mohades, Mohammad Ali Abam, and Mohammad Reza Kazemi. Efficiently approximating color-spanning balls. Theoretical Computer Science, 634(??): 120–126, June 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300615.

Anonymous:2016:EBbb

[856] Anonymous. Editorial Board. Theoretical Computer Science, 635(??): iii–ix, July 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301712.

Wang:2016:WRA

[857] Qichao Wang and Friedrich Otto. Weighted restarting automata and pushdown relations. Theoretical Computer Science, 635(??):1–15, July 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516301116.

Fici:2016:APR

[858] Gabriele Fici, Alessio Langiu, Thierry Lecroq, Arnaud Lefebvre, Filippo Mignosi, Jarkko Peltomäki, and Élise Prieur-Gaston. Abelian powers and repetitions in Sturmian words. Theoretical Computer Science, 635(??): 16–34, July 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301128.

Rute:2016:WDR

[859] Jason Rute. When does randomness come from randomness? Theoretical Computer Science, 635(??): 35–50, July 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301141.

Barnabei:2016:SPD

[860] Marilena Barnabei, Flavio Bonetti, Niccolò Castronuovo, and Robert Cori. Some permutations on Dyck words.

Theoretical Computer Science, 635 (??):51–63, July 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301207.

Hu:2016:SMP

[861] Xiaomin Hu, Yingzhi Tian, Xiaodong Liang, and Jixiang Meng. Strong matching preclusion for ndimensional torus networks. Theoretical Computer Science, 635(??): 64–73, July 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301219.

Dourado:2016:CRA

[862] Mitre Costa Dourado. Computing role assignments of split graphs. Theoretical Computer Science, 635 (??):74–84, July 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301244.

He:2016:CRP

[863] Yong He, Zhenhe Cui, and Zihan Yuan. Conjugacy relations of prefix codes. Theoretical Computer Science, 635(??):85–93, July 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301578.

Lewenstein:2016:DRO

[864] Moshe Lewenstein, J. Ian Munro, Yakov Nekrich, and Sharma V. Thankachan. Document retrieval with one wildcard. Theoretical Computer Science, 635(??):94– 101, July 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301608.

Anonymous:2016:EBbc

[865] Anonymous. Editorial Board. Theoretical Computer Science, 636(??): iii–ix, July 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301906.

Brandenburg:2016:RDI

[866] Franz J. Brandenburg, Walter Didimo, William S. Evans, Philipp Kindermann, Giuseppe Liotta, and Fabrizio Montecchiani. Recognizing and drawing IC-planar graphs. Theoretical Computer Science, 636(??):1– 16, July 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300652.

Dolev:2016:MCG

[867] Shlomi Dolev, Ephraim Korach, Ximing Li, Yin Li, and Galit Uzan. Magnifying computing gaps: Establishing encrypted communication over unidirectional channels. Theoretical Computer Science, 636(??): 17–26, July 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300718.

Serafino:2016:HFL

[868] Paolo Serafino and Carmine Ventre. Heterogeneous facility location without money. Theoretical Computer Science, 636(??):27–46, July 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300925.

DiGiacomo:2016:LUB

[869] Emilio Di Giacomo, Giuseppe Liotta, and Tamara Mchedlidze. Lower and upper bounds for long induced paths in 3-connected planar graphs. Theoretical Computer Science, 636 (??):47–55, July 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300937.

Kosolobov:2016:FLC

[870] Dmitry Kosolobov. Finding the leftmost critical factorization on unordered alphabet. Theoretical Computer Science, 636(??):56– 65, July 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301104.

Brandenburg:2016:RCS

[871] Franz J. Brandenburg and Andreas Gleißner. Ranking chain sum orders. Theoretical Computer Science, 636(??):66–76, July 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301621.

Groshaus:2016:TLB

[872] Marina Groshaus and Leandro Montero. Tight lower bounds on the number of bicliques in false-twin-free graphs. Theoretical Computer Science, 636(??):77–84, July 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301633.

Bang-Jensen:2016:FGPa

[873] J. Bang-Jensen and Frédéric Havet. Finding good 2-partitions of digraphs I. Hereditary properties. Theoretical Computer Science, 636(??):85– 94, July 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301657.

Prodinger:2016:BOE

[874] Helmut Prodinger. Batcher’s oddeven exchange revisited: a generating functions approach. Theoretical Computer Science, 636(??): 95–100, July 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301414.

Anonymous:2016:EBbd

[875] Anonymous. Editorial Board. Theoretical Computer Science, 637(??): iii–ix, July 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302031.

Chambart:2016:FAM

[876] Pierre Chambart, Alain Finkel, and Sylvain Schmitz. Forward analysis and model checking for trace bounded WSTS. Theoretical Computer Science, 637(??):1–29, July 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300597.

Courtault:2016:LSM

[877] Jean-René Courtault, Didier Galmiche, and David Pym. A logic of separating modalities. Theoretical Computer Science, 637(??):30– 58, July 18, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630113X.

Anonymous:2016:EBbe

[878] Anonymous. Editorial Board. Theoretical Computer Science, 638(??): iii–ix, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302122.

Benson:2016:SIH

[879] Gary Benson, Martin Farach-Colton, Moshe Lewenstein, and Ely Porat. Special issue in honor of the 60th birthday of Amihood Amir. Theoretical Computer Science, 638(??): 1–2, July 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302195.

Shiftan:2016:SIS

[880] Ariel Shiftan and Ely Porat. Set intersection and sequence matching with mismatch counting. Theoretical Computer Science, 638(??):3– 10, July 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000049.

Benson:2016:LRS

[881] G. Benson, A. Levy, S. Maimoni, D. Noifeld, and B. R. Shalom. LCSk: a refined similarity measure. Theoretical Computer Science, 638(??): 11–26, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010877.

Butman:2016:PSM

[882] Ayelet Butman, Noa Lewenstein, and J. Ian Munro. Permuted scaled matching. Theoretical Computer Science, 638 (??):27–32, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751600178X.

Brodal:2016:TDR

[883] Gerth Stølting Brodal, Pooya Davoodi, Moshe Lewenstein, Rajeev Raman, and Srinivasa Rao Satti. Two dimensional range minimum queries and Fibonacci lattices. Theoretical Computer Science, 638(??):33–43, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001389.

Kopelowitz:2016:PST

[884] Tsvi Kopelowitz. The property suffix tree with dynamic properties. Theoretical Computer Science, 638(??): 44–51, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001754.

Klein:2016:CMF

[885] Shmuel T. Klein and Dana Shapira. Compressed matching for feature vectors. Theoretical Computer Science, 638(??):52–62, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011822.

Beigel:2016:SDP

[886] Richard Beigel and William Gasarch. On the sizes of DPDAs, PDAs, LBAs. Theoretical Computer Science, 638 (??):63–75, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007756.

Apostolico:2016:SSM

[887] Alberto Apostolico, Concettina Guerra, Gad M. Landau, and Cinzia Pizzi. Sequence similarity measures based on bounded Hamming distance. Theoretical Computer Science, 638(??): 76–90, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000475.

Munro:2016:FCW

[888] J. Ian Munro, Yakov Nekrich, and Jeffrey S. Vitter. Fast construction of wavelet trees. Theoretical Computer Science, 638(??):91– 97, July 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010087.

Bille:2016:LCE

[889] Philip Bille, Paweł Gawrychowski, Inge Li Gørtz, Gad M. Landau, and Oren Weimann. Longest common extensions in trees. Theoretical Computer Science, 638(??): 98–107, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751500729X.

Navarro:2016:RCS

[890] Gonzalo Navarro and Sharma V. Thankachan. Reporting consecutive substring occurrences under bounded gap constraints. Theoretical Computer Science, 638(??): 108–111, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001158.

Babenko:2016:CMM

[891] Maxim Babenko, Paweł Gawrychowski, Tomasz Kociumaka, Ignat Kolesnichenko, and Tatiana Starikovskaya. Computing minimal and maximal suffixes of a substring. Theoretical Computer Science, 638(??):112– 121, July 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007707.

Crochemore:2016:OPI

[892] Maxime Crochemore, Costas S. Iliopoulos, Tomasz Kociumaka, Marcin Kubica, Alessio Langiu, Solon P. Pissis, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. Order-preserving indexing. Theoretical Computer Science, 638(??): 122–135, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515005605.

Gawrychowski:2016:OPP

[893] Paweł Gawrychowski and Przemysław Uznański. Order-preserving pattern matching with k mismatches. Theoretical Computer Science, 638(??): 136–144, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007690.

Kucherov:2016:ASM

[894] Gregory Kucherov, Kamil Salikhov, and Dekel Tsur. Approximate string matching using a bidirectional index. Theoretical Computer Science, 638(??): 145–158, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009494.

Na:2016:FIA

[895] Joong Chae Na, Hyunjoon Kim, Heejin Park, Thierry Lecroq, Martine Léonard, Laurent Mouchard, and Kunsoo Park. FM-index of alignment: a compressed index for similar strings. Theoretical Computer Science, 638(??): 159–170, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007288.

Crochemore:2016:LSS

[896] Maxime Crochemore, Chiara Epifanio, Roberto Grossi, and Filippo Mignosi. Linear-size suffix tries. Theoretical Computer Science, 638(??): 171–178, July 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300305.

Anonymous:2016:EBbf

[897] Anonymous. Editorial Board. Theoretical Computer Science, 639(??): iii–ix, August 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302626.

Brewster:2016:DTC

[898] Richard C. Brewster, Sean McGuinness, Benjamin Moore, and Jonathan A. Noel. A dichotomy theorem for circular colouring reconfiguration. Theoretical Computer Science, 639(??): 1–13, August 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301426.

Araujo:2016:POC

[899] Julio Araujo, Frédéric Havet, Claudia Linhares Sales, and Ana Silva. Proper orientation of cacti. Theoretical Computer Science, 639(??): 14–25, August 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301438.

Dehkordi:2016:CRA

[900] Hooman Reisi Dehkordi, Peter Eades, Seok-Hee Hong, and Quan Nguyen. Circular right-angle crossing drawings in linear time. Theoretical Computer Science, 639(??):26– 41, August 1, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630144X.

Lu:2016:PSC

[901] Yang Lu and Jiguo Li. Provably secure certificateless proxy signature scheme in the standard model. Theoretical Computer Science, 639(??): 42–59, August 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301554.

Fernandes:2016:KCP

[902] Cristina G. Fernandes and Marcio T. I. Oshiro. Kinetic clustering of points on the line. Theoretical Computer Science, 639(??): 60–71, August 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301566.

Coudert:2016:DCI

[903] David Coudert and Guillaume Ducoffe. Data center interconnection networks are not hyperbolic. Theoretical Computer Science, 639(??): 72–90, August 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630161X.

Riechert:2016:ADP

[904] Maik Riechert, Christian Höner zu Siederdissen, and Peter F. Stadler. Algebraic dynamic programming for multiple context-free grammars. Theoretical Computer Science, 639(??): 91–109, August 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301797.

Anonymous:2016:EBbg

[905] Anonymous. Editorial Board. Theoretical Computer Science, 640(??): iii–ix, August 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302857.

Bang-Jensen:2016:FGPb

[906] J. Bang-Jensen, Nathann Cohen, and Frédéric Havet. Finding good 2partitions of digraphs II. Enumerable properties. Theoretical Computer Science, 640(??):1–19, August 9, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301815.

Piotrow:2016:FMN

[907] Marek Piotrów. Faster merging networks with a small constant period. Theoretical Computer Science, 640(??): 20–40, August 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301852.

Blanchet-Sadri:2016:CAC

[908] F. Blanchet-Sadri, Daniel Seita, and David Wise. Computing abelian complexity of binary uniform morphic words. Theoretical Computer Science, 640(??):41–51, August 9, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630233X.

Dourado:2016:GHN

[909] Mitre C. Dourado, Lucia D. Penso, and Dieter Rautenbach. On the geodetic hull number of P_{k}-free graphs. Theoretical Computer Science, 640(??): 52–60, August 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302341.

Jiang:2016:EDP

[910] Minghui Jiang, Ge Xia, and Yong Zhang. Edge-disjoint packing of stars and cycles. Theoretical Computer Science, 640(??):61–69, August 9, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302377.

Golovach:2016:IDP

[911] Petr A. Golovach, Daniël Paulusma, and Erik Jan van Leeuwen. Induced disjoint paths in circular-arc graphs in linear time. Theoretical Computer Science, 640(??):70– 83, August 9, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302390.

Liao:2016:AAC

[912] Chung-Shou Liao, Chih-Wei Liang, and Sheung Hung Poon. Approximation algorithms on consistent dynamic map labeling. Theoretical Computer Science, 640(??): 84–93, August 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302420.

Bujtas:2016:NMG

[913] Cs. Bujtás, Gy. Dósa, Cs. Imreh, J. Nagy-György, and Zs. Tuza. New models of graph-bin packing. Theoretical Computer Science, 640(??): 94–103, August 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302444.

Lee:2016:AGA

[914] Hyung Tae Lee, San Ling, and Huaxiong Wang. Analysis of Gong et al.’s CCA2-secure homomorphic encryption. Theoretical Computer Science, 640(??):104–114, August 9, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630250X.

Zhao:2016:CCH

[915] Shuli Zhao, Weihua Yang, and Shurong Zhang. Component connectivity of hypercubes. Theoretical Computer Science, 640(??):115– 118, August 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301827.

Li:2016:ACP

[916] Shuangshuang Li, Zhao Zhang, and Hong-Jian Lai. Algorithm for constraint partial inverse matroid problem with weight increase forbidden. Theoretical Computer Science, 640(??): 119–124, August 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302389.

Song:2016:NNS

[917] Hong Jie Song, Wen Hua Pan, Xiang Nan Gong, and Chang Qing Xu. A note on the neighbor sum distinguishing total coloring of planar graphs. Theoretical Computer Science, 640(??): 125–129, August 9, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302432.

Anonymous:2016:EBbh

[918] Anonymous. Editorial Board. Theoretical Computer Science, 641(??): iii–ix, August 16, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303061.

Kari:2016:Ed

[919] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 641(??):1, August 16, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303139.

Meunier:2016:USE

[920] Pierre-Étienne Meunier. Unraveling simplicity in elementary cellular automata. Theoretical Computer Science, 641(??):2–10, August 16, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000050.

Gratie:2016:CCF

[921] Cristian Gratie and Ion Petre. Complete characterization for the fitpreserving data refinement of massaction reaction networks. Theoretical Computer Science, 641(??):11– 24, August 16, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300160.

Olarte:2016:PTV

[922] C. Olarte, D. Chiarugi, M. Falaschi, and D. Hermith. A proof theoretic view of spatial and temporal dependencies in biochemical systems. Theoretical Computer Science, 641(??): 25–42, August 16, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300184.

Song:2016:CPT

[923] Bosheng Song and Linqiang Pan. The computational power of tissuelike P systems with promoters. Theoretical Computer Science, 641(??): 43–52, August 16, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630158X.

Correia:2016:NIN

[924] Marisol B. Correia. New insights on neutral binary representations for evolutionary optimization. Theoretical Computer Science, 641(??): 53–74, August 16, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301803.

Tan:2016:IEL

[925] Jian-Ding Tan, Mahidzal Dahari, Siaw-Paw Koh, Ying-Ying Koay, and Issa-Ahmed Abed. An improved electromagnetism-like algorithm for numerical optimization. Theoretical Computer Science, 641(??):75– 84, August 16, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302328.

Zhang:2016:NSM

[926] Zhiqiang Zhang, Tingfang Wu, Andrei Paun, and Linqiang Pan. Numerical P systems with migrating variables. Theoretical Computer Science, 641(??): 85–108, August 16, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302407.

Anonymous:2016:EBbi

[927] Anonymous. Editorial Board. Theoretical Computer Science, 642(??): iii–ix, August 23, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303309.

Navarro-Lopez:2016:DHD

[928] Eva M. Navarro-López and Rebekah Carter. Deadness and how to disprove liveness in hybrid dynamical systems. Theoretical Computer Science, 642(??): 1–23, August 23, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302456.

Bujtor:2016:NMI

[929] Ferenc Bujtor, Sascha Fendrich, Gerald Lüttgen, and Walter Vogler. Nondeterministic modal interfaces. Theoretical Computer Science, 642(??): 24–53, August 23, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630247X.

Anonymous:2016:EBbj

[930] Anonymous. Editorial Board. Theoretical Computer Science, 643(??): iii–ix, August 30, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303425.

Jaskolka:2016:MCC

[931] Jason Jaskolka and Ridha Khedri. Mitigating covert channels based on analysis of the potential for communication. Theoretical Computer Science, 643(??): 1–37, August 30, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302675.

Rossi:2016:TLM

[932] Matteo Rossi, Dino Mandrioli, Angelo Morzenti, and Luca Ferrucci. A temporal logic for micro- and macrostep-based real-time systems: Foundations and applications. Theoretical Computer Science, 643(??):38– 64, August 30, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302985.

Anonymous:2016:EBbk

[933] Anonymous. Editorial Board. Theoretical Computer Science, 644(??): iii–ix, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303644.

vandenHerik:2016:RAC

[934] H. Jaap van den Herik, Walter A. Kosters, and Aske Plaat. Recent advances in computer games. Theoretical Computer Science, 644(??): 1, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303693.

Bonnet:2016:CCG

[935] Édouard Bonnet, Florian Jamain, and Abdallah Saffidine. On the complexity of connection games. Theoretical Computer Science, 644(??): 2–28, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630278X.

Browne:2016:AIS

[936] Cameron Browne. Algorithms for interactive Sprouts. Theoretical Computer Science, 644(??):29–42, September 6, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302778.

Cazenave:2016:PPA

[937] Tristan Cazenave. Playout policy adaptation with move features. Theoretical Computer Science, 644(??): 43–52, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302699.

Graf:2016:APO

[938] Tobias Graf and Marco Platzner. Adaptive playouts for online learning of policies during Monte Carlo Tree Search. Theoretical Computer Science, 644(??):53–62, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516302742.

Hsueh:2016:ASI

[939] Chu-Hsuan Hsueh, I-Chen Wu, Wen-Jie Tseng, Shi-Jim Yen, and Jr-Chang Chen. An analysis for strength improvement of an MCTS-based program playing Chinese dark chess. Theoretical Computer Science, 644(??): 63–75, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302705.

Janko:2016:PPC

[940] Vito Janko and Matej Guid. A program for Progressive chess. Theoretical Computer Science, 644(??): 76–91, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302730.

Liu:2016:MIU

[941] Yun-Ching Liu and Yoshimasa Tsuruoka. Modification of improved upper confidence bounds for regulating exploration in Monte-Carlo tree search. Theoretical Computer Science, 644(??): 92–105, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302791.

Lorentz:2016:UEF

[942] Richard Lorentz. Using evaluation functions in Monte-Carlo Tree Search. Theoretical Computer Science, 644(??): 106–113, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302717.

Mandai:2016:LAM

[943] Yusaku Mandai and Tomoyuki Kaneko. LinUCB applied to Monte Carlo tree search. Theoretical Computer Science, 644(??):114–126, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397516302808.

Pawlewicz:2016:CNS

[944] Jakub Pawlewicz and Ryan B. Hayward. Conspiracy number search with relative sibling scores. Theoretical Computer Science, 644(??):127– 142, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302729.

Uiterwijk:2016:PCS

[945] Jos W. H. M. Uiterwijk. Polymerization and crystallization of Snowflake molecules in Domineering. Theoretical Computer Science, 644(??):143– 158, September 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302754.

Anonymous:2016:EBbl

[946] Anonymous. Editorial Board. Theoretical Computer Science, 645(??): iii–ix, September 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304005.

Libert:2016:BRD

[947] Benoît Libert, Marc Joye, and Moti Yung. Born and raised distributively: Fully distributed noninteractive adaptively-secure threshold signatures with short shares. Theoretical Computer Science, 645(??):1– 24, September 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001626.

Bonamy:2016:KPF

[948] Marthe Bonamy and Łukasz Kowalik. A 13k-kernel for planar feedback vertex set via region decomposition. Theoretical Computer Science, 645(??): 25–40, September 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301785.

Halman:2016:DFP

[949] Nir Halman. A deterministic fully polynomial time approximation scheme for counting integer knapsack solutions made easy. Theoretical Computer Science, 645(??):41– 47, September 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302511.

Konstantinidis:2016:SMC

[950] G. Konstantinidis and Ath. Kehagias. Simultaneously moving cops and robbers. Theoretical Computer Science, 645(??):48–59, September 13, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630295X.

Kammer:2016:ATD

[951] Frank Kammer and Torsten Tholey. Approximate tree decompositions of planar graphs in linear time. Theoretical Computer Science, 645(??):60– 90, September 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302961.

Guo:2016:ECB

[952] Jia Guo and Mei Lu. The extra connectivity of bubble-sort star graphs. Theoretical Computer Science, 645(??): 91–99, September 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302997.

Evans:2016:SVR

[953] William S. Evans, Giuseppe Liotta, and Fabrizio Montecchiani. Simultaneous visibility representations of plane st-graphs using L-shapes. Theoretical Computer Science, 645(??):100– 111, September 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303012.

Bauer:2016:SSS

[954] Reinhard Bauer, Tobias Columbus, Ignaz Rutter, and Dorothea Wagner. Search-space size in contraction hierarchies. Theoretical Computer Science, 645(??):112–127, September 13, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303176.

Mazo:2016:NLE

[955] Loïc Mazo and Étienne Baudrier. Nonlocal estimators: a new class of multigrid convergent length estimators. Theoretical Computer Science, 645(??): 128–146, September 13, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303218.

Anonymous:2016:EBbm

[956] Anonymous. Editorial Board. Theoretical Computer Science, 646(??): iii–ix, September 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304133.

Kelk:2016:RRM

[957] Steven Kelk, Mareike Fischer, Vincent Moulton, and Taoyang Wu. Reduction rules for the maximum parsimony distance on phylogenetic trees. Theoretical Computer Science, 646(??): 1–15, September 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303243.

Beretta:2016:PTM

[958] Stefano Beretta, Mauro Castelli, and Riccardo Dondi. Parameterized tractability of the maximum-duo preservation string mapping problem. Theoretical Computer Science, 646(??): 16–25, September 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303255. See

corrigendum [1023].

Avni:2016:CSS

[959] Guy Avni and Tami Tamir. Costsharing scheduling games on restricted unrelated machines. Theoretical Computer Science, 646(??):26– 39, September 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303358.

Kacem:2016:SOS

[960] Imed Kacem and Hans Kellerer. Semionline scheduling on a single machine with unexpected breakdown. Theoretical Computer Science, 646(??):40– 48, September 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303371.

Ho:2016:TPS

[961] Nhan Bao Ho. Three-pile Sharing Nim and the quadratic time winning strategy. Theoretical Computer Science, 646(??):49–60, September 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303498.

Harland:2016:BBM

[962] James Harland. Busy beaver machines and the observant otter heuristic (or how to tame dreadful dragons). Theoretical Computer Science, 646(??):61– 85, September 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303504.

Glasser:2016:EAM

[963] Christian Glaßer, Heinz Schmitz, and Victor Selivanov. Efficient algorithms for membership in boolean hierarchies of regular languages. Theoretical Computer Science, 646(??):86– 108, September 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303516.

Anonymous:2016:EBbn

[964] Anonymous. Editorial Board. Theoretical Computer Science, 647(??): iii–ix, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304297.

Chen:2016:EME

[965] Yangjun Chen and Weixin Shen. An efficient method to evaluate intersections on big data sets. Theoretical Computer Science, 647(??):1– 21, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303528.

Cicalese:2016:TSP

[966] Ferdinando Cicalese, Balázs Keszegh, Bernard Lidický, Dömötör Pálvölgyi, and Tomás Valla. On the tree search problem with non-uniform costs. Theoretical Computer Science, 647(??):22– 32, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630353X.

Atminas:2016:SPC

[967] Aistis Atminas, Marcin Kamiński, and Jean-Florent Raymond. Scattered packings of cycles. Theoretical Computer Science, 647(??):33– 42, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303553.

deParga:2016:MCD

[968] Manuel Vázquez de Parga, Pedro García, and Damián López. Minimal consistent DFA revisited. Theoretical Computer Science, 647(??):43– 49, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303577.

Xiao:2016:NDF

[969] Lin Xiao. A new design formula exploited for accelerating Zhang neural network and its application to timevarying matrix inversion. Theoretical Computer Science, 647(??):50– 58, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303589.

Elmasry:2016:DRM

[970] Amr Elmasry, Meng He, J. Ian Munro, and Patrick K. Nicholson. Dynamic range majority data structures. Theoretical Computer Science, 647(??):59– 73, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630384X.

Belazzougui:2016:IQC

[971] Djamal Belazzougui, Roman Kolpakov, and Mathieu Raffinot. Indexing and querying color sets of images. Theoretical Computer Science, 647(??):74– 84, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303863.

Balkova:2016:APN

[972] L’ubomíra Balková, Michelangelo Bucci, Alessandro De Luca, Jiří Hladký, and Svetlana Puzynina. Aperiodic pseudorandom number generators based on infinite words. Theoretical Computer Science, 647(??):85– 100, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303875.

Vignatti:2016:MVC

[973] André L. Vignatti and Murilo V. G. da Silva. Minimum vertex cover in generalized random graphs with power law degree distribution. Theoretical Computer Science, 647(??):101– 111, September 27, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303899.

Anonymous:2016:EBbo

[974] Anonymous. Editorial Board. Theoretical Computer Science, 648(??): iii–ix, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304509.

Bresar:2016:CGD

[975] Bostjan Bresar, Paul Dorbec, Sandi Klavzar, Gasper Kosmrlj, and Gabriel Renault. Complexity of the game domination problem. Theoretical Computer Science, 648(??):1–7, October 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303590.

Boldi:2016:EOL

[976] Paolo Boldi and Sebastiano Vigna. Efficient optimally lazy algorithms for minimal-interval semantics. Theoretical Computer Science, 648(??): 8–25, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303814.

Le:2016:UAR

[977] Van Bang Le, Andrea Oversberg, and Oliver Schaudt. A unified approach to recognize squares of split graphs. Theoretical Computer Science, 648(??): 26–33, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303826.

Dutta:2016:LCG

[978] Tridib Dutta, Lenwood S. Heath, V. S. Anil Kumar, and Madhav V. Marathe. Labeled cuts in graphs. Theoretical Computer Science, 648(??): 34–39, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303851.

Zehavi:2016:PAA

[979] Meirav Zehavi. Parameterized approximation algorithms for packing problems. Theoretical Computer Science, 648(??):40–55, October 4, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303917.

Alvarez:2016:CG

[980] Carme Àlvarez, Maria J. Blesa, Amalia Duch, Arnau Messegué, and Maria Serna. Celebrity games. Theoretical Computer Science, 648(??): 56–71, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303929.

Zhang:2016:LCP

[981] Peng Zhang and Bin Fu. The label cut problem with respect to path length and label frequency. Theoretical Computer Science, 648(??): 72–83, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303930.

Kociumaka:2016:MND

[982] Tomasz Kociumaka, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. Maximum number of distinct and nonequivalent nonstandard squares in a word. Theoretical Computer Science, 648(??): 84–95, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304066.

Ferraioli:2016:DDF

[983] Diodato Ferraioli, Paul W. Goldberg, and Carmine Ventre. Decentralized dynamics for finite opinion games. Theoretical Computer Science, 648(??): 96–115, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304078.

Feigenblat:2016:CFI

[984] Guy Feigenblat, Ofra Itzhaki, and Ely Porat. Corrigendum to “The frequent items problem, under polynomial decay, in the streaming model” [Theoret. Comput. Sci. 411(34-36) (2010) 3048– 3054]. Theoretical Computer Science, 648(??):116, October 4, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630456X. See

[1].

Anonymous:2016:EBbp

[985] Anonymous. Editorial Board. Theoretical Computer Science, 649(??): iii–ix, October 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304613.

Jegourel:2016:CBI

[986] Cyrille Jegourel, Axel Legay, and Sean Sedwards. Command-based importance sampling for statistical model checking. Theoretical Computer Science, 649(??):1–24, October 11, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303966.

Ferrarotti:2016:NTC

[987] Flavio Ferrarotti, Klaus-Dieter Schewe, Loredana Tec, and Qing Wang. A new thesis concerning synchronised parallel computing — simplified parallel ASM thesis. Theoretical Computer Science, 649(??):25– 53, October 11, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304212.

Anonymous:2016:EBbq

[988] Anonymous. Editorial Board. Theoretical Computer Science, 650(??): iii–ix, October 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304728.

Auer:2016:GEF

[989] Peter Auer, Alexander Clark, and Thomas Zeugmann. Guest Editors’ foreword. Theoretical Computer Science, 650(??):1–3, October 18, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303723.

Kotzing:2016:MUC

[990] Timo Kötzing and Raphaela Palenta. A map of update constraints in inductive inference. Theoretical Computer Science, 650(??): 4–24, October 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303735.

Jain:2016:PLA

[991] Sanjay Jain and Efim Kinber. Parallel learning of automatic classes of languages. Theoretical Computer Science, 650(??):25–44, October 18, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303747.

Abasi:2016:LBH

[992] Hasan Abasi, Ali Z. Abdi, and Nader H. Bshouty. Learning boolean halfspaces with small weights from membership queries. Theoretical Computer Science, 650(??):45–56, October 18, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303759.

Angluin:2016:LRO

[993] Dana Angluin and Dana Fisman. Learning regular omega languages. Theoretical Computer Science, 650(??): 57–72, October 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303760.

Hutter:2016:ESA

[994] Marcus Hutter. Extreme state aggregation beyond Markov decision processes. Theoretical Computer Science, 650(??): 73–91, October 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303772.

Ailon:2016:BOO

[995] Nir Ailon, Kohei Hatano, and Eiji Takimoto. Bandit online optimization over the permutahedron. Theoretical Computer Science, 650(??): 92–108, October 18, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303784.

Maurer:2016:CRE

[996] Andreas Maurer. A chain rule for the expected suprema of Gaussian processes. Theoretical Computer Science, 650(??):109–122, October 18, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303796.

Anonymous:2016:EBbr

[997] Anonymous. Editorial Board. Theoretical Computer Science, 651(??): iii–ix, October 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304947.

Elkin:2016:SEP

[998] Michael Elkin, Ofer Neiman, and Christian Wulff-Nilsen. Spaceefficient path-reporting approximate distance oracles. Theoretical Computer Science, 651(??):1–10, October 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303838.

DArco:2016:SCC

[999] Paolo D’Arco and Roberto De Prisco. Secure computation without computers. Theoretical Computer Science, 651 (??):11–36, October 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303905.

Haddadan:2016:CDS

[1000] Arash Haddadan, Takehiro Ito, Amer E. Mouawad, Naomi Nishimura, Hirotaka Ono, Akira Suzuki, and Youcef Tebbal. The complexity of dominating set reconfiguration. Theoretical Computer Science, 651(??): 37–49, October 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304248.

Wehmuth:2016:MG

[1001] Klaus Wehmuth, Éric Fleury, and Artur Ziviani. On MultiAspect graphs. Theoretical Computer Science, 651(??): 50–61, October 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304352.

Hermelin:2016:PCC

[1002] Danny Hermelin, Moshe Kaspi, Christian Komusiewicz, and Barak Navon. Parameterized complexity of critical node cuts. Theoretical Computer Science, 651(??):62–75, October 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304364.

Deng:2016:ARC

[1003] Xingchao Deng, Hengzhe Li, and Guiying Yan. Algorithm on rainbow connection for maximal outerplanar graphs. Theoretical Computer Science, 651(??): 76–86, October 25, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304388.

Pai:2016:CIC

[1004] Kung-Jui Pai, Jou-Ming Chang, Jinn-Shyong Yang, and Ro-Yu Wu. Corrigendum to “Incidence coloring on hypercubes” [Theoret. Comput. Sci. 557 (2014) 59–65]. Theoretical Computer Science, 651(??):87–88, October 25, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303887. See

[6].

Anonymous:2016:EBbs

[1005] Anonymous. Editorial Board. Theoretical Computer Science, 652(??): iii–ix, November 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305254.

Cao:2016:SOS

[1006] Qian Cao and Zhaohui Liu. Semionline scheduling with bounded job sizes on two uniform machines. Theoretical Computer Science, 652(??): 1–17, November 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304406.

He:2016:STI

[1007] Kai He, Jian Weng, Robert H. Deng, and Joseph K. Liu. On the security of two identity-based conditional proxy re-encryption schemes. Theoretical Computer Science, 652(??): 18–27, November 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304443.

Pai:2016:CTC

[1008] Kung-Jui Pai and Jou-Ming Chang.

Constructing two completely independent spanning trees in hypercubevariant networks. Theoretical Computer Science, 652(??):28–37, November 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http://www.

sciencedirect.com/science/article/
pii/S0304397516304455.

Karthick:2016:SSE

[1009] T. Karthick. Structure of squares and efficient domination in graph classes. Theoretical Computer Science, 652(??): 38–46, November 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304662.

Bell:2016:DCP

[1010] Paul C. Bell, Shang Chen, and Lisa Jackson. On the decidability and complexity of problems for restricted hierarchical hybrid systems. Theoretical Computer Science, 652(??): 47–63, November 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304674.

Altman:2016:ICR

[1011] Harry Altman. Integer complexity: Representing numbers of bounded defect. Theoretical Computer Science, 652(??):64–85, November 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397516304789.

Bilo:2016:EAA

[1012] Davide Bilò, Luciano Gualà, Stefano Leucci, and Guido Proietti. Exact and approximate algorithms for movement problems on (special classes of) graphs. Theoretical Computer Science, 652(??): 86–101, November 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304807.

Berend:2016:SCR

[1013] Daniel Berend and Aryeh Kontorovich. The state complexity of random DFAs. Theoretical Computer Science, 652(??): 102–108, November 1, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304881.

Anonymous:2016:EBbt

[1014] Anonymous. Editorial Board. Theoretical Computer Science, 653(??): iii–ix, November 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305370.

Yero:2016:SNS

[1015] Ismael González Yero, Marko Jakovac, and Dorota Kuziak. The security number of strong grid-like graphs. Theoretical Computer Science, 653(??): 1–14, November 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304893.

Cechlarova:2016:SMT

[1016] Katarína Cechlárová, Tamás Fleiner, David F. Manlove, and Iain McBride. Stable matchings of teachers to schools. Theoretical Computer Science, 653(??): 15–25, November 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305011.

Sureson:2016:MLR

[1017] Claude Sureson. Π_{1}^{1}–Martin-Löf random reals as measures of natural open sets. Theoretical Computer Science, 653(??):26–41, November 15, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305059.

Damaschke:2016:DVR

[1018] Peter Damaschke. Deterministic versus randomized adaptive test cover. Theoretical Computer Science, 653(??):

42–52, November 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305060.

Dagdelen:2016:RLS

[1019] Özgür Dagdelen, Payman Mohassel, and Daniele Venturi. Rate-limited secure function evaluation. Theoretical Computer Science, 653(??):53– 78, November 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305072.

Li:2016:ECR

[1020] Xiang-Jun Li, Qi-Qi Dong, Zheng Yan, and Jun-Ming Xu. Embedded connectivity of recursive networks. Theoretical Computer Science, 653(??): 79–86, November 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305096.

Bourne:2016:SHS

[1021] Tom Bourne and Nik Ruskuc. On the star-height of subword counting languages and their relationship to Rees zero-matrix semigroups. Theoretical Computer Science, 653(??): 87–96, November 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305114.

Reis:2016:IRL

[1022] Rogério Reis and Emanuele Rodaro. Ideal regular languages and strongly connected synchronizing automata. Theoretical Computer Science, 653(??): 97–107, November 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305138.

Beretta:2016:CPT

[1023] Stefano Beretta, Mauro Castelli, and Riccardo Dondi. Corrigendum to “Parameterized tractability of the maximum-duo preservation string mapping problem” [Theoret. Comput. Sci. 646 (2016) 16–25]. Theoretical Computer Science, 653(??):108– 110, November 15, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305023. See

[958].

Anonymous:2016:EBbu

[1024] Anonymous. Editorial Board. Theoretical Computer Science, 654(??): iii–ix, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305588.

Xu:2016:ECC

[1025] Dachuan Xu, Donglei Du, and Ding-Zhu Du. Editorial for Computing and Combinatorics Conference. Theoretical Computer Science, 654(??): 1, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305643.

Chen:2016:CBS

[1026] Ruiwen Chen and Valentine Kabanets. Correlation bounds and #SAT algorithms for small linear-size circuits. Theoretical Computer Science, 654(??): 2–10, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301189.

Zhang:2016:MFC

[1027] Nan Zhang, Zhenhua Duan, and Cong Tian. A mechanism of function calls in MSVL. Theoretical Computer Science, 654(??):11– 25, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751600181X.

He:2016:NOM

[1028] Dayu He and Xin He. Nearly optimal monotone drawing of trees. Theoretical Computer Science, 654(??): 26–32, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000207.

Dang:2016:QCS

[1029] Zhe Dang, Thomas R. Fischer, William J. Hutton III, Oscar H. Ibarra, and Qin Li. Quantifying communication in synchronized languages. Theoretical Computer Science, 654(??): 33–44, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000864.

Yu:2016:IAA

[1030] Wei Yu and Zhaohui Liu. Improved approximation algorithms for some min-max and minimum cycle cover problems. Theoretical Computer Science, 654(??):45–58, November 22, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000852.

Massberg:2016:ERS

[1031] Jens Maßberg. Embedding rectilinear Steiner trees with length restrictions. Theoretical Computer Science, 654(??): 59–69, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302547.

Meesum:2016:RRA

[1032] S. M. Meesum, Pranabendu Misra, and Saket Saurabh. Reducing rank of the adjacency matrix by graph modification. Theoretical Computer Science, 654(??):70–79, November 22, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001420.

Jo:2016:SER

[1033] Seungbum Jo and Srinivasa Rao Satti. Simultaneous encodings for range and next/previous larger/smaller value queries. Theoretical Computer Science, 654(??):80–91, November 22, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000876.

Chakraborty:2016:UBF

[1034] Sourav Chakraborty, Raghav Kulkarni, Satyanarayana V. Lokam, and Nitin Saurabh. Upper bounds on Fourier entropy. Theoretical Computer Science, 654(??):92–112, November 22, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301190.

Su:2016:SSP

[1035] Shenghui Su, Shuwang Lü, Maozhi Xu, and Tao Xie. A semantically secure public key cryptoscheme using bit-pair shadows. Theoretical Computer Science, 654(??):113– 127, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300561.

Su:2016:PSN

[1036] Shenghui Su, Tao Xie, and Shuwang Lü. A provably secure non-iterative hash function resisting birthday attack. Theoretical Computer Science, 654(??): 128–142, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001547.

Balas:2016:NAR

[1037] Kevin Balas and Csaba D. Tóth. On the number of anchored rectangle packings for a planar point set. Theoretical Computer Science, 654(??):143– 154, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001973.

Uno:2016:MPS

[1038] Takeaki Uno and Yushi Uno. Mining preserving structures in a graph sequence. Theoretical Computer Science, 654(??):155–163, November 22, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011469.

Wang:2016:AAC

[1039] Yin Wang and Yinfeng Xu. The approximation algorithms for a class of multiple-choice problem. Theoretical Computer Science, 654(??):164– 174, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302948.

Nip:2016:SSC

[1040] Kameng Nip, Zhenbo Wang, and Wenxun Xing. A study on several combination problems of classic shop scheduling and shortest path. Theoretical Computer Science, 654(??):175– 187, November 22, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011883.

Ye:2016:ACT

[1041] Deshi Ye and Guochuan Zhang. Approximate composable truthful mechanism design. Theoretical Computer Science, 654(??):188–198, November 22, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001560.

Anonymous:2016:EBbv

[1042] Anonymous. Editorial Board. Theoretical Computer Science, 655 (part A) (??):iii–ix, December 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305746.

Fomin:2016:FSI

[1043] Fedor V. Fomin, Pierre Fraigniaud, Nicolas Nisse, and Dimitrios M. Thilikos. Forewords: Special issue on Theory and Applications of Graph Searching Problems. Theoretical Computer Science, 655 (part A) (??):1, December 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305813.

Bonato:2016:PVG

[1044] Anthony Bonato, Dieter Mitsche, Xavier Pérez-Giménez, and Paweł Pralat. A probabilistic version of the game of Zombies and Survivors on graphs. Theoretical Computer Science, 655 (part A)(??):2– 14, December 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011731.

Foerster:2016:LUC

[1045] Klaus-Tycho Foerster and Roger Wattenhofer. Lower and upper competitive bounds for online directed graph exploration. Theoretical Computer Science, 655 (part A)(??):15– 29, December 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515010166.

Lehner:2016:PEI

[1046] Florian Lehner. Pursuit evasion on infinite graphs. Theoretical Computer Science, 655 (part A)(??):30– 40, December 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300639.

Luccio:2016:MAM

[1047] Fabrizio Luccio and Linda Pagli.

More agents may decrease global work: a case in butterfly decontamination. Theoretical Computer Science, 655 (part A)(??):41–57, December 6, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305151.

Puchala:2016:JRD

[1048] Bernd Puchala and Roman Rabinovich.

Jumping robbers in digraphs. Theoretical Computer Science, 655 (part A) (??):58–77, December 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305035.

Amiri:2016:DWP

[1049] Saeed Akhoondian Amiri, Stephan Kreutzer, and Roman Rabinovich. DAG-width is PSPACE-complete. Theoretical Computer Science, 655 (part A)(??):78–89, December 6, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439751630487X.

Anonymous:2016:EBbw

[1050] Anonymous. Editorial Board. Theoretical Computer Science, 656 (part A)(??):iii–ix, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397516305862.

Carotenuto:2016:OMS

[1051] Dario Carotenuto, Aniello Murano, and Adriano Peron. Ordered multi-stack visibly pushdown automata. Theoretical Computer Science, 656 (part A)(??):1–26, December 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304200.

Fu:2016:DES

[1052] Xiang Fu. On detecting environment sensitivity using slicing. Theoretical Computer Science, 656 (part A)(??): 27–45, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304686.

Diaconescu:2016:FSF

[1053] Razvan Diaconescu. Functorial semantics of first-order views. Theoretical Computer Science, 656 (part A)(??): 46–59, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304844.

Basu:2016:DSA

[1054] Samik Basu and Tevfik Bultan.

On deciding synchronizability for asynchronously communicating systems. Theoretical Computer Science, 656 (part A)(??):60–75, December 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305102.

Lellmann:2016:HRR

[1055] Björn Lellmann. Hypersequent rules with restricted contexts for propositional modal logics. Theoretical Computer Science, 656 (part A)(??):76– 105, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305187.

Anonymous:2016:EBbx

[1056] Anonymous. Editorial Board. Theoretical Computer Science, 656 (Part B)(??):iii–ix, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397516306296.

Iliopoulos:2016:F

[1057] Costas S. Iliopoulos and Simon J. Pugisi. Foreword. Theoretical Computer Science, 656 (Part B)(??): 107, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306880.

Heliou:2016:EDR

[1058] A. Heliou, M. Léonard, L. Mouchard, and M. Salson. Efficient dynamic range minimum query. Theoretical Computer Science, 656 (Part B)(??):108– 117, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303164.

Daykin:2016:BBO

[1059] Jacqueline W. Daykin, Richard Groult, Yannick Guesnet, Thierry Lecroq, Arnaud Lefebvre, Martine Léonard, and Élise Prieur-Gaston. Binary block order Rouen Transform. Theoretical Computer Science, 656 (Part B)(??): 118–134, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301645.

Cordova:2016:SEF

[1060] Joshimar Cordova and Gonzalo Navarro. Simple and efficient fully-functional succinct trees. Theoretical Computer Science, 656 (Part B)(??):135– 145, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300706.

Amir:2016:AJI

[1061] Amihood Amir, Alberto Apostolico, Tirza Hirst, Gad M. Landau, Noa Lewenstein, and Liat Rozenberg. Algorithms for Jumbled Indexing, Jumbled Border and Jumbled Square on run-length encoded strings. Theoretical Computer Science, 656 (Part B)(??): 146–159, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630069X.

Barton:2016:LTC

[1062] Carl Barton, Chang Liu, and Solon P.

Pissis. Linear-time computation of prefix table for weighted strings & applications. Theoretical Computer Science, 656 (Part B)(??):160– 172, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300688.

Kim:2016:SEA

[1063] Youngho Kim, Joong Chae Na, Heejin Park, and Jeong Seop Sim. A spaceefficient alphabet-independent Four-Russians’ lookup table and a multithreaded Four-Russians’ edit distance algorithm. Theoretical Computer Science, 656 (Part B)(??):173– 179, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300676.

Rytter:2016:TFC

[1064] Wojciech Rytter. Two fast constructions of compact representations of binary words with given set of periods. Theoretical Computer Science, 656 (Part B)(??):180–187, December 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300664.

Zohora:2016:EAD

[1065] Fatema Tuz Zohora and M. Sohel Rahman. An efficient algorithm to detect common ancestor genes for non-overlapping inversion and applications. Theoretical Computer Science, 656 (Part B)(??):188– 214, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516001961.

I:2016:FLF

[1066] Tomohiro I., Yuto Nakashima, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Faster Lyndon factorization algorithms for SLP and LZ78 compressed text. Theoretical Computer Science, 656 (Part B)(??):215– 224, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751600195X.

Matsuoka:2016:GPM

[1067] Yoshiaki Matsuoka, Takahiro Aoki, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Generalized pattern matching and periodicity under substring consistent equivalence relations. Theoretical Computer Science, 656 (Part B)(??):225– 233, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001390.

Crochemore:2016:DLR

[1068] Maxime Crochemore and Robert Mercas. On the density of Lyndon roots in factors. Theoretical Computer Science, 656 (Part B)(??):234– 240, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001377.

Badkobeh:2016:ECM

[1069] Golnaz Badkobeh, Maxime Crochemore, Manal Mohamed, and Chalita Toopsuwan. Efficient computation of maximal anti-exponent in palindrome-free strings. Theoretical Computer Science, 656 (Part B)(??):241–248, December 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516001365.

Simpson:2016:APL

[1070] Jamie Simpson. An abelian periodicity lemma. Theoretical Computer Science, 656 (Part B)(??):249– 255, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011755.

Fici:2016:FCA

[1071] Gabriele Fici, Tomasz Kociumaka, Thierry Lecroq, Arnaud Lefebvre, and Élise Prieur-Gaston. Fast computation of abelian runs. Theoretical Computer Science, 656 (Part B)(??):256– 264, December 20, 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011718.

Karkkainen:2016:TBS

[1072] Juha Kärkkäinen, Dominik Kempa, and Marcin Pia̧tkowski. Tighter bounds for the sum of irreducible LCP values. Theoretical Computer Science, 656 (Part B)(??):265–278, December 20, 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011706.

Anonymous:2016:EBby

[1073] Anonymous. Editorial Board. Theoretical Computer Science, 657 (Part B)(??):iii–ix, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306454.

Mohan:2016:CFA

[1074] Usha Mohan, Sivaramakrishnan Ramani, and Sounaka Mishra. Constant factor approximation algorithm for TSP satisfying a biased triangle inequality. Theoretical Computer Science, 657 (Part B)(??):111– 126, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305047.

Anselmo:2016:NEN

[1075] Marcella Anselmo, Dora Giammarresi, and Maria Madonia. Non-expandable non-overlapping sets of pictures. Theoretical Computer Science, 657 (Part B)(??):127–136, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305126.

Meira:2016:CTC

[1076] Luis A. A. Meira, Flávio K. Miyazawa, and Lehilton L. C. Pedrosa. Clustering through Continuous Facility Location Problems. Theoretical Computer Science, 657 (Part B)(??):137– 145, ???? 2016. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630514X.

Du:2016:DAF

[1077] Chen Fei Du, Hamoon Mousavi, Eric Rowland, Luke Schaeffer, and Jeffrey Shallit. Decision algorithms for Fibonacci-automatic words, II: Related sequences and avoidability. Theoretical Computer Science, 657 (Part B) (??):146–162, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305199.

Guo:2016:CDR

[1078] Jia Guo and Mei Lu. Conditional diagnosability of the round matching composition networks. Theoretical Computer Science, 657 (Part B) (??):163–172, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630531X.

Kawahara:2016:BBO

[1079] Jun Kawahara, Koji M. Kobayashi, and Shuichi Miyazaki. Better bounds for online k-frame throughput maximization in network switches. Theoretical Computer Science, 657 (Part B)(??):173–190, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305321.

Zhang:2016:MLS

[1080] Haixia Zhang and Yi Wang. The minimal Laplacian spectral radius of trees with diameter 4. Theoretical Computer Science, 657 (Part B) (??):191–195, ???? 2016. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305163.

Anonymous:2017:EBa

[1081] Anonymous. Editorial Board. Theoretical Computer Science, 657 (part A) (??):iii–ix, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306375.

Wang:2017:P

[1082] Jianxin Wang and Chee K. Yap.

Preface. Theoretical Computer Science, 657 (part A)(??):1–2, January 2, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306867.

Furer:2017:ECC

[1083] Martin Fürer. Efficient computation of the characteristic polynomial of a threshold graph. Theoretical Computer Science, 657 (part A)(??): 3–10, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630336X.

Lin:2017:PTS

[1084] Mugang Lin, Qilong Feng, Jianer Chen, and Wenjun Li. Partition on trees with supply and demand: Kernelization and algorithms. Theoretical Computer Science, 657 (part A)(??): 11–19, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303000.

Li:2017:IKR

[1085] Wenjun Li, Qilong Feng, Jianer Chen, and Shuai Hu. Improved kernel results for some FPT problems based on simple observations. Theoretical Computer Science, 657 (part A)(??):

20–27, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302481.

Wang:2017:TAA

[1086] Jinyan Wang, Minghao Yin, and Jingli Wu. Two approximate algorithms for model counting. Theoretical Computer Science, 657 (part A) (??):28–37, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302316.

Liu:2017:FQA

[1087] Yang Liu and Shengyu Zhang. Fast quantum algorithms for least squares regression and statistic leverage scores. Theoretical Computer Science, 657 (part A)(??):38–47, January 2, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302304.

Cheng:2017:IAI

[1088] Jie Cheng, Daming Zhu, and Binhai Zhu. Improved algorithms for intermediate dataset storage in a cloud-based dataflow. Theoretical Computer Science, 657 (part A)(??): 48–53, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302286.

Xian:2017:AMN

[1089] Aiyong Xian, Kaiyuan Zhu, Daming Zhu, Lianrong Pu, and Hong Liu. Approximating Max NAE-k-SAT by anonymous local search. Theoretical Computer Science, 657 (part A) (??):54–63, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302262.

Dong:2017:FPT

[1090] Jianming Dong, Weitian Tong, Taibo Luo, Xueshi Wang, Jueliang Hu, Yinfeng Xu, and Guohui Lin. An FP-TAS for the parallel two-stage flowshop problem. Theoretical Computer Science, 657 (part A)(??):64– 72, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302250. See

corrigendum [1354].

Ivanyos:2017:SSD

[1091] Gábor Ivanyos and Miklos Santha.

Solving systems of diagonal polynomial equations over finite fields. Theoretical Computer Science, 657 (part A) (??):73–85, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302249.

Xiao:2017:EAM

[1092] Mingyu Xiao and Shaowei Kou. Exact algorithms for the maximum dissociation set and minimum 3-path vertex cover problems. Theoretical Computer Science, 657 (part A)(??): 86–97, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302225.

Zhang:2017:PSG

[1093] Peng Zhang and Yong Gao. A probabilistic study of generalized solution concepts in satisfiability testing and constraint programming. Theoretical Computer Science, 657 (part A)(??): 98–110, January 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302201.

Anonymous:2017:EBb

[1094] Anonymous. Editorial Board. Theoretical Computer Science, 658 (part A) (??):iii–ix, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306612.

Giammarresi:2017:P

[1095] Dora Giammarresi, Sabrina Mantaci, Marinella Sciortino, and Filippo Mignosi. Preface. Theoretical Computer Science, 658 (part A)(??):1– 3, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306958.

Anselmo:2017:TDC

[1096] Marcella Anselmo and Maria Madonia. Two-dimensional comma-free and cylindric codes. Theoretical Computer Science, 658 (part A)(??): 4–17, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302523.

Beal:2017:SNS

[1097] Marie-Pierre Béal and Pavel Heller.

Shifts of k-nested sequences. Theoretical Computer Science, 658 (part A) (??):18–26, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302973.

Badkobeh:2017:CME

[1098] Golnaz Badkobeh, Maxime Crochemore, and Robert Mercas. Counting maximal-exponent factors in words. Theoretical Computer Science, 658 (part A)(??):27–35, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516001778.

Barcucci:2017:NOM

[1099] Elena Barcucci, Antonio Bernini, Stefano Bilotta, and Renzo Pinzani. Nonoverlapping matrices. Theoretical Computer Science, 658 (part A)(??): 36–45, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301220.

Bersani:2017:LCT

[1100] Marcello M. Bersani, Matteo Rossi, and Pierluigi San Pietro. A logical characterization of timed regular languages. Theoretical Computer Science, 658 (part A)(??):46– 59, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303541.

Bonizzoni:2017:CGA

[1101] Paola Bonizzoni, Anna Paola Carrieri, Gianluca Della Vedova, Raffaella Rizzi, and Gabriella Trucco. A colored graph approach to perfect phylogeny with persistent characters. Theoretical Computer Science, 658 (part A) (??):60–73, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304236.

Castiglione:2017:CLH

[1102] Giusi Castiglione and Paolo Massazza. On a class of languages with holonomic generating functions. Theoretical Computer Science, 658 (part A) (??):74–84, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303565.

Choffrut:2017:SWD

[1103] Christian Choffrut. Sequences of words defined by two-way transducers. Theoretical Computer Science, 658 (part A) (??):85–96, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301177.

Cori:2017:SB

[1104] R. Cori and C. Reutenauer. On Sillke’s bijection. Theoretical Computer Science, 658 (part A)(??):97– 104, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300548.

Reghizzi:2017:TTI

[1105] Stefano Crespi Reghizzi, Violetta Lonati, Dino Mandrioli, and Matteo Pradella. Toward a theory of input-driven locally parsable languages. Theoretical Computer Science, 658 (part A)(??):105–121, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301165.

DAniello:2017:CSW

[1106] Alma D’Aniello, Aldo de Luca, and Alessandro De Luca. On Christoffel and standard words and their derivatives. Theoretical Computer Science, 658 (part A)(??):122–147, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301232.

DeFelice:2017:USC

[1107] Clelia De Felice, Rocco Zaccagnino, and Rosalba Zizza. Unavoidable sets and circular splicing languages. Theoretical Computer Science, 658 (part A)(??):148–158, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304819.

Dolce:2017:NTS

[1108] Francesco Dolce and Dominique Perrin. Neutral and tree sets of arbitrary characteristic. Theoretical Computer Science, 658 (part A)(??): 159–174, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000190.

Guaiana:2017:LTW

[1109] Giovanna Guaiana. Local testability from words to traces, a suitable definition. Theoretical Computer Science, 658 (part A)(??):175– 189, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302687.

Karhumaki:2017:CAE

[1110] Juhani Karhumäki, Svetlana Puzynina, Michaël Rao, and Markus A. Whiteland. On cardinalities of kabelian equivalence classes. Theoretical Computer Science, 658 (part A)(??): 190–204, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302468.

Kleijn:2017:AR

[1111] Jetty Kleijn, Maciej Koutny, Marta Pietkiewicz-Koutny, and Grzegorz Rozenberg. Applying regions. Theoretical Computer Science, 658 (part A)(??):205–215, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516000840.

Marsault:2017:SRL

[1112] Victor Marsault and Jacques Sakarovitch.

The signature of rational languages. Theoretical Computer Science, 658 (part A)(??):216–234, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516300627.

Moreira:2017:OSR

[1113] Nelma Moreira, Giovanni Pighizzini, and Rogério Reis. Optimal state reductions of automata with partially specified behaviors. Theoretical Computer Science, 658 (part A)(??):235– 245, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301153.

Pin:2017:UCF

[1114] Jean-Éric Pin and Pedro V. Silva.

On uniformly continuous functions for some profinite topologies. Theoretical Computer Science, 658 (part A)(??): 246–262, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302493.

Anonymous:2017:EBc

[1115] Anonymous. Editorial Board. Theoretical Computer Science, 658 (Part B) (??):iii–ix, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306697.

Anonymous:2017:EBd

[1116] Anonymous. Editorial Board. Theoretical Computer Science, 658 (part B) (??):iii–ix, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306697.

Adaricheva:2017:F

[1117] Kira Adaricheva, Giuseppe F. Italiano, Hans Kleine Büning, and György Turán. Foreword. Theoretical Computer Science, 658 (Part B)(??): 263, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306934.

Wild:2017:JIA

[1118] Marcel Wild. The joy of implications, aka pure Horn formulas: Mainly a survey. Theoretical Computer Science, 658 (Part B)(??):264– 292, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516002206.

Ausiello:2017:DHI

[1119] Giorgio Ausiello and Luigi Laura. Directed hypergraphs: Introduction and fundamental algorithms — a survey. Theoretical Computer Science, 658 (Part B)(??):293–306, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002097.

Adaricheva:2017:DBB

[1120] K. Adaricheva and J. B. Nation.

Discovery of the D-basis in binary tables based on hypergraph dualization. Theoretical Computer Science, 658 (Part B)(??):307–315, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011172.

Babin:2017:DLG

[1121] Mikhail A. Babin and Sergei O.

Kuznetsov. Dualization in lattices given by ordered sets of irreducibles. Theoretical Computer Science, 658 (Part B)(??):316–326, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516000165.

Rudolph:2017:STC

[1122] Sebastian Rudolph. Succinctness and tractability of closure operator representations. Theoretical Computer Science, 658 (Part B)(??):327– 345, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011986.

Arias:2017:LDH

[1123] Marta Arias, José L. Balcázar, and Cristina Tîrnauca. Learning definite Horn formulas from closure queries. Theoretical Computer Science, 658 (Part B)(??):346–356, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011809.

Chardin:2017:RQL

[1124] Brice Chardin, Emmanuel Coquery, Marie Pailloux, and Jean-Marc Petit. RQL: a query language for rule discovery in databases. Theoretical Computer Science, 658 (Part B)(??): 357–374, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305941.

Obiedkov:2017:PCP

[1125] Sergei Obiedkov. Parameterized ceteris paribus preferences over atomic conjunctions under conservative semantics. Theoretical Computer Science, 658 (Part B)(??):375–390, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000591.

Beaudou:2017:AMS

[1126] Laurent Beaudou, Arnaud Mary, and Lhouari Nourine. Algorithms for kmeet-semidistributive lattices. Theoretical Computer Science, 658 (Part B)(??):391–398, January 7, 2017.

CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397515009354.

Kucera:2017:HCG

[1127] Petr Kucera. Hydras: Complexity on general graphs and a subclass of trees. Theoretical Computer Science, 658 (Part B)(??):399– 416, January 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301840.

Sloan:2017:HDH

[1128] Robert H. Sloan, Despina Stasi, and György Turán. Hydras: Directed hypergraphs and Horn formulas. Theoretical Computer Science, 658 (Part B)(??):417–428, January 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301839.

Anonymous:2017:EBe

[1129] Anonymous. Editorial Board. Theoretical Computer Science, 659(??): iii–ix, January 10, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307241.

Burcsi:2017:PNW

[1130] Péter Burcsi, Gabriele Fici, Zsuzsanna Lipták, Frank Ruskey, and Joe Sawada. On prefix normal words and prefix normal forms. Theoretical Computer Science, 659(??):

1–13, January 10, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305680.

Liu:2017:MPS

[1131] Wen An Liu and Ming Yue Wang. Multi-player subtraction games. Theoretical Computer Science, 659(??): 14–35, January 10, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305692.

Bacquey:2017:LET

[1132] Nicolas Bacquey. Leader election on two-dimensional periodic cellular automata. Theoretical Computer Science, 659(??):36–52, January 10, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305990.

Xu:2017:GND

[1133] Xiang Xu, Xiaowang Li, Shuming Zhou, Rong-Xia Hao, and Mei-Mei Gu. The g-good-neighbor diagnosability of (n, k)-star graphs. Theoretical Computer Science, 659(??): 53–63, January 10, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306028.

Kellerer:2017:AIU

[1134] Hans Kellerer, Rebecca Sarto Basso, and Vitaly A. Strusevich. Approximability issues for unconstrained and constrained maximization of halfproduct related functions. Theoretical Computer Science, 659(??):64– 71, January 10, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306041.

Xiao:2017:CKB

[1135] Mingyu Xiao and Hiroshi Nagamochi.

Complexity and kernels for bipartition into degree-bounded induced graphs. Theoretical Computer Science, 659(??): 72–82, January 10, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306065.

Kintali:2017:DWP

[1136] Shiva Kintali. Directed width parameters and circumference of digraphs. Theoretical Computer Science, 659(??): 83–87, January 10, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305436.

Karpinski:2017:VCM

[1137] Michal Karpiński. Vertex 2-coloring without monochromatic cycles of fixed size is NP-complete. Theoretical Computer Science, 659(??): 88–94, January 10, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305448.

Kamiyama:2017:NSV

[1138] Naoyuki Kamiyama. A note on the submodular vertex cover problem with submodular penalties. Theoretical Computer Science, 659(??): 95–97, January 10, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305928.

Anonymous:2017:EBf

[1139] Anonymous. Editorial Board. Theoretical Computer Science, 660(??): iii–ix, January 17, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307344.

Rusu:2017:LLT

[1140] Irena Rusu. log-Lists and their applications to sorting by transpositions, reversals and block-interchanges. Theoretical Computer Science, 660(??): 1–15, January 17, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306077.

Barmpalias:2017:CHP

[1141] George Barmpalias and Andrew Lewis-Pye. Computing halting probabilities from other halting probabilities. Theoretical Computer Science, 660(??): 16–22, January 17, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306089.

Faonio:2017:FLR

[1142] Antonio Faonio, Jesper Buus Nielsen, and Daniele Venturi. Fully leakageresilient signatures revisited: Graceful degradation, noisy leakage, and construction in the bounded-retrieval model. Theoretical Computer Science, 660(??):23–56, January 17, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516306119.

Akutsu:2017:PCA

[1143] Tatsuya Akutsu, Jesper Jansson, Atsuhiro Takasu, and Takeyuki Tamura. On the parameterized complexity of associative and commutative unification. Theoretical Computer Science, 660(??): 57–74, January 17, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306892.

Guo:2017:NLF

[1144] Siyao Guo and Ilan Komargodski.

Negation-limited formulas. Theoretical Computer Science, 660(??):75– 85, January 17, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306909.

Cheng:2017:NCA

[1145] Maggie X. Cheng, Yi Ling, and Brian M. Sadler. Network connectivity assessment and improvement through relay node deployment. Theoretical Computer Science, 660(??): 86–101, January 17, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306922.

Shitov:2017:CFZ

[1146] Yaroslav Shitov. On the complexity of failed zero forcing. Theoretical Computer Science, 660(??):102– 104, January 17, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306971.

Anonymous:2017:EBg

[1147] Anonymous. Editorial Board. Theoretical Computer Science, 661(??): iii–ix, January 24, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300178.

Wang:2017:NSD

[1148] Jihui Wang, Jiansheng Cai, and Baojian Qiu. Neighbor sum distinguishing total choosability of planar graphs without adjacent triangles. Theoretical Computer Science, 661(??): 1–7, January 24, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630593X.

Correa:2017:LTC

[1149] Ricardo C. Corrêa and Pablo M. S. Farias. Linear time computation of the maximal linear and circular sums of multiple independent insertions into a sequence. Theoretical Computer Science, 661(??):

8–17, January 24, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305953.

Birks:2017:TAO

[1150] Martin Birks and Stanley P. Y. Fung.

Temperature aware online algorithms for minimizing flow time. Theoretical Computer Science, 661(??):18– 34, January 24, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306004.

Aubrun:2017:NES

[1151] Nathalie Aubrun, Sebastián Barbieri, and Mathieu Sablik. A notion of effectiveness for subshifts on finitely generated groups. Theoretical Computer Science, 661(??):35– 55, January 24, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306983.

Ashok:2017:PCS

[1152] Pradeesha Ashok, Sudeshna Kolay, S. M. Meesum, and Saket Saurabh. Parameterized complexity of Strip Packing and Minimum Volume Packing. Theoretical Computer Science, 661(??): 56–64, January 24, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307022.

Khodakarami:2017:TGP

[1153] Farnoosh Khodakarami, Farzad Didehvar, and Ali Mohades. 1.5D terrain guarding problem parameterized by guard range. Theoretical Computer Science, 661(??):65– 69, January 24, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306107.

Anonymous:2017:EBh

[1154] Anonymous. Editorial Board. Theoretical Computer Science, 662(??): iii–ix, February 1, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300270.

Czygrinow:2017:IDL

[1155] A. Czygrinow, M. Hanćkowiak, E. Szymańska, W. Wawrzyniak, and M. Witkowski. Improved distributed local approximation algorithm for minimum 2dominating set in planar graphs. Theoretical Computer Science, 662(??): 1–8, February 1, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307071.

Bilo:2017:ARM

[1156] Vittorio Bilò, Michele Flammini, and Gianpiero Monaco. Approximating the revenue maximization problem with sharp demands. Theoretical Computer Science, 662(??): 9–30, February 1, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307083.

Montejano:2017:CCR

[1157] Luis Pedro Montejano and Ignasi Sau.

On the complexity of computing the krestricted edge-connectivity of a graph. Theoretical Computer Science, 662(??): 31–39, February 1, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307125.

Kintali:2017:FDM

[1158] Shiva Kintali and Qiuyi Zhang. Forbidden directed minors and Kelly-width. Theoretical Computer Science, 662(??): 40–47, February 1, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307149.

Hwang:2017:SSE

[1159] Jung Yeon Hwang, Boyeon Song, Daeseon Choi, Seung-Hun Jin, Hyun Sook Cho, and Mun-Kyu Lee. Simplified small exponent test for batch verification. Theoretical Computer Science, 662(??):48–58, February 1, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516307459.

Lenhart:2017:PEP

[1160] William J. Lenhart, Giuseppe Liotta, and Fabrizio Montecchiani. On partitioning the edges of 1-plane graphs. Theoretical Computer Science, 662(??): 59–65, February 1, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307101.

Anonymous:2017:EBi

[1161] Anonymous. Editorial Board. Theoretical Computer Science, 663(??): iii–ix, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300452.

Fiadeiro:2017:HAN

[1162] José L. Fiadeiro and Antónia Lopes. Heterogeneous and asynchronous networks of timed systems. Theoretical Computer Science, 663(??): 1–33, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307423.

Seldin:2017:SRC

[1163] Jonathan P. Seldin. The search for a reduction in combinatory logic equivalent to λβ-reduction, Part II. Theoretical Computer Science, 663 (??):34–58, ???? 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307447.

Anonymous:2017:EBj

[1164] Anonymous. Editorial Board. Theoretical Computer Science, 664(??): iii–ix, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300580.

Campadelli:2017:P

[1165] Paola Campadelli, Giovanni Pighizzini, and Massimiliano Goldwurm. Preface. Theoretical Computer Science, 664 (??):1, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300129.

Campadelli:2017:ABS

[1166] Paola Campadelli, Massimiliano Goldwurm, and Giovanni Pighizzini. Alberto Bertoni: a scientist and a friend. Theoretical Computer Science, 664(??): 2–11, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300336.

Adamo:2017:SDI

[1167] Alessandro Adamo, Giuliano Grossi, Raffaella Lanzarotti, and Jianyi Lin. Sparse decomposition by iterating Lipschitzian-type mappings. Theoretical Computer Science, 664(??):12– 28, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516300640.

Barcucci:2017:CBF

[1168] Elena Barcucci, Antonio Bernini, Stefano Bilotta, and Renzo Pinzani. Cross-bifix-free sets in two dimensions. Theoretical Computer Science, 664(??): 29–38, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515007793.

Bianchi:2017:QFA

[1169] Maria Paola Bianchi, Carlo Mereghetti, and Beatrice Palano. Quantum finite automata: Advances on Bertoni’s ideas. Theoretical Computer Science, 664(??):39–53, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439751600089X.

Brocchi:2017:EGC

[1170] Stefano Brocchi, Giusi Castiglione, and Paolo Massazza. On the exhaustive generation of k-convex polyominoes. Theoretical Computer Science, 664(??): 54–66, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751600116X.

Carpi:2017:ISF

[1171] Arturo Carpi and Flavio D’Alessandro. On incomplete and synchronizing finite sets. Theoretical Computer Science, 664(??):67–77, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397515007896.

Choffrut:2017:HOR

[1172] Christian Choffrut. An Hadamard operation on rational relations. Theoretical Computer Science, 664(??): 78–90, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515009482.

Reghizzi:2017:CMP

[1173] Stefano Crespi Reghizzi and Pierluigi San Pietro. Counter machines, Petri Nets, and consensual computation. Theoretical Computer Science, 664(??): 91–116, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011433.

DeFelice:2017:DPC

[1174] Clelia De Felice, Sabrina Mantaci, and Antonio Restivo. On the decomposition of prefix codes. Theoretical Computer Science, 664(??):117– 130, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751501141X.

Dennunzio:2017:CCF

[1175] Alberto Dennunzio, Enrico Formenti, Luca Manzoni, Giancarlo Mauri, and Antonio E. Porreca. Computational complexity of finite asynchronous cellular automata. Theoretical Computer Science, 664(??):131– 143, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515011421.

Sabadini:2017:GAN

[1176] N. Sabadini, F. Schiavio, and R. F. C.

Walters. On the geometry and algebra of networks with state. Theoretical Computer Science, 664(??): 144–163, February 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516000529.

Anonymous:2017:EBk

[1177] Anonymous. Editorial Board. Theoretical Computer Science, 665(??): iii–ix, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300762.

Golovach:2017:GEG

[1178] Petr A. Golovach and George B. Mertzios. Graph editing to a given degree sequence. Theoretical Computer Science, 665(??):1– 12, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307137.

Jez:2017:UCG

[1179] Artur Jez and Alexander Okhotin. Unambiguous conjunctive grammars over a one-symbol alphabet. Theoretical Computer Science, 665(??): 13–39, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307150.

Giotis:2017:AEC

[1180] Ioannis Giotis, Lefteris Kirousis, Kostas I. Psaromiligkos, and Dimitrios M. Thilikos. Acyclic edge coloring through the Lovász Local Lemma.

Theoretical Computer Science, 665 (??):40–50, ???? 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307174.

Datta:2017:SSS

[1181] Ajoy K. Datta, Stéphane Devismes, and Lawrence L. Larmore. Selfstabilizing silent disjunction in an anonymous network. Theoretical Computer Science, 665(??):51– 72, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307186.

Lancia:2017:SSS

[1182] Giuseppe Lancia, Luke Mathieson, and Pablo Moscato. Separating sets of strings by finding matching patterns is almost always hard. Theoretical Computer Science, 665(??): 73–86, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307460.

Al-Bataineh:2017:FMM

[1183] Omar Al-Bataineh, Mark Reynolds, and Tim French. Finding minimum and maximum termination time of timed automata models with cyclic behaviour. Theoretical Computer Science, 665(??):87–104, ???? 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307484.

Anonymous:2017:EBl

[1184] Anonymous. Editorial Board. Theoretical Computer Science, 666(??): iii–ix, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300932.

Mishra:2017:MAC

[1185] Apoorva Mishra and Anupam Shukla. Mathematical analysis of the cumulative effect of novel ternary crossover operator and mutation on probability of survival of a schema. Theoretical Computer Science, 666(??): 1–11, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303802.

Teh:2017:IRS

[1186] Wen Chean Teh and Adrian Atanasiu. Irreducible reaction systems and reaction system rank. Theoretical Computer Science, 666(??): 12–20, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630439X.

Zaitsev:2017:GNC

[1187] Dmitry A. Zaitsev. A generalized neighborhood for cellular automata. Theoretical Computer Science, 666 (??):21–35, ???? 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305916.

Durand:2017:ATE

[1188] Bruno Durand, Guilhem Gamard, and Anaël Grandjean. Aperiodic tilings and entropy. Theoretical Computer Science, 666(??):36–47, ???? 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307198.

Zheng:2017:PPS

[1189] Shenggen Zheng, Lvzhou Li, Daowen Qiu, and Jozef Gruska. Promise problems solved by quantum and classical finite automata. Theoretical Computer Science, 666(??):48– 64, ???? 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307538.

Anonymous:2017:EBm

[1190] Anonymous. Editorial Board. Theoretical Computer Science, 667(??): iii–ix, March 8, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301147.

Chalk:2017:CIR

[1191] Cameron T. Chalk, Bin Fu, Eric Martinez, Robert Schweller, and Tim Wylie. Concentration independent random number generation in tile self-assembly. Theoretical Computer Science, 667(??): 1–15, March 8, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307496.

Datta:2017:SFH

[1192] Pratish Datta, Ratna Dutta, and Sourav Mukhopadhyay. Strongly fullhiding inner product encryption. Theoretical Computer Science, 667(??): 16–50, March 8, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307526.

Lee:2017:SUE

[1193] Kwangsu Lee, Seung Geol Choi, Dong Hoon Lee, Jong Hwan Park, and Moti Yung. Self-updatable encryption: Time constrained access control with hidden attributes and better efficiency. Theoretical Computer Science, 667(??):51–92, March 8, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307551.

Gentner:2017:DMD

[1194] Michael Gentner and Dieter Rautenbach. Dynamic monopolies for degree proportional thresholds in connected graphs of girth at least five and trees. Theoretical Computer Science, 667(??): 93–100, March 8, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307563.

Flammini:2017:NMG

[1195] M. Flammini, V. Gallotti, G. Melideo, G. Monaco, and L. Moscardelli. Network movement games. Theoretical Computer Science, 667(??):101– 118, March 8, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307575.

Anonymous:2017:EBn

[1196] Anonymous. Editorial Board. Theoretical Computer Science, 668(??): iii–ix, March 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301445.

Rao:2017:CFA

[1197] Y. Sreenivasa Rao and Ratna Dutta.

Computational friendly attributebased encryptions with short ciphertext. Theoretical Computer Science, 668(??):1–26, March 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307587.

Charalambidis:2017:ETF

[1198] Angelos Charalambidis, Panos Rondogiannis, and Ioanna Symeonidou. Equivalence of two fixed-point semantics for definitional higher-order logic programs. Theoretical Computer Science, 668(??):27–42, March 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397517300373.

Foucaud:2017:ILD

[1199] Florent Foucaud, George B. Mertzios, Reza Naserasr, Aline Parreau, and Petru Valicov. Identification, locationdomination and metric dimension on interval and permutation graphs. I. Bounds. Theoretical Computer Science, 668(??):43–58, March 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397517300385.

Dekking:2017:CCF

[1200] F. Michel Dekking and Michael S. Keane. On the conjugacy class of the Fibonacci dynamical system. Theoretical Computer Science, 668(??): 59–69, March 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300506.

Dvorak:2017:GGC

[1201] Tomás Dvorák, Petr Gregor, and Václav Koubek. Generalized Gray codes with prescribed ends. Theoretical Computer Science, 668(??): 70–94, March 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300518.

Anonymous:2017:EBo

[1202] Anonymous. Editorial Board. Theoretical Computer Science, 669(??): iii–ix, March 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301718.

Jovanovic:2017:SOE

[1203] Aleksandra Jovanović, Marta Kwiatkowska, Gethin Norman, and Quentin Peyras. Symbolic optimal expected time reachability computation and controller synthesis for probabilistic timed automata.

Theoretical Computer Science, 669 (??):1–21, March 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300646.

Xi:2017:LCC

[1204] Xiaoyong Xi, Qingyu He, and Lingyun Yang. On the largest Cartesian closed category of stable domains. Theoretical Computer Science, 669(??): 22–32, March 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301007.

Bocchi:2017:MNT

[1205] Laura Bocchi, Tzu-Chun Chen, Romain Demangeon, Kohei Honda, and Nobuko Yoshida. Monitoring networks through multiparty session types. Theoretical Computer Science, 669(??): 33–58, March 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301263.

Anonymous:2017:EBp

[1206] Anonymous. Editorial Board. Theoretical Computer Science, 670(??): iii–ix, March 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301809.

Gourves:2017:BOM

[1207] Laurent Gourvès, Jérôme Monnot, Fanny Pascual, and Daniel Vanderpooten. Bi-objective matchings with the triangle inequality. Theoretical Computer Science, 670(??): 1–10, March 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300531.

Angelopoulos:2017:ILP

[1208] Spyros Angelopoulos, Diogo Arsénio, and Christoph Dürr. Infinite linear programming and online searching with turn cost. Theoretical Computer Science, 670(??):11– 22, March 29, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300622.

Wang:2017:EIS

[1209] Yan Wang, Hong Shen, and Jianxi Fan. Edge-independent spanning trees in augmented cubes. Theoretical Computer Science, 670(??): 23–32, March 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300658.

Chen:2017:DVR

[1210] Jianer Chen, Chao Xu, and Jianxin Wang. Dealing with 4-variables by resolution: an improved MaxSAT algorithm. Theoretical Computer Science, 670(??):33–44, March 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397517300695.

Kung:2017:ESR

[1211] Tzu-Liang Kung and Chun-Nan Hung.

Estimating the subsystem reliability of bubblesort networks. Theoretical Computer Science, 670(??): 45–55, March 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300701.

Martin:2017:CCQ

[1212] Barnaby Martin, András Pongrácz, and Michal Wrona. The complexity of counting quantifiers on equality languages. Theoretical Computer Science, 670(??):56–67, March 29, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300713.

Meszaros-Karkus:2017:HRS

[1213] Zsuzsa Mészáros-Karkus. Hardness results for stable exchange problems. Theoretical Computer Science, 670(??): 68–78, March 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300993.

Lan:2017:FPR

[1214] Yan Lan, Xin Han, Yinling Wang, Min Ge, He Guo, and Xin Chen. Flowshop problem F 2 → D|v = 1, c ≥ 1|C_{max} revisited. Theoretical Computer Science, 670(??): 79–85, March 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300671.

Anonymous:2017:EBq

[1215] Anonymous. Editorial Board. Theoretical Computer Science, 671(??): iii–ix, April 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301949.

Kari:2017:Ea

[1216] Lila Kari.  Editorial.  Theo-

retical Computer Science, 671(??): 1, April 6, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517303122.

Patitz:2017:TSI

[1217] Matthew J. Patitz. TCS Special Issue on Computational Self-Assembly. Theoretical Computer Science, 671 (??):2–3, April 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302232.

Demaine:2017:NGA

[1218] Erik D. Demaine, Sándor P. Fekete, Christian Scheffer, and Arne Schmidt. New geometric algorithms for fully connected staged self-assembly. Theoretical Computer Science, 671(??): 4–18, April 6, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630679X.

Ipparthi:2017:DCA

[1219] Dhananjay Ipparthi, Massimo Mastrangeli, and Andrew Winslow. Dipole codes attractively encode glue functions. Theoretical Computer Science, 671(??):19–25, April 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002000.

Ota:2017:RSD

[1220] Makoto Ota and Shinnosuke Seki. Rule set design problems for oritatami systems. Theoretical Computer Science, 671(??):26–35, April 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304868.

Michail:2017:CPN

[1221] Othon Michail and Paul G. Spirakis.

Connectivity preserving network transformers. Theoretical Computer Science, 671(??):36–55, April 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002024.

Derakhshandeh:2017:UCP

[1222] Zahra Derakhshandeh, Robert Gmyr, Andréa W. Richa, Christian Scheideler, and Thim Strothmann. Universal coating for programmable matter. Theoretical Computer Science, 671 (??):56–68, April 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516002012.

Ellis-Monaghan:2017:DTR

[1223] Joanna A. Ellis-Monaghan, Greta Pangborn, Nadrian C. Seeman, Sam Blakeley, Conor Disher, Mary Falcigno, Brianna Healy, Ada Morse, Bharti Singh, and Melissa Westland. Design tools for reporter strands and DNA origami scaffold strands. Theoretical Computer Science, 671(??): 69–78, April 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305308.

Anonymous:2017:EBr

[1224] Anonymous. Editorial Board. Theoretical Computer Science, 672(??): iii–ix, April 11, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302141.

Meseguer:2017:SCC

[1225] José Meseguer. Strict coherence of conditional rewriting modulo axioms. Theoretical Computer Science, 672 (??):1–35, April 11, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630754X.

Bonelli:2017:ANB

[1226] Eduardo Bonelli, Delia Kesner, Carlos Lombardi, and Alejandro Ríos. On abstract normalisation beyond neededness. Theoretical Computer Science, 672(??):36–63, April 11, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301019.

Ayala-Rincon:2017:IDP

[1227] Mauricio Ayala-Rincón, Maribel Fernández, and Daniele Nantes-Sobrinho. Intruder deduction problem for locally stable theories with normal forms and inverses. Theoretical Computer Science, 672(??):64–100, April 11, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301214.

Anonymous:2017:EBs

[1228] Anonymous. Editorial Board. Theoretical Computer Science, 673(??): iii–ix, April 18, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302402.

Barmpalias:2017:RNP

[1229] George Barmpalias, Douglas Cenzer, and Christopher P. Porter. Random numbers as probabilities of machine behavior. Theoretical Computer Science, 673(??):1–18, April 18, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301032.

Furusawa:2017:UCR

[1230] Hitoshi Furusawa. Uniform continuity of relations and nondeterministic cellular automata. Theoretical Computer Science, 673(??): 19–29, April 18, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301068.

Pan:2017:NSP

[1231] Linqiang Pan, Zhiqiang Zhang, Tingfang Wu, and Jinbang Xu. Numerical P systems with production thresholds. Theoretical Computer Science, 673(??):30–41, April 18, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301597.

Leporati:2017:TSA

[1232] Alberto Leporati, Luca Manzoni, Giancarlo Mauri, Antonio E. Porreca, and Claudio Zandron. A toolbox for simpler active membrane algorithms. Theoretical Computer Science, 673 (??):42–57, April 18, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302098.

Anonymous:2017:EBt

[1233] Anonymous. Editorial Board. Theoretical Computer Science, 674(??): iii–ix, April 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302682.

Brough:2017:ASN

[1234] Tara Brough and Alan J. Cain. Automaton semigroups: New constructions results and examples of non-automaton semigroups. Theoretical Computer Science, 674(??): 1–15, April 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301056.

Ding:2017:ISS

[1235] Wei Ding and Ke Qiu. Incremental single-source shortest paths in digraphs with arbitrary positive arc weights. Theoretical Computer Science, 674 (??):16–31, April 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301093.

Kortsarz:2017:ASL

[1236] Guy Kortsarz and Zeev Nutov. Approximating source location and star survivable network problems. Theoretical Computer Science, 674(??): 32–42, April 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301251.

Xu:2017:CDC

[1237] Min Xu, Krishnaiyan Thulasiraman, and Qiang Zhu. Conditional diagnosability of a class of matching composition networks under the comparison model. Theoretical Computer Science, 674(??):43–52, April 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301275.

Li:2017:GNC

[1238] Desai Li and Mei Lu. The g-good-neighbor conditional diagnosability of star graphs under the PMC and MM* model. Theoretical Computer Science, 674(??):53–59, April 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301287.

Ahadi:2017:ACW

[1239] Arash Ahadi, Ali Dehghan, and Mohsen Mollahajiaghaei. Algorithmic complexity of weakly semiregular partitioning and the representation number. Theoretical Computer Science, 674 (??):60–72, April 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301305.

Kalimullin:2017:ASC

[1240] Iskander Kalimullin, Alexander Melnikov, and Keng Meng Ng. Algebraic structures computable without delay. Theoretical Computer Science, 674(??):73–98, April 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301317.

Fraser:2017:WSD

[1241] Robert Fraser and Alejandro López-Ortiz. The within-strip discrete unit disk cover problem. Theoretical Computer Science, 674(??): 99–115, April 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301329.

Shitov:2017:TNC

[1242] Yaroslav Shitov. A tractable NPcompleteness proof for the twocoloring without monochromatic cycles of fixed length. Theoretical Computer Science, 674(??):116– 118, April 25, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730107X.

Anonymous:2017:EBu

[1243] Anonymous. Editorial Board. Theoretical Computer Science, 675(??): iii–ix, May 2, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303018.

Kadelka:2017:MNC

[1244] Claus Kadelka, Yuan Li, Jack Kuipers, John O. Adeyeye, and Reinhard Laubenbacher. Multistate nested canalizing functions and their networks. Theoretical Computer Science, 675(??):1–14, May 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301330.

Yoon:2017:RRL

[1245] Sang Duk Yoon, Hee-Kap Ahn, and Jessica Sherette. Realistic roofs without local minimum edges over a rectilinear polygon. Theoretical Computer Science, 675(??):15– 26, May 2, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301378.

Kobayashi:2017:CBM

[1246] Koji M. Kobayashi, Shuichi Miyazaki, and Yasuo Okabe. Competitive buffer management for multi-queue switches in QoS networks using packet buffering algorithms. Theoretical Computer Science, 675(??): 27–42, May 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730138X.

Baffier:2017:HNH

[1247] Jean-François Baffier, Man-Kwun Chiu, Yago Diez, Matias Korman, Valia Mitsou, André van Renssen, Marcel Roeloffzen, and Yushi Uno. Hanabi is NP-hard, even for cheaters who look at their cards. Theoretical Computer Science, 675(??):43– 55, May 2, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301573.

Reimers:2017:MES

[1248] Arne C. Reimers, Alexandra-M. Reimers, and Yaron Goldstein. Minimal equivalent subgraphs containing a given set of arcs. Theoretical Computer Science, 675(??):56– 63, May 2, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301585.

Li:2017:MMS

[1249] Huan Li and Zhongzhi Zhang. Maximum matchings in scale-free networks with identical degree distribution. Theoretical Computer Science, 675(??):64–81, May 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301603.

Lin:2017:LAP

[1250] Yuqing Lin and Fuji Zhang. A linear algorithm for a perfect matching in polyomino graphs. Theoretical Computer Science, 675(??):82– 88, May 2, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301615.

Barmpalias:2017:KC

[1251] George Barmpalias and Rodney G.

Downey. Kobayashi compressibility. Theoretical Computer Science, 675 (??):89–100, May 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301640.

Mamino:2017:SRS

[1252] Marcello Mamino. Strategy recovery for stochastic mean payoff games. Theoretical Computer Science, 675 (??):101–104, May 2, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301391.

Anonymous:2017:EBv

[1253] Anonymous. Editorial Board. Theoretical Computer Science, 676(??): iii–ix, May 9, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303213.

Abello:2017:MRC

[1254] James Abello, Pavel Klavík, Jan Kratochvíl, and Tomás Vyskocil. MSOL restricted contractibility to planar graphs. Theoretical Computer Science, 676(??):1–14, May 9, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301652.

Dey:2017:FBV

[1255] Palash Dey, Neeldhara Misra, and Y. Narahari. Frugal bribery in voting. Theoretical Computer Science, 676(??):15–32, May 9, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301664.

Chapelle:2017:EEA

[1256] Mathieu Chapelle, Manfred Cochefert, Dieter Kratsch, Romain Letourneur, and Mathieu Liedloff. Exact exponential algorithms to find tropical connected sets of minimum size. Theoretical Computer Science, 676 (??):33–41, May 9, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301883.

Chlebikova:2017:FPF

[1257] Janka Chlebíková and Morgan Chopin.

The firefighter problem: Further steps in understanding its complexity. Theoretical Computer Science, 676 (??):42–51, May 9, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301895.

Huang:2017:SPS

[1258] Zhenyu Huang and Dongdai Lin. Solving polynomial systems with noise over F_{2}: Revisited. Theoretical Computer Science, 676(??):52– 68, May 9, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301998.

Ham:2017:GTR

[1259] Lucy Ham. Gap theorems for robust satisfiability: Boolean CSPs and beyond. Theoretical Computer Science, 676(??):69–91, May 9, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302001.

Badano:2017:CCF

[1260] M. Badano and D. Vaggione. Characterization of context-free languages. Theoretical Computer Science, 676 (??):92–96, May 9, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301871.

Anonymous:2017:EBw

[1261] Anonymous. Editorial Board. Theoretical Computer Science, 677(??): iii–ix, May 16, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303365.

Li:2017:PDP

[1262] Jing Li, Guoren Wang, and Lichao Chen. Paired 2-disjoint path covers of multi-dimensional torus networks with 2n − 3 faulty edges. Theoretical Computer Science, 677 (??):1–11, May 16, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302025.

Shan:2017:DNM

[1263] Liren Shan, Huan Li, and Zhongzhi Zhang. Domination number and minimum dominating sets in pseudofractal scale-free web and Sierpiński graph. Theoretical Computer Science, 677 (??):12–30, May 16, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302037.

He:2017:SPC

[1264] Weihua He, Ziwen Liu, and Chao Yang. Snowman is PSPACE-complete. Theoretical Computer Science, 677 (??):31–40, May 16, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302050.

Taubenfeld:2017:CSD

[1265] Gadi Taubenfeld. Contention-sensitive data structures and algorithms. Theoretical Computer Science, 677(??): 41–55, May 16, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730230X.

Mathieson:2017:GEP

[1266] Luke Mathieson. Graph editing problems with extended regularity constraints. Theoretical Computer Science, 677(??):56–68, May 16, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302323.

Ito:2017:ESC

[1267] Takehiro Ito, Naonori Kakimura, Naoyuki Kamiyama, Yusuke Kobayashi, and Yoshio Okamoto. Efficient stabilization of cooperative matching games. Theoretical Computer Science, 677(??):69–82, May 16, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302335.

Anonymous:2017:EBx

[1268] Anonymous. Editorial Board. Theoretical Computer Science, 678(??): iii–ix, May 23, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730347X.

Antoniadis:2017:CSS

[1269] Antonios Antoniadis, Peter Kling, Sebastian Ott, and Sören Riechers. Continuous speed scaling with variability: a simple and direct approach. Theoretical Computer Science, 678 (??):1–13, May 23, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302347.

Hemaspaandra:2017:CCC

[1270] Edith Hemaspaandra, Lane A. Hemaspaandra, and Jörg Rothe. The complexity of controlling candidatesequential elections. Theoretical Computer Science, 678(??):14–21, May 23, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302608.

Louza:2017:IES

[1271] Felipe A. Louza, Simon Gog, and Guilherme P. Telles. Inducing enhanced suffix arrays for string collections. Theoretical Computer Science, 678(??):22–39, May 23, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302621.

Markakis:2017:IBC

[1272] Evangelos Markakis and Orestis Telelis. Item bidding for combinatorial public projects. Theoretical Computer Science, 678(??):40– 55, May 23, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302724.

Pilsniak:2017:EMD

[1273] Monika Pilśniak. Edge motion and the distinguishing index. Theoretical Computer Science, 678(??):56– 62, May 23, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302736.

Liu:2017:GRE

[1274] Jing Liu, Minmin Liu, Changji Wang, and Shaowen Yao. Group rekeying in the exclusive subset-cover framework. Theoretical Computer Science, 678(??):63–77, May 23, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302852.

Anonymous:2017:EBy

[1275] Anonymous. Editorial Board. Theoretical Computer Science, 679(??): iii–ix, May 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303584.

Drewes:2017:P

[1276] Frank Drewes.  Preface.  Theo-

retical Computer Science, 679(??): 1, May 30, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303845.

Maletti:2017:SFS

[1277] Andreas Maletti. Survey: Finite-state technology in natural language processing. Theoretical Computer Science, 679(??):2–17, May 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301669.

Gruber:2017:MDN

[1278] Hermann Gruber, Markus Holzer, and Sebastian Jakobi. More on deterministic and nondeterministic finite cover automata. Theoretical Computer Science, 679(??):18– 30, May 30, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305205.

Fernau:2017:CCR

[1279] Henning Fernau, Meenakshi Paramasivan, Markus L. Schmid, and Vojtech Vorel. Characterization and complexity results on jumping finite automata. Theoretical Computer Science, 679(??):31–52, May 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303206.

Mohri:2017:DAW

[1280] Mehryar Mohri and Michael D. Riley. A disambiguation algorithm for weighted automata. Theoretical Computer Science, 679(??):53– 68, May 30, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304376.

Berglund:2017:SRE

[1281] Martin Berglund and Brink van der Merwe. On the semantics of regular expression parsing in the wild. Theoretical Computer Science, 679 (??):69–82, May 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304790.

Kutrib:2017:CFL

[1282] Martin Kutrib and Matthias Wendlandt. Concatenation-free languages. Theoretical Computer Science, 679 (??):83–94, May 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516304224.

Caron:2017:HGO

[1283] Pascal Caron, Ludovic Mignot, and Clément Miklarz. On the hierarchy of generalizations of oneunambiguous regular languages. Theoretical Computer Science, 679(??): 95–106, May 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516303188.

Ng:2017:SCP

[1284] Timothy Ng, David Rappaport, and Kai Salomaa. State complexity of prefix distance. Theoretical Computer Science, 679(??):107– 117, May 30, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516301451.

Prusa:2017:UEP

[1285] Daniel Prusa and Klaus Reinhardt. Undecidability of the emptiness problem for context-free picture languages. Theoretical Computer Science, 679(??): 118–125, May 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300147.

Adiga:2017:ILT

[1286] Abhijin Adiga, Chris J. Kuhlman, Madhav V. Marathe, S. S. Ravi, Daniel J. Rosenkrantz, and Richard E. Stearns. Inferring local transition functions of discrete dynamical systems from observations of system behavior. Theoretical Computer Science, 679 (??):126–144, May 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630322X.

Hoffmann:2017:SIT

[1287] Stefan Hoffmann, Sibylle Schwarz, and Ludwig Staiger. Shift-invariant topologies for the Cantor space X^{ω}. Theoretical Computer Science, 679(??): 145–161, May 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305175.

Anonymous:2017:EBz

[1288] Anonymous. Editorial Board. Theoretical Computer Science, 680(??): iii–ix, June 5, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303705.

Fijalkow:2017:PTP

[1289] Nathanaël Fijalkow. Profinite techniques for probabilistic automata and the Markov Monoid algorithm. Theoretical Computer Science, 680(??): 1–14, June 5, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730316X.

Higgins:2017:LMD

[1290] Peter M. Higgins and Suhear Alwan. Languages with membership determined by single letter factors. Theoretical Computer Science, 680 (??):15–24, June 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303158.

Kehagias:2017:SCP

[1291] Ath. Kehagias and G. Konstantinidis. Selfish cops and passive robber: Qualitative games. Theoretical Computer Science, 680(??):25– 35, June 5, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303146.

Anonymous:2017:EBba

[1292] Anonymous. Editorial Board. Theoretical Computer Science, 681(??): iii–ix, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517303808.

Verschelde:2017:TSP

[1293] Jan Verschelde, Stephen M. Watt, and Lihong Zhi. TCS SNC preface. Theoretical Computer Science, 681(??):1–2, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302451.

Barkatou:2017:CIA

[1294] Moulay Barkatou, Paola Boito, and Esteban Segura Ugalde. A contour integral approach to the computation of invariant pairs. Theoretical Computer Science, 681(??): 3–26, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302463.

Batenkov:2017:ASN

[1295] Dmitry Batenkov. Accurate solution of near-colliding Prony systems via decimation and homotopy continuation. Theoretical Computer Science, 681(??):27–40, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302487.

Elkadi:2017:MPB

[1296] Mohamed Elkadi and André Galligo. On mixed polynomials of bidegree (n, 1). Theoretical Computer Science, 681(??):41–53, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302499.

Helmer:2017:DAC

[1297] Martin Helmer. A direct algorithm to compute the topological Euler characteristic and Chern–Schwartz– MacPherson class of projective complete intersection varieties. Theoretical Computer Science, 681(??): 54–74, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302517.

Li:2017:WNT

[1298] Yi Li. Witness to non-termination of linear programs. Theoretical Computer Science, 681(??):75–100, June 12, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302591.

Pan:2017:RPR

[1299] Victor Y. Pan and Liang Zhao. Real polynomial root-finding by means of matrix and polynomial iterations. Theoretical Computer Science, 681(??): 101–116, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302554.

Pan:2017:NOC

[1300] Victor Y. Pan and Elias P. Tsigaridas. Nearly optimal computations with structured matrices. Theoretical Computer Science, 681(??): 117–137, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302530.

Pan:2017:AAC

[1301] Victor Y. Pan and Elias Tsigaridas. Accelerated approximation of the complex roots and factors of a univariate polynomial. Theoretical Computer Science, 681(??):138– 145, June 12, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302529.

Schleicher:2017:NMP

[1302] Dierk Schleicher and Robin Stoll. Newton’s method in practice: Finding all roots of polynomials of degree one million efficiently. Theoretical Computer Science, 681(??): 146–166, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302475.

Sekigawa:2017:NPM

[1303] Hiroshi Sekigawa. The nearest polynomial to multiple given polynomials with a given zero in the real case. Theoretical Computer Science, 681(??): 167–175, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302566.

Usevich:2017:VPM

[1304] Konstantin Usevich and Ivan Markovsky.

Variable projection methods for approximate (greatest) common divisor computations. Theoretical Computer Science, 681(??):176–198, June 12, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302505.

Wang:2017:MCC

[1305] Wei guo Wang and Yimin Wei. Mixed and componentwise condition numbers for matrix decompositions. Theoretical Computer Science, 681(??): 199–216, June 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302578.

Wu:2017:CRW

[1306] Wenyuan Wu, Greg Reid, and Yong Feng. Computing real witness points of positive dimensional polynomial systems. Theoretical Computer Science, 681(??):217–231, June 12, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730258X.

Anonymous:2017:EBbb

[1307] Anonymous. Editorial Board. Theoretical Computer Science, 682(??): iii–ix, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303894.

Kari:2017:Eb

[1308] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 682(??):1, June 19, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304139.

Manea:2017:TSIa

[1309] Florin Manea, Bianca Truthe, and György Vaszil. TCS Special Issue on Languages and Combinatorics in Theory and Nature. Theoretical Computer Science, 682(??):2– 4, June 19, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304127.

Alhazov:2017:CAG

[1310] Artiom Alhazov, Henning Fernau, Rudolf Freund, Sergiu Ivanov, Rani Siromoney, and K. G. Subramanian. Contextual array grammars with matrix control, regular control languages, and tissue P systems control. Theoretical Computer Science, 682(??): 5–21, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302062.

Arroyo:2017:TBN

[1311] Fernando Arroyo, Sandra Gómez-Canaval, Victor Mitrana, and José Ramón Sánchez-Couso. Transducers based on networks of polarized evolutionary processors. Theoretical Computer Science, 682(??):22–29, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305679.

Bjorklund:2017:FBV

[1312] Johanna Björklund, Frank Drewes, and Anna Jonsson. Finding the N best vertices in an infinite weighted hypergraph. Theoretical Computer Science, 682(??):30–41, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302049.

Bordihn:2017:ASG

[1313] Henning Bordihn. Active symbols in grammars with valuations. Theoretical Computer Science, 682(??): 42–48, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751630603X.

Bottoni:2017:PSS

[1314] Paolo Bottoni, Anna Labella, and Victor Mitrana. Prefix-suffix square reduction. Theoretical Computer Science, 682(??):49–56, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307113.

Chen:2017:TLG

[1315] Haiming Chen, Mihai Ionescu, Andrei Paun, and Gheorghe Paun. On trace languages generated by (small) spiking neural P systems. Theoretical Computer Science, 682(??): 57–66, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305667.

Cho:2017:SCP

[1316] Da-Jung Cho, Daniel Goc, Yo-Sub Han, Sang-Ki Ko, Alexandros Palioudakis, and Kai Salomaa. State complexity of permutation on finite languages over a binary alphabet. Theoretical Computer Science, 682 (??):67–78, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302013.

Ehrenfeucht:2017:ERS

[1317] Andrzej Ehrenfeucht, Jetty Kleijn, Maciej Koutny, and Grzegorz Rozenberg. Evolving reaction systems. Theoretical Computer Science, 682(??): 79–99, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730035X.

Fernau:2017:CCG

[1318] Henning Fernau, Lakshmanan Kuppusamy, and Indhumathi Raman. On the computational completeness of graph-controlled insertion-deletion systems with binary sizes. Theoretical Computer Science, 682(??): 100–121, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300683.

Holzer:2017:CL

[1319] Markus Holzer, Sebastian Jakobi, and Martin Kutrib. The chop of languages. Theoretical Computer Science, 682(??): 122–137, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301044.

Jurgensen:2017:HLC

[1320] Helmut Jürgensen. Higher-level constructs for families and multisets. Theoretical Computer Science, 682(??): 138–148, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301366.

Kutrib:2017:OWR

[1321] Martin Kutrib and Andreas Malcher.

One-way reversible multi-head finite automata. Theoretical Computer Science, 682(??):149–164, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516306016.

Masopust:2017:BCF

[1322] Tomás Masopust and Michaël Thomazo.

On Boolean combinations forming piecewise testable languages. Theoretical Computer Science, 682(??): 165–179, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730066X.

Mercas:2017:AAB

[1323] Robert Mercas. On the aperiodic avoidability of binary patterns with variables and reversals. Theoretical Computer Science, 682(??): 180–189, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307502.

Otto:2017:RVL

[1324] Friedrich Otto and Frantisek Mráz. Regulated variants of limited context restarting automata. Theoretical Computer Science, 682(??): 190–207, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307162.

Salomaa:2017:MRS

[1325] Arto Salomaa. Minimal reaction systems: Duration and blips. Theoretical Computer Science, 682(??): 208–216, June 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301342.

Anonymous:2017:EBbc

[1326] Anonymous. Editorial Board. Theoretical Computer Science, 683(??): iii–ix, June 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304012.

Herlihy:2017:WFA

[1327] Maurice Herlihy, Sergio Rajsbaum, Michel Raynal, and Julien Stainer. From wait-free to arbitrary concurrent solo executions in colorless distributed computing. Theoretical Computer Science, 683(??): 1–21, June 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303298.

Cormode:2017:SLC

[1328] Graham Cormode and Hossein Jowhari.

A second look at counting triangles in graph streams (corrected). Theoretical Computer Science, 683(??): 22–30, June 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302560.

Cormode:2017:CSL

[1329] Graham Cormode and Hossein Jowhari.

Corrigendum to “A second look at counting triangles in graph streams” [Theoret. Comput. Sci. 552 (2014) 44– 51]. Theoretical Computer Science, 683 (??):31–32, June 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302924.

Anonymous:2017:EBbd

[1330] Anonymous. Editorial Board. Theoretical Computer Science, 684(??): iii–ix, July 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304279.

Manea:2017:TSIb

[1331] Florin Manea and Dirk Nowotka. TCS Special Issue: Combinatorics on Words — WORDS 2015. Theoretical Computer Science, 684(??): 1–2, July 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304693.

Berthe:2017:SS

[1332] Valérie Berthé, Clelia De Felice, Vincent Delecroix, Francesco Dolce, Julien Leroy, Dominique Perrin, Christophe Reutenauer, and Giuseppina Rindone. Specular sets. Theoretical Computer Science, 684(??): 3–28, July 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730186X.

Fici:2017:ASR

[1333] Gabriele Fici, Filippo Mignosi, and Jeffrey Shallit. Abelian-square-rich words. Theoretical Computer Science, 684(??):29–42, July 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301299.

Holub:2017:PFL

[1334] Stepán Holub. Prefix frequency of lost positions. Theoretical Computer Science, 684(??):43–52, July 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397517301020.

Holub:2017:FBW

[1335] Stepán Holub and Mike Müller.

Fully bordered words. Theoretical Computer Science, 684(??):53– 58, July 7, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305989.

Anonymous:2017:EBbe

[1336] Anonymous. Editorial Board. Theoretical Computer Science, 685(??): iii–ix, July 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304413.

Ayala-Rincon:2017:LSF

[1337] Mauricio Ayala-Rincón, Ian Mackie, and Ugo Montanari. Logical and semantic frameworks with applications. Theoretical Computer Science, 685 (??):1–2, July 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304711.

Alves:2017:GBF

[1338] Sandra Alves and Maribel Fernández.

A graph-based framework for the analysis of access control policies. Theoretical Computer Science, 685 (??):3–22, July 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305965.

Benevides:2017:BLE

[1339] Mario Benevides. Bisimilar and logically equivalent programs in PDL with parallel operator. Theoretical Computer Science, 685(??):23– 45, July 15, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302785.

Olarte:2017:CBF

[1340] Carlos Olarte and Elaine Pimentel. On concurrent behaviors and focusing in linear logic. Theoretical Computer Science, 685(??):46– 64, July 15, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304832.

Copello:2017:FML

[1341] Ernesto Copello, Nora Szasz, and Álvaro Tasistro. Formal metatheory of the Lambda calculus using Stoughton’s substitution. Theoretical Computer Science, 685(??):65– 82, July 15, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516304820.

Veloso:2017:GCM

[1342] Paulo A. S. Veloso, Sheila R. M. Veloso, and Mario R. F. Benevides. On a graph calculus for modalities. Theoretical Computer Science, 685(??): 83–103, July 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301238.

Anonymous:2017:EBbf

[1343] Anonymous. Editorial Board. Theoretical Computer Science, 686(??): iii–ix, July 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304504.

Gaina:2017:FLP

[1344] Daniel Gaina. Foundations of logic programming in hybrid logics with user-defined sharing. Theoretical Computer Science, 686(??):1– 24, July 22, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303316.

Cavalheiro:2017:TPG

[1345] Simone André da Costa Cavalheiro, Luciana Foss, and Leila Ribeiro. Theorem proving graph grammars with attributes and negative application conditions. Theoretical Computer Science, 686(??):25–77, July 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303419.

Anonymous:2017:EBbg

[1346] Anonymous. Editorial Board. Theoretical Computer Science, 687(??): iii–ix, July 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304929.

Lin:2017:AFT

[1347] Shangwei Lin, Chanchan Zhou, and Chunfang Li. Arc fault tolerance of Kautz digraphs. Theoretical Computer Science, 687(??):1– 10, July 29, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517303420.

Bulatov:2017:FCE

[1348] Andrei Bulatov, Leslie Ann Goldberg, Mark Jerrum, David Richerby, and Stanislav Zivný. Functional clones and expressibility of partition functions. Theoretical Computer Science, 687(??):11–39, July 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517303535.

Hu:2017:MPD

[1349] Xiaomin Hu, Yingzhi Tian, Xiaodong Liang, and Jixiang Meng. Matching preclusion for n-dimensional torus networks. Theoretical Computer Science, 687(??):40–47, July 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730364X.

Vesti:2017:RSF

[1350] Jetro Vesti. Rich square-free words. Theoretical Computer Science, 687 (??):48–61, July 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517303651.

Gorzkowska:2017:PBD

[1351] Aleksandra Gorzkowska and Monika Pilśniak. Precise bounds for the distinguishing index of the Cartesian product. Theoretical Computer Science, 687 (??):62–69, July 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303821.

Arvind:2017:FFP

[1352] V. Arvind. Finding fixed point free elements and small bases in permutation groups. Theoretical Computer Science, 687(??):70–78, July 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303948.

Zhu:2017:SEA

[1353] Daxin Zhu, Lei Wang, Tinran Wang, and Xiaodong Wang. A space efficient algorithm for the longest common subsequence in k-length substrings. Theoretical Computer Science, 687(??):79–92, July 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304206.

Dong:2017:CFP

[1354] Jianming Dong, Jueliang Hu, Mikhail Y.

Kovalyov, Guohui Lin, Taibo Luo, Weitian Tong, Xueshi Wang, and Yinfeng Xu. Corrigendum to “An FPTAS for the parallel two-stage flowshop problem” [Theoret. Comput. Sci. 657 (2017) 64–72]. Theoretical Computer Science, 687(??):

93–94, July 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304218. See

[1090].

Anonymous:2017:EBbh

[1355] Anonymous. Editorial Board. Theoretical Computer Science, 688(??): iii–ix, August 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305017.

Kothapalli:2017:P

[1356] Kishore Kothapalli and Sergio Rajsbaum. Preface. Theoretical Computer Science, 688(??):1, August 6, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305157.

Abraham:2017:TPS

[1357] Uri Abraham and Gal Amram. Twoprocess synchronization. Theoretical Computer Science, 688(??):2– 23, August 6, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516300500.

Natarajan:2017:EAA

[1358] Aravind Natarajan, Himanshu Chauhan, Neeraj Mittal, and Vijay K. Garg. Efficient abstraction algorithms for predicate detection. Theoretical Computer Science, 688(??):24–48, August 6, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012074.

Altisen:2017:PSS

[1359] Karine Altisen and Stéphane Devismes.

On probabilistic snap-stabilization. Theoretical Computer Science, 688(??): 49–76, August 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515007161.

Feinerman:2017:FRC

[1360] Ofer Feinerman, Amos Korman, Shay Kutten, and Yoav Rodeh. Fast rendezvous on a cycle by agents with different speeds. Theoretical Computer Science, 688(??):77– 85, August 6, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012050.

Hirve:2017:HHP

[1361] Sachin Hirve, Roberto Palmieri, and Binoy Ravindran. HiperTM: High performance, fault-tolerant transactional memory. Theoretical Computer Science, 688(??):86–102, August 6, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516301591.

Kuznetsov:2017:NIL

[1362] Petr Kuznetsov and Sathya Peri.

Non-interference and local correctness in transactional memory. Theoretical Computer Science, 688(??): 103–116, August 6, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302572.

Zhong:2017:PMS

[1363] Xueling Zhong and Jinwen Ou. Parallel machine scheduling with restricted job rejection. Theoretical Computer Science, 690(??):1–11, August 22, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304772.

Hon:2017:PAE

[1364] Wing-Kai Hon, Sharma V. Thankachan, and Bojian Xu. In-place algorithms for exact and approximate shortest unique substring problems. Theoretical Computer Science, 690(??): 12–25, August 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304760.

Keshavarz-Kohjerdi:2017:LTA

[1365] Fatemeh Keshavarz-Kohjerdi and Alireza Bagheri. A linear-time algorithm for finding Hamiltonian (s, t)- paths in even-sized rectangular grid graphs with a rectangular hole. Theoretical Computer Science, 690(??): 26–58, August 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304759.

Gu:2017:ERB

[1366] Mei-Mei Gu, Rong-Xia Hao, Jun-Ming Xu, and Yan-Quan Feng. Equal relation between the extra connectivity and pessimistic diagnosability for some regular graphs. Theoretical Computer Science, 690(??): 59–72, August 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304802.

Mazo:2017:NLL

[1367] Loïc Mazo and Étienne Baudrier.

Non-local length estimators and concave functions. Theoretical Computer Science, 690(??):73–90, August 22, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305121.

Gozupek:2017:PCM

[1368] Didem Gözüpek, Sibel Özkan, Christophe Paul, Ignasi Sau, and Mordechai Shalom. Parameterized complexity of the MINCCA problem on graphs of bounded decomposability. Theoretical Computer Science, 690(??): 91–103, August 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730525X.

Mishra:2017:AOP

[1369] Sounaka Mishra, Shijin Rajakrishnan, and Saket Saurabh. On approximability of optimization problems related to Red/Blue-split graphs. Theoretical Computer Science, 690(??): 104–113, August 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305200.

Cohen:2017:ACD

[1370] Nathann Cohen, David Coudert, Guillaume Ducoffe, and Aurélien Lancin. Applying clique-decomposition for computing Gromov hyperbolicity. Theoretical Computer Science, 690(??): 114–139, August 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730508X.

Anonymous:2017:EBbi

[1371] Anonymous. Editorial Board. Theoretical Computer Science, 690(??): ifc, August 22, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305881.

Zhang:2017:CCE

[1372] Zhongxi Zhang and Qingguo Li. On cartesian closed extensions of non-pointed domains. Theoretical Computer Science, 691(??):1–9, August 29, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305558.

Bilokon:2017:DTA

[1373] Paul Bilokon and Abbas Edalat. A domain-theoretic approach to Brownian motion and general continuous stochastic processes. Theoretical Computer Science, 691(??):10– 26, August 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305546.

Vasile:2017:TWT

[1374] Cristian-Ioan Vasile, Derya Aksaray, and Calin Belta. Time window temporal logic. Theoretical Computer Science, 691(??):27–54, August 29, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305509.

Rabe:2017:MA

[1375] Florian Rabe. Morphism axioms. Theoretical Computer Science, 691(??): 55–80, August 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305443.

Xu:2017:CBF

[1376] Kevin H. Xu. A class of bounded functions, a database language and an extended lambda calculus. Theoretical Computer Science, 691(??): 81–106, August 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305406.

Anonymous:2017:EBbj

[1377] Anonymous. Editorial Board. Theoretical Computer Science, 691(??): ifc, August 29, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306047.

Terrier:2017:RPS

[1378] Véronique Terrier. Recognition of poly-slender context-free languages by trellis automata. Theoretical Computer Science, 692(??):1–24, September 5, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304875.

Liskiewicz:2017:SLS

[1379] Maciej Liśkiewicz, Rüdiger Reischuk, and Ulrich Wölfel. Security levels in steganography — insecurity does not imply detectability. Theoretical Computer Science, 692(??): 25–45, September 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305194.

Munaro:2017:BCG

[1380] Andrea Munaro. Boundary classes for graph problems involving nonlocal properties. Theoretical Computer Science, 692(??):46–71, September 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397517305248.

Anta:2017:APS

[1381] Antonio Fernández Anta, Chryssis Georgiou, Dariusz R. Kowalski, and Elli Zavou. Adaptive packet scheduling over a wireless channel under constrained jamming. Theoretical Computer Science, 692(??):72– 89, September 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305327.

Biagioni:2017:NYP

[1382] Silvio Biagioni, Daniel Masny, and Daniele Venturi. Naor–Yung paradigm with shared randomness and applications. Theoretical Computer Science, 692(??):90–113, September 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397517305315.

Anonymous:2017:EBbk

[1383] Anonymous. Editorial Board. Theoretical Computer Science, 692(??): ifc, September 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730614X.

vanDitmarsch:2017:UAA

[1384] Hans van Ditmarsch, Wiebe van der Hoek, and Louwe B. Kuijer. The undecidability of arbitrary arrow update logic. Theoretical Computer Science, 693(??):1–12, September 12, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730556X.

Bujtor:2017:AMI

[1385] Ferenc Bujtor and Walter Vogler. ACTL for Modal Interface Automata. Theoretical Computer Science, 693(??): 13–34, September 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305236.

Nigam:2017:SFM

[1386] Vivek Nigam, Carlos Olarte, and Elaine Pimentel. On subexponentials, focusing and modalities in concurrent systems. Theoretical Computer Science, 693(??):35– 58, September 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305212.

Anonymous:2017:EBbl

[1387] Anonymous. Editorial Board. Theoretical Computer Science, 693(??): ifc, September 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306321.

Li:2017:CPM

[1388] Weian Li, Wenjing Liu, Tiantian Chen, Xiaoying Qu, Qizhi Fang, and Ker-I Ko. Competitive profit maximization in social networks. Theoretical Computer Science, 694(??): 1–9, September 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305388.

Anderson:2017:GEP

[1389] Jaclyn Anderson, Kristina C. Garrett, and Brian Hopkins. Garden of Eden partitions in the sand pile and related models. Theoretical Computer Science, 694(??):10– 20, September 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305364.

Lukasiewicz:2017:NCC

[1390] Thomas Lukasiewicz and Enrico Malizia. A novel characterization of the complexity class Θ_{k}^{P} based on counting and comparison. Theoretical Computer Science, 694(??):21– 33, September 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305352.

Jayaram:2017:DSB

[1391] B. Jayaram, S. Arumugam, and K. Thulasiraman. Dominator sequences in bipartite graphs. Theoretical Computer Science, 694(??):34– 41, September 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305340.

Epstein:2017:SSJ

[1392] Leah Epstein and Elena Kleiman.

Scheduling selfish jobs on multidimensional parallel machines. Theoretical Computer Science, 694(??):42– 59, September 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305303.

Ma:2017:VDE

[1393] Meijie Ma, Douglas B. West, and Jun-Ming Xu. The vulnerability of the diameter of the enhanced hypercubes. Theoretical Computer Science, 694(??): 60–65, September 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305431.

Dragan:2017:MES

[1394] Feodor F. Dragan and Arne Leitert. On the minimum eccentricity shortest path problem. Theoretical Computer Science, 694(??):66– 78, September 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730542X.

Anonymous:2017:EBbm

[1395] Anonymous. Editorial Board. Theoretical Computer Science, 694(??): ifc, September 19, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306412.

Halffmann:2017:GAM

[1396] Pascal Halffmann, Stefan Ruzika, Clemens Thielen, and David Willems. A general approximation method for bicriteria minimization problems. Theoretical Computer Science, 695(??):1– 15, September 26, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305418.

Rusu:2017:GMP

[1397] Irena Rusu. Graph matching problems and the NP-hardness of sortedness constraints. Theoretical Computer Science, 695(??):16–27, September 26, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305674.

Durr:2017:MDA

[1398] Christoph Dürr, Łukasz Jeż, and Óscar C. Vásquez. Mechanism design for aggregating energy consumption and quality of service in speed scaling scheduling. Theoretical Computer Science, 695(??):28– 41, September 26, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305686.

Bacher:2017:ERS

[1399] Axel Bacher, Olivier Bodini, and Alice Jacquot. Efficient random sampling of binary and unary-binary trees via holonomic equations. Theoretical Computer Science, 695(??):42– 53, September 26, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305479.

Glantz:2017:FCC

[1400] Roland Glantz and Henning Meyerhenke. On finding convex cuts in general, bipartite and plane graphs. Theoretical Computer Science, 695(??):54– 73, September 26, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305741.

Zhang:2017:FTP

[1401] Shurong Zhang, Lin Chen, and Weihua Yang. On fault-tolerant path optimization under QoS constraint in multi-channel wireless networks. Theoretical Computer Science, 695(??):74– 82, September 26, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305753.

Anonymous:2017:EBbn

[1402] Anonymous. Editorial Board. Theoretical Computer Science, 695(??): ifc, September 26, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306540.

Wei:2017:IFI

[1403] Jianxin Wei, Yujun Yang, and Guangfu Wang. Infinite families of 2-isometric and not 3-isometric binary words. Theoretical Computer Science, 696(??): 1–10, October 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305765.

Bodei:2017:SAB

[1404] C. Bodei, L. Brodo, R. Gori, F. Levi, A. Bernini, and D. Hermith. A static analysis for Brane Calculi providing global occurrence counting information. Theoretical Computer Science, 696(??):11–51, October 5, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305467.

Cho:2017:PGO

[1405] Da-Jung Cho, Yo-Sub Han, Timothy Ng, and Kai Salomaa. Pseudoknotgenerating operation. Theoretical Computer Science, 696(??):52– 64, October 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730539X.

Ning:2017:CEC

[1406] Wantao Ning. The h-connectivity of exchanged crossed cube. Theoretical Computer Science, 696(??): 65–68, October 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305716. See

erratum [1496].

Anonymous:2017:EBbo

[1407] Anonymous. Editorial Board. Theoretical Computer Science, 696(??): ifc, October 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306680.

Elkin:2017:TE

[1408] Michael Elkin, Arnold Filtser, and Ofer Neiman. Terminal embeddings. Theoretical Computer Science, 697(??): 1–36, October 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305339.

Fuerst:2017:DLR

[1409] Carlo Fuerst, Maciej Pacut, and Stefan Schmid. Data locality and replica aware virtual cluster embeddings. Theoretical Computer Science, 697(??): 37–57, October 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305376.

Sakai:2017:IEA

[1410] Takayuki Sakai, Kazuhisa Seto, Suguru Tamaki, and Junichi Teruyama. Improved exact algorithms for mildly sparse instances of Max SAT. Theoretical Computer Science, 697(??): 58–68, October 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305492.

Abu-Khzam:2017:CVP

[1411] Faisal N. Abu-Khzam, Édouard Bonnet, and Florian Sikora. On the complexity of various parameterizations of common induced subgraph isomorphism. Theoretical Computer Science, 697(??):69–78, October 12, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305480.

Wei:2017:GNC

[1412] Yulong Wei and Min Xu. On ggood-neighbor conditional diagnosability of (n, k)-star networks. Theoretical Computer Science, 697(??): 79–90, October 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305790.

Ren:2017:GND

[1413] Yunxia Ren and Shiying Wang.

The g-good-neighbor diagnosability of locally twisted cubes. Theoretical Computer Science, 697(??):91– 97, October 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305789.

Navarra:2017:OKU

[1414] Alfredo Navarra and Cristina M.

Pinotti. Online knapsack of unknown capacity: How to optimize energy consumption in smartphones. Theoretical Computer Science, 697(??): 98–109, October 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305777.

Anonymous:2017:EBbp

[1415] Anonymous. Editorial Board. Theoretical Computer Science, 697(??): ifc, October 12, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306928.

Rombo:2017:FAS

[1416] Simona E. Rombo and Filippo Utro. Foreword: Algorithms, strings and theoretical approaches in the big data era — special issue in honor of the 60th birthday of Professor Raffaele Giancarlo. Theoretical Computer Science, 698(??):1–3, October 25, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517307661.

Amir:2017:SC

[1417] Amihood Amir, Alberto Apostolico, Travis Gagie, and Gad M. Landau. String cadences. Theoretical Computer Science, 698(??):4–8, October 25, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730467X.

Jamshidpey:2017:AMR

[1418] Arash Jamshidpey and David Sankoff. Asymptotic medians of random permutations sampled from reversal random walks. Theoretical Computer Science, 698(??):9–13, October 25, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305261.

Lim:2017:FAA

[1419] Jihyuk Lim and Kunsoo Park. A fast algorithm for the all-pairs suffix– prefix problem. Theoretical Computer Science, 698(??):14–24, October 25, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305510.

Crochemore:2017:CPP

[1420] Maxime Crochemore, Costas S. Iliopoulos, Tomasz Kociumaka, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. Covering problems for partial words and for indeterminate strings. Theoretical Computer Science, 698(??):25–39, October 25, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304681.

Georgiadis:2017:SCC

[1421] Loukas Georgiadis, Giuseppe F. Italiano, Aikaterini Karanasiou, Charis Papadopoulos, and Nikos Parotsidis. Sparse certificates for 2connectivity in directed graphs. Theoretical Computer Science, 698(??): 40–66, October 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305273.

Gagie:2017:WGF

[1422] Travis Gagie, Giovanni Manzini, and Jouni Sirén. Wheeler graphs: a framework for BWT-based data structures.

Theoretical Computer Science, 698(??): 67–78, October 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305285.

Mantaci:2017:MCE

[1423] Sabrina Mantaci, Antonio Restivo, Giovanna Rosone, Marinella Sciortino, and Luca Versari. Measuring the clustering effect of BWT via RLE. Theoretical Computer Science, 698(??): 79–87, October 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305534.

Girotto:2017:MRB

[1424] Samuele Girotto, Matteo Comin, and Cinzia Pizzi. Metagenomic reads binning with spaced seeds. Theoretical Computer Science, 698(??): 88–99, October 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304632.

Petrillo:2017:NDA

[1425] Umberto Ferraro Petrillo, Concettina Guerra, and Cinzia Pizzi. A new distributed alignment-free approach to compare whole proteomes. Theoretical Computer Science, 698(??):100– 112, October 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305297.

Anonymous:2017:EBbq

[1426] Anonymous. Editorial Board. Theoretical Computer Science, 698(??): ifc, October 25, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307120.

Whidden:2017:ROC

[1427] Chris Whidden and Frederick A.

Matsen IV. Ricci–Ollivier curvature of the rooted phylogenetic subtree–prune–regraft graph. Theoretical Computer Science, 699(??): 1–20, November 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301081.

Bhakta:2017:SWP

[1428] Prateek Bhakta and Dana Randall.

Sampling weighted perfect matchings on the square–octagon lattice. Theoretical Computer Science, 699(??): 21–32, November 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300634.

Durant:2017:DBC

[1429] Kevin Durant and Stephan Wagner.

On the distribution of betweenness centrality in random trees. Theoretical Computer Science, 699(??): 33–52, November 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307514.

Kobayashi:2017:RSC

[1430] Yusuke Kobayashi and Kenjiro Takazawa. Randomized strategies for cardinality robustness in the knapsack problem. Theoretical Computer Science, 699(??): 53–62, November 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307472.

Shiraga:2017:TVD

[1431] Takeharu Shiraga, Yukiko Yamauchi, Shuji Kijima, and Masafumi Yamashita. Total variation discrepancy of deterministic random walks for ergodic Markov chains. Theoretical Computer Science, 699(??): 63–74, November 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306120.

Anonymous:2017:EBbr

[1432] Anonymous. Editorial Board. Theoretical Computer Science, 699(??): ifc, November 7, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307223.

Fuentes-Sepulveda:2017:PCS

[1433] José Fuentes-Sepúlveda, Leo Ferres, Meng He, and Norbert Zeh. Parallel construction of succinct trees. Theoretical Computer Science, 700(??): 1–22, November 14, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730573X.

Lo:2017:OCD

[1434] Shih-Ming Lo, Wu-Hsiung Lin, Chiuyuan Chen, and Yu-Chee Tseng. Optimal coloring for data collection in tree-based wireless sensor networks. Theoretical Computer Science, 700(??): 23–36, November 14, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305728.

Wang:2017:OAS

[1435] Qian Wang, Ji Tian, Ruyan Fu, and Xiangjuan Yao. Online algorithms for scheduling on batch processing machines with interval graph compatibilities between jobs. Theoretical Computer Science, 700(??):37– 44, November 14, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305704.

Amit:2017:LMA

[1436] Mika Amit, Maxime Crochemore, Gad M. Landau, and Dina Sokol. Locating maximal approximate runs in a string. Theoretical Computer Science, 700(??):45–62, November 14, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305698.

Lozin:2017:MRW

[1437] Vadim Lozin, Dmitriy Malyshev, Raffaele Mosca, and Viktor Zamaraev. More results on weighted independent domination. Theoretical Computer Science, 700(??):63– 74, November 14, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306023.

Ganguly:2017:STT

[1438] Arnab Ganguly, Wing-Kai Hon, Rahul Shah, and Sharma V. Thankachan. Space-time trade-offs for finding shortest unique substrings and maximal unique matches. Theoretical Computer Science, 700(??):75– 88, November 14, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305832.

Jain:2017:EIL

[1439] Sanjay Jain and Jason Teutsch. Enumerations including laconic enumerators. Theoretical Computer Science, 700(??):89–95, November 14, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397517305819.

Anonymous:2017:EBbs

[1440] Anonymous. Editorial Board. Theoretical Computer Science, 700(??): ifc, November 14, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517307296.

Kari:2017:Ec

[1441] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 701(??):1, November 21, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307594.

Paun:2017:FSG

[1442] Gheorghe Paun, Ion Petre, Grzegorz Rozenberg, and Arto Salomaa. From finite state grammars to natural computing — in memory of Solomon Marcus. Theoretical Computer Science, 701 (??):2–7, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307570.

Marcus:2017:THN

[1443] Solomon Marcus. Ten human needs that education should take into account. Theoretical Computer Science, 701(??):8–11, November 21, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307582.

Alhazov:2017:SAS

[1444] Artiom Alhazov and Svetlana Cojocaru. Small asynchronous P systems with inhibitors defining nonsemilinear sets. Theoretical Computer Science, 701(??):12–19, November 21, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306667.

Azimi:2017:SSC

[1445] Sepinoud Azimi. Steady states of constrained reaction systems. Theoretical Computer Science, 701(??): 20–26, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302864.

Boykett:2017:FGS

[1446] Tim Boykett, Jarkko Kari, and Ville Salo. Finite generating sets for reversible gate sets under general conservation laws. Theoretical Computer Science, 701(??):27– 39, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301226.

Brijder:2017:SRT

[1447] Robert Brijder. Sorting by reversals and the theory of 4-regular graphs. Theoretical Computer Science, 701(??): 40–53, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302748.

Calude:2017:QFG

[1448] Cristian S. Calude, Michael J. Dinneen, and Richard Hua. QUBO formulations for the graph isomorphism problem and related problems. Theoretical Computer Science, 701(??): 54–69, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304590.

Cho:2017:OGI

[1449] Da-Jung Cho, Yo-Sub Han, Timothy Ng, and Kai Salomaa. Outfixguided insertion. Theoretical Computer Science, 701(??):70–84, November 21, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302797.

Ciobanu:2017:DSM

[1450] Gabriel Ciobanu and Eneia Nicolae Todoran. Denotational semantics of membrane systems by using complete metric spaces. Theoretical Computer Science, 701(??):85– 108, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306382.

Genova:2017:GIC

[1451] Daniela Genova, Hendrik Jan Hoogeboom, and Natasa Jonoska. A graph isomorphism condition and equivalence of reaction systems. Theoretical Computer Science, 701(??):109– 119, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304541.

Halava:2017:WTP

[1452] Vesa Halava and Tero Harju. Walks on tilings of polygons. Theoretical Computer Science, 701(??):120– 124, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730275X.

Honkala:2017:DWC

[1453] Juha Honkala. Discrete Watson– Crick dynamical systems. Theoretical Computer Science, 701(??):125– 131, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730124X.

Kleijn:2017:SST

[1454] Jetty Kleijn, Maciej Koutny, and Marta Pietkiewicz-Koutny. Signal set tissue systems and overlapping localities. Theoretical Computer Science, 701(??):132–145, November 21, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302773.

Krishna:2017:FRG

[1455] Shankara Narayanan Krishna, Marian Gheorghe, Florentin Ipate, Erzsébet Csuhaj-Varjú, and Rodica Ceterchi. Further results on generalised communicating P systems. Theoretical Computer Science, 701(??):146– 160, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304553.

Leporati:2017:CPS

[1456] Alberto Leporati, Luca Manzoni, Giancarlo Mauri, Antonio E. Porreca, and Claudio Zandron. The counting power of P systems with antimatter.

Theoretical Computer Science, 701(??): 161–173, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302840.

Li:2017:ICB

[1457] Yang Li, Andrei Paun, and Mihaela Paun. Improvements on contours based segmentation for DNA microarray image processing. Theoretical Computer Science, 701(??):174– 189, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304565.

Manca:2017:PIG

[1458] Vincenzo Manca. The principles of informational genomics. Theoretical Computer Science, 701(??):190– 202, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302761.

Mohammed:2017:CMK

[1459] Abdulmelik Mohammed, Elena Czeizler, and Eugen Czeizler. Computational modelling of the kinetic Tile Assembly Model using a rule-based approach. Theoretical Computer Science, 701(??):203– 215, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305522.

Perez-Jimenez:2017:FRS

[1460] Mario J. Pérez-Jiménez, Carmen Graciani, David Orellana-Martín, Agustín Riscos-Núñez, Álvaro Romero-Jiménez, and Luis Valencia-Cabrera. Fuzzy reasoning spiking neural P systems revisited: a formalization. Theoretical Computer Science, 701(??):216– 225, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304577.

Valencia-Cabrera:2017:RET

[1461] Luis Valencia-Cabrera, David Orellana-Martín, Miguel A. Martínez del Amor, Agustín Riscos-Núñez, and Mario J. Pérez-Jiménez. Reaching efficiency through collaboration in membrane systems: Dissolution, polarization and cooperation. Theoretical Computer Science, 701(??):226– 234, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304589.

Anonymous:2017:EBbt

[1462] Anonymous. Editorial Board. Theoretical Computer Science, 701(??): ifc, November 21, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730751X.

Borlido:2017:WP

[1463] Célia Borlido. The κ-word problem over DRH. Theoretical Computer Science, 702(??):1–22, November 30, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306102.

Downey:2017:LLD

[1464] Rod Downey, Michael McInerney, and Keng Meng Ng. Lowness and logical depth. Theoretical Computer Science, 702(??):23–33, November 30, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306126.

Walter:2017:PRC

[1465] Tobias Walter. Parikh-reducing Church–Rosser representations for some classes of regular languages. Theoretical Computer Science, 702(??): 34–47, November 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306114.

Asaeedi:2017:CHG

[1466] Saeed Asaeedi, Farzad Didehvar, and Ali Mohades. α-Concave hull, a generalization of convex hull. Theoretical Computer Science, 702(??):48– 59, November 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306230.

Antunes:2017:RDL

[1467] L. F. Antunes, A. Souto, and P. M. B. Vitányi. On the rate of decrease in logical depth. Theoretical Computer Science, 702(??):60– 64, November 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306217. See

corrigendum [2024].

Zhou:2017:TCT

[1468] Hao Zhou, Yiwei Jiang, Ping Zhou, Min Ji, and Yun Zhao. Total completion time minimization scheduling on two hierarchical uniform machines. Theoretical Computer Science, 702(??): 65–76, November 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306254.

Cabello:2017:ISS

[1469] Sergio Cabello and Pablo Pérez-Lantero. Interval selection in the streaming model. Theoretical Computer Science, 702(??):77– 96, November 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306242.

Anonymous:2017:EBbu

[1470] Anonymous. Editorial Board. Theoretical Computer Science, 702(??): ifc, November 30, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307818.

Bose:2017:GFS

[1471] Prosenjit Bose and Jean-Lou De Carufel. A general framework for searching on a line. Theoretical Computer Science, 703(??):1–17, December 5, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306394.

Glasser:2017:CSC

[1472] Christian Glaßer, Peter Jonsson, and Barnaby Martin. Circuit satisfiability and constraint satisfaction around Skolem Arithmetic. Theoretical Computer Science, 703(??): 18–36, December 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306485.

Jin:2017:MMM

[1473] Yujia Jin, Huan Li, and Zhongzhi Zhang. Maximum matchings and minimum dominating sets in Apollonian networks and extended Tower of Hanoi graphs. Theoretical Computer Science, 703(??):37–54, December 5, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306473.

Garg:2017:MEU

[1474] Jugal Garg. Market equilibrium under piecewise Leontief concave utilities. Theoretical Computer Science, 703(??): 55–65, December 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306497.

Li:2017:TAP

[1475] Dawei Li and Xiwen Lu. Twoagent parallel-machine scheduling with rejection. Theoretical Computer Science, 703(??):66–75, December 5, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306527.

Bulyovszky:2017:PFF

[1476] Balázs Bulyovszky and Gábor Horváth. Polynomial functions over finite commutative rings. Theoretical Computer Science, 703(??):76–86, December 5, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306503.

Anonymous:2017:EBbv

[1477] Anonymous. Editorial Board. Theoretical Computer Science, 703(??): ifc, December 5, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517307880.

Kang:2017:WPU

[1478] Dong Yeap Kang, O joung Kwon, Torstein J. F. Strømme, and Jan Arne Telle. A width parameter useful for chordal and co-comparability graphs. Theoretical Computer Science, 704(??): 1–17, December 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306618.

Abam:2017:FTS

[1479] M. A. Abam, F. Baharifard, M. S.

Borouny, and H. Zarrabi-Zadeh. Faulttolerant spanners in networks with symmetric directional antennas. Theoretical Computer Science, 704(??): 18–27, December 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306606.

Jayapaul:2017:FME

[1480] Varukumar Jayapaul, J. Ian Munro, Venkatesh Raman, and Srinivas Rao Satti. Finding modes with equality comparisons. Theoretical Computer Science, 704(??):28–41, December 15, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730662X.

Blanchet-Sadri:2017:CGB

[1481] F. Blanchet-Sadri and Sinziana Munteanu.

Computing generalized de Bruijn sequences. Theoretical Computer Science, 704(??):42–61, December 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397517306655.

Liu:2017:ECD

[1482] Aixia Liu, Shiying Wang, Jun Yuan, and Jing Li. On g-extra conditional diagnosability of hypercubes and folded hypercubes. Theoretical Computer Science, 704(??):62– 73, December 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306989.

Bunder:2017:GAR

[1483] Martin Bunder, Abderrahmane Nitaj, Willy Susilo, and Joseph Tonien. A generalized attack on RSA type cryptosystems. Theoretical Computer Science, 704(??):74–81, December 15, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306643.

Li:2017:FTS

[1484] Xiang-Jun Li, Yong-Ni Guan, Zheng Yan, and Jun-Ming Xu. On fault tolerance of (n, k)-star networks. Theoretical Computer Science, 704(??): 82–86, December 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305984.

Sepasian:2017:UMM

[1485] Ali Reza Sepasian and Ehsan Monabbati. Upgrading min–max spanning tree problem under various cost functions. Theoretical Computer Science, 704(??):87–91, December 15, 2017. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306011.

Penso:2017:CCA

[1486] Lucia Draque Penso, Fábio Protti, Dieter Rautenbach, and Uéverton dos Santos Souza. Corrigendum to “Complexity analysis of P_{3}-convexity problems on bounded-degree and planar graphs” [Theoret. Comput. Sci. 607 Part 1 (2015) 83–95]. Theoretical Computer Science, 704(??):92– 93, December 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306229. See

[521].

Anonymous:2017:EBbw

[1487] Anonymous. Editorial Board. Theoretical Computer Science, 704(??): ifc, December 15, 2017. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307995.

Leiss:2018:WPH

[1488] Ernst L. Leiss and Isaac Mackey. The worst parallel Hanoi graphs. Theoretical Computer Science, 705(??): 1–8, January 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306746.

Shibata:2018:MOP

[1489] Masahiro Shibata, Fukuhito Ooshita, Hirotsugu Kakugawa, and Toshimitsu Masuzawa. Move-optimal partial gathering of mobile agents in asynchronous trees. Theoretical Computer Science, 705(??):9–30, January 1, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306783.

Hackl:2018:RBT

[1490] Benjamin Hackl, Clemens Heuberger, and Helmut Prodinger. Reductions of binary trees and lattice paths induced by the register function. Theoretical Computer Science, 705(??): 31–57, January 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306771.

Yu:2018:MSC

[1491] Hung-I Yu, Cheng-Chung Li, and D. T. Lee. The multi-service center problem. Theoretical Computer Science, 705(??): 58–74, January 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307028.

Chiarelli:2018:MCT

[1492] Nina Chiarelli, Tatiana R. Hartinger, Matthew Johnson, Martin Milanic, and Daniël Paulusma. Minimum connected transversals in graphs: New hardness results and tractable cases using the price of connectivity. Theoretical Computer Science, 705(??): 75–83, January 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307016.

Kobayashi:2018:ILB

[1493] Koji M. Kobayashi. Improved lower bounds for online scheduling to minimize total stretch. Theoretical Computer Science, 705(??): 84–98, January 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307004.

Greenberg:2018:DSC

[1494] Noam Greenberg, Joseph S. Miller, Alexander Shen, and Linda Brown Westrick. Dimension 1 sequences are close to randoms. Theoretical Computer Science, 705(??):99– 112, January 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306990.

Chatterjee:2018:NRU

[1495] Kingshuk Chatterjee and Kumar Sankar Ray. Non-regular unary language and parallel communicating Watson-Crick automata systems. Theoretical Computer Science, 705(??):113– 117, January 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306631.

Ning:2018:ECC

[1496] Wantao Ning. Erratum and corrections to “The h-connectivity of exchanged crossed cube” [Theoret. Comput. Sci. 696 (2017) 65–68]. Theoretical Computer Science, 705(??): 118–121, January 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730703X. See

[1406].

Anonymous:2018:EBa

[1497] Anonymous. Editorial Board. Theoretical Computer Science, 705(??): ifc, January 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308101.

Ebrahimi:2018:SSI

[1498] M. M. Ebrahimi, Kh. Keshvardoost, and M. Mahmoudi. Simple and subdirectly irreducible finitely supported Cb-sets. Theoretical Computer Science, 706(??):1–21, January 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517305844.

Zhao:2018:FFO

[1499] Chunna Zhao and Shanshan Li. Formalization of fractional order PD control systems in HOL4. Theoretical Computer Science, 706(??): 22–34, January 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306205.

Dimovski:2018:VAP

[1500] Aleksandar S. Dimovski. Verifying annotated program families using symbolic game semantics. Theoretical Computer Science, 706(??):

35–53, January 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306977.

Baldan:2018:MTI

[1501] Paolo Baldan and Alessandro Beggiato. Multilevel transitive and intransitive non-interference, causally. Theoretical Computer Science, 706(??): 54–82, January 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307090.

Leitsch:2018:PCI

[1502] Alexander Leitsch and Michael Lettmann. The problem of Π_{2}-cut-introduction. Theoretical Computer Science, 706(??):83–116, January 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307089.

Anonymous:2018:EBb

[1503] Anonymous. Editorial Board. Theoretical Computer Science, 706(??): ifc, January 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308216.

Anonymous:2018:EBc

[1504] Anonymous. Editorial Board. Theoretical Computer Science, 707(??): ii, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308745.

Kayaaslan:2018:SSP

[1505] Enver Kayaaslan, Thomas Lambert, Loris Marchal, and Bora Uçar. Scheduling series-parallel task graphs to minimize peak memory. Theoretical Computer Science, 707(??): 1–23, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307053.

Godin:2018:BMA

[1506] Thibault Godin and Ines Klimann. On bireversible Mealy automata and the Burnside problem. Theoretical Computer Science, 707(??):24– 35, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307107.

Alizadeh:2018:OAI

[1507] Behrooz Alizadeh and Roghayeh Etemad. Optimal algorithms for inverse vertex obnoxious center location problems on graphs. Theoretical Computer Science, 707(??):36– 45, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307065.

Alipour:2018:RAA

[1508] Sharareh Alipour, Mohammad Ghodsi, and Amir Jafari. Randomized approximation algorithms for planar visibility counting problem. Theoretical Computer Science, 707(??): 46–55, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307375.

Li:2018:FTS

[1509] Pingshan Li and Min Xu. Faulttolerant strong Menger (edge) connectivity and 3-extra edge-connectivity of balanced hypercubes. Theoretical Computer Science, 707(??):56– 68, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307454.

Martins:2018:LIC

[1510] Nicolas Martins and Rudini Sampaio. Locally identifying coloring of graphs with few P4s. Theoretical Computer Science, 707(??):69– 76, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307430.

Karpinski:2018:ECC

[1511] Michal Karpiński. Encoding cardinality constraints using standard encoding of generalized selection networks preserves arc-consistency. Theoretical Computer Science, 707(??): 77–81, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307041.

Zhang:2018:R

[1512] Jie-Meng Zhang, Ying-Jun Guo, and Zhi-Xiong Wen. On the regularity of {⌊log_{b}(αn + β)⌋}_{n≥0}. Theoretical Computer Science, 707(??): 82–88, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517307387.

Bienkowski:2018:LPB

[1513] Marcin Bienkowski, Martin Böhm, Łukasz Jeż, Paweł Laskoś-Grabowski, Jan Marcinkowski, Jirí Sgall, Aleksandra Spyra, and Pavel Veselý. Logarithmic price of buffer downscaling on line metrics. Theoretical Computer Science, 707(??):89– 93, January 10, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517307363.

Li:2018:TPA

[1514] Wangwei Li and Xianhua Li. On two problems of almost synchronizing groups. Theoretical Computer Science, 707(??):94–95, January 10, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517307351.

Guo:2018:RAT

[1515] Litao Guo. Reliability analysis of twisted cubes. Theoretical Computer Science, 707(??):96–101, January 10, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307442.

Anonymous:2018:EBd

[1516] Anonymous. Editorial Board. Theoretical Computer Science, 708(??): ii, January 17, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308836.

Marcilon:2018:MTN

[1517] Thiago Marcilon and Rudini Sampaio.

The maximum time of 2-neighbor bootstrap percolation: Complexity results. Theoretical Computer Science, 708(??): 1–17, January 17, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307429.

Rai:2018:BCA

[1518] Ashutosh Rai and Saket Saurabh.

Bivariate complexity analysis of Almost Forest Deletion. Theoretical Computer Science, 708(??):18– 33, January 17, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307491.

Mahajan:2018:SRO

[1519] Meena Mahajan and Anuj Tawari.

Sums of read-once formulas: How many summands are necessary? Theoretical Computer Science, 708(??): 34–45, January 17, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307478.

Ma:2018:IMC

[1520] Fei Ma and Bing Yao. An iteration method for computing the total number of spanning trees and its applications in graph theory. Theoretical Computer Science, 708(??): 46–57, January 17, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730779X.

Talmon:2018:SPR

[1521] Nimrod Talmon. Structured proportional representation. Theoretical Computer Science, 708(??):58– 74, January 17, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307776.

Tseng:2018:EML

[1522] Kuo-Tsung Tseng, De-Sheng Chan, Chang-Biau Yang, and Shou-Fu Lo. Efficient merged longest common subsequence algorithms for similar sequences. Theoretical Computer Science, 708(??):75–90, January 17, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307764.

Litow:2018:TMB

[1523] B. Litow. A transfer method from bounded existential Diophantine equations to Tarski algebra formulas. Theoretical Computer Science, 708(??): 91–95, January 17, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307417.

Mikos:2018:NLB

[1524] Patryk Mikos. A new lower bound for the on-line coloring of intervals with bandwidth. Theoretical Computer Science, 708(??):96– 100, January 17, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307788.

Anonymous:2018:EBe

[1525] Anonymous. Editorial Board. Theoretical Computer Science, 709(??): ii, January 24, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308952.

Pandurangan:2018:SII

[1526] Gopal Pandurangan and Peter Robinson. Special issue of ICDCN 2016 (Distributed Computing Track). Theoretical Computer Science, 709(??): 1, January 24, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308988.

Castaneda:2018:CRM

[1527] Armando Castañeda, Danny Dolev, and Amitabh Trehan. Compact routing messages in self-healing trees. Theoretical Computer Science, 709(??):

2–19, January 24, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306818.

Czyzowicz:2018:ETR

[1528] Jurek Czyzowicz, Stefan Dobrev, Konstantinos Georgiou, Evangelos Kranakis, and Fraser MacQuarrie. Evacuating two robots from multiple unknown exits in a circle. Theoretical Computer Science, 709(??): 20–30, January 24, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306788.

He:2018:GME

[1529] Yuan He, K. Gopalakrishnan, and Eli Gafni. Group mutual exclusion in linear time and space. Theoretical Computer Science, 709(??): 31–47, January 24, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304747.

Foerster:2018:LCN

[1530] Klaus-Tycho Foerster, Thomas Luedi, Jochen Seidel, and Roger Wattenhofer. Local checkability, no strings attached: (a)cyclicity, reachability, loop free updates in SDNs. Theoretical Computer Science, 709(??): 48–63, January 24, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306776.

Bonomi:2018:OSS

[1531] Silvia Bonomi, Antonella Del Pozzo, and Maria Potop-Butucaru. Optimal self-stabilizing synchronous mobile Byzantine-tolerant atomic register. Theoretical Computer Science, 709 (??):64–79, January 24, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306291.

Mostefaoui:2018:RSA

[1532] Achour Mostéfaoui, Hamouma Moumen, and Michel Raynal. Randomized k-set agreement in crash-prone and Byzantine asynchronous systems. Theoretical Computer Science, 709(??): 80–97, January 24, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302311.

Anonymous:2018:EBf

[1533] Anonymous. Editorial Board. Theoretical Computer Science, 710(??): ii, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300100.

Crochemore:2018:AAC

[1534] Maxime Crochemore and Solon P. Pissis. Advances in algorithms & combinatorics on strings (honoring 60th birthday for Prof. Costas S. Iliopoulos). Theoretical Computer Science, 710(??):1, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300069.

Amir:2018:PRS

[1535] Amihood Amir, Mika Amit, Gad M. Landau, and Dina Sokol. Period recovery of strings over the Hamming and edit distances. Theoretical Computer Science, 710(??): 2–18, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517307752.

Aronica:2018:OPL

[1536] Salvatore Aronica, Alessio Langiu, Francesca Marzi, Salvatore Mazzola, and Filippo Mignosi. On optimal parsing for LZ78-like compressors. Theoretical Computer Science, 710(??): 19–28, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301524.

Bae:2018:FLC

[1537] Sang Won Bae and Inbok Lee. On finding a longest common palindromic subsequence. Theoretical Computer Science, 710(??):29–34, February 1, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301512.

Cho:2018:TAB

[1538] Sukhyeun Cho, Joong Chae Na, and Jeong Seop Sim. An O(n^{2} log m)-time algorithm for the boxed-mesh permutation pattern matching problem. Theoretical Computer Science, 710(??): 35–43, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730155X.

Daykin:2018:RSL

[1539] Jacqueline W. Daykin, Frantisek Franek, Jan Holub, A. S. M. Sohidull Islam, and W. F. Smyth. Reconstructing a string from its Lyndon arrays. Theoretical Computer Science, 710(??): 44–51, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303304.

Daykin:2018:SSO

[1540] Jacqueline W. Daykin, Richard Groult, Yannick Guesnet, Thierry Lecroq, Arnaud Lefebvre, Martine Léonard, and Élise Prieur-Gaston. A survey of string orderings and their application to the Burrows–Wheeler transform. Theoretical Computer Science, 710(??): 52–65, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301548.

Gawrychowski:2018:NCN

[1541] Paweł Gawrychowski, Gad M. Landau, Shay Mozes, and Oren Weimann. The nearest colored node in a tree. Theoretical Computer Science, 710(??): 66–73, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306308.

Grossi:2018:MTE

[1542] Roberto Grossi, Giulia Menconi, Nadia Pisanti, Roberto Trani, and Søren Vind. Motif trie: an efficient text index for pattern discovery with don’t cares. Theoretical Computer Science, 710(??): 74–87, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517303523.

Helling:2018:CIS

[1543] Joel Helling, P. J. Ryan, W. F. Smyth, and Michael Soltys. Constructing an indeterminate string from its associated graph. Theoretical Computer Science, 710(??): 88–96, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301494.

Hoffmann:2018:ENL

[1544] Michael Hoffmann, John Iacono, Patrick K. Nicholson, and Rajeev Raman. Encoding nearest larger values. Theoretical Computer Science, 710(??): 97–115, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301500.

Iakovidou:2018:SDA

[1545] Nantia D. Iakovidou, Nikos A. Laskaris, Costas Tsichlas, Yannis Manolopoulos, Manolis Christodoulakis, Eleftherios S. Papathanasiou, Savvas S. Papacostas, and Georgios D. Mitsis. A symbolic dynamics approach to Epileptic Chronnectomics: Employing strings to predict crisis onset. Theoretical Computer Science, 710(??): 116–125, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517301561.

Kociumaka:2018:SCP

[1546] Tomasz Kociumaka, Jakub W. Pachocki, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. On the string consensus problem and the Manhattan sequence consensus problem. Theoretical Computer Science, 710(??): 126–138, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302359.

Kociumaka:2018:EAS

[1547] Tomasz Kociumaka, Solon P. Pissis, Jakub Radoszewski, Wojciech Rytter, and Tomasz Waleń. Efficient algorithms for shortest partial seeds in words. Theoretical Computer Science, 710(??):139–147, February 1, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307034.

Na:2018:FIA

[1548] Joong Chae Na, Hyunjoon Kim, Seunghwan Min, Heejin Park, Thierry Lecroq, Martine Léonard, Laurent Mouchard, and Kunsoo Park. FMindex of alignment with gaps. Theoretical Computer Science, 710(??): 148–157, February 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517301536.

Anonymous:2018:EBg

[1549] Anonymous. Editorial Board. Theoretical Computer Science, 711(??): ii, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300239.

Guo:2018:PSC

[1550] Yuyan Guo, Jiguo Li, Yang Lu, Yichen Zhang, and Futai Zhang. Provably secure certificate-based encryption with leakage resilience. Theoretical Computer Science, 711(??): 1–10, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730748X.

Mozes:2018:FSP

[1551] Shay Mozes, Yahav Nussbaum, and Oren Weimann. Faster shortest paths in dense distance graphs, with applications. Theoretical Computer Science, 711(??):11–35, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397517307971.

Hu:2018:SMP

[1552] Xiaomin Hu, Yingzhi Tian, Xiaodong Liang, and Jixiang Meng. Strong matching preclusion for kcomposition networks. Theoretical Computer Science, 711(??):36– 43, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730796X.

Sabir:2018:SFT

[1553] Eminjan Sabir and Jixiang Meng.

Structure fault tolerance of hypercubes and folded hypercubes. Theoretical Computer Science, 711(??): 44–55, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307958.

Karpinski:2018:QBN

[1554] Marek Karpinski, Andrzej Lingas, and Dzmitry Sledneu. A QPTAS for the base of the number of crossing-free structures on a planar point set. Theoretical Computer Science, 711(??): 56–65, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308307.

Garnero:2018:FIC

[1555] Valentin Garnero, Konstanty Junosza-Szaniawski, Mathieu Liedloff, Pedro Montealegre, and Paweł Rza̧żewski. Fixing improper colorings of graphs. Theoretical Computer Science, 711(??): 66–78, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308435.

Ebrahimnejad:2018:GBS

[1556] Farzam Ebrahimnejad. On the gap between separating words and separating their reversals. Theoretical Computer Science, 711(??):79– 91, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308423.

Seebold:2018:SIN

[1557] Patrice Séébold. Sturmian images of non Sturmian words and standard morphisms. Theoretical Computer Science, 711(??):92–104, February 8, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308411.

Kantor:2018:TWC

[1558] Erez Kantor, Zvi Lotker, Merav Parter, and David Peleg. The topology of wireless communication on a line. Theoretical Computer Science, 711(??): 105–108, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517307077.

Masopust:2018:SPT

[1559] Tomáš Masopust. Separability by piecewise testable languages is PTimecomplete. Theoretical Computer Science, 711(??):109–114, February 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397517308319.

Anonymous:2018:EBh

[1560] Anonymous. Editorial Board. Theoretical Computer Science, 712(??): ii, February 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300367.

Lipton:2018:KSH

[1561] James Lipton and Susana Nieva.

Kripke semantics for higher-order type theory applied to constraint logic programming languages. Theoretical Computer Science, 712(??):1– 37, February 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308356.

Kalorkoti:2018:SLT

[1562] K. Kalorkoti. Solving Łukasiewicz μ-terms. Theoretical Computer Science, 712(??):38–49, February 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397517308289.

Avni:2018:SCL

[1563] Guy Avni and Orna Kupferman.

Synthesis from component libraries with costs. Theoretical Computer Science, 712(??):50–72, February 15, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308277.

Anonymous:2018:EBi

[1564] Anonymous. Editorial Board. Theoretical Computer Science, 713(??): ii, February 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300471.

Jones:2018:CGL

[1565] Átila A. Jones, Fábio Protti, and Renata R. Del-Vecchio. Cograph generation with linear delay. Theoretical Computer Science, 713(??): 1–10, February 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300033.

Mao:2018:SMP

[1566] Yaping Mao, Zhao Wang, Eddie Cheng, and Christopher Melekian. Strong matching preclusion number of graphs. Theoretical Computer Science, 713(??): 11–20, February 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830001X.

Friedman:2018:DNG

[1567] Eric J. Friedman and Adam S. Landsberg. Duality and nonlinear graph Laplacians. Theoretical Computer Science, 713(??):21– 30, February 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309295.

Rutanen:2018:MCN

[1568] Kalle Rutanen. Minimal characterization of O-notation in algorithm analysis. Theoretical Computer Science, 713 (??):31–41, February 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309210.

You:2018:SSC

[1569] Lantao You, Jianxi Fan, and Yuejuan Han. Super spanning connectivity on WK-recursive networks. Theoretical Computer Science, 713(??): 42–55, February 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309180.

Goeckner:2018:UPW

[1570] Bennet Goeckner, Corbin Groothuis, Cyrus Hettle, Brian Kell, Pamela Kirkpatrick, Rachel Kirsch, and Ryan Solava. Universal partial words over non-binary alphabets. Theoretical Computer Science, 713(??):56– 65, February 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309179.

Bille:2018:TST

[1571] Philip Bille, Mikko Berggren Ettienne, Inge Li Gørtz, and Hjalte Wedel Vildhøj. Time-space trade-offs for Lempel–Ziv compressed indexing. Theoretical Computer Science, 713(??): 66–77, February 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309167.

DiGiacomo:2018:NRE

[1572] Emilio Di Giacomo, Walter Didimo, William S. Evans, Giuseppe Liotta, Henk Meijer, Fabrizio Montecchiani, and Stephen K. Wismath. New results on edge partitions of 1-plane graphs. Theoretical Computer Science, 713(??): 78–84, February 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309192.

Anonymous:2018:EBj

[1573] Anonymous. Editorial Board. Theoretical Computer Science, 714(??): ii, March 1, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300562.

Kayibi:2018:TTM

[1574] K. K. Kayibi and S. Pirzada. T - tetrominoes tiling’s Markov chain mixes fast. Theoretical Computer Science, 714(??):1–14, March 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397517309155.

Lai:2018:EOT

[1575] Jianchang Lai, Yi Mu, Fuchun Guo, Rongmao Chen, and Sha Ma. Efficient k-out-of-n oblivious transfer scheme with the ideal communication cost. Theoretical Computer Science, 714(??):15–26, March 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309143.

Cao:2018:OOD

[1576] Huabin Cao, Bicheng Zhang, and Zhiheng Zhou. One-to-one disjoint path covers in digraphs. Theoretical Computer Science, 714(??): 27–35, March 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308708.

Docker:2018:ECP

[1577] Janosch Döcker and Simone Linz. On the existence of a cherry-picking sequence. Theoretical Computer Science, 714(??):36–50, March 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308691.

DiGiacomo:2018:DSP

[1578] Emilio Di Giacomo, Giuseppe Liotta, and Fabrizio Montecchiani. Drawing subcubic planar graphs with four slopes and optimal angular resolution. Theoretical Computer Science, 714(??):51–73, March 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730868X.

Gregor:2018:TGG

[1579] Petr Gregor and Torsten Mütze. Trimming and gluing Gray codes. Theoretical Computer Science, 714(??): 74–95, March 1, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308678.

Anonymous:2018:EBk

[1580] Anonymous. Editorial Board. Theoretical Computer Science, 715(??): ii, March 8, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300732.

LeGuiban:2018:CPL

[1581] Kaourintin Le Guiban, Arpad Rimmel, Marc-Antoine Weisser, and Joanna Tomasik. Completion of partial latin hypercube designs: NPcompleteness and inapproximability. Theoretical Computer Science, 715 (??):1–20, March 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300409.

Jonsson:2018:TCN

[1582] Peter Jonsson and Johan Thapper. Tractability conditions for numeric CSPs. Theoretical Computer Science, 715(??):21–34, March 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300392.

Lintzmayer:2018:SPB

[1583] Carla Negri Lintzmayer, Guillaume Fertin, and Zanoni Dias. Sorting permutations and binary strings by length-weighted rearrangements. Theoretical Computer Science, 715(??):

35–59, March 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830032X.

Yoon:2018:GMA

[1584] Sang Duk Yoon, Min-Gyu Kim, Wanbin Son, and Hee-Kap Ahn. Geometric matching algorithms for two realistic terrains. Theoretical Computer Science, 715(??):60– 70, March 8, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300318.

Duchene:2018:SOP

[1585] Eric Duchêne, Marc Heinrich, Urban Larsson, and Aline Parreau. The switch operators and push-the-button games: A sequential compound over rulesets. Theoretical Computer Science, 715(??):71–85, March 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300148.

Capdevielle:2018:CBA

[1586] Claire Capdevielle, Colette Johnen, and Alessia Milani. On the complexity of basic abstractions to implement consensus. Theoretical Computer Science, 715(??):86–96, March 8, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300057.

Anonymous:2018:EBl

[1587] Anonymous. Editorial Board. Theoretical Computer Science, 716(??): ii, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300902.

Chaudhuri:2018:SIA

[1588] Kamalika Chaudhuri and Claudio Gentile. Special issue on ALT 2015: Guest Editors’ introduction. Theoretical Computer Science, 716(??): 1–3, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300860.

Hermo:2018:ELM

[1589] Montserrat Hermo and Ana Ozaki. Exact learning of multivalued dependency formulas. Theoretical Computer Science, 716(??):4–14, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397517308484.

Abasi:2018:NAL

[1590] Hasan Abasi, Nader H. Bshouty, and Hanna Mazzawi. Non-adaptive learning of a hidden hypergraph. Theoretical Computer Science, 716(??): 15–27, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308496.

Leike:2018:CSI

[1591] Jan Leike and Marcus Hutter. On the computability of Solomonoff induction and AIXI. Theoretical Computer Science, 716(??):28– 49, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308502.

Orabona:2018:SFO

[1592] Francesco Orabona and Dávid Pál.

Scale-free online learning. Theoretical Computer Science, 716(??):50– 69, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308514.

Darnstadt:2018:HDF

[1593] Malte Darnstädt, Christoph Ries, and Hans Ulrich Simon. Hierarchical design of fast Minimum Disagreement algorithms. Theoretical Computer Science, 716(??):70–88, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308526.

Balle:2018:GBL

[1594] Borja Balle and Mehryar Mohri.

Generalization bounds for learning weighted automata. Theoretical Computer Science, 716(??):89– 106, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308538.

Gao:2018:TCL

[1595] Ziyuan Gao, Hans Ulrich Simon, and Sandra Zilles. On the teaching complexity of linear sets. Theoretical Computer Science, 716(??): 107–123, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730854X.

Yang:2018:BMR

[1596] Liu Yang, Steve Hanneke, and Jaime Carbonell. Bounds on the minimax rate for estimating a prior over a VC class from independent learning tasks. Theoretical Computer Science, 716(??):124–140, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308551.

Anonymous:2018:EBm

[1597] Anonymous. Editorial Board. Theoretical Computer Science, 717(??): ii, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301051.

Dondi:2018:E

[1598] Riccardo Dondi, Guillaume Fertin, and Giancarlo Mauri. Editorial. Theoretical Computer Science, 717 (??):1, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301014.

Bazgan:2018:MFU

[1599] Cristina Bazgan, Ljiljana Brankovic, Katrin Casel, Henning Fernau, Klaus Jansen, Kim-Manuel Klein, Michael Lampis, Mathieu Liedloff, Jérôme Monnot, and Vangelis Th. Paschos. The many facets of upper domination. Theoretical Computer Science, 717(??):2–25, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305145.

Bredereck:2018:PCT

[1600] Robert Bredereck, Jiehua Chen, Falk Hüffner, and Stefan Kratsch. Parameterized complexity of team formation in social networks. Theoretical Computer Science, 717(??): 26–36, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304644.

Goldwurm:2018:CCR

[1601] Massimiliano Goldwurm, Jianyi Lin, and Francesco Saccà. On the complexity of clustering with relaxed size constraints in fixed dimension. Theoretical Computer Science, 717 (??):37–46, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517304619.

Georgiou:2018:KWP

[1602] Konstantinos Georgiou, George Karakostas, Evangelos Kranakis, and Danny Krizanc. Know when to persist: Deriving value from a stream buffer.

Theoretical Computer Science, 717 (??):47–61, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517304607.

Kaklamanis:2018:NFG

[1603] Christos Kaklamanis, Panagiotis Kanellopoulos, and Sophia Tsokana. On network formation games with heterogeneous players and basic network creation games. Theoretical Computer Science, 717(??):62– 72, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302803.

Bellitto:2018:SCT

[1604] Thomas Bellitto. Separating codes and traffic monitoring. Theoretical Computer Science, 717(??):73– 85, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517302839.

Anonymous:2018:EBn

[1605] Anonymous. Editorial Board. Theoretical Computer Science, 718(??): ii, March 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301129.

Petreschi:2018:P

[1606] Rossella Petreschi and Mohammad Kaykobad. Preface. Theoretical Computer Science, 718(??):1, March 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301154.

Bae:2018:CMW

[1607] Sang Won Bae. Computing a minimum-width square annulus in arbitrary orientation. Theoretical Computer Science, 718(??):2–13, March 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306053.

Bae:2018:CPC

[1608] Sang Won Bae, Hwan-Gue Cho, William Evans, Noushin Saeedi, and Chan-Su Shin. Covering points with convex sets of minimum size. Theoretical Computer Science, 718(??): 14–23, March 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516306090.

Czibula:2018:LRV

[1609] Oliver G. Czibula, Hanyu Gu, and Yakov Zinder. Lagrangian relaxation versus genetic algorithm based metaheuristic for a large partitioning problem. Theoretical Computer Science, 718(??):24– 36, March 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517300348.

Gragera:2018:RTI

[1610] Alonso Gragera and Vorapong Suppakitpaisarn. Relaxed triangle inequality ratio of the Sørensen–Dice and Tversky indexes. Theoretical Computer Science, 718(??):37– 45, March 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300361.

Calamoneri:2018:DTG

[1611] Tiziana Calamoneri, Angelo Monti, and Rossella Petreschi. On dynamic threshold graphs and related classes. Theoretical Computer Science, 718(??): 46–57, March 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300397.

Hasegawa:2018:OOA

[1612] Shun Hasegawa and Toshiya Itoh. Optimal online algorithms for the multiobjective time series search problem. Theoretical Computer Science, 718(??): 58–66, March 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517300403.

Anonymous:2018:EBo

[1613] Anonymous. Editorial Board. Theoretical Computer Science, 719(??): ii, April 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830135X.

Greenhill:2018:SMC

[1614] Catherine Greenhill and Matteo Sfragara. The switch Markov chain for sampling irregular graphs and digraphs. Theoretical Computer Science, 719 (??):1–20, April 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730840X.

Schridde:2018:CGP

[1615] Christian Schridde. On the construction of graphs with a planar bipartite double cover from boolean formulas and its application to counting satisfying solutions. Theoretical Computer Science, 719(??):21– 45, April 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308393.

Golovnev:2018:GEC

[1616] Alexander Golovnev, Alexander S. Kulikov, Alexander V. Smal, and Suguru Tamaki. Gate elimination: Circuit size lower bounds and #SAT upper bounds. Theoretical Computer Science, 719(??):46–63, April 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308381.

Bang-Jensen:2018:DRP

[1617] J. Bang-Jensen, S. Bessy, F. Havet, and A. Yeo. Out-degree reducing partitions of digraphs. Theoretical Computer Science, 719(??): 64–72, April 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730837X.

Barash:2018:LSR

[1618] Mikhail Barash and Alexander Okhotin. Linear-space recognition for grammars with contexts. Theoretical Computer Science, 719(??):73– 85, April 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308368.

Teh:2018:SUP

[1619] Wen Chean Teh, Adrian Atanasiu, and Ghajendran Poovanandran. On strongly M-unambiguous prints and Serbănută’s conjecture for Parikh matrices. Theoretical Computer Science, 719(??):86–93, April 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308460.

Anonymous:2018:EBp

[1620] Anonymous. Editorial Board. Theoretical Computer Science, 720(??): ii, April 11, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301518.

Sima:2018:QPE

[1621] Jirí Síma and Petr Savický. Quasiperiodic β-expansions and cut languages. Theoretical Computer Science, 720(??):1–23, April 11, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301257.

Sultana:2018:RGT

[1622] Shaheena Sultana and Md. Saidur Rahman. Realizability of graphs as triangle cover contact graphs. Theoretical Computer Science, 720(??): 24–35, April 11, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301245.

Boyar:2018:MCV

[1623] Joan Boyar and Magnus Gausdal Find. Multiplicative complexity of vector valued Boolean functions. Theoretical Computer Science, 720(??): 36–46, April 11, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301208.

Shan:2018:INN

[1624] Liren Shan, Huan Li, and Zhongzhi Zhang. Independence number and the number of maximum independent sets in pseudofractal scale-free web and Sierpiński gasket. Theoretical Computer Science, 720(??): 47–54, April 11, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301191.

Repke:2018:SPB

[1625] Damian Repke and Wojciech Rytter. On semi-perfect de Bruijn words. Theoretical Computer Science, 720 (??):55–63, April 11, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300859.

Furst:2018:LSL

[1626] Maximilian Fürst, Marilena Leichter, and Dieter Rautenbach. Locally searching for large induced matchings. Theoretical Computer Science, 720(??): 64–72, April 11, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300835.

Anonymous:2018:EBq

[1627] Anonymous. Editorial Board. Theoretical Computer Science, 721(??): ii, April 18, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301579.

Moulton:2018:UNE

[1628] Vincent Moulton, Andreas Spillner, and Taoyang Wu. UPGMA and the normalized equidistant minimum evolution problem. Theoretical Computer Science, 721(??): 1–15, April 18, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300690.

Cheng:2018:BCM

[1629] Chu-Han Cheng, Po-An Chen, and Wing-Kai Hon. Budget-constrained multi-battle contests: A new perspective and analysis. Theoretical Computer Science, 721(??):16– 26, April 18, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300653.

Liu:2018:TAP

[1630] Zhiqiang Liu, Yukiko Yamauchi, Shuji Kijima, and Masafumi Yamashita. Team assembling problem for asynchronous heterogeneous mobile robots. Theoretical Computer Science, 721 (??):27–41, April 18, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830029X.

Bezrukov:2018:NIF

[1631] Sergei Bezrukov, Pavle Bulatovic, and Nikola Kuzmanovski. New infinite family of regular edge-isoperimetric graphs. Theoretical Computer Science, 721(??):42–53, April 18, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300021.

Altman:2018:ISA

[1632] Harry Altman. Internal structure of addition chains: Well-ordering. Theoretical Computer Science, 721(??): 54–69, April 18, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308666.

Ouazene:2018:IPM

[1633] Yassine Ouazene and Farouk Yalaoui.

Identical parallel machine scheduling with time-dependent processing times. Theoretical Computer Science, 721 (??):70–77, April 18, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308654.

Anonymous:2018:EBr

[1634] Anonymous. Editorial Board. Theoretical Computer Science, 722(??): ii, April 25, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301634.

Bizjak:2018:MGR

[1635] Ales Bizjak and Lars Birkedal. A model of guarded recursion via generalised equilogical spaces. Theoretical Computer Science, 722(??): 1–18, April 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300951.

Bravetti:2018:BBD

[1636] Mario Bravetti, Marco Carbone, and Gianluigi Zavattaro. On the boundary between decidability and undecidability of asynchronous session subtyping. Theoretical Computer Science, 722 (??):19–51, April 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300938.

Zhang:2018:HVU

[1637] Min Zhang and Kazuhiro Ogata. From hidden to visible: A unified framework for transforming behavioral theories into rewrite theories. Theoretical Computer Science, 722(??): 52–75, April 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300173.

Anonymous:2018:EBs

[1638] Anonymous. Editorial Board. Theoretical Computer Science, 723(??): ii, May 2, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301877.

Adil:2018:PAS

[1639] Deeksha Adil, Sushmita Gupta, Sanjukta Roy, Saket Saurabh, and Meirav Zehavi. Parameterized algorithms for stable matching with ties and incomplete lists. Theoretical Computer Science, 723(??):1– 10, May 2, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301762.

Kolpakov:2018:NGR

[1640] Roman Kolpakov. On the number of gapped repeats with arbitrary gap. Theoretical Computer Science, 723 (??):11–22, May 2, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301464.

Angelini:2018:FBP

[1641] Patrizio Angelini, Michael A. Bekos, Michael Kaufmann, Philipp Kindermann, and Thomas Schneck. 1fan-bundle-planar drawings of graphs. Theoretical Computer Science, 723 (??):23–50, May 2, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301440.

Canard:2018:NTC

[1642] S. Canard, D. H. Phan, D. Pointcheval, and V. C. Trinh. A new technique for compacting ciphertext in multi-channel broadcast encryption and attribute-based encryption. Theoretical Computer Science, 723(??): 51–72, May 2, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301427.

Yang:2018:OBM

[1643] Yi-Hua Yang, Chung-Shou Liao, Xin Han, and Louxin Zhang. Online buffer management for transmitting packets with processing cycles. Theoretical Computer Science, 723(??): 73–83, May 2, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301385.

Anonymous:2018:EBt

[1644] Anonymous. Editorial Board. Theoretical Computer Science, 724(??): ii, May 9, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301932.

Kari:2018:Ea

[1645] Lila Kari.  Editorial.  Theo-

retical Computer Science, 724(??): iii, May 9, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302020.

Martinez-del-Amor:2018:P

[1646] Miguel Á. Martínez del Amor, Agustín Riscos-Núñez, and Luis Valencia-Cabrera. Preface. Theoretical Computer Science, 724(??):1– 2, May 9, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730912X.

Zhang:2018:CPE

[1647] Zhiqiang Zhang, Yansen Su, and Linqiang Pan. The computational power of enzymatic numerical P systems working in the sequential mode. Theoretical Computer Science, 724 (??):3–12, May 9, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309088.

Wu:2018:SNS

[1648] Tingfang Wu, Yanfeng Wang, Suxia Jiang, Yansen Su, and Xiaolong Shi. Spiking neural P systems with rules on synapses and anti-spikes. Theoretical Computer Science, 724(??): 13–27, May 9, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309076.

Pan:2018:LGA

[1649] Linqiang Pan, Bosheng Song, Atulya K. Nagar, and K. G. Subramanian. Language generating alphabetic flat splicing P systems. Theoretical Computer Science, 724(??): 28–34, May 9, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309064.

Kantor:2018:CLC

[1650] Kristóf Kántor and György Vaszil. On the classes of languages characterized by generalized P colony automata. Theoretical Computer Science, 724(??):35–44, May 9, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309027.

Gheorghe:2018:KSM

[1651] Marian Gheorghe, Rodica Ceterchi, Florentin Ipate, Savas Konur, and Raluca Lefticaru. Kernel P systems: From modelling to verification and testing. Theoretical Computer Science, 724(??):45–60, May 9, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309015.

Ciencialova:2018:GCP

[1652] Lucie Ciencialová, Ludek Cienciala, and Petr Sosík. Generalized P colonies with passive environment. Theoretical Computer Science, 724 (??):61–68, May 9, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730899X.

Wu:2018:SRA

[1653] Zijun Wu, Rolf H. Möhring, and Jianhui Lai. Stochastic runtime analysis of a cross-entropy algorithm for traveling salesman problems. Theoretical Computer Science, 724(??): 69–86, May 9, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307405.

Anonymous:2018:EBu

[1654] Anonymous. Editorial board. Theoretical Computer Science, 725(??): ii, May 16, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302135.

Cohen:2018:SGG

[1655] Nathann Cohen, Nícolas A. Martins, Fionn Mc Inerney, Nicolas Nisse, Stéphane Pérennes, and Rudini Sampaio. Spy-game on graphs: Complexity and simple topologies. Theoretical Computer Science, 725(??): 1–15, May 16, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308459.

Acharyya:2018:MWC

[1656] Ankush Acharyya, Subhas C. Nandy, and Sasanka Roy. Minimum width color spanning annulus. Theoretical Computer Science, 725(??):16– 30, May 16, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308617.

Flores-Lamas:2018:AFM

[1657] Alejandro Flores-Lamas, José Alberto Fernández-Zepeda, and Joel Antonio Trejo-Sánchez. Algorithm to find a maximum 2-packing set in a cactus. Theoretical Computer Science, 725 (??):31–51, May 16, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308605.

Larsson:2018:GCT

[1658] Urban Larsson, Richard J. Nowakowski, and Carlos P. Santos. Game comparison through play. Theoretical Computer Science, 725(??):52– 63, May 16, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517308599.

Jin:2018:NLT

[1659] Kai Jin, Jian Li, Haitao Wang, Bowei Zhang, and Ningye Zhang. Nearlinear time approximation schemes for geometric maximum coverage. Theoretical Computer Science, 725(??): 64–78, May 16, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308563.

Hong:2018:SAT

[1660] Seok-Hee Hong and Hiroshi Nagamochi. Simpler algorithms for testing two-page book embedding of partitioned graphs. Theoretical Computer Science, 725(??):79– 98, May 16, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012207.

Anonymous:2018:EBv

[1661] Anonymous. Editorial Board. Theoretical Computer Science, 726(??): ii, May 23, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302263.

Gamard:2018:ACF

[1662] Guilhem Gamard, Pascal Ochem, Gwenaël Richomme, and Patrice Séébold. Avoidability of circular formulas. Theoretical Computer Science, 726(??):1–4, May 23, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308447.

Chen:2018:SMM

[1663] Yangjun Chen and Yujia Wu. On the string matching with k mismatches. Theoretical Computer Science, 726 (??):5–29, May 23, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300781.

Fox:2018:EES

[1664] Kyle Fox, Sungjin Im, and Benjamin Moseley. Energy efficient scheduling of parallelizable jobs. Theoretical Computer Science, 726(??): 30–40, May 23, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830118X.

Biely:2018:GDC

[1665] Martin Biely, Peter Robinson, Ulrich Schmid, Manfred Schwarz, and Kyrill Winkler. Gracefully degrading consensus and k-set agreement in directed dynamic networks. Theoretical Computer Science, 726(??): 41–77, May 23, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301166.

Dey:2018:CMP

[1666] Palash Dey, Neeldhara Misra, and Y. Narahari. Complexity of manipulation with partial information in voting. Theoretical Computer Science, 726 (??):78–99, May 23, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301737.

Anonymous:2018:EBw

[1667] Anonymous. Editorial board. Theoretical Computer Science, 727(??): ii, May 30, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518302482.

Alqarni:2018:ISP

[1668] Mohammad Alqarni and Ryszard Janicki. Interval semantics for Petri nets with inhibitor arcs. Theoretical Computer Science, 727(??):1– 23, May 30, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518301415.

Chen:2018:PTA

[1669] Taolue Chen, Tingting Han, and Yongzhi Cao. Polynomial-time algorithms for computing distances of fuzzy transition systems. Theoretical Computer Science, 727(??):24– 36, May 30, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518301403.

Enqvist:2018:CMC

[1670] Sebastian Enqvist, Fatemeh Seifan, and Yde Venema. Completeness for the modal μ-calculus: Separating the combinatorics from the dynamics. Theoretical Computer Science, 727(??): 37–100, May 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397518301397.

Anonymous:2018:EBx

[1671] Anonymous. Editorial board. Theoretical Computer Science, 728(??): ii, June 5, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518302548.

Cheng:2018:CEF

[1672] Qi Cheng, Pingshan Li, and Min Xu. Conditional (edge-)fault-tolerant strong Menger (edge) connectivity of folded hypercubes. Theoretical Computer Science, 728(??):1– 8, June 5, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518301725.

Huang:2018:RAG

[1673] Yanze Huang, Limei Lin, and Dajin Wang. On the reliability of alternating group graph-based networks. Theoretical Computer Science, 728 (??):9–28, June 5, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397518301713.

Engelfriet:2018:MCF

[1674] Joost Engelfriet, Andreas Maletti, and Sebastian Maneth. Multiple context-free tree grammars: Lexicalization and characterization. Theoretical Computer Science, 728(??): 29–99, June 5, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518301750.

Anonymous:2018:EBy

[1675] Anonymous. Editorial board. Theoretical Computer Science, 729(??): ii, June 12, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302822.

Yamanaka:2018:SCT

[1676] Katsuhisa Yamanaka, Takashi Horiyama, J. Mark Keil, David Kirkpatrick, Yota Otachi, Toshiki Saitoh, Ryuhei Uehara, and Yushi Uno. Swapping colored tokens on graphs. Theoretical Computer Science, 729(??): 1–10, June 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301774.

Araujo:2018:RFA

[1677] Júlio Araújo, Julien Baste, and Ignasi Sau. Ruling out FPT algorithms for weighted coloring on forests. Theoretical Computer Science, 729 (??):11–19, June 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301749.

Giraudo:2018:AAA

[1678] Samuele Giraudo and Stéphane Vialette.

Algorithmic and algebraic aspects of unshuffling permutations. Theoretical Computer Science, 729(??): 20–41, June 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300847.

Buning:2018:FRO

[1679] Hans Kleine Büning, Piotr Wojciechowski, and K. Subramani. Finding read-once resolution refutations in systems of 2CNF clauses. Theoretical Computer Science, 729(??): 42–56, June 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300793.

Sheikhi:2018:PMB

[1680] Farnaz Sheikhi and Ali Mohades. Planar maximum-box problem revisited. Theoretical Computer Science, 729 (??):57–67, June 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300045.

Bereg:2018:OSC

[1681] Sergey Bereg, Binay Bhattacharya, Sandip Das, Tsunehiko Kameda, Priya Ranjan Sinha Mahapatra, and Zhao Song. Optimizing squares covering a set of points. Theoretical Computer Science, 729(??):68– 83, June 12, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397515011159.

Anonymous:2018:EBz

[1682] Anonymous. Editorial board. Theoretical Computer Science, 730(??): ii, June 19, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302883.

Ling:2018:LBG

[1683] San Ling, Khoa Nguyen, Adeline Roux-Langlois, and Huaxiong Wang. A lattice-based group signature scheme with verifier-local revocation. Theoretical Computer Science, 730(??): 1–20, June 19, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302056.

Basavaraju:2018:KCS

[1684] Manu Basavaraju, Fahad Panolan, Ashutosh Rai, M. S. Ramanujan, and Saket Saurabh. On the kernelization complexity of string problems. Theoretical Computer Science, 730 (??):21–31, June 19, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301993.

Bresar:2018:CID

[1685] Bostjan Bresar, Tanja Gologranc, and Tim Kos. Convex and isometric domination of (weak) dominating pair graphs. Theoretical Computer Science, 730(??):32–43, June 19, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301981.

Quaglia:2018:SVA

[1686] Elizabeth A. Quaglia and Ben Smyth.

Secret, verifiable auctions from elections. Theoretical Computer Science, 730(??):44–92, June 19, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830183X.

Anonymous:2018:EBba

[1687] Anonymous. Editorial board. Theoretical Computer Science, 731(??): ii, June 30, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303207.

Salem:2018:SOS

[1688] Iosif Salem, Elad M. Schiller, Marina Papatriantafilou, and Philippas Tsigas. Shared-object system equilibria: Delay and throughput analysis. Theoretical Computer Science, 731 (??):1–27, June 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302081.

Wang:2018:TPM

[1689] Fan Wang and Wuyang Sun. A type of perfect matchings extend to Hamiltonian cycles in k-ary n-cubes. Theoretical Computer Science, 731 (??):28–35, June 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830207X.

Dey:2018:MEN

[1690] Palash Dey. Manipulative elicitation — a new attack on elections with incomplete preferences. Theoretical Computer Science, 731(??): 36–49, June 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302068.

He:2018:SME

[1691] Shengjie He, Rong-Xia Hao, and Eddie Cheng. Strongly Menger-edge-connectedness and strongly Mengervertex-connectedness of regular networks. Theoretical Computer Science, 731(??):50–67, June 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302093.

Case:2018:MDR

[1692] Adam Case and Jack H. Lutz. Mutual dimension and random sequences. Theoretical Computer Science, 731 (??):68–87, June 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302172.

Araujo:2018:DIC

[1693] João Araújo, Michael Kinyon, Janusz Konieczny, and António Malheiro. Decidability and independence of conjugacy problems in finitely presented monoids. Theoretical Computer Science, 731(??):88–98, June 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397518302160.

Kelk:2018:TDP

[1694] Steven Kelk, Georgios Stamoulis, and Taoyang Wu. Treewidth distance on phylogenetic trees. Theoretical Computer Science, 731(??):

99–117, June 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302184.

Anonymous:2018:EBbb

[1695] Anonymous. Editorial board. Theoretical Computer Science, 732(??): ii, July 7, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303268.

Masse:2018:LRP

[1696] Alexandre Blondin Massé, Julien de Carufel, Alain Goupil, Mélodie Lapointe, Émile Nadeau, and Élise Vandomme. Leaf realization problem, caterpillar graphs and prefix normal words. Theoretical Computer Science, 732(??):1–13, July 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830241X.

vanEe:2018:ACM

[1697] Martijn van Ee and René Sitters. Approximation and complexity of multitarget graph search and the Canadian traveler problem. Theoretical Computer Science, 732(??):14– 25, July 7, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302445.

Dang:2018:CET

[1698] Chuangyin Dang and Yinyu Ye. On the complexity of an expanded Tarski’s fixed point problem under the componentwise ordering. Theoretical Computer Science, 732(??): 26–45, July 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302433. See

erratum/correction [2417].

Farr:2018:SGF

[1699] Graham Farr and Nhan Bao Ho.

The Sprague–Grundy function for some nearly disjunctive sums of Nim and Silver Dollar games. Theoretical Computer Science, 732(??):46– 59, July 7, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302615.

Vagvolgyi:2018:DRT

[1700] Sándor Vágvölgyi. Descendants of a recognizable tree language for prefix constrained linear monadic term rewriting with position cutting strategy. Theoretical Computer Science, 732(??):60–72, July 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302603.

Dekking:2018:FPH

[1701] Michel Dekking. The Frobenius problem for homomorphic embeddings of languages into the integers. Theoretical Computer Science, 732(??): 73–79, July 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302573.

Demaine:2018:SPP

[1702] Erik D. Demaine and Mikhail Rudoy. A simple proof that the (n^{2} − 1)-puzzle is hard. Theoretical Computer Science, 732(??):80–84, July 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302652.

Halava:2018:FPR

[1703] Vesa Halava, Tero Harju, and Esa Sahla. On fixed points of rational transductions. Theoretical Computer Science, 732(??):85– 88, July 7, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302640.

Anonymous:2018:EBbc

[1704] Anonymous. Editorial board. Theoretical Computer Science, 733(??): ii, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303463.

Zeugmann:2018:GEF

[1705] Thomas Zeugmann. Guest Editor’s foreword [Preface: special issue on learning theory and complexity]. Theoretical Computer Science, 733(??):1–3, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302718.

Bshouty:2018:ELH

[1706] Nader H. Bshouty. Exact learning from an honest teacher that answers membership queries. Theoretical Computer Science, 733(??): 4–43, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830272X.

Fernau:2018:RSA

[1707] Henning Fernau, Florin Manea, Robert Mercas, and Markus L. Schmid. Revisiting Shinohara’s algorithm for computing descriptive patterns. Theoretical Computer Science, 733(??): 44–54, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302731.

Case:2018:EQK

[1708] John Case, Sanjay Jain, and Frank Stephan. Effectivity questions for Kleene’s recursion theorem. Theoretical Computer Science, 733(??): 55–70, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302743.

Hutter:2018:TBS

[1709] Marcus Hutter. Tractability of batch to sequential conversion. Theoretical Computer Science, 733(??):71– 82, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302755.

Kawachi:2018:CLB

[1710] Akinori Kawachi. Circuit lower bounds from learning-theoretic approaches. Theoretical Computer Science, 733 (??):83–98, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302767.

Anonymous:2018:EBbd

[1711] Anonymous. Editorial board. Theoretical Computer Science, 734(??): ii, July 22, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303530.

Zhu:2018:P

[1712] Daming Zhu and Sergey Bereg. Preface. Theoretical Computer Science, 734(??):1–2, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303499.

Chen:2018:NGA

[1713] Zihe Chen, Danyang Chen, Xiangyu Wang, Robert J. Damiano, Hui Meng, and Jinhui Xu. Novel geometric approach for virtual coiling. Theoretical Computer Science, 734(??): 3–14, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300987.

Zou:2018:FDD

[1714] Peng Zou, Hui Li, Wencheng Wang, Chunlin Xin, and Binhai Zhu. Finding disjoint dense clubs in a social network. Theoretical Computer Science, 734(??):15–23, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307466.

Tong:2018:ASM

[1715] Weitian Tong, Eiji Miyano, Randy Goebel, and Guohui Lin. An approximation scheme for minimizing the makespan of the parallel identical multi-stage flow-shops. Theoretical Computer Science, 734(??): 24–31, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306801.

Li:2018:SMI

[1716] Xingfu Li, Haodi Feng, Haotao Jiang, and Binhai Zhu. Solving the maximum internal spanning tree problem on interval graphs in polynomial time. Theoretical Computer Science, 734(??):32–37, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306795.

Pu:2018:CBG

[1717] Lianrong Pu, Yu Lin, Daming Zhu, and Haitao Jiang. Can a breakpoint graph be decomposed into none other than 2cycles? Theoretical Computer Science, 734(??):38–45, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306813.

Mei:2018:MDO

[1718] Lili Mei, Deshi Ye, and Guochuan Zhang. Mechanism design for onefacility location game with obnoxious effects on a line. Theoretical Computer Science, 734(??):46– 57, July 22, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306825.

Ma:2018:NPD

[1719] Yuzhe Ma, Kun He, John Hopcroft, and Pan Shi. Neighbourhoodpreserving dimension reduction via localised multidimensional scaling. Theoretical Computer Science, 734(??): 58–71, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306837.

Yang:2018:KSG

[1720] Yongjie Yang, Yash Raj Shrestha, Wenjun Li, and Jiong Guo. On the kernelization of split graph problems. Theoretical Computer Science, 734(??):72–82, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517306850.

Liu:2018:PCM

[1721] Yunlong Liu, Shaokai Wang, and Jianxin Wang. Parameterized counting matching and packing: A family of hard problems that admit FP-TRAS. Theoretical Computer Science, 734(??):83–93, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306849.

Feng:2018:DSP

[1722] Qilong Feng, Neng Huang, Xiong Jiang, and Jianxin Wang. Dealing with several parameterized problems by random methods. Theoretical Computer Science, 734(??): 94–104, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306862.

Feng:2018:PAE

[1723] Qilong Feng, Shaohua Li, Zeyang Zhou, and Jianxin Wang. Parameterized algorithms for edge biclique and related problems. Theoretical Computer Science, 734(??): 105–118, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306898.

Fallat:2018:CCG

[1724] Shaun Fallat, Karen Meagher, Abolghasem Soltani, and Boting Yang. Compressed cliques graphs, clique coverings and positive zero forcing. Theoretical Computer Science, 734(??): 119–130, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306278.

Chen:2018:AAS

[1725] Zhi-Zhong Chen, Youta Harada, Fei Guo, and Lusheng Wang. Approximation algorithms for the scaffolding problem and its generalizations. Theoretical Computer Science, 734(??): 131–141, July 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302815.

Anonymous:2018:EBbe

[1726] Anonymous. Editorial board. Theoretical Computer Science, 735(??): ii, July 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303591.

Ouaknine:2018:RPS

[1727] Joël Ouaknine, Igor Potapov, and James Worrell. Reachability problems 2014: Special issue. Theoretical Computer Science, 735(??): 1, July 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303621.

Demri:2018:EBM

[1728] Stéphane Demri, Amit Kumar Dhar, and Arnaud Sangnier. Equivalence between model-checking flat counter systems and Presburger arithmetic. Theoretical Computer Science, 735 (??):2–23, July 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305455.

Geeraerts:2018:SSS

[1729] Gilles Geeraerts, Joël Goossens, Thi-Van-Anh Nguyen, and Amélie Stainer. Synthesising succinct strategies in safety games with an application to real-time scheduling. Theoretical Computer Science, 735(??):24– 49, July 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730511X.

Hofman:2018:TIO

[1730] Piotr Hofman and Patrick Totzke.

Trace inclusion for one-counter nets revisited. Theoretical Computer Science, 735(??):50–63, July 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303961.

Jovanovic:2018:PSP

[1731] Aleksandra Jovanović and Marta Kwiatkowska. Parameter synthesis for probabilistic timed automata using stochastic game abstractions. Theoretical Computer Science, 735(??): 64–81, July 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517303833.

Hunter:2018:MPG

[1732] Paul Hunter, Arno Pauly, Guillermo A.

Pérez, and Jean-François Raskin. Mean-payoff games with partial observation. Theoretical Computer Science, 735(??):82–110, July 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751730261X.

Haddad:2018:IIA

[1733] Serge Haddad and Benjamin Monmege. Interval iteration algorithm for MDPs and IMDPs. Theoretical Computer Science, 735(??):111– 131, July 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516307095.

Bazille:2018:CBT

[1734] Hugo Bazille, Olivier Bournez, Walid Gomaa, and Amaury Pouly. On the complexity of bounded time and precision reachability for piecewise affine systems. Theoretical Computer Science, 735(??):132– 146, July 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516305084.

Chistikov:2018:CFC

[1735] Dmitry Chistikov, Christoph Haase, and Simon Halfon. Context-free commutative grammars with integer counters and resets. Theoretical Computer Science, 735(??): 147–161, July 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302535.

Kari:2018:Eb

[1736] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 736(??):ii, August 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518303797.

Buno:2018:SQP

[1737] Kelvin C. Buño, Francis George C.

Cabarle, Marj Darrel Calabia, and Henry N. Adorna. Solving the N-Queens problem using dP systems with active membranes. Theoretical Computer Science, 736(??): iii, August 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397517309040.

Valencia-Cabrera:2018:DRC

[1738] Luis Valencia-Cabrera, David Orellana-Martín, Miguel Á. Martínez del Amor, Agustín Riscos-Núñez, and Mario J. Pérez-Jiménez. From distribution to replication in cooperative systems with active membranes: a frontier of the efficiency. Theoretical Computer Science, 736(??):1–14, August 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397517309039.

Anonymous:2018:EBbf

[1739] Anonymous. Editorial Board. Theoretical Computer Science, 736(??): 1–102, August 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397518303761.

Rowe:2018:LMO

[1740] Jonathan E. Rowe. Linear multiobjective drift analysis. Theoretical Computer Science, 736(??):15– 24, August 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S030439751830094X.

Regnault:2018:LSS

[1741] Damien Regnault and Éric Rémila. Lost in self-stabilization: a local process that aligns connected cells. Theoretical Computer Science, 736(??): 25–40, August 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301002.

Grillo:2018:QQS

[1742] S. A. Grillo and F. L. Marquezino. Quantum query as a state decomposition. Theoretical Computer Science, 736(??):41–61, August 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S0304397518301786.

Paun:2018:DRT

[1743] Gheorghe Paun. A dozen of research topics in membrane computing. Theoretical Computer Science, 736(??): 62–75, August 6, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302226.

Barbuti:2018:PFM

[1744] Roberto Barbuti, Roberta Gori, and Paolo Milazzo. Predictors for flat membrane systems. Theoretical Computer Science, 736(??):76– 78, August 6, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518302214.

Anonymous:2018:EBbg

[1745] Anonymous. Editorial board. Theoretical Computer Science, 737(??): ii, August 15, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303839.

Polyvyanyy:2018:IIE

[1746] Artem Polyvyanyy, Jan Sürmeli, and Matthias Weidlich. Interleaving isotactics — an equivalence notion on behaviour abstractions. Theoretical Computer Science, 737(??):1–18, August 15, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300161.

Choffrut:2018:TET

[1747] Ch. Choffrut and Z. Ésik. Two equational theories of partial words. Theoretical Computer Science, 737(??): 19–39, August 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302329.

DAgostino:2018:CAD

[1748] Giovanna D’Agostino and Giacomo Lenzi. The μ-calculus alternation depth hierarchy is infinite over finite planar graphs. Theoretical Computer Science, 737(??):40–61, August 15, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302317.

Ayala-Rincon:2018:NEI

[1749] Maurício Ayala-Rincón, Maribel Fernández, Ana Cristina Rocha-Oliveira, and Daniel Lima Ventura. Nominal essential intersection types. Theoretical Computer Science, 737(??): 62–80, August 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303141.

Badia:2018:FCG

[1750] Guillermo Badia and Carles Noguera. Fraïssé classes of graded relational structures. Theoretical Computer Science, 737(??):81–90, August 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397518303165.

Ibarra:2018:VCS

[1751] Oscar H. Ibarra and Ian McQuillan. Variations of checking stack automata: Obtaining unexpected decidability properties. Theoretical Computer Science, 738(??):ii, August 22, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518302585.

Sun:2018:CBO

[1752] Lei Sun and Fang-Wei Fu. Constructions of balanced odd-variable rotation symmetric Boolean functions with optimal algebraic immunity and high nonlinearity. Theoretical Computer Science, 738(??): 1–12, August 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397518302780.

Anonymous:2018:EBbh

[1753] Anonymous. Editorial Board. Theoretical Computer Science, 738(??): 1–80, August 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S030439751830389X.

Currie:2018:ABF

[1754] James D. Currie, Lucas Mol, and Narad Rampersad. Avoidance bases for formulas with reversal. Theoretical Computer Science, 738(??): 13–24, August 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397518302779.

Izumi:2018:TCC

[1755] Taisuke Izumi, Daichi Kaino, Maria Gradinariu Potop-Butucaru, and Sébastien Tixeuil. On time complexity for connectivity-preserving scattering of mobile robots. Theoretical Computer Science, 738(??):25–41, August 22, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S0304397518302974.

Cornelissen:2018:BPM

[1756] Kamiel Cornelissen and Bodo Manthey. Belief propagation for the maximumweight independent set and minimum spanning tree problems. Theoretical Computer Science, 738(??): 42–52, August 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397518302962.

Anthony:2018:LWN

[1757] Martin Anthony and Joel Ratsaby. Large width nearest prototype classification on general distance spaces. Theoretical Computer Science, 738(??): 53–64, August 22, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S0304397518302950.

Anonymous:2018:EBbi

[1758] Anonymous. Editorial board. Theoretical Computer Science, 739(??): ii, August 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303955.

Dondi:2018:PCA

[1759] Riccardo Dondi and Florian Sikora. Parameterized complexity and approximation issues for the colorful components problems. Theoretical Computer Science, 739(??): 1–12, August 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302949.

Chen:2018:RLF

[1760] Yu Chen, Baodong Qin, and Haiyang Xue. Regular lossy functions and their applications in leakage-resilient cryptography. Theoretical Computer Science, 739(??):13–38, August 29, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302937.

Noori:2018:LMG

[1761] Narges Noori and Volkan Isler. The lion and man game on polyhedral surfaces with obstacles. Theoretical Computer Science, 739(??):39– 58, August 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302913.

Zhang:2018:ECP

[1762] Wenwen Zhang and Jian-Liang Wu.

Edge coloring of planar graphs without adjacent 7-cycles. Theoretical Computer Science, 739(??):59– 64, August 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303128.

Hatanaka:2018:PCL

[1763] Tatsuhiko Hatanaka, Takehiro Ito, and Xiao Zhou. Parameterized complexity of the list coloring reconfiguration problem with graph parameters. Theoretical Computer Science, 739(??): 65–79, August 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303116.

Li:2018:KAV

[1764] Wenjun Li and Binhai Zhu. A 2kkernelization algorithm for vertex cover based on crown decomposition. Theoretical Computer Science, 739(??): 80–85, August 29, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303104.

Anonymous:2018:EBbj

[1765] Anonymous. Editorial board. Theoretical Computer Science, 740(??): ii, September 5, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304250.

Kostolanyi:2018:AWA

[1766] Peter Kostolányi and Filip Misún. Alternating weighted automata over commutative semirings. Theoretical Computer Science, 740(??):1– 27, September 5, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303098.

Mittelbach:2018:FSH

[1767] Arno Mittelbach and Daniele Venturi. Fiat–Shamir for highly sound protocols is instantiable. Theoretical Computer Science, 740(??):28– 62, September 5, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303074.

Honkala:2018:ESB

[1768] Juha Honkala. Equality sets of binary D0L sequences. Theoretical Computer Science, 740(??):63– 67, September 5, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303311.

Li:2018:BPL

[1769] Wenhua Li, Xing Chai, and Yali Song. A best possible on-line algorithm for scheduling on uniform parallel-batch machines. Theoretical Computer Science, 740(??):68– 75, September 5, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303347.

Konstantinidis:2018:STC

[1770] Athanasios L. Konstantinidis, Stavros D.

Nikolopoulos, and Charis Papadopoulos. Strong triadic closure in cographs and graphs of low maximum degree. Theoretical Computer Science, 740(??): 76–84, September 5, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303323.

Hofmann:2018:F

[1771] Martin Hofmann, David Aspinall, Brian Campbell, Ian Stark, and Perdita Stevens. Foreword. Theoretical Computer Science, 741(??): ii, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304481.

Kennedy:2018:TGC

[1772] Andrew J. Kennedy and Claudio V. Russo. Transposing G to C^{♯}: Expressivity of generalized algebraic data types in an object-oriented language. Theoretical Computer Science, 741 (??):1–2, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301270.

Anonymous:2018:EBbk

[1773] Anonymous. Editorial Board. Theoretical Computer Science, 741(??): 1–70, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304808.

Orejas:2018:INL

[1774] Fernando Orejas, Elvira Pino, Marisa Navarro, and Leen Lambers. Institutions for navigational logics for graphical structures. Theoretical Computer Science, 741(??):3– 18, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301282.

Harper:2018:ETO

[1775] Robert Harper. Exception tracking in an open world. Theoretical Computer Science, 741(??):19– 24, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301294.

Hennicker:2018:BAS

[1776] Rolf Hennicker, Alexandre Madeira, and Martin Wirsing. Behavioural and abstractor specifications revisited. Theoretical Computer Science, 741(??):25–31, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301300.

Mosses:2018:CDS

[1777] Peter D. Mosses. CoFI with Don Sannella. Theoretical Computer Science, 741(??):32–43, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301312.

Ghani:2018:CTI

[1778] Neil Ghani, Clemens Kupke, Alasdair Lambert, and Fredrik Nordvall Forsberg. A compositional treatment of iterated open games. Theoretical Computer Science, 741(??): 44–47, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303645.

Mossakowski:2018:PPS

[1779] Till Mossakowski and Bernd Krieg-Brückner. Partial pushout semantics of generics in DOL. Theoretical Computer Science, 741(??): 48–57, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518301828.

Ortner:2018:GEF

[1780] Ronald Ortner and Hans Ulrich Simon. Guest Editors’ Foreword. Theoretical Computer Science, 742(??): ii, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309222.

Kaufmann:2018:SAA

[1781] Emilie Kaufmann, Thomas Bonald, and Marc Lelarge. A spectral algorithm with additive clustering for the recovery of overlapping communities in networks. Theoretical Computer Science, 742(??):1–2, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309234.

Anonymous:2018:EBbl

[1782] Anonymous. Editorial Board. Theoretical Computer Science, 742(??): 1–114, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305085.

Zhivotovskiy:2018:LVC

[1783] N. Zhivotovskiy and S. Hanneke. Localization of VC classes: Beyond local Rademacher complexities. Theoretical Computer Science, 742(??): 3–26, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309246.

Kotlowski:2018:MFL

[1784] Wojciech Kotlowski. On minimaxity of Follow the Leader strategy in the stochastic setting. Theoretical Computer Science, 742(??):27– 49, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309258.

Holzl:2018:LPL

[1785] Rupert Hölzl, Sanjay Jain, and Frank Stephan. Learning pattern languages over groups. Theoretical Computer Science, 742(??):50– 65, ???? 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730926X.

Bshouty:2018:ELJ

[1786] Nader H. Bshouty and Areej Costa. Exact learning of juntas from membership queries. Theoretical Computer Science, 742(??):66–81, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309271.

Sabato:2018:SLC

[1787] Sivan Sabato. Submodular learning and covering with response-dependent costs. Theoretical Computer Science, 742(??):82–97, ???? 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517309283.

Ferenczi:2018:AWE

[1788] Sébastien Ferenczi, Christian Mauduit, and Carlos Gustavo Moreira. An algorithm for the word entropy. Theoretical Computer Science, 743(??): ii, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304572.

Gabric:2018:CBS

[1789] Daniel Gabric and Joe Sawada. Constructing de Bruijn sequences by concatenating smaller universal cycles. Theoretical Computer Science, 743(??): 1–11, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304559.

Anonymous:2018:EBbm

[1790] Anonymous. Editorial Board. Theoretical Computer Science, 743(??):1– 166, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305140.

Klouda:2018:FPS

[1791] Karel Klouda, Katerina Medková, Edita Pelantová, and Stepán Starosta. Fixed points of Sturmian morphisms and their derivated words. Theoretical Computer Science, 743(??):12– 22, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304535.

Cicalese:2018:BFN

[1792] Ferdinando Cicalese, Zsuzsanna Lipták, and Massimiliano Rossi. Bubbleflip — a new generation algorithm for prefix normal words. Theoretical Computer Science, 743(??):23– 37, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304377.

Carey:2018:PFC

[1793] Norman Carey and David Clampitt.

Partitioned factors in Christoffel and Sturmian words. Theoretical Computer Science, 743(??):38– 52, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304353.

Petersen:2018:TST

[1794] Karl Petersen and Ibrahim Salama.

Tree shift topological entropy. Theoretical Computer Science, 743(??):53– 63, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303724.

Currie:2018:UPU

[1795] James Currie, Florin Manea, Dirk Nowotka, and Kamellia Reshadi. Unary patterns under permutations. Theoretical Computer Science, 743(??): 64–71, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303712.

Teh:2018:OWR

[1796] Wen Chean Teh, K. G. Subramanian, and Somnath Bera. Order of weak Mrelation and Parikh matrices. Theoretical Computer Science, 743(??):72– 82, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302925.

Bertet:2018:LCS

[1797] Karell Bertet, Christophe Demko, Jean-François Viaud, and Clément Guérin. Lattices, closures systems and implication bases: a survey of structural aspects and algorithms. Theoretical Computer Science, 743(??):83– 92, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306806.

Ge:2018:CEV

[1798] Cunjing Ge, Feifei Ma, Peng Zhang, and Jian Zhang. Computing and estimating the volume of the solution space of SMT(LA) constraints. Theoretical Computer Science, 743(??):93– 109, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516305977.

Bender:2018:RQR

[1799] Michael A. Bender, Rezaul A. Chowdhury, Pramod Ganapathi, Samuel Mc-Cauley, and Yuan Tang. The range 1 query (R1Q) problem. Theoretical Computer Science, 743(??):110– 129, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397515012219.

Hendricks:2018:PDS

[1800] Jacob Hendricks, Matthew J. Patitz, Trent A. Rogers, and Scott M. Summers. The power of duples (in selfassembly): It’s not so hip to be square. Theoretical Computer Science, 743(??): 130–147, September 26, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751501169X.

Anonymous:2018:EBbn

[1801] Anonymous. Editorial Board. Theoretical Computer Science, 744(??): ii, October 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305450.

Sampaio:2018:TAC

[1802] Augusto Sampaio and Farn Wang. Theoretical aspects of computing. Theoretical Computer Science, 744 (??):1–2, October 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304705.

Babari:2018:WRA

[1803] Parvaneh Babari, Manfred Droste, and Vitaly Perevoshchikov. Weighted register automata and weighted logic on data words. Theoretical Computer Science, 744(??):3– 21, October 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830015X.

Benevides:2018:TRA

[1804] Mario Benevides, Bruno Lopes, and Edward Hermann Haeusler. Towards reasoning about Petri nets: a Propositional Dynamic Logic based approach. Theoretical Computer Science, 744 (??):22–36, October 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518300185.

Guttmann:2018:AFM

[1805] Walter Guttmann. An algebraic framework for minimum spanning tree problems. Theoretical Computer Science, 744(??):37–55, October 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302342.

Guzman:2018:CRI

[1806] Michell Guzmán, Salim Perchy, Camilo Rueda, and Frank D. Valencia. Characterizing right inverses for spatial constraint systems with applications to modal logic. Theoretical Computer Science, 744(??):56– 77, October 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303426.

Madeira:2018:LSD

[1807] Alexandre Madeira, Luís S. Barbosa, Rolf Hennicker, and Manuel A. Martins. A logic for the stepwise development of reactive systems. Theoretical Computer Science, 744(??): 78–96, October 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301439.

McIver:2018:SFG

[1808] Annabelle McIver, Tahiry Rabehaja, and Georg Struth. Schedulers and finishers: On generating and filtering the behaviours of an event structure. Theoretical Computer Science, 744 (??):97–112, October 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300410.

Neves:2018:LMH

[1809] Renato Neves and Luís S. Barbosa.

Languages and models for hybrid automata: a coalgebraic perspective. Theoretical Computer Science, 744(??):

113–142, October 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397517307193.

Schwammberger:2018:AMP

[1810] Maike Schwammberger. An abstract model for proving safety of autonomous urban traffic. Theoretical Computer Science, 744(??): 143–169, October 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303669.

Buhmann:2018:PAL

[1811] Joachim M. Buhmann, Julien Dumazert, Alexey Gronskiy, and Wojciech Szpankowski. Posterior agreement for large parameter-rich optimization problems. Theoretical Computer Science, 745(??):ii, October 12, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302378.

Hanneke:2018:TPF

[1812] Steve Hanneke and Liu Yang. Testing piecewise functions. Theoretical Computer Science, 745(??):1– 22, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303396.

Anonymous:2018:EBbo

[1813] Anonymous. Editorial Board. Theoretical Computer Science, 745(??):

1–224, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305516.

Bae:2018:GPG

[1814] Sang Won Bae, Jean-François Baffier, Jinhee Chun, Peter Eades, Kord Eickmeyer, Luca Grilli, Seok-Hee Hong, Matias Korman, Fabrizio Montecchiani, Ignaz Rutter, and Csaba D. Tóth. Gap-planar graphs. Theoretical Computer Science, 745(??): 23–35, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303670.

Reddy:2018:PAC

[1815] I. Vinod Reddy. Parameterized algorithms for conflict-free colorings of graphs. Theoretical Computer Science, 745(??):36–52, October 12, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303633.

Mezzini:2018:PTA

[1816] Mauro Mezzini. Polynomial time algorithm for computing a minimum geodetic set in outerplanar graphs. Theoretical Computer Science, 745(??): 53–62, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303700.

Cao:2018:VDP

[1817] Yixin Cao, Yuping Ke, Yota Otachi, and Jie You. Vertex deletion problems on chordal graphs. Theoretical Computer Science, 745(??):63– 74, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304067.

Thorncharoensri:2018:PCS

[1818] Pairat Thorncharoensri, Willy Susilo, and Yi Mu. Policy controlled system with anonymity. Theoretical Computer Science, 745(??):75– 86, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304055.

Ibarra:2018:SLL

[1819] Oscar H. Ibarra and Ian McQuillan. On store languages of language acceptors. Theoretical Computer Science, 745(??): 87–113, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304031.

Garraffa:2018:EEB

[1820] Michele Garraffa, Lei Shang, Federico Della Croce, and Vincent T’kindt. An exact exponential branch-and-merge algorithm for the single machine total tardiness problem. Theoretical Computer Science, 745(??): 114–132, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304079.

Gatti:2018:TBM

[1821] N. Gatti, M. Rocco, P. Serafino, and C. Ventre. Towards better models of externalities in sponsored search auctions. Theoretical Computer Science, 745(??):133– 149, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304195.

Cai:2018:CGH

[1822] Jin-Yi Cai, Heng Guo, and Tyson Williams. Clifford gates in the Holant framework. Theoretical Computer Science, 745(??):150–162, October 12, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304183.

Escamocher:2018:PFM

[1823] Guillaume Escamocher and Barry O’Sullivan. Pushing the frontier of minimality. Theoretical Computer Science, 745(??):163–171, October 12, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830416X.

Watel:2018:TSP

[1824] Dimitri Watel and Alain Faye. Taxisharing: Parameterized complexity and approximability of the dial-a-ride problem with money as an incentive. Theoretical Computer Science, 745(??):

172–201, October 12, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304146.

Anonymous:2018:EBbp

[1825] Anonymous. Editorial Board. Theoretical Computer Science, 746(??): ii, October 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305632.

Xu:2018:BGN

[1826] Rongxing Xu and Xuding Zhu. Bounded greedy Nim. Theoretical Computer Science, 746(??):1–5, October 25, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304316.

Komargodski:2018:LRO

[1827] Ilan Komargodski. Leakage resilient one-way functions: the auxiliaryinput setting. Theoretical Computer Science, 746(??):6–18, October 25, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304304.

Beaudou:2018:OGG

[1828] Laurent Beaudou, Pierre Coupechoux, Antoine Dailly, Sylvain Gravier, Julien Moncel, Aline Parreau, and Éric Sopena. Octal games on graphs: the game 0.33 on subdivided stars and bistars. Theoretical Computer Science, 746(??):19– 35, October 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304341.

Alves:2018:PCR

[1829] Sancrey Rodrigues Alves, Konrad K.

Dabrowski, Luerbio Faria, Sulamita Klein, Ignasi Sau, and Uéverton S. Souza. On the (parameterized) complexity of recognizing well-covered (r, l)-graph. Theoretical Computer Science, 746(??):36–48, October 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397518304407.

Diner:2018:CDB

[1830] Öznur Yasar Diner, Daniël Paulusma, Christophe Picouleau, and Bernard Ries. Contraction and deletion blockers for perfect graphs and H-free graphs. Theoretical Computer Science, 746(??): 49–72, October 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304390.

Bilo:2018:OFG

[1831] Vittorio Bilò, Angelo Fanelli, and Luca Moscardelli. Opinion formation games with dynamic social influences. Theoretical Computer Science, 746(??): 73–87, October 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304419.

Kobayashi:2018:NHF

[1832] Yusuke Kobayashi. NP-hardness and fixed-parameter tractability of the minimum spanner problem. Theoretical Computer Science, 746(??): 88–97, October 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830447X.

Biswas:2018:RDR

[1833] Sudip Biswas, Arnab Ganguly, Rahul Shah, and Sharma V. Thankachan. Ranked document retrieval for multiple patterns. Theoretical Computer Science, 746(??):98–111, October 25, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304456.

Bang-Jensen:2018:DCP

[1834] Jørgen Bang-Jensen and Tilde My Christiansen. Degree constrained 2partitions of semicomplete digraphs. Theoretical Computer Science, 746(??): 112–123, October 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304444.

Mitsche:2018:BNG

[1835] Dieter Mitsche, Paweł Pralat, and Elham Roshanbin. Burning number of graph products. Theoretical Computer Science, 746(??):124– 135, October 25, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304523.

Ibarra:2018:GCN

[1836] Oscar H. Ibarra. Grammatical characterizations of NPDAs and VPDAs with counters. Theoretical Computer Science, 746(??):136–150, October 25, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304511.

Anonymous:2018:EBbq

[1837] Anonymous. Editorial Board. Theoretical Computer Science, 747(??): ii, November 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305814.

Matt:2018:TAT

[1838] Christian Matt, Ueli Maurer, Christopher Portmann, Renato Renner, and Björn Tackmann. Toward an algebraic theory of systems. Theoretical Computer Science, 747(??):1– 25, November 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304092.

Ishii:2018:PLR

[1839] Katsumasa Ishii. A proof of the leftmost reduction theorem for λβηcalculus. Theoretical Computer Science, 747(??):26–32, November 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397518304110.

Kojima:2018:GHP

[1840] Kensuke Kojima, Minoru Kinoshita, and Kohei Suenaga. Generalized homogeneous polynomials for efficient template-based nonlinear invariant synthesis. Theoretical Computer Science, 747(??):33–47, November 7, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304134.

Bart:2018:RPI

[1841] Anicet Bart, Benoît Delahaye, Paulin Fournier, Didier Lime, Éric Monfroy, and Charlotte Truchet. Reachability in parametric Interval Markov Chains using constraints. Theoretical Computer Science, 747(??):48– 74, November 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304328.

Moggi:2018:SRR

[1842] Eugenio Moggi, Amin Farjudian, Adam Duracz, and Walid Taha. Safe and robust reachability analysis of hybrid systems. Theoretical Computer Science, 747(??):75– 99, November 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304365.

Weinert:2018:VLD

[1843] Alexander Weinert and Martin Zimmermann. Visibly linear dynamic logic. Theoretical Computer Science, 747(??): 100–117, November 7, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304468.

Anonymous:2018:EBbr

[1844] Anonymous. Editorial Board. Theoretical Computer Science, 748(??): ii, November 14, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830625X.

Demaine:2018:EF

[1845] Erik D. Demaine and Fabrizio Grandoni. Editorial fun. Theoretical Computer Science, 748(??): 1, November 14, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304833.

Bekos:2018:AIR

[1846] Michael A. Bekos, Till Bruckdorfer, Henry Förster, Michael Kaufmann, Simon Poschenrieder, and Thomas Stüber. Algorithms and insights for RaceTrack. Theoretical Computer Science, 748(??):2– 16, November 14, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302627.

Langerman:2018:TFH

[1847] Stefan Langerman and Yushi Uno. Threes!, fives, 1024!, and 2048 are hard. Theoretical Computer Science, 748(??): 17–27, November 14, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301798.

Demaine:2018:FCP

[1848] Erik D. Demaine, Fermi Ma, Ariel Schvartzman, Erik Waingarten, and Scott Aaronson. The fewest clues problem. Theoretical Computer Science, 748(??):28–39, November 14, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397518300677.

Herter:2018:LGC

[1849] Felix Herter and Günter Rote. Loopless Gray code enumeration and the Tower of Bucharest. Theoretical Computer Science, 748(??):40– 54, November 14, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308472.

DeBiasi:2018:CSU

[1850] Marzio De Biasi and Tim Ophelders. The complexity of snake and undirected NCL variants. Theoretical Computer Science, 748(??):55– 65, November 14, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307946.

Almanza:2018:TNH

[1851] Matteo Almanza, Stefano Leucci, and Alessandro Panconesi. Trainyard is NP-Hard. Theoretical Computer Science, 748(??):66–76, November 14, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730720X.

Anonymous:2018:EBbs

[1852] Anonymous. Editorial Board. Theoretical Computer Science, 749(??): ii, November 21, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306182.

Anonymous:2018:P

[1853] Anonymous.  Preface.  Theoret-

ical Computer Science, 749(??):1, November 21, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306212.

Yang:2018:CMA

[1854] Kai Yang, Zhenhua Duan, Cong Tian, and Nan Zhang. A compiler for MSVL and its applications. Theoretical Computer Science, 749(??):2– 16, November 21, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305807.

Miao:2018:SVD

[1855] Dongjing Miao, Zhipeng Cai, and Yingshu Li. SEF view deletion under bounded condition. Theoretical Computer Science, 749(??):17– 25, November 21, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305820.

Liu:2018:TSR

[1856] Xianmin Liu, Zhipeng Cai, Dongjing Miao, and Jianzhong Li. Tree size reduction with keeping distinguishability. Theoretical Computer Science, 749(??): 26–35, November 21, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306266.

Gu:2018:AAR

[1857] Qian-Ping Gu, Jiajian Leo Liang, and Guochuan Zhang. Algorithmic analysis for ridesharing of personal vehicles. Theoretical Computer Science, 749(??): 36–46, November 21, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730628X.

Zhang:2018:AHR

[1858] Peng Zhang, Chenchen Wu, and Dachuan Xu. Approximation and hardness results for the Max k-Uncut problem. Theoretical Computer Science, 749(??):47–58, November 21, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306515.

Bai:2018:CCP

[1859] Chunsong Bai, Liying Kang, and Erfang Shan. The connected p-center problem on cactus graphs. Theoretical Computer Science, 749(??):59– 65, November 21, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517306904.

Gao:2018:RAN

[1860] Xiaofeng Gao, Yueyang Xianzang, Xiaotian You, Yaru Dang, Guihai Chen, and Xinglong Wang. Reachability for airline networks: fast algorithm for shortest path problem with time windows. Theoretical Computer Science, 749(??):66– 79, November 21, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830063X.

Wu:2018:AAM

[1861] Chenchen Wu, Donglei Du, and Dachuan Xu. An approximation algorithm for the k-median problem with uniform penalties via pseudo-solution. Theoretical Computer Science, 749(??): 80–92, November 21, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301233.

Anonymous:2018:EBbt

[1862] Anonymous. Editorial Board. Theoretical Computer Science, 750(??): ii, November 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306613.

Larsen:2018:RPS

[1863] Kim G. Larsen, Igor Potapov, and Jirí Srba. Reachability problems: Special issue. Theoretical Computer Science, 750(??):1, November 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830656X.

Jones:2018:WPG

[1864] Sam A. M. Jones and Richard M. Thomas. Word problems of groups: Formal languages, characterizations and decidability. Theoretical Computer Science, 750(??):2–23, November 30, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830313X.

Hague:2018:DMI

[1865] Matthew Hague and Anthony W. Lin. Decidable models of integermanipulating programs with recursive parallelism. Theoretical Computer Science, 750(??):24–37, November 30, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303049.

Geffroy:2018:ORA

[1866] Thomas Geffroy, Jérôme Leroux, and Grégoire Sutre. Occam’s Razor applied to the Petri net coverability problem. Theoretical Computer Science, 750(??):

38–52, November 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302366.

LeCoent:2018:CSS

[1867] Adrien Le Coënt, Laurent Fribourg, Nicolas Markey, Florian De Vuyst, and Ludovic Chamoin. Compositional synthesis of state-dependent switching control. Theoretical Computer Science, 750(??):53–68, November 30, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300689.

Alechina:2018:CRB

[1868] N. Alechina, N. Bulling, S. Demri, and B. Logan. On the complexity of resource-bounded logics. Theoretical Computer Science, 750(??):69– 100, November 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300665.

Best:2018:SCM

[1869] Eike Best, Thomas Hujsa, and Harro Wimmel. Sufficient conditions for the marked graph realisability of labelled transition systems. Theoretical Computer Science, 750(??):101– 116, November 30, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307181.

Anonymous:2018:EBbu

[1870] Anonymous. Editorial Board. Theoretical Computer Science, 751(??): ii, December 3, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830673X.

Scheideler:2018:P

[1871] Christian Scheideler. Preface. Theoretical Computer Science, 751(??): 1, December 3, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306698.

Barenboim:2018:FND

[1872] Leonid Barenboim, Michael Elkin, and Cyril Gavoille. A fast networkdecomposition algorithm and its applications to constant-time distributed computation. Theoretical Computer Science, 751(??):2–23, December 3, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751630319X.

Bury:2018:ROB

[1873] Marc Bury. Randomized OBDDbased graph algorithms. Theoretical Computer Science, 751(??):24– 45, December 3, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516306910.

Moses:2018:MEM

[1874] Yoram Moses and Katia Patkin. Mutual exclusion as a matter of priority. Theoretical Computer Science, 751(??): 46–60, December 3, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516307435.

Fraigniaud:2018:NLL

[1875] Pierre Fraigniaud, Juho Hirvonen, and Jukka Suomela. Node labels in local decision. Theoretical Computer Science, 751(??):61– 73, December 3, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751730052X.

Censor-Hillel:2018:FRI

[1876] Keren Censor-Hillel and Tariq Toukan.

On fast and robust information spreading in the Vertex-Congest model. Theoretical Computer Science, 751(??): 74–90, December 3, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305224.

Anonymous:2018:EBbv

[1877] Anonymous. Editorial Board. Theoretical Computer Science, 752(??): ii, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306868.

Kari:2018:Ec

[1878] Lila Kari. Editorial. Theoretical Computer Science, 752(??):iii, December 15, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306893.

Aerts:2018:SIQ

[1879] Diederik Aerts, Massimo Melucci, Massimiliano Sassoli de Bianchi, Sandro Sozzo, and Tomas Veloz. Special issue: Quantum structures in computer science: Language, semantics, retrieval. Theoretical Computer Science, 752(??): 1–4, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306819.

DiBuccio:2018:USE

[1880] Emanuele Di Buccio. Utilizing sources of evidence in relevance feedback through geometry. Theoretical Computer Science, 752(??):5– 20, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303657.

Zhang:2018:QIM

[1881] Yazhou Zhang, Dawei Song, Peng Zhang, Panpan Wang, Jingfei Li, Xiang Li, and Benyou Wang. A quantum-inspired multimodal sentiment analysis framework. Theoretical Computer Science, 752(??):21– 40, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302639.

Busemeyer:2018:DFU

[1882] Jerome Busemeyer and Zheng Wang.

Data fusion using Hilbert space multi-dimensional models. Theoretical Computer Science, 752(??):41– 55, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308915.

Bruza:2018:MCP

[1883] Peter D. Bruza. Modelling contextuality by probabilistic programs with hypergraph semantics. Theoretical Computer Science, 752(??):56– 70, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308587.

Yan:2018:FRM

[1884] Fei Yan, Abdullah M. Iliyasu, Yiming Guo, and Huamin Yang. Flexible representation and manipulation of audio signals on quantum computers. Theoretical Computer Science, 752(??): 71–85, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309209.

Leporini:2018:QSQ

[1885] R. Leporini and C. Bertini. Quantum structures in qudit spaces. Theoretical Computer Science, 752(??): 86–96, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517308575.

Danilov:2018:PQB

[1886] V. I. Danilov and A. Lambert-Mogiliansky. Preparing a (quantum) belief system. Theoretical Computer Science, 752(??):97–103, December 15, 2018. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301087.

Coecke:2018:GRL

[1887] Bob Coecke, Fabrizio Genovese, Martha Lewis, Dan Marsden, and Alex Toumi. Generalized relations in linguistics and cognition. Theoretical Computer Science, 752(??):104– 115, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301476.

Aerts:2018:TQW

[1888] Diederik Aerts, Jonito Aerts Arguëlles, Lester Beltran, Lyneth Beltran, Isaac Distrito, Massimiliano Sassoli de Bianchi, Sandro Sozzo, and Tomas Veloz. Towards a quantum World Wide Web. Theoretical Computer Science, 752(??):116– 131, December 15, 2018. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301804.

Anonymous:2019:EBa

[1889] Anonymous. Editorial Board. Theoretical Computer Science, 753(??): ii, January 1, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306996.

Tomohiro:2019:IUB

[1890] Tomohiro I and Dominik Köppl. Improved upper bounds on all maximal αgapped repeats and palindromes. Theoretical Computer Science, 753(??): 1–15, January 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304493.

Sakai:2019:SSL

[1891] Yoshifumi Sakai. A substringsubstring LCS data structure. Theoretical Computer Science, 753(??): 16–34, January 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830450X.

Datta:2019:SSS

[1892] Ajoy K. Datta, Stéphane Devismes, and Lawrence L. Larmore. A silent self-stabilizing algorithm for the generalized minimal k-dominating set problem. Theoretical Computer Science, 753(??):35–63, January 1, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304560.

Ileri:2019:SSA

[1893] Can Umut Ileri and Orhan Dagdeviren. A self-stabilizing algorithm for b-matching. Theoretical Computer Science, 753(??):64–75, January 1, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304584.

Ryzhikov:2019:ARB

[1894] Andrew Ryzhikov. On automata recognizing birecurrent sets. Theoretical Computer Science, 753(??): 76–79, January 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304596.

Bampas:2019:ARG

[1895] Evangelos Bampas, Lélia Blin, Jurek Czyzowicz, David Ilcinkas, Arnaud Labourel, Maria Potop-Butucaru, and Sébastien Tixeuil. On asynchronous rendezvous in general graphs. Theoretical Computer Science, 753(??): 80–90, January 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304614.

Anonymous:2019:EBb

[1896] Anonymous. Editorial Board. Theoretical Computer Science, 754(??): ii, January 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306935.

Fotakis:2019:PSI

[1897] Dimitris Fotakis, Aris Pagourtzis, and Vangelis Paschos. Preface to special issue on algorithms and complexity. Theoretical Computer Science, 754(??):1–2, January 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306820.

Golovach:2019:EMI

[1898] Petr A. Golovach, Dieter Kratsch, and Mohamed Yosri Sayadi. Enumeration of maximal irredundant sets for claw-free graphs. Theoretical Computer Science, 754(??):3– 15, January 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300999.

Hagerup:2019:SEE

[1899] Torben Hagerup, Frank Kammer, and Moritz Laudahn. Space-efficient Euler partition and bipartite edge coloring. Theoretical Computer Science, 754(??):16–34, January 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300197.

DiLuna:2019:PPF

[1900] Giuseppe A. Di Luna, Paola Flocchini, Taisuke Izumi, Tomoko Izumi, Nicola Santoro, and Giovanni Viglietta. Population protocols with faulty interactions: the impact of a leader. Theoretical Computer Science, 754 (??):35–49, January 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830570X.

Gourves:2019:MSA

[1901] Laurent Gourvès and Jérôme Monnot. On maximin share allocations in matroids. Theoretical Computer Science, 754(??):50–64, January 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303384.

Kleer:2019:TIB

[1902] Pieter Kleer and Guido Schäfer. Tight inefficiency bounds for perceptionparameterized affine congestion games. Theoretical Computer Science, 754 (??):65–87, January 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518302597.

Brauer:2019:CSS

[1903] Sascha Brauer. Complexity of singleswap heuristics for metric facility location and related problems. Theoretical Computer Science, 754(??): 88–106, January 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303025.

Abed:2019:SMJ

[1904] Fidaa Abed, Lin Chen, Yann Disser, Martin Groß, Nicole Megow, Julie Meißner, Alexander T. Richter, and Roman Rischke. Scheduling maintenance jobs in networks. Theoretical Computer Science, 754(??): 107–121, January 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301178.

Emek:2019:MCP

[1905] Yuval Emek, Yaacov Shapiro, and Yuyi Wang. Minimum cost perfect matching with delays for two sources. Theoretical Computer Science, 754(??): 122–129, January 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304857.

Anonymous:2019:EBc

[1906] Anonymous. Editorial Board. Theoretical Computer Science, 755(??): ii, January 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307400.

Wallner:2019:BPI

[1907] Michael Wallner. A bijection of plane increasing trees with relaxed binary trees of right height at most one. Theoretical Computer Science, 755(??): 1–12, January 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304699.

Nader:2019:PSD

[1908] Rafic Nader, Alain Bretto, Bassam Mourad, and Hassan Abbas. On the positive semi-definite property of similarity matrices. Theoretical Computer Science, 755(??): 13–28, January 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304687.

Guan:2019:LWB

[1909] Albert Guan and Wen-Guey Tzeng. A light-weight bit commitment protocol based on unpredictable channel noise. Theoretical Computer Science, 755(??): 29–37, January 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304675.

Wang:2019:MPNa

[1910] Zhao Wang, Christopher Melekian, Eddie Cheng, and Yaping Mao. Matching preclusion number in product graphs. Theoretical Computer Science, 755(??): 38–47, January 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304663.

Leitert:2019:PAA

[1911] Arne Leitert and Feodor F. Dragan. Parameterized approximation algorithms for some location problems in graphs. Theoretical Computer Science, 755(??):48–64, January 10, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304651.

Zhang:2019:EPK

[1912] Kai Zhang, Jie Chen, Hyung Tae Lee, Haifeng Qian, and Huaxiong Wang. Efficient public key encryption with equality test in the standard model. Theoretical Computer Science, 755(??): 65–80, January 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830464X.

Guo:2019:GNC

[1913] Jia Guo, Desai Li, and Mei Lu. The g-good-neighbor conditional diagnosability of the crossed cubes under the PMC and MM* model. Theoretical Computer Science, 755(??): 81–88, January 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304626.

Jing:2019:PTA

[1914] Rui-Juan Jing, Chun-Ming Yuan, and Xiao-Shan Gao. A polynomialtime algorithm to compute generalized Hermite normal forms of matrices over Z[x]. Theoretical Computer Science, 755(??):89–109, January 10, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304845.

Anonymous:2019:EBd

[1915] Anonymous. Editorial Board. Theoretical Computer Science, 756(??): ii, January 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307461.

Baghoolizadeh:2019:UBA

[1916] Shirin Baghoolizadeh, Ali Movaghar, and Negin Majidi. A uniformizationbased algorithm for continuous-time stochastic games model checking. Theoretical Computer Science, 756(??): 1–18, January 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306145.

Ribeiro:2019:APC

[1917] Pedro Ribeiro and Ana Cavalcanti. Angelic processes for CSP via the UTP. Theoretical Computer Science, 756(??): 19–63, January 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306133.

Marinkovic:2019:PPC

[1918] Bojan Marinković, Paola Glavan, and Zoran Ognjanović. Proving properties of the Chord protocol using the ASM formalism. Theoretical Computer Science, 756(??):64– 93, January 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306467.

Anonymous:2019:EBe

[1919] Anonymous. Editorial Board. Theoretical Computer Science, 757(??): ii, January 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830759X.

Czygrinow:2019:DCC

[1920] Andrzej Czygrinow, Michał Hanćkowiak, Wojciech Wawrzyniak, and Marcin Witkowski. Distributed CONGESTBC constant approximation of MDS in bounded genus graphs. Theoretical Computer Science, 757(??):1– 10, January 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304894.

Chang:2019:HRF

[1921] Yi-Jun Chang. Hardness of RNA folding problem with four symbols. Theoretical Computer Science, 757(??): 11–26, January 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304912.

Bartal:2019:ANN

[1922] Yair Bartal and Lee-Ad Gottlieb. Approximate nearest neighbor search for l p-spaces (2 < p < ∞) via embeddings. Theoretical Computer Science, 757(??): 27–35, January 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304924.

daSilva:2019:OSW

[1923] Rodrigo Ferreira da Silva, Sebastián Urrutia, and Vinícius Fernandes dos Santos. One-sided weak dominance drawing. Theoretical Computer Science, 757(??):36–43, January 24, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304936.

Lv:2019:RSN

[1924] Mengjie Lv, Shuming Zhou, Xueli Sun, Guanqin Lian, and Jiafei Liu. Reliability of (n, k)-star network based on g-extra conditional fault. Theoretical Computer Science, 757(??): 44–55, January 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305000.

Cheng:2019:GAD

[1925] Eddie Cheng, Ke Qiu, and Zhizhang Shen. A general approach to deriving the g-good-neighbor conditional diagnosability of interconnection networks. Theoretical Computer Science, 757(??): 56–67, January 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305012.

Anonymous:2019:EBf

[1926] Anonymous. Editorial Board. Theoretical Computer Science, 758(??): ii, February 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930012X.

Zhu:2019:HFD

[1927] Qiang Zhu, Lili Li, Sanyang Liu, and Xing Zhang. Hybrid fault diagnosis capability analysis of hypercubes under the PMC model and MM∗ model. Theoretical Computer Science, 758(??): 1–8, February 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305024.

Sun:2019:EFT

[1928] Gaoxing Sun, Shuang Zhao, and Jixiang Meng. Edge fault tolerance of interconnection networks with respect to maximally edge-connectivity. Theoretical Computer Science, 758(??): 9–16, February 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305036.

Bonomi:2019:AAU

[1929] Silvia Bonomi, Antonella Del Pozzo, Maria Potop-Butucaru, and Sébastien Tixeuil. Approximate agreement under mobile Byzantine faults. Theoretical Computer Science, 758(??): 17–29, February 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305176.

Escoffier:2019:SCM

[1930] Bruno Escoffier. Saving colors and max coloring: Some fixedparameter tractability results. Theoretical Computer Science, 758(??): 30–41, February 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305188.

Stipulanti:2019:CPL

[1931] Manon Stipulanti. Convergence of Pascal-like triangles in Parry– Bertrand numeration systems. Theoretical Computer Science, 758(??): 42–60, February 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830519X.

Li:2019:ACM

[1932] Shimin Li and Haitao Wang. Algorithms for covering multiple barriers. Theoretical Computer Science, 758(??): 61–72, February 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305206.

Goldberg:2019:AOA

[1933] Leslie Ann Goldberg, John Lapinskas, Johannes Lengler, Florian Meier, Konstantinos Panagiotou, and Pascal Pfister. Asymptotically optimal amplifiers for the Moran process. Theoretical Computer Science, 758(??): 73–93, February 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305218.

Anonymous:2019:EBg

[1934] Anonymous. Editorial Board. Theoretical Computer Science, 759(??): ii, February 8, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300374.

Huang:2019:LRR

[1935] Jianye Huang, Qiong Huang, and Willy Susilo. Leakage-resilient ring signature schemes. Theoretical Computer Science, 759(??):1– 13, February 8, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300210.

Crespelle:2019:FDR

[1936] Christophe Crespelle. Fully dynamic representations of interval graphs. Theoretical Computer Science, 759(??): 14–49, February 8, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300209.

Lajou:2019:ANS

[1937] Dimitri Lajou. On the achromatic number of signed graphs. Theoretical Computer Science, 759(??): 50–60, February 8, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300192.

Wang:2019:MPNb

[1938] Zhao Wang, Yaping Mao, Eddie Cheng, and Jinyu Zou. Matching preclusion number of graphs. Theoretical Computer Science, 759(??):61– 71, February 8, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300180.

Libert:2019:ZKA

[1939] Benoît Libert, San Ling, Fabrice Mouhartem, Khoa Nguyen, and Huaxiong Wang. Zero-knowledge arguments for matrix-vector relations and lattice-based group encryption. Theoretical Computer Science, 759(??): 72–97, February 8, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300167.

Faonio:2019:CNM

[1940] Antonio Faonio, Jesper Buus Nielsen, Mark Simkin, and Daniele Venturi. Continuously non-malleable codes with split-state refresh. Theoretical Computer Science, 759(??): 98–132, February 8, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300064.

Anonymous:2019:EBh

[1941] Anonymous. Editorial Board. Theoretical Computer Science, 760(??): ii, February 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930060X.

Wei:2019:HFD

[1942] Yulong Wei and Min Xu. Hybrid fault diagnosis capability analysis of regular graphs. Theoretical Computer Science, 760(??):1– 14, February 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305231.

Lafond:2019:CCM

[1943] M. Lafond, N. El-Mabrouk, K. T. Huber, and V. Moulton. The complexity of comparing multiply-labelled trees by extending phylogenetic-tree metrics. Theoretical Computer Science, 760(??): 15–34, February 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830522X.

Bringmann:2019:GIR

[1944] Karl Bringmann, Ralph Keusch, and Johannes Lengler. Geometric inhomogeneous random graphs. Theoretical Computer Science, 760(??):35– 54, February 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305309.

Yang:2019:CBD

[1945] Yongjie Yang, Yash Raj Shrestha, and Jiong Guo. On the complexity of bribery with distance restrictions. Theoretical Computer Science, 760(??): 55–71, February 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305292.

Dosa:2019:RAS

[1946] Gyorgy Dosa, Hans Kellerer, and Zsolt Tuza. Restricted assignment scheduling with resource constraints. Theoretical Computer Science, 760(??): 72–87, February 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305322.

Anonymous:2019:EBi

[1947] Anonymous. Editorial Board. Theoretical Computer Science, 761(??): ii, February 21, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300714.

Li:2019:CFT

[1948] Xiang-Jun Li, Min Liu, Zheng Yan, and Jun-Ming Xu. On conditional fault tolerance of hierarchical cubic networks. Theoretical Computer Science, 761(??): 1–6, February 21, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305048.

Liu:2019:MPE

[1949] Wen An Liu and Tao Wu. Multiplayer End-Nim games. Theoretical Computer Science, 761(??):7– 22, February 21, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305346.

Lu:2019:HPP

[1950] Huazhong Lü and Fan Wang. Hamiltonian paths passing through prescribed edges in balanced hypercubes. Theoretical Computer Science, 761(??): 23–33, February 21, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305334.

Khonji:2019:CDS

[1951] Majid Khonji, Areg Karapetyan, Khaled Elbassioni, and Sid Chi-Kin Chau. Complex-demand scheduling problem with application in smart grid. Theoretical Computer Science, 761(??): 34–50, February 21, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305395.

Takayasu:2019:PKE

[1952] Atsushi Takayasu and Noboru Kunihiro. Partial key exposure attacks on RSA: Achieving the Boneh– Durfee bound. Theoretical Computer Science, 761(??):51–77, February 21, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305371.

Gu:2019:CQT

[1953] Qian-Ping Gu and Gengchun Xu. Constant query time (1 + ϵ)-approximate distance oracle for planar graphs. Theoretical Computer Science, 761(??): 78–88, February 21, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305401.

Anonymous:2019:EBj

[1954] Anonymous. Editorial Board. Theoretical Computer Science, 762(??): ii, March 1, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930088X.

He:2019:TRB

[1955] Kun He, Qian Li, and Xiaoming Sun. A tighter relation between sensitivity complexity and certificate complexity. Theoretical Computer Science, 762 (??):1–12, March 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305413.

Sabeti:2019:PEN

[1956] Rostam Sabeti. Polynomial expressions for non-binomial structures. Theoretical Computer Science, 762(??): 13–24, March 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306418.

Kawachi:2019:GPE

[1957] Akinori Kawachi, Mitsunori Ogihara, and Kei Uchizawa. Generalized predecessor existence problems for Boolean finite dynamical systems on directed graphs. Theoretical Computer Science, 762(??):25–40, March 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305589.

Navarro:2019:UCT

[1958] Gonzalo Navarro and Nicola Prezza. Universal compressed text indexing. Theoretical Computer Science, 762 (??):41–50, March 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305723.

Kucera:2019:LBC

[1959] Petr Kucera, Petr Savický, and Vojtech Vorel. A lower bound on CNF encodings of the at-most-one constraint. Theoretical Computer Science, 762 (??):51–73, March 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305681.

Anonymous:2019:EBk

[1960] Anonymous. Editorial Board. Theoretical Computer Science, 763(??): ii, April 4, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300994.

He:2019:WAI

[1961] Qingyu He and Luoshan Xu. Weak algebraic information systems and a new equivalent category of DOM of domains. Theoretical Computer Science, 763(??):1–11, April 4, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300830.

deSilva:2019:HLS

[1962] Lavindra de Silva, Lin Padgham, and Sebastian Sardina. HTN-like solutions for classical planning problems: an application to BDI agent systems. Theoretical Computer Science, 763 (??):12–37, April 4, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300672.

Bodei:2019:FAO

[1963] Chiara Bodei, Linda Brodo, and Roberto Bruni. A formal approach to open multiparty interactions. Theoretical Computer Science, 763(??): 38–65, April 4, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300660.

Adamek:2019:FPC

[1964] Jirí Adámek and Stefan Milius. On functors preserving coproducts and algebras with iterativity. Theoretical Computer Science, 763(??):66– 87, April 4, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930043X.

Anonymous:2019:EBl

[1965] Anonymous. Editorial Board. Theoretical Computer Science, 764(??): ii, April 11, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301100.

Bilo:2019:E

[1966] Vittorio Bilò and Antonio Caruso. Editorial. Theoretical Computer Science, 764(??):1, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301069.

Calamoneri:2019:SLT

[1967] Tiziana Calamoneri, Matteo Dell’Orefice, and Angelo Monti. A simple linear time algorithm for the locally connected spanning tree problem on maximal planar chordal graphs. Theoretical Computer Science, 764(??): 2–14, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301221.

Cordasco:2019:AIS

[1968] Gennaro Cordasco, Luisa Gargano, and Adele A. Rescigno. Active influence spreading in social networks. Theoretical Computer Science, 764 (??):15–29, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830121X.

DAngelo:2019:RLT

[1969] Gianlorenzo D’Angelo, Lorenzo Severini, and Yllka Velaj. Recommending links through influence maximization. Theoretical Computer Science, 764(??):30–41, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300641.

Das:2019:GRR

[1970] Shantanu Das, Riccardo Focardi, Flaminia L. Luccio, Euripides Markou, and Marco Squarcina. Gathering of robots in a ring with mobile faults. Theoretical Computer Science, 764 (??):42–60, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518303086.

Vinci:2019:NAO

[1971] Cosimo Vinci. Non-atomic one-round walks in congestion games. Theoretical Computer Science, 764(??): 61–79, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304547.

Bernardo:2019:CLC

[1972] Marco Bernardo and Marino Miculan. Constructive logical characterizations of bisimilarity for reactive probabilistic systems. Theoretical Computer Science, 764(??): 80–99, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307217.

Bodei:2019:MSI

[1973] Chiara Bodei, Stefano Chessa, and Letterio Galletta. Measuring security in IoT communications. Theoretical Computer Science, 764(??): 100–124, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307205.

Bozzelli:2019:WFI

[1974] Laura Bozzelli, Alberto Molinari, Angelo Montanari, Adriano Peron, and Pietro Sala. Which fragments of the interval temporal logic HS are tractable in model checking? Theoretical Computer Science, 764(??): 125–144, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302330.

Giannini:2019:FRU

[1975] Paola Giannini, Marco Servetto, Elena Zucca, and James Cone. Flexible recovery of uniqueness and immutability. Theoretical Computer Science, 764 (??):145–172, April 11, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305668.

Anonymous:2019:EBm

[1976] Anonymous. Editorial Board. Theoretical Computer Science, 765(??): ii, April 18, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301434.

Kari:2019:Ea

[1977] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 765(??):iii, April 18, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930146X.

Feret:2019:P

[1978] Jérôme Feret, Loïc Paulevé, and David Safránek. Preface. Theoretical Computer Science, 765(??):1– 2, April 18, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302464.

Johnson:2019:VCR

[1979] Robert Johnson, Qing Dong, and Erik Winfree. Verifying chemical reaction network implementations: a bisimulation approach. Theoretical Computer Science, 765(??):3– 46, April 18, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300136.

Cardelli:2019:CCR

[1980] Luca Cardelli, Mirco Tribastone, Max Tschaikowski, and Andrea Vandin. Comparing chemical reaction networks: a categorical and algorithmic perspective. Theoretical Computer Science, 765(??):47–66, April 18, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517309131.

Shin:2019:VCR

[1981] Seung Woo Shin, Chris Thachuk, and Erik Winfree. Verifying chemical reaction network implementations: a pathway decomposition approach. Theoretical Computer Science, 765(??): 67–96, April 18, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517307399.

Honorato-Zimmer:2019:CLP

[1982] Ricardo Honorato-Zimmer, Andrew J.

Millar, Gordon D. Plotkin, and Argyris Zardilis. Chromar, a language of parameterised agents. Theoretical Computer Science, 765(??): 97–119, April 18, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517305996.

Kolcak:2019:PSA

[1983] Juraj Kolcák, David Safránek, Stefan Haar, and Loïc Paulevé. Parameter space abstraction and unfolding semantics of discrete regulatory networks. Theoretical Computer Science, 765(??): 120–144, April 18, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518301701.

Bernot:2019:GMH

[1984] G. Bernot, J.-P. Comet, Z. Khalis, A. Richard, and O. Roux. A genetically modified Hoare logic. Theoretical Computer Science, 765(??): 145–157, April 18, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830080X.

Islam:2019:PRM

[1985] Md. Ariful Islam, Rance Cleaveland, Flavio H. Fenton, Radu Grosu, Paul L.

Jones, and Scott A. Smolka. Probabilistic reachability for multi-parameter bifurcation analysis of cardiac alternans. Theoretical Computer Science, 765(??): 158–169, April 18, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518300823.

Anonymous:2019:EBn

[1986] Anonymous. Editorial Board. Theoretical Computer Science, 766(??): ii, April 25, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301598.

Bazgan:2019:PAC

[1987] Cristina Bazgan, Florent Foucaud, and Florian Sikora. Parameterized and approximation complexity of Partial VC Dimension. Theoretical Computer Science, 766(??): 1–15, April 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830584X.

Li:2019:ECE

[1988] Xiaoyan Li, Jianxi Fan, Cheng-Kuan Lin, Baolei Cheng, and Xiaohua Jia. The extra connectivity, extra conditional diagnosability and t/k-diagnosability of the data center network DCell. Theoretical Computer Science, 766(??): 16–29, April 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305851.

Chitturi:2019:SPT

[1989] Bhadrachalam Chitturi and Priyanshu Das. Sorting permutations with transpositions in O(n^{3}) amortized time. Theoretical Computer Science, 766 (??):30–37, April 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305863.

Chang:2019:CCA

[1990] Jou-Ming Chang, Kung-Jui Pai, Ro-Yu Wu, and Jinn-Shyong Yang. The 4-component connectivity of alternating group networks. Theoretical Computer Science, 766(??):38– 45, April 25, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305899.

Henning:2019:AAS

[1991] Michael A. Henning and Arti Pandey.

Algorithmic aspects of semitotal domination in graphs. Theoretical Computer Science, 766(??):46– 57, April 25, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305905.

Anonymous:2019:EBo

[1992] Anonymous. Editorial Board. Theoretical Computer Science, 767(??): i, May 3, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301744.

Anonymous:2019:EBp

[1993] Anonymous. Editorial Board. Theoretical Computer Science, 767(??): ii, May 3, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301756.

Yang:2019:EFF

[1994] Yuxing Yang, Jing Li, and Shiying Wang. Embedding fault-free Hamiltonian paths with prescribed linear forests into faulty ternary n-cubes. Theoretical Computer Science, 767 (??):1–15, May 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305917.

Xiao:2019:IPT

[1995] Yuzhi Xiao, Haixing Zhao, Yaping Mao, and Guanrong Chen. Invulnerability of planar two-tree networks. Theoretical Computer Science, 767 (??):16–25, May 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305929.

Elbassioni:2019:GPA

[1996] Khaled Elbassioni, Imran Rauf, and Saurabh Ray. A global parallel algorithm for enumerating minimal transversals of geometric hypergraphs. Theoretical Computer Science, 767 (??):26–33, May 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830598X.

Zhang:2019:RNV

[1997] Zhen Zhang, Shu qiang Huang, Dong Guo, and Yong-Hui Li. RVCCC: a new variational network of cube-connected cycles and its topological properties. Theoretical Computer Science, 767 (??):34–50, May 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305991.

Jiang:2019:EIB

[1998] Peng Jiang, Fuchun Guo, and Yi Mu.

Efficient identity-based broadcast encryption with keyword search against insider attacks for database systems. Theoretical Computer Science, 767 (??):51–72, May 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306017.

Ghosal:2019:RMM

[1999] Pratik Ghosal, Meghana Nasre, and Prajakta Nimbhorkar. Rank-maximal matchings — structure and algorithms. Theoretical Computer Science, 767 (??):73–82, May 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306042.

Hudry:2019:UOS

[2000] Olivier Hudry and Antoine Lobstein.

Unique (optimal) solutions: Complexity results for identifying and locating-dominating codes. Theoretical Computer Science, 767(??): 83–102, May 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306054.

Anonymous:2019:EBq

[2001] Anonymous. Editorial Board. Theoretical Computer Science, 768(??): ii, May 10, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301847.

Howe:2019:ICS

[2002] Jacob M. Howe, Andy King, and Axel Simon. Incremental closure for systems of two variables per inequality. Theoretical Computer Science, 768 (??):1–42, May 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307199.

Vatandoost:2019:GML

[2003] Mehdi Vatandoost, Ali Akbar Estaji, and Rahimeh Pourkhandani. A generalized modal logic in causal structures. Theoretical Computer Science, 768(??):43–53, May 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301045.

Cintula:2019:SHT

[2004] Petr Cintula, Denisa Diaconescu, and George Metcalfe. Skolemization and Herbrand theorems for lattice-valued logics. Theoretical Computer Science, 768(??):54–75, May 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301057.

Jay:2019:ICH

[2005] Barry Jay. Intensional computation with higher-order functions. Theoretical Computer Science, 768(??): 76–90, May 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301227.

Strassburger:2019:DPM

[2006] Lutz Straßburger. On the decision problem for MELL. Theoretical Computer Science, 768(??): 91–98, May 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301288.

Hirsch:2019:AFQ

[2007] Robin Hirsch, Marcel Jackson, and Tomasz Kowalski. Algebraic foundations for qualitative calculi and networks. Theoretical Computer Science, 768(??):99–116, May 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301471.

Anonymous:2019:EBr

[2008] Anonymous. Editorial Board. Theoretical Computer Science, 769(??): ii, May 17, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302051.

Lafond:2019:WQC

[2009] Manuel Lafond and Celine Scornavacca. On the Weighted Quartet Consensus problem. Theoretical Computer Science, 769(??):1– 17, May 17, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306108.

deBerg:2019:CDS

[2010] Mark de Berg, Sándor Kisfaludi-Bak, and Gerhard Woeginger. The complexity of Dominating Set in geometric intersection graphs. Theoretical Computer Science, 769(??): 18–31, May 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306121.

Bazgan:2019:FPC

[2011] Cristina Bazgan, Thomas Pontoizeau, and Zsolt Tuza. Finding a potential community in networks. Theoretical Computer Science, 769(??): 32–42, May 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306285.

Dolev:2019:PSM

[2012] Shlomi Dolev, Juan A. Garay, Niv Gilboa, Vladimir Kolesnikov, and Muni Venkateswarlu Kumaramangalam. Perennial secure multi-party computation of universal Turing machine. Theoretical Computer Science, 769(??):43–62, May 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306297.

Sadhu:2019:LTA

[2013] Sanjib Sadhu, Sasanka Roy, Subhas C. Nandy, and Suchismita Roy. Linear time algorithm to cover and hit a set of line segments optimally by two axis-parallel squares. Theoretical Computer Science, 769(??): 63–74, May 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306303. See

corrigendum [2415].

Anonymous:2019:EBs

[2014] Anonymous. Editorial Board. Theoretical Computer Science, 770(??): ii, May 24, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302221.

Pena:2019:EDO

[2015] N. Pena and A. Borodin. On extensions of the deterministic online model for bipartite matching and maxsat. Theoretical Computer Science, 770 (??):1–24, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306327.

Manoussakis:2019:NDT

[2016] George Manoussakis. A new decomposition technique for maximal clique enumeration for sparse graphs. Theoretical Computer Science, 770 (??):25–33, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306315.

Gagie:2019:PQF

[2017] Travis Gagie, Meng He, and Gonzalo Navarro. Path queries on functions. Theoretical Computer Science, 770(??):34–50, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830642X.

Durocher:2019:SLS

[2018] Stephane Durocher and Robby Singh. A simple linear-space data structure for constant-time range minimum query. Theoretical Computer Science, 770 (??):51–61, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306406.

Telle:2019:FAD

[2019] J. A. Telle and Y. Villanger. FPT algorithms for domination in sparse graphs and beyond. Theoretical Computer Science, 770(??):62– 68, May 24, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306510.

Le:2019:CDM

[2020] Hoang-Oanh Le and Van Bang Le. A complexity dichotomy for matching cut in (bipartite) graphs of fixed diameter. Theoretical Computer Science, 770 (??):69–78, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306509.

Kowaluk:2019:FDD

[2021] Miroslaw Kowaluk and Andrzej Lingas.

A fast deterministic detection of small pattern graphs in graphs without large cliques. Theoretical Computer Science, 770(??):79–87, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306492.

Lazard:2019:ELR

[2022] Sylvain Lazard, William J. Lenhart, and Giuseppe Liotta. On the edgelength ratio of outerplanar graphs. Theoretical Computer Science, 770 (??):88–94, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306078. See

corrigendum [2416].

Harju:2019:SFA

[2023] Tero Harju. On square-free arithmetic progressions in infinite words. Theoretical Computer Science, 770 (??):95–100, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306030.

Vitanyi:2019:CRD

[2024] P. M. B. Vitányi. Corrigendum to “On the rate of decrease in logical depth” [Theor. Comput. Sci. 702 (2017) 60–64] by L. F. Antunes, A. Souto, and P. M. B. Vitányi. Theoretical Computer Science, 770 (??):101, May 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304900. See

[1467].

Anonymous:2019:EBt

[2025] Anonymous. Editorial Board. Theoretical Computer Science, 771(??): ii, June 1, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302282.

Gharahi:2019:OLS

[2026] Motahhareh Gharahi and Shahram Khazaei. Optimal linear secret sharing schemes for graph access structures on six participants. Theoretical Computer Science, 771(??): 1–8, June 1, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306789.

Ravelo:2019:PMC

[2027] Santiago Valdés Ravelo and Carlos Eduardo Ferreira. A PTAS for the metric case of the optimum weighted source-destination communication spanning tree problem. Theoretical Computer Science, 771 (??):9–22, June 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306790.

Rusu:2019:MCF

[2028] Irena Rusu. Min (a)cyclic feedback vertex sets and min ones monotone 3-SAT. Theoretical Computer Science, 771(??):23–38, June 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306807.

Araujo:2019:WPO

[2029] Júlio Araújo, Cláudia Linhares Sales, Ignasi Sau, and Ana Silva. Weighted proper orientations of trees and graphs of bounded treewidth. Theoretical Computer Science, 771(??):39– 48, June 1, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307035.

Aung:2019:FHE

[2030] Khin Mi Mi Aung, Hyung Tae Lee, Benjamin Hong Meng Tan, and Huaxiong Wang. Fully homomorphic encryption over the integers for non-binary plaintexts without the sparse subset sum problem. Theoretical Computer Science, 771 (??):49–70, June 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307047.

Dinesh:2019:ASS

[2031] Krishnamoorthy Dinesh and Jayalal Sarma. Alternation, sparsity and sensitivity: Bounds and exponential gaps. Theoretical Computer Science, 771(??):71–82, June 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307059.

Farbstein:2019:DT

[2032] Boaz Farbstein and Asaf Levin. Deadline TSP. Theoretical Computer Science, 771(??):83–92, June 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397518307060.

Dong:2019:HMT

[2033] Qiang Dong and Xi Wang. How many triangles and quadrilaterals are there in an n-dimensional augmented cube? Theoretical Computer Science, 771(??):93–98, June 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306686.

Anonymous:2019:EBu

[2034] Anonymous. Editorial Board. Theoretical Computer Science, 772(??): ii, June 7, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930235X.

Zhou:2019:GCC

[2035] Yanwei Zhou, Bo Yang, and Yi Mu. The generic construction of continuous leakage-resilient identity-based cryptosystems. Theoretical Computer Science, 772(??):1–45, June 7, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397518307096.

Ehard:2019:VYT

[2036] Stefan Ehard and Dieter Rautenbach. Vaccinate your trees! Theoretical Computer Science, 772(??): 46–57, June 7, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307084.

Navarro:2019:DLR

[2037] Gonzalo Navarro. Document listing on repetitive collections with guaranteed performance. Theoretical Computer Science, 772(??):58– 72, June 7, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307126.

Alizadeh:2019:IOM

[2038] Behrooz Alizadeh, Esmaeil Afrashteh, and Fahimeh Baroughi. Inverse obnoxious p-median location problems on trees with edge length modifications under different norms. Theoretical Computer Science, 772(??): 73–87, June 7, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307102.

Bournat:2019:SSR

[2039] Marjorie Bournat, Ajoy K. Datta, and Swan Dubois. Self-stabilizing robots in highly dynamic environments. Theoretical Computer Science, 772(??):88–110, June 7, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307163.

Ito:2019:RCS

[2040] Takehiro Ito and Yota Otachi. Reconfiguration of colorable sets in classes of perfect graphs. Theoretical Computer Science, 772(??): 111–122, June 7, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830714X.

Xue:2019:ACG

[2041] Jie Xue, Huiqiu Lin, and Jinlong Shu. The algebraic connectivity of graphs with given circumference. Theoretical Computer Science, 772(??): 123–131, June 7, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307266.

Kolay:2019:HCP

[2042] Sudeshna Kolay, Ragukumar Pandurangan, Fahad Panolan, Venkatesh Raman, and Prafullkumar Tale. Harmonious coloring: Parameterized algorithms and upper bounds. Theoretical Computer Science, 772(??): 132–142, June 7, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307321.

Le:2019:MGH

[2043] Hoang-Oanh Le and Van Bang Le. Map graphs having witnesses of large girth. Theoretical Computer Science, 772(??): 143–148, June 7, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830731X.

Anonymous:2019:EBv

[2044] Anonymous. Editorial Board. Theoretical Computer Science, 773(??): ii, June 14, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302427.

Espitia:2019:CPC

[2045] Helbert Eduardo Espitia and Jorge Iván Sofrony. Considerations for parameter configuration on Vortex Particle Swarm Optimization. Theoretical Computer Science, 773(??): 1–26, June 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303402.

Dinh:2019:CCD

[2046] Hai Q. Dinh, Abhay Kumar Singh, Sukhamoy Pattanayak, and Songsak Sriboonchitta. Construction of cyclic DNA codes over the ring Z_{4}[u]/⟨u_{2} − 1⟩ based on the deletion distance. Theoretical Computer Science, 773 (??):27–42, June 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304109.

PourAliAkbar:2019:EDR

[2047] Ehsan PourAliAkbar and Mohammad Mosleh. An efficient design for reversible Wallace unsigned multiplier. Theoretical Computer Science, 773 (??):43–52, June 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304158.

Oliveto:2019:BRU

[2048] Pietro S. Oliveto, Dirk Sudholt, and Christine Zarges. On the benefits and risks of using fitness sharing for multimodal optimisation. Theoretical Computer Science, 773(??): 53–70, June 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304882.

Sartakhti:2019:CPL

[2049] Javad Salimi Sartakhti and Saeed Jalili. On the computational power of the light: a plan for breaking Data Encryption Standard. Theoretical Computer Science, 773(??): 71–78, June 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305310.

Becker:2019:TRS

[2050] Ruben Becker, Vincenzo Bonifaci, Andreas Karrenbauer, Pavel Kolev, and Kurt Mehlhorn. Two results on slime mold computations. Theoretical Computer Science, 773(??): 79–106, June 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305590.

Wang:2019:GNE

[2051] Shiying Wang and Mujiangshan Wang.

The g-good-neighbor and g-extra diagnosability of networks. Theoretical Computer Science, 773(??): 107–114, June 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830567X.

Doerr:2019:ARS

[2052] Benjamin Doerr. Analyzing randomized search heuristics via stochastic domination. Theoretical Computer Science, 773(??):115–137, June 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397518305954.

Pan:2019:EDG

[2053] Minghua Pan, Daowen Qiu, Paulo Mateus, and Jozef Gruska. Entangling and disentangling in Grover’s search algorithm. Theoretical Computer Science, 773(??):138–152, June 14, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397518306066.

Anonymous:2019:EBw

[2054] Anonymous. Editorial Board. Theoretical Computer Science, 774(??): ii, June 25, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302518.

Zhang:2019:DDL

[2055] Kejia Zhang, Qilong Han, Zhipeng Cai, Guisheng Yin, and Junyu Lin. DOAMI: a distributed on-line algorithm to minimize interference for routing in wireless sensor networks. Theoretical Computer Science, 774 (??):1–13, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302298.

Mahmoody:2019:OSA

[2056] Ahmad Mahmoody and Eli Upfal. Optimizing static and adaptive probing schedules for rapid event detection. Theoretical Computer Science, 774(??):14–30, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302274.

Ding:2019:ARC

[2057] Wei Ding and Ke Qiu. Approximating the restricted 1-center in graphs. Theoretical Computer Science, 774 (??):31–43, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302237.

Huang:2019:MDC

[2058] Fei Huang, Xueliang Li, Zhongmei Qin, and Colton Magnant. Minimum degree condition for proper connection number 2. Theoretical Computer Science, 774(??):44–50, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302213.

Kawase:2019:PCB

[2059] Yasushi Kawase, Xin Han, and Kazuhisa Makino. Proportional cost buyback problem with weight bounds. Theoretical Computer Science, 774 (??):51–64, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302419.

Tian:2019:MCO

[2060] Cong Tian and Zhenhua Duan. Model checking open systems with alternating projection temporal logic. Theoretical Computer Science, 774(??): 65–81, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302559.

Zhao:2019:MCP

[2061] Liang Zhao, Xiaobing Wang, and Zhenhua Duan. Model checking of pushdown systems for projection temporal logic. Theoretical Computer Science, 774(??):82–94, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397516302766.

Xu:2019:DMM

[2062] Yi Xu, Jigen Peng, Yinfeng Xu, and Binhai Zhu. The discrete and mixed minimax 2-center problems. Theoretical Computer Science, 774(??): 95–102, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516302936.

Miao:2019:VCC

[2063] Dongjing Miao, Xianmin Liu, Yingshu Li, and Jianzhong Li. Vertex cover in conflict graphs. Theoretical Computer Science, 774(??): 103–112, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303231.

Xu:2019:RLC

[2064] Chao Xu, Jianer Chen, and Jianxin Wang. Resolution and linear CNF formulas: Improved (n, 3)- MaxSAT algorithms. Theoretical Computer Science, 774(??):113–123, June 25, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303954.

Takaoka:2019:ATS

[2065] Tadao Takaoka. Algebraic theory on shortest paths for all flows. Theoretical Computer Science, 774(??): 124–132, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397516303942.

Armaselu:2019:DMB

[2066] Bogdan Armaselu and Ovidiu Daescu.

Dynamic minimum bichromatic separating circle. Theoretical Computer Science, 774(??):133–142, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397516307046.

Xu:2019:IAA

[2067] Yicheng Xu, Dachuan Xu, Donglei Du, and Chenchen Wu. Improved approximation algorithm for universal facility location problem with linear penalties. Theoretical Computer Science, 774 (??):143–151, June 25, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397517302086.

Anonymous:2019:EBx

[2068] Anonymous. Editorial Board. Theoretical Computer Science, 775(??): ii, July 5, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302609.

Khoshkhah:2019:CAE

[2069] Kaveh Khoshkhah, Mehdi Khosravian Ghadikolaei, Jérôme Monnot, and Dirk Oliver Theis. Complexity and approximability of extended Spanning Star Forest problems in general and complete graphs. Theoretical Computer Science, 775(??): 1–15, July 5, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307151.

Colucci:2019:TPC

[2070] Lucas Colucci, Péter L. Erdős, Ervin Györi, and Tamás Róbert Mezei. Terminal-pairability in complete bipartite graphs with non-bipartite demands: Edge-disjoint paths in complete bipartite graphs. Theoretical Computer Science, 775(??):16– 25, July 5, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751830728X.

Ahadi:2019:CGM

[2071] Arash Ahadi and Alireza Zarei. Connecting guards with minimum Steiner points inside simple polygons. Theoretical Computer Science, 775(??): 26–31, July 5, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307291.

Zhang:2019:GCT

[2072] Xiao Zhang, Shengli Liu, Dawu Gu, and Joseph K. Liu. A generic construction of tightly secure signatures in the multi-user setting. Theoretical Computer Science, 775(??):32– 52, July 5, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307333.

Chatterjee:2019:CCT

[2073] Tanima Chatterjee, Bhaskar Das-Gupta, Nasim Mobasheri, Venkatkumar Srinivasan, and Ismael G. Yero. On the computational complexities of three problems related to a privacy measure for large networks under active attack. Theoretical Computer Science, 775(??):53–67, July 5, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307345.

Bonardo:2019:NVF

[2074] Pierre Bonardo, Anna E. Frid, and Jeffrey Shallit. The number of valid factorizations of Fibonacci prefixes. Theoretical Computer Science, 775 (??):68–75, July 5, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307497.

Genc:2019:CSR

[2075] Begum Genc, Mohamed Siala, Gilles Simonin, and Barry O’Sullivan. Complexity study for the robust stable marriage problem. Theoretical Computer Science, 775(??):76– 92, July 5, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307503.

Gurski:2019:KPP

[2076] Frank Gurski, Carolin Rehs, and Jochen Rethmann. Knapsack problems: a parameterized point of view. Theoretical Computer Science, 775 (??):93–108, July 5, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307527.

Anonymous:2019:EBy

[2077] Anonymous. Editorial Board. Theoretical Computer Science, 776(??): ii, July 12, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302671.

Russo:2019:SST

[2078] Luís M. S. Russo. A study on splay trees. Theoretical Computer Science, 776(??):1–18, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307539.

Goyal:2019:BAG

[2079] Navin Goyal and Manoj Gupta. Better analysis of GREEDY binary search tree on decomposable sequences. Theoretical Computer Science, 776(??): 19–42, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307540.

Jain:2019:ICP

[2080] Sanjay Jain and Efim Kinber. Intrinsic complexity of partial learning. Theoretical Computer Science, 776 (??):43–63, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307552.

Bang-Jensen:2019:DCP

[2081] Jørgen Bang-Jensen and Stéphane Bessy. Degree-constrained 2-partitions of graphs. Theoretical Computer Science, 776(??):64–74, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397519300015.

Lintzmayer:2019:OCS

[2082] Carla Negri Lintzmayer, Flávio Keidi Miyazawa, and Eduardo Candido Xavier. Online circle and sphere packing. Theoretical Computer Science, 776 (??):75–94, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300179.

Bohm:2019:OPS

[2083] Martin Böhm, Marek Chrobak, Łukasz Jeż, Fei Li, Jiří Sgall, and Pavel Veselý. Online packet scheduling with bounded delay and lookahead. Theoretical Computer Science, 776(??): 95–113, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930026X.

Litow:2019:EDD

[2084] B. Litow. Existential Diophantine definability of string length. Theoretical Computer Science, 776(??): 114–116, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300271.

Wu:2019:STS

[2085] Guangwei Wu, Jianer Chen, and Jianxin Wang. Scheduling two-stage jobs on multiple flowshops. Theoretical Computer Science, 776(??): 117–124, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300428.

Park:2019:DPC

[2086] Jung-Heum Park, Jae-Hoon Kim, and Hyeong-Seok Lim. Disjoint path covers joining prescribed source and sink sets in interval graphs. Theoretical Computer Science, 776(??): 125–137, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300441.

Gu:2019:FDD

[2087] Mei-Mei Gu, Rong-Xia Hao, and Shuming Zhou. Fault diagnosability of data center networks. Theoretical Computer Science, 776(??): 138–147, July 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300453.

Anonymous:2019:EBz

[2088] Anonymous. Editorial Board. Theoretical Computer Science, 777(??): ii, July 19, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302762.

Ausiello:2019:P

[2089] Giorgio Ausiello, Lila Kari, Grzegorz Rozenberg, Donald Sannella, Paul Spirakis, and Pierre-Louis Curien. Preface. Theoretical Computer Science, 777(??):1–2, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302725.

Aceto:2019:WPF

[2090] Luca Aceto, Dario Della Monica, Ignacio Fábregas, and Anna Ingólfsdóttir. When are prime formulae characteristic? Theoretical Computer Science, 777(??):3–31, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307254.

Alvim:2019:AIF

[2091] Mário S. Alvim, Konstantinos Chatzikokolakis, Annabelle McIver, Carroll Morgan, Catuscia Palamidessi, and Geoffrey Smith. An axiomatization of information flow measures. Theoretical Computer Science, 777(??): 32–54, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306376.

Anselmo:2019:FSP

[2092] Marcella Anselmo, Dora Giammarresi, and Maria Madonia. Full sets of pictures to encode pictures. Theoretical Computer Science, 777(??): 55–68, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306522.

Beaton:2019:EFF

[2093] Nicholas R. Beaton, Mathilde Bouvel, Veronica Guerrini, and Simone Rinaldi. Enumerating five families of patternavoiding inversion sequences; and introducing the powered Catalan numbers. Theoretical Computer Science, 777(??):69–92, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300933.

Berthe:2019:BCS

[2094] Valérie Berthé and Paulina Cecchi Bernales. Balancedness and coboundaries in symbolic systems. Theoretical Computer Science, 777(??): 93–110, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305772.

Blanchette:2019:QAS

[2095] Benjamin Blanchette, Christian Choffrut, and Christophe Reutenauer. Quasi-automatic semigroups. Theoretical Computer Science, 777(??): 111–120, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300155.

Brzozowski:2019:SCP

[2096] Janusz A. Brzozowski, Sylvie Davies, and Abhishek Madan. State complexity of pattern matching in regular languages. Theoretical Computer Science, 777(??):121–131, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397518307357.

Cardelli:2019:SCD

[2097] Luca Cardelli, Mirco Tribastone, Max Tschaikowski, and Andrea Vandin. Symbolic computation of differential equivalences. Theoretical Computer Science, 777(??):132– 154, July 19, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301719.

Chen:2019:ASE

[2098] Jing Chen and Silvio Micali. Algorand:

a secure and efficient distributed ledger. Theoretical Computer Science, 777(??): 155–183, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930091X.

Coquand:2019:CND

[2099] Thierry Coquand. Canonicity and normalization for dependent type theory. Theoretical Computer Science, 777 (??):184–191, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300325.

Dassow:2019:OCI

[2100] Jürgen Dassow. On the orbit of closure-involution operations — the case of formal languages. Theoretical Computer Science, 777(??): 192–203, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830608X.

Dershowitz:2019:DCA

[2101] Nachum Dershowitz and Jean-Pierre Jouannaud. Drags: a compositional algebraic framework for graph rewriting. Theoretical Computer Science, 777 (??):204–231, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300544.

Dolce:2019:GLW

[2102] Francesco Dolce, Antonio Restivo, and Christophe Reutenauer. On generalized Lyndon words. Theoretical Computer Science, 777(??): 232–242, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307369.

Dondi:2019:TFD

[2103] Riccardo Dondi, Giancarlo Mauri, and Italo Zoppis. On the tractability of finding disjoint clubs in a network. Theoretical Computer Science, 777(??): 243–251, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930218X.

Droste:2019:WSR

[2104] Manfred Droste, Sven Dziadek, and Werner Kuich. Weighted simple reset pushdown automata. Theoretical Computer Science, 777(??): 252–259, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300337.

Endrullis:2019:BTR

[2105] Jörg Endrullis and Jan Willem Klop.

Braids via term rewriting. Theoretical Computer Science, 777(??): 260–295, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307278.

Engeler:2019:NAH

[2106] E. Engeler. Neural algebra on “how does the brain think?”. Theoretical Computer Science, 777(??): 296–307, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302117.

Etessami:2019:RSG

[2107] Kousha Etessami, Dominik Wojtczak, and Mihalis Yannakakis. Recursive stochastic games with positive rewards. Theoretical Computer Science, 777(??): 308–328, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307515.

Frosini:2019:TRC

[2108] Andrea Frosini and Laurent Vuillon. Tomographic reconstruction of 2-convex polyominoes using dual Horn clauses. Theoretical Computer Science, 777(??):329–337, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397519300076.

Gerard:2019:RSC

[2109] Yan Gerard. Regular switching components. Theoretical Computer Science, 777(??):338–355, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397519300234.

Ibarra:2019:CFS

[2110] Oscar H. Ibarra, Ian McQuillan, and Bala Ravikumar. On counting functions and slenderness of languages. Theoretical Computer Science, 777(??): 356–378, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518307114.

Kari:2019:NCP

[2111] Jarkko Kari and Etienne Moutot. Nivat’s conjecture and pattern complexity in algebraic subshifts. Theoretical Computer Science, 777(??): 379–386, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300088.

Kimoto:2019:MCR

[2112] Nobuya Kimoto, Ken Komiya, Kenzo Fujimoto, and Satoshi Kobayashi. Monotonically controlling right linear grammars with unknown behaviors to output a target string. Theoretical Computer Science, 777(??): 387–408, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300222.

Kobayashi:2019:IBF

[2113] Naoki Kobayashi. Inclusion between the frontier language of a non-deterministic recursive program scheme and the Dyck language is undecidable. Theoretical Computer Science, 777(??):409–416, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397519300283.

Okhotin:2019:EDN

[2114] Alexander Okhotin and Kai Salomaa.

Edit distance neighbourhoods of inputdriven pushdown automata. Theoretical Computer Science, 777(??): 417–430, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301525.

Okubo:2019:DFC

[2115] Fumiya Okubo and Takashi Yokomori. Decomposition and factorization of chemical reaction transducers. Theoretical Computer Science, 777 (??):431–442, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300659.

Orellana-Martin:2019:PCE

[2116] David Orellana-Martín, Luis Valencia-Cabrera, Agustín Riscos-Núñez, and Mario J. Pérez-Jiménez. A path to computational efficiency through membrane computing. Theoretical Computer Science, 777(??):443– 453, July 19, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300027.

Rampersad:2019:CEI

[2117] Narad Rampersad, Jeffrey Shallit, and Élise Vandomme. Critical exponents of infinite balanced words. Theoretical Computer Science, 777(??): 454–463, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306388.

vanLeeuwen:2019:QAH

[2118] Jan van Leeuwen and Jirí Wiedermann. Question answering by humans and machines: a complexity-theoretic view. Theoretical Computer Science, 777(??): 464–473, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305280.

Wu:2019:CPA

[2119] Tingfang Wu, Linqiang Pan, and Artiom Alhazov. Computation power of asynchronous spiking neural P systems with polarizations. Theoretical Computer Science, 777(??): 474–489, July 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306455.

Anonymous:2019:EBba

[2120] Anonymous. Editorial Board. Theoretical Computer Science, 778(??): ii, July 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302932.

Zhao:2019:PCG

[2121] Xiao Zhao and Sheng Chen. Proper 2-coloring game on some trees. Theoretical Computer Science, 778(??): 1–18, July 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300465.

Giannakopoulos:2019:ASM

[2122] Yiannis Giannakopoulos, Elias Koutsoupias, and Maria Kyropoulou. The anarchy of scheduling without money. Theoretical Computer Science, 778 (??):19–32, July 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300477.

Kline:2019:GSH

[2123] Jeffery Kline. Geometric search for Hadamard matrices. Theoretical Computer Science, 778(??):33– 46, July 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300507.

Mohan:2019:PDC

[2124] Chris Monica Mohan, Samivel Santhakumar, Micheal Arockiaraj, and Jia-Bao Liu. Partition dimension of certain classes of series parallel graphs. Theoretical Computer Science, 778 (??):47–60, July 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300519.

Kao:2019:TAP

[2125] Mong-Jen Kao, Jia-Yau Shiau, Ching-Chi Lin, and D. T. Lee. Tight approximation for partial vertex cover with hard capacities. Theoretical Computer Science, 778(??):61– 72, July 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300520.

Heckenberger:2019:PLB

[2126] I. Heckenberger and J. Sawada. A Pascal-like bound for the number of necklaces with fixed density. Theoretical Computer Science, 778(??): 73–77, July 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300556.

Vitanyi:2019:LDR

[2127] Paul M. B. Vitányi. Logical depth for reversible Turing machines with an application to the rate of decrease in logical depth for general Turing machines. Theoretical Computer Science, 778(??):78–80, July 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300568.

Anonymous:2019:EBbb

[2128] Anonymous. Editorial Board. Theoretical Computer Science, 779(??): ii, August 2, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303305.

Abam:2019:VTC

[2129] Mohammad Ali Abam, Sharareh Alipour, Mohammad Ghodsi, and Mohammad Mahdian. Visibility testing and counting for uncertain segments. Theoretical Computer Science, 779(??):1–7, August 2, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300763.

Zhang:2019:PCF

[2130] Guozhen Zhang and Shangwei Lin.

Path and cycle fault tolerance of bubble-sort graph networks. Theoretical Computer Science, 779(??): 8–16, August 2, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300775.

Cicerone:2019:AAD

[2131] Serafino Cicerone and Gabriele Di Stefano. Approximation algorithms for decomposing octilinear polygons. Theoretical Computer Science, 779(??): 17–36, August 2, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300787.

Mohabeddine:2019:NRT

[2132] Amine Mohabeddine and Mourad Boudhar. New results in two identical machines scheduling with agreement graphs. Theoretical Computer Science, 779(??):37–46, August 2, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300799.

Han:2019:QAD

[2133] Shuai Han, Shengli Liu, and Lin Lyu. QANIZK for adversary-dependent languages and their applications. Theoretical Computer Science, 779(??): 47–71, August 2, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300829.

Bhattacharya:2019:NAC

[2134] Sayan Bhattacharya, Monika Henzinger, and Stefan Neumann. New amortized cell-probe lower bounds for dynamic problems. Theoretical Computer Science, 779(??):72– 87, August 2, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300842.

Anonymous:2019:EBbc

[2135] Anonymous. Editorial Board. Theoretical Computer Science, 780(??): ii, August 9, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303391.

Damaschke:2019:CST

[2136] Peter Damaschke. Combinatorial search in two and more rounds. Theoretical Computer Science, 780(??): 1–11, August 9, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300945.

Tsur:2019:SDS

[2137] Dekel Tsur. Succinct data structure for dynamic trees with faster queries. Theoretical Computer Science, 780(??): 12–19, August 9, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300957.

Casteigts:2019:COH

[2138] A. Casteigts, Y. Métivier, J. M. Robson, and A. Zemmari. Counting in one-hop beeping networks. Theoretical Computer Science, 780(??): 20–28, August 9, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930115X.

Florian:2019:NTG

[2139] Josef Florian, Tereza Veselá, and Lubomíra Dvoráková. Normalization of ternary generalized pseudostandard words. Theoretical Computer Science, 780(??):29–53, August 9, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301161.

Etesami:2019:WOD

[2140] Omid Etesami, Narges Ghareghani, Michel Habib, Mohammadreza Hooshmandasl, Reza Naserasr, and Pouyeh Sharifani. When an optimal dominating set with given constraints exists. Theoretical Computer Science, 780 (??):54–65, August 9, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301185.

Chen:2019:ACG

[2141] Jin Chen and Zhi-Xiong Wen. On the abelian complexity of generalized Thue–Morse sequences. Theoretical Computer Science, 780(??): 66–73, August 9, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301203.

Ambroz:2019:PLW

[2142] Petr Ambroz, Ondrej Kadlec, Zuzana Masáková, and Edita Pelantová. Palindromic length of words and morphisms in class P. Theoretical Computer Science, 780(??):74– 83, August 9, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301306.

Konstantinidis:2019:SCA

[2143] G. Konstantinidis and A. Kehagias.

Selfish cops and active robber: Multiplayer pursuit evasion on graphs. Theoretical Computer Science, 780(??): 84–102, August 9, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301318.

Anonymous:2019:EBbd

[2144] Anonymous. Editorial Board. Theoretical Computer Science, 781(??): ii, August 16, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303482.

Nigam:2019:LSF

[2145] Vivek Nigam and René Thiemann.

Logical and semantic frameworks with applications. Theoretical Computer Science, 781(??):1–2, August 16, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303445.

Ayala-Rincon:2019:FNE

[2146] Mauricio Ayala-Rincón, Washington de Carvalho-Segundo, Maribel Fernández, Daniele Nantes-Sobrinho, and Ana Cristina Rocha-Oliveira. A formalisation of nominal α-equivalence with A, C, and AC function symbols. Theoretical Computer Science, 781(??): 3–23, August 16, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301264.

Chaudhuri:2019:FMT

[2147] Kaustuv Chaudhuri, Leonardo Lima, and Giselle Reis. Formalized metatheory of sequent calculi for linear logics. Theoretical Computer Science, 781 (??):24–38, August 16, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930129X.

Caleiro:2019:GPS

[2148] Carlos Caleiro, Filipe Casal, and Andreia Mordido. Generalized probabilistic satisfiability and applications to modelling attackers with side-channel capabilities. Theoretical Computer Science, 781(??):39– 62, August 16, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301276.

Arieli:2019:LAD

[2149] Ofer Arieli and Christian Straßer. Logical argumentation by dynamic proof systems. Theoretical Computer Science, 781(??):63–91, August 16, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301252.

Lombardi:2019:PIR

[2150] Carlos Lombardi, Alejandro Ríos, and Roel de Vrijer. Projections for infinitary rewriting (extended version). Theoretical Computer Science, 781(??): 92–110, August 16, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301239.

Ayala-Rincon:2019:TPP

[2151] Mauricio Ayala-Rincón, Eduardo Bonelli, Juan Edi, and Andrés Viso. Typed path polymorphism. Theoretical Computer Science, 781(??): 111–130, August 16, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301240.

Anonymous:2019:EBbe

[2152] Anonymous. Editorial Board. Theoretical Computer Science, 782(??): ii, August 23, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930355X.

Shraibman:2019:NCC

[2153] Adi Shraibman. Nondeterministic communication complexity with help and graph functions. Theoretical Computer Science, 782(??): 1–10, August 23, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930132X.

Bentz:2019:WLB

[2154] Cédric Bentz. Weighted and locally bounded list-colorings in split graphs, cographs, and partial k-trees. Theoretical Computer Science, 782(??): 11–29, August 23, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301355.

Bergougnoux:2019:FEA

[2155] Benjamin Bergougnoux and Mamadou Kanté. Fast exact algorithms for some connectivity problems parameterized by clique-width. Theoretical Computer Science, 782(??):30– 53, August 23, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301367.

Cohen:2019:FPS

[2156] Johanne Cohen, Jonas Lefèvre, Khaled Maâmra, George Manoussakis, and Laurence Pilard. The first polynomial self-stabilizing 1-maximal matching algorithm for general graphs. Theoretical Computer Science, 782(??): 54–78, August 23, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301379.

Das:2019:SPF

[2157] Bireswar Das, Murali Krishna Enduri, Masashi Kiyomi, Neeldhara Misra, Yota Otachi, I. Vinod Reddy, and Shunya Yoshimura. On structural parameterizations of firefighting. Theoretical Computer Science, 782(??):

79–90, August 23, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301392.

Vakhania:2019:FSS

[2158] Nodari Vakhania. Fast solution of single-machine scheduling problem with embedded jobs. Theoretical Computer Science, 782(??):91– 106, August 23, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301483.

Chatterjee:2019:UWC

[2159] Kingshuk Chatterjee and Kumar Sankar Ray. Unary Watson–Crick automata. Theoretical Computer Science, 782(??): 107–112, August 23, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301628.

Moldagaliyev:2019:RFL

[2160] Birzhan Moldagaliyev. Randomness of formal languages via automatic martingales. Theoretical Computer Science, 782(??):113–128, August 23, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930163X.

Xie:2019:NCB

[2161] Ning Xie, Shuai Xu, and Yekun Xu. A new coding-based algorithm for finding closest pair of vectors. Theoretical Computer Science, 782(??):

129–144, August 23, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301641.

Louza:2019:ACB

[2162] Felipe A. Louza, Guilherme P. Telles, Simon Gog, and Liang Zhao. Algorithms to compute the Burrows– Wheeler similarity distribution. Theoretical Computer Science, 782(??): 145–156, August 23, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301653.

Anonymous:2019:EBbf

[2163] Anonymous. Editorial Board. Theoretical Computer Science, 783(??): ii, September 6, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303627.

Tsur:2019:PAP

[2164] Dekel Tsur. Parameterized algorithm for 3-path vertex cover. Theoretical Computer Science, 783(??): 1–8, September 6, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301665.

Gheibi:2019:WMB

[2165] Amin Gheibi, Anil Maheshwari, and Jörg-Rüdiger Sack. Weighted minimum backward Fréchet distance. Theoretical Computer Science, 783(??): 9–21, September 6, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301689.

Kim:2019:NAP

[2166] Sungwook Kim, Jinsu Kim, and Jae Hong Seo. A new approach to practical function-private inner product encryption. Theoretical Computer Science, 783(??):22–40, September 6, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301690.

Golovach:2019:EMN

[2167] Petr A. Golovach, Dieter Kratsch, Mathieu Liedloff, and Mohamed Yosri Sayadi. Enumeration and maximum number of minimal dominating sets for chordal graphs. Theoretical Computer Science, 783(??): 41–52, September 6, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301707.

Dey:2019:PDC

[2168] Palash Dey, Neeldhara Misra, and Y. Narahari. Parameterized dichotomy of choosing committees based on approval votes in the presence of outliers. Theoretical Computer Science, 783(??): 53–70, September 6, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301872.

Ling:2019:LBG

[2169] San Ling, Khoa Nguyen, Huaxiong Wang, and Yanhong Xu. Lattice-based group signatures: Achieving full dynamicity (and deniability) with ease. Theoretical Computer Science, 783(??): 71–94, September 6, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301884.

Zhang:2019:EFT

[2170] Yaoyao Zhang, Shuang Zhao, and Jixiang Meng. Edge fault tolerance of graphs with respect to λ_{2}-optimal property. Theoretical Computer Science, 783(??):95–104, September 6, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301896.

Anonymous:2019:EBbg

[2171] Anonymous. Editorial Board. Theoretical Computer Science, 784(??): ii, September 13, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303688.

Budinich:2019:BSP

[2172] Marco Budinich. The Boolean SATisfiability Problem in Clifford algebra. Theoretical Computer Science, 784(??): 1–10, September 13, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301938.

Angelini:2019:VRG

[2173] Patrizio Angelini, Michael A. Bekos, Michael Kaufmann, and Fabrizio Montecchiani. On 3D visibility representations of graphs with few crossings per edge. Theoretical Computer Science, 784(??):11–20, September 13, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301951.

Liu:2019:ECA

[2174] Aixia Liu, Shiying Wang, Jun Yuan, and Xue Ma. The h-extra connectivity of k-ary n-cubes. Theoretical Computer Science, 784(??):21– 45, September 13, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301963.

Chan:2019:DOS

[2175] T.-H. Hubert Chan, Zhihao Gavin Tang, Xiaowei Wu, and Chenzi Zhang. Diffusion operator and spectral analysis for directed hypergraph Laplacian. Theoretical Computer Science, 784(??): 46–64, September 13, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301987.

Pai:2019:TST

[2176] Kung-Jui Pai, Ruay-Shiung Chang, Ro-Yu Wu, and Jou-Ming Chang. A two-stages tree-searching algorithm for finding three completely independent spanning trees. Theoretical Computer Science, 784(??):65– 74, September 13, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302087.

Balister:2019:ANP

[2177] Paul Balister and Stefanie Gerke. The asymptotic number of prefix normal words. Theoretical Computer Science, 784(??):75–80, September 13, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302099.

Merta:2019:CIV

[2178] Łukasz Merta. Composition inverses of the variations of the Baum–Sweet sequence. Theoretical Computer Science, 784(??):81– 98, September 13, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302142.

Said:2019:ADD

[2179] Mouhammad Said and Jacques-Olivier Lachaud. Analytical description of digital intersections: Minimal parameters and multiscale representation. Theoretical Computer Science, 784(??):99– 112, September 13, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302154.

Kaminski:2019:PES

[2180] Bogumil Kamiński, Tomasz Olczak, and Pawel Pralat. Parallel execution of schedules with random dependency graph. Theoretical Computer Science, 784(??):113– 132, September 13, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302166.

Carmi:2019:ACC

[2181] Paz Carmi, Anil Maheshwari, Saeed Mehrabi, and Luís Fernando Schultz Xavier da Silveira. Approximability of covering cells with line segments. Theoretical Computer Science, 784(??):133– 141, September 13, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302804.

Anonymous:2019:EBbh

[2182] Anonymous. Editorial Board. Theoretical Computer Science, 785(??): ii, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304232.

Kari:2019:Eb

[2183] Lila Kari.  Editorial.  Theoret-

ical Computer Science, 785(??):iii, September 20, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304268.

Noordman:2019:ENS

[2184] C. R. Noordman and G. A. W. Vreeswijk. Evolving novelty strategies for the Iterated Prisoner’s Dilemma in deceptive tournaments. Theoretical Computer Science, 785(??):1– 16, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306479.

Breik:2019:CPS

[2185] Keenan Breik, Chris Thachuk, Marijn Heule, and David Soloveichik. Computing properties of stable configurations of thermodynamic binding networks. Theoretical Computer Science, 785(??): 17–29, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306480.

Spettel:2019:ASA

[2186] Patrick Spettel and Hans-Georg Beyer. Analysis of the (1, λ)-σ-Self-Adaptation Evolution Strategy with repair by projection applied to a conically constrained problem. Theoretical Computer Science, 785(??):30– 45, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306571.

Friedrich:2019:UED

[2187] Tobias Friedrich, Timo Kötzing, and Martin S. Krejca. Unbiasedness of estimation-of-distribution algorithms. Theoretical Computer Science, 785(??):

46–59, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306662.

Waldmajer:2019:TTA

[2188] Jacek Waldmajer, Zbigniew Bonikowski, and Sebastian Sakowski. Theory of tailor automata. Theoretical Computer Science, 785(??):60– 82, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300921.

Abrahao:2019:ANC

[2189] Felipe S. Abrahão, Klaus Wehmuth, and Artur Ziviani. Algorithmic networks: Central time to trigger expected emergent open-endedness. Theoretical Computer Science, 785(??):83– 116, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301550.

Ellingham:2019:EOG

[2190] M. N. Ellingham and Joanna A. Ellis-Monaghan. Edge-outer graph embedding and the complexity of the DNA reporter strand problem. Theoretical Computer Science, 785(??):117– 127, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301781.

Han:2019:ROI

[2191] Yo-Sub Han and Hwee Kim. Ruleset optimization on isomorphic oritatami systems. Theoretical Computer Science, 785(??):128–139, September 20, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301793.

Liu:2019:NPS

[2192] Liucheng Liu, Wenmei Yi, Qian Yang, Hong Peng, and Jun Wang. Numerical P systems with Boolean condition. Theoretical Computer Science, 785(??): 140–149, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930180X.

Einarsson:2019:LHS

[2193] Hafsteinn Einarsson, Marcelo Matheus Gauy, Johannes Lengler, Florian Meier, Asier Mujika, Angelika Steger, and Felix Weissenberger. The linear hidden subset problem for the (1 + 1) EA with scheduled and adaptive mutation rates. Theoretical Computer Science, 785(??): 150–170, September 20, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303421.

Anonymous:2019:EBbi

[2194] Anonymous. Editorial Board. Theoretical Computer Science, 786(??): ii, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304736.

Xiao:2019:FA

[2195] Mingyu Xiao and Frances A. Rosamond. Frontiers in algorithmics. Theoretical Computer Science, 786(??): 1, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307023.

Abu-Khzam:2019:EPA

[2196] Faisal N. Abu-Khzam, Shouwei Li, Christine Markarian, Friedhelm Meyer auf der Heide, and Pavel Podlipyan. Efficient parallel algorithms for parameterized problems. Theoretical Computer Science, 786(??):2– 12, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306777.

Luo:2019:LSM

[2197] Kelin Luo, Yinfeng Xu, and Huili Zhang. On-line scheduling with monotone subsequence constraints. Theoretical Computer Science, 786(??):13– 25, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305693.

Bereg:2019:SMP

[2198] Sergey Bereg, Feifei Ma, Wencheng Wang, Jian Zhang, and Binhai Zhu. On some matching problems under the color-spanning model. Theoretical Computer Science, 786(??):26– 31, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305243.

Yuasa:2019:DIA

[2199] Shota Yuasa, Zhi-Zhong Chen, Bin Ma, and Lusheng Wang. Designing and implementing algorithms for the closest string problem. Theoretical Computer Science, 786(??):32– 43, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303372.

Wang:2019:PSZ

[2200] Lusheng Wang and Boting Yang. Positive semidefinite zero forcing numbers of two classes of graphs. Theoretical Computer Science, 786(??):44– 54, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303153.

Li:2019:ILK

[2201] Wenjun Li, Haiyan Liu, Jianxin Wang, Lingyun Xiang, and Yongjie Yang. An improved linear kernel for complementary maximal strip recovery: Simpler and smaller. Theoretical Computer Science, 786(??):55– 66, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302421.

Wu:2019:SIJ

[2202] Guangwei Wu, Jianer Chen, and Jianxin Wang. On scheduling inclined jobs on multiple two-stage flowshops. Theoretical Computer Science, 786(??): 67–77, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302196.

Chen:2019:CSC

[2203] Wei Chen and Hanrui Zhang. Complete submodularity characterization in the comparative independent cascade model. Theoretical Computer Science, 786(??):78– 87, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302019.

Han:2019:OKP

[2204] Xin Han, Qinyang Chen, and Kazuhisa Makino. Online knapsack problem under concave functions. Theoretical Computer Science, 786(??):88– 95, September 27, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518302007.

Anonymous:2019:EBbj

[2205] Anonymous. Editorial Board. Theoretical Computer Science, 787(??): ii, October 1, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304797.

Carayol:2019:SII

[2206] Arnaud Carayol and Cyril Nicaud. Special issue — Implementation and Application of Automata (CIAA 2017). Theoretical Computer Science, 787 (??):1, October 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303731.

Brzozowski:2019:CPP

[2207] Janusz A. Brzozowski and Corwin Sinnamon. Complexity of proper prefix-convex regular languages. Theoretical Computer Science, 787(??): 2–13, October 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518304961.

Ferens:2019:CBF

[2208] Robert Ferens and Marek Szykuła. Complexity of bifix-free regular languages. Theoretical Computer Science, 787(??):14–27, October 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397518305875.

Dando:2019:HEW

[2209] Louis-Marie Dando and Sylvain Lombardy. From Hadamard expressions to weighted rotating automata and back. Theoretical Computer Science, 787(??): 28–44, October 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305887.

Szykula:2019:SCB

[2210] Marek Szykuła and John Wittnebel.

Syntactic complexity of bifix-free regular languages. Theoretical Computer Science, 787(??):45–76, October 1, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300039.

Ryzhikov:2019:SPA

[2211] Andrew Ryzhikov. Synchronization problems in automata without non-trivial cycles. Theoretical Computer Science, 787(??):77– 88, October 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300040.

Hospodar:2019:NCS

[2212] Michal Hospodár, Galina Jirásková, and Peter Mlynárcik. Nondeterministic complexity in subclasses of convex languages. Theoretical Computer Science, 787(??):89– 110, October 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300052.

Kutrib:2019:TRF

[2213] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Transducing reversibly with finite state machines. Theoretical Computer Science, 787(??): 111–126, October 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304128.

Han:2019:ADR

[2214] Yo-Sub Han and Sang-Ki Ko. Alignment distance of regular tree languages. Theoretical Computer Science, 787(??): 127–137, October 1, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930413X.

Anonymous:2019:EBbk

[2215] Anonymous. Editorial Board. Theoretical Computer Science, 788(??): ii, October 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304852.

Gao:2019:PSI

[2216] Xiaofeng Gao and Dongdong Ge. Preface: Special issue on the Annual International Conference on Combinatorial Optimization and Applications (COCOA). Theoretical Computer Science, 788(??):1, October 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303743.

Miao:2019:TED

[2217] Dongjing Miao, Zhipeng Cai, Jiguo Yu, and Yingshu Li. Triangle edge deletion on planar glasses-free RGB-digraphs. Theoretical Computer Science, 788 (??):2–11, October 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306558.

Angel:2019:AFR

[2218] Eric Angel, Thang Nguyen Kim, and Shikha Singh. Approximating kforest with resource augmentation: a primal-dual approach. Theoretical Computer Science, 788(??):12– 20, October 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307229.

Duan:2019:ISE

[2219] Zhenhua Duan, Cong Tian, Nan Zhang, Qian Ma, and Hongwei Du. Index set expressions can represent temporal logic formulas. Theoretical Computer Science, 788(??):21– 38, October 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307230.

Huang:2019:ACS

[2220] Jiapeng Huang, Zhenzao Wen, Linghe Kong, Li Ge, Min-You Wu, and Guihai Chen. Accelerate the classification statistics in RFID systems. Theoretical Computer Science, 788 (??):39–52, October 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518307242.

Guo:2019:EAA

[2221] Longkun Guo, Min Li, and Dachuan Xu. Efficient approximation algorithms for maximum coverage with group budget constraints. Theoretical Computer Science, 788(??):53– 65, October 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301215.

Erlebach:2019:COA

[2222] Thomas Erlebach, Fu-Hong Liu, Hsiang-Hsuan Liu, Mordechai Shalom, Prudence W. H. Wong, and Shmuel Zaks. Complexity and online algorithms for minimum skyline coloring of intervals. Theoretical Computer Science, 788(??):66– 78, October 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302853.

Gu:2019:EAR

[2223] Qian-Ping Gu, Jiajian Leo Liang, and Guochuan Zhang. Efficient algorithms for ridesharing of personal vehicles. Theoretical Computer Science, 788(??):79–94, October 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303767.

Anonymous:2019:EBbl

[2224] Anonymous. Editorial Board. Theoretical Computer Science, 789(??): ii, October 15, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305171.

Yen:2019:SIS

[2225] Hsu-Chun Yen, Md. Saidur Rahman, and Sheung-Hung Poon. Special issue on selected papers from the 11th International Conference and Workshops on Algorithms and Computation (WALCOM 2017). Theoretical Computer Science, 789(??): 1, October 15, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305134.

Oh:2019:CCR

[2226] Eunjin Oh and Hee-Kap Ahn. Computing the center region and its variants. Theoretical Computer Science, 789(??): 2–12, October 15, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304420.

Bahoo:2019:TST

[2227] Yeganeh Bahoo, Bahareh Banyassady, Prosenjit K. Bose, Stephane Durocher, and Wolfgang Mulzer. A time-space trade-off for computing the k-visibility region of a point in a polygon. Theoretical Computer Science, 789(??): 13–21, October 15, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830433X.

Vaezi:2019:VEM

[2228] Arash Vaezi and Mohammad Ghodsi.

Visibility extension via mirror-edges to cover invisible segments. Theoretical Computer Science, 789(??):

22–33, October 15, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301173.

Jansen:2019:IAT

[2229] Klaus Jansen and Malin Rau. Improved approximation for two dimensional Strip Packing with polynomial bounded width. Theoretical Computer Science, 789(??):34– 49, October 15, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302555.

Akitaya:2019:MWC

[2230] Hugo A. Akitaya, Rajasekhar Inkulu, Torrie L. Nichols, Diane L. Souvaine, Csaba D. Tóth, and Charles R. Winston. Minimum weight connectivity augmentation for planar straight-line graphs. Theoretical Computer Science, 789(??):50–63, October 15, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303694.

Jiao:2019:AAR

[2231] Yang Jiao, R. Ravi, and Wolfgang Gatterbauer. Algorithms for automatic ranking of participants and tasks in an anonymized contest. Theoretical Computer Science, 789(??): 64–76, October 15, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830495X.

Cano:2019:SDL

[2232] Rafael G. Cano, Cid C. de Souza, and Pedro J. de Rezende. Solving dynamic labeling problems to optimality using solution space reductions. Theoretical Computer Science, 789(??): 77–92, October 15, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305267.

Anonymous:2019:EBbm

[2233] Anonymous. Editorial Board. Theoretical Computer Science, 790(??): ii, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305419.

Aziz:2019:ERU

[2234] Haris Aziz, Péter Biró, Jérôme Lang, Julien Lesca, and Jérôme Monnot. Efficient reallocation under additive and responsive preferences. Theoretical Computer Science, 790(??): 1–15, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303196.

Jahannia:2019:PZL

[2235] Marieh Jahannia, Morteza Mohammadnoori, Narad Rampersad, and Manon Stipulanti. Palindromic Ziv–Lempel and Crochemore factorizations of m-Bonacci infinite words. Theoretical Computer Science, 790(??):16– 40, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303184.

Covanov:2019:IMF

[2236] Svyatoslav Covanov. Improved method for finding optimal formulas for bilinear maps in a finite field. Theoretical Computer Science, 790(??): 41–65, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303172.

Liu:2019:ECD

[2237] Huiqing Liu, Shunzhe Zhang, and Dong Li. On g-extra conditional diagnosability of hierarchical cubic networks. Theoretical Computer Science, 790(??): 66–79, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303160.

Gerdjikov:2019:SEB

[2238] Stefan Gerdjikov, Stoyan Mihov, and Klaus U. Schulz. Space-efficient bimachine construction based on the equalizer accumulation principle. Theoretical Computer Science, 790(??): 80–95, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303159.

Tsur:2019:FDP

[2239] Dekel Tsur. Faster deterministic parameterized algorithm for k-path. Theoretical Computer Science, 790(??): 96–104, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303123.

Bae:2019:SPQ

[2240] Sang Won Bae and Haitao Wang.

L_{1} shortest path queries in simple polygons. Theoretical Computer Science, 790(??):105–116, October 22, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303111.

Sun:2019:FTA

[2241] Xueli Sun, Qingfeng Dong, Shuming Zhou, Mengjie Lv, Guanqin Lian, and Jiafei Liu. Fault tolerance analysis of hierarchical folded cube. Theoretical Computer Science, 790(??): 117–130, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930310X.

Klouda:2019:CCD

[2242] Karel Klouda and Stepán Starosta.

Characterization of circular D0Lsystems. Theoretical Computer Science, 790(??):131–137, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0304397519303093.

Machida:2019:SRC

[2243] Motoya Machida and Alexander Y.

Shibakov. Stack and register complexity of radix conversions. Theoretical Computer Science, 790(??):

138–151, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303081.

Chen:2019:KPC

[2244] Jianer Chen, Henning Fernau, Peter Shaw, Jianxin Wang, and Zhibiao Yang. Kernels for packing and covering problems. Theoretical Computer Science, 790(??):152– 166, October 22, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303068.

Anonymous:2019:EBbn

[2245] Anonymous. Editorial Board. Theoretical Computer Science, 791(??): ii, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305353.

Jo:2019:MCP

[2246] Changseong Jo, Jihoon Choi, Suh-Ryung Kim, and Yoshio Sano. On the minimum clique partitioning problem on weighted chordal graphs. Theoretical Computer Science, 791(??): 1–9, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302841.

Kante:2019:PCG

[2247] Mamadou Moustapha Kanté, Thiago Marcilon, and Rudini Sampaio. On the parameterized complexity of the geodesic hull number. Theoretical Computer Science, 791(??): 10–27, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302828.

Kueng:2019:FRH

[2248] Richard Kueng, Dustin G. Mixon, and Soledad Villar. Fair redistricting is hard. Theoretical Computer Science, 791(??):28–35, October 29, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302798.

Hartman:2019:RUF

[2249] Patrick Hartman and Joe Sawada.

Ranking and unranking fixed-density necklaces and Lyndon words. Theoretical Computer Science, 791(??): 36–47, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302865.

Fukunaga:2019:CTH

[2250] Takuro Fukunaga and Takanori Maehara. Computing a tree having a small vertex cover. Theoretical Computer Science, 791(??):48– 61, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303263.

Baiocchi:2019:SCG

[2251] Pierluigi Baiocchi, Tiziana Calamoneri, Angelo Monti, and Rossella Petreschi. Some classes of graphs that are not PCGs. Theoretical Computer Science, 791(??):62– 75, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303251.

Bahoo:2019:PSM

[2252] Yeganeh Bahoo, Stephane Durocher, J. Mark Keil, Debajyoti Mondal, Saeed Mehrabi, and Sahar Mehrpour. Polygon simplification by minimizing convex corners. Theoretical Computer Science, 791(??):76–86, October 29, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930324X.

BenBasat:2019:GMS

[2253] Ran Ben Basat, Gil Einziger, and Roy Friedman. Give me some slack: Efficient network measurements. Theoretical Computer Science, 791(??): 87–108, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303202.

Bruneaux:2019:NAW

[2254] Margot Bruneaux. A note about words which coincide except in one position. Theoretical Computer Science, 791(??): 109–111, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302816.

Recasens:2019:CCP

[2255] J. Recasens. On a conjecture concerning positive semi-definiteness. Theoretical Computer Science, 791(??): 112–115, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303238.

Li:2019:MSB

[2256] Yuan Li and John O. Adeyeye. Maximal sensitivity of Boolean nested canalizing functions. Theoretical Computer Science, 791(??):116–122, October 29, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303135.

Fleischmann:2019:RAP

[2257] Pamela Fleischmann, Pascal Ochem, and Kamellia Reshadi. Repetition avoidance in products of factors. Theoretical Computer Science, 791(??): 123–126, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303019.

Alm:2019:FCT

[2258] Jeremy F. Alm and Andrew Ylvisaker.

A fast coset-translation algorithm for computing the cycle structure of Comer relation algebras over Z/pZ. Theoretical Computer Science, 791(??): 127–131, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303342.

Melekian:2019:NGM

[2259] Christopher Melekian and Eddie Cheng. A note on generalized matching preclusion in bipartite graphs. Theoretical Computer Science, 791(??): 132–140, October 29, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303433.

Anonymous:2019:EBbo

[2260] Anonymous. Editorial Board. Theoretical Computer Science, 792(??): ii, November 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305717.

Crochemore:2019:SIH

[2261] Maxime Crochemore and Jakub Radoszewski. Special issue in honor of the 70th birthday of Prof. Wojciech Rytter. Theoretical Computer Science, 792(??):1, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305894.

Czumaj:2019:LEM

[2262] Artur Czumaj and Peter Davies. Leader election in multi-hop radio networks. Theoretical Computer Science, 792(??):2–11, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301331.

Alzamel:2019:LLA

[2263] Mai Alzamel, Costas S. Iliopoulos, W. F. Smyth, and Wing-Kin Sung. Off-line and on-line algorithms for closed string factorization. Theoretical Computer Science, 792(??): 12–19, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306546.

Plandowski:2019:PGS

[2264] Wojciech Plandowski. On PSPACE generation of a solution set of a word equation and its applications. Theoretical Computer Science, 792(??): 20–61, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306443.

Plandowski:2019:CCM

[2265] Wojciech Plandowski and Aleksy Schubert. On the complexity of computation maximal exponent of periodicity of word equations and expressible relations (note). Theoretical Computer Science, 792(??): 62–68, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306431.

Czyzowicz:2019:GSP

[2266] Jurek Czyzowicz, Maxime Godon, Evangelos Kranakis, and Arnaud Labourel. Group search of the plane with faulty robots. Theoretical Computer Science, 792(??): 69–84, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306005.

Almirantis:2019:OWT

[2267] Yannis Almirantis, Panagiotis Charalampopoulos, Jia Gao, Costas S. Iliopoulos, Manal Mohamed, Solon P. Pissis, and Dimitris Polychronopoulos. On overabundant words and their application to biological sequence analysis. Theoretical Computer Science, 792 (??):85–95, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305760.

Shallit:2019:SCP

[2268] Jeffrey Shallit and Arseny Shur. Subword complexity and power avoidance. Theoretical Computer Science, 792(??): 96–116, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305759.

Giancarlo:2019:DCM

[2269] Raffaele Giancarlo, Simona E. Rombo, and Filippo Utro. DNA combinatorial messages and epigenomics: the case of chromatin organization and nucleosome occupancy in eukaryotic genomes. Theoretical Computer Science, 792(??): 117–130, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304638.

Nakashima:2019:SSA

[2270] Yuto Nakashima, Takuya Takagi, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. On the size of the smallest alphabet for Lyndon trees. Theoretical Computer Science, 792(??): 131–143, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304602.

Fici:2019:MFF

[2271] Gabriele Fici, Antonio Restivo, and Laura Rizzo. Minimal forbidden factors of circular words. Theoretical Computer Science, 792(??): 144–153, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304043.

Avgustinovich:2019:ASI

[2272] Sergey Avgustinovich, Julien Cassaigne, Juhani Karhumäki, Svetlana Puzynina, and Aleksi Saarela. On abelian saturated infinite words. Theoretical Computer Science, 792(??): 154–160, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303335.

Hendrian:2019:EDD

[2273] Diptarama Hendrian, Shunsuke Inenaga, Ryo Yoshinaka, and Ayumi Shinohara. Efficient dynamic dictionary matching with DAWGs and AC-automata. Theoretical Computer Science, 792(??):161–172, November 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S030439751830238X.

Anonymous:2019:EBbp

[2274] Anonymous. Editorial Board. Theoretical Computer Science, 793(??): ii, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305997.

Gelle:2019:OGO

[2275] Kitti Gelle and Szabolcs Iván. The ordinal generated by an ordinal grammar is computable. Theoretical Computer Science, 793(??):1– 13, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303044.

Wang:2019:SSR

[2276] Bin Wang and Guangyan Zhou. Super solutions of random (3 + p)-SAT. Theoretical Computer Science, 793(??): 14–27, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303032.

Banik:2019:DVG

[2277] Aritra Banik, Sandip Das, Anil Maheshwari, and Michiel Smid. The discrete Voronoi game in a simple polygon. Theoretical Computer Science, 793(??):28–35, November 12, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303007.

Faal:2019:MVD

[2278] Hossein Teimoori Faal. A multiset version of determinants and the Coin arrangements lemma. Theoretical Computer Science, 793(??):36– 43, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302981.

Yu:2019:NAA

[2279] Wei Yu, Zhaohui Liu, and Xiaoguang Bao. New approximation algorithms for the minimum cycle cover problem. Theoretical Computer Science, 793(??): 44–58, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930297X.

Amir:2019:ACS

[2280] A. Amir, A. Levy, R. Lubin, and E. Porat. Approximate cover of strings. Theoretical Computer Science, 793(??): 59–69, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303354.

Kononov:2019:MMA

[2281] A. V. Kononov, M. Y. Kovalyov, and B. M. T. Lin. Minimizing machine assignment costs over Δapproximate solutions of the scheduling problem P ||C_{max}. Theoretical Computer Science, 793(??):70– 78, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303512.

Wang:2019:QSS

[2282] Hui Wang, Cees H. Elzinga, Zhiwei Lin, and Jordan Vincent. Quantifying sequential subsumption. Theoretical Computer Science, 793(??): 79–99, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930372X.

Charron-Bost:2019:FSP

[2283] Bernadette Charron-Bost and Shlomo Moran. The firing squad problem revisited. Theoretical Computer Science, 793(??):100–112, November 12, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930458X.

Lv:2019:PDC

[2284] Mengjie Lv, Shuming Zhou, Xueli Sun, Guanqin Lian, Jiafei Liu, and Dajin Wang. Probabilistic diagnosis of clustered faults for hypercubebased multiprocessor system. Theoretical Computer Science, 793(??):113– 131, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304141.

Sakai:2019:MCS

[2285] Yoshifumi Sakai. Maximal common subsequence algorithms. Theoretical Computer Science, 793(??):132– 139, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304074.

Kuo:2019:AHM

[2286] Tung-Wei Kuo. On the approximability and hardness of the minimum connected dominating set with routing cost constraint. Theoretical Computer Science, 793(??):140– 151, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304062.

Cho:2019:BDD

[2287] Da-Jung Cho, Yo-Sub Han, and Hwee Kim. Bound-decreasing duplication system. Theoretical Computer Science, 793(??):152–168, November 12, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304050.

Brauner:2019:DBW

[2288] Nadia Brauner, Yves Crama, Etienne Delaporte, Vincent Jost, and Luc Libralesso. Do balanced words have a short period? Theoretical Computer Science, 793(??):169– 180, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304049.

Li:2019:DSM

[2289] Pingshan Li and Min Xu. The t/kdiagnosability and strong Menger connectivity on star graphs with conditional faults. Theoretical Computer Science, 793(??):181–192, November 12, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304037.

Blanchette:2019:QAG

[2290] Benjamin Blanchette. Quasi-automatic groups are asynchronously automatic. Theoretical Computer Science, 793(??): 193–195, November 12, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303755.

Anonymous:2019:EBbq

[2291] Anonymous. Editorial Board. Theoretical Computer Science, 794(??): ii, November 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306115.

Angelopoulos:2019:PSI

[2292] Spyros Angelopoulos, Nicolas Nisse, and Dimitrios M. Thilikos. Preface to special issue on theory and applications of graph searching. Theoretical Computer Science, 794(??): 1–2, November 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306073.

Pralat:2019:HMZ

[2293] Paweł Prałat. How many zombies are needed to catch the survivor on toroidal grids? Theoretical Computer Science, 794(??):3– 11, November 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518303414.

Bonato:2019:BBN

[2294] Anthony Bonato and Thomas Lidbetter. Bounds on the burning numbers of spiders and path-forests. Theoretical Computer Science, 794(??): 12–19, November 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751830402X.

Alpern:2019:SIH

[2295] Steve Alpern. Search for an immobile Hider in a known subset of a network. Theoretical Computer Science, 794(??): 20–26, November 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518304389.

Lamprou:2019:EDL

[2296] Ioannis Lamprou, Russell Martin, and Sven Schewe. Eternally dominating large grids. Theoretical Computer Science, 794(??):27–46, November 19, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518305735.

Dereniowski:2019:CFR

[2297] Dariusz Dereniowski, Tomás Gavenciak, and Jan Kratochvíl. Cops, a fast robber and defensive domination on interval graphs. Theoretical Computer Science, 794(??):47– 58, November 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306029.

Bonato:2019:HCR

[2298] A. Bonato, N. E. Clarke, D. Cox, S. Finbow, F. McInerney, and M. E. Messinger. Hyperopic cops and robbers. Theoretical Computer Science, 794(??):59–68, November 19, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397518306091.

Coupechoux:2019:FT

[2299] Pierre Coupechoux, Marc Demange, David Ellison, and Bertrand Jouve. Firefighting on trees. Theoretical Computer Science, 794(??):69– 84, November 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519300817.

Dereniowski:2019:FSW

[2300] Dariusz Dereniowski, Dorota Osula, and Paweł Rzazewski. Finding smallwidth connected path decompositions in polynomial time. Theoretical Computer Science, 794(??):85– 100, November 19, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302129.

Anonymous:2019:EBbr

[2301] Anonymous. Editorial Board. Theoretical Computer Science, 795(??): ii, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306358.

Minh:2019:BNM

[2302] Do Le Paul Minh and Truong Thi Thuy Duong. Bayesian networks: the minimal triangulations of a graph. Theoretical Computer Science, 795(??): 1–8, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303792.

delaMaza:2019:CKP

[2303] Sebastián González Hermosillo de la Maza and César Hernández-Cruz. On the complexity of the k-kernel problem on cyclically k-partite digraphs. Theoretical Computer Science, 795(??): 9–19, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303809.

Zeng:2019:PKE

[2304] Ming Zeng, Jie Chen, Kai Zhang, and Haifeng Qian. Public key encryption with equality test via hash proof system. Theoretical Computer Science, 795(??):20–35, November 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303822.

Huang:2019:EDG

[2305] Yanze Huang, Limei Lin, Li Xu, and Xiaoding Wang. Extra diagnosability and good-neighbor diagnosability of ndimensional alternating group graph AG_{n} under the PMC model. Theoretical Computer Science, 795(??): 36–49, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303834.

Akhtar:2019:GCN

[2306] Muhammad S. Akhtar, Usman Ali, Ghulam Abbas, and Mutahira Batool. On the game chromatic number of splitting graphs of path and cycle. Theoretical Computer Science, 795(??): 50–56, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303846.

Lyu:2019:SPP

[2307] Lin Lyu, Shengli Liu, and Dawu Gu. Structure-preserving publickey encryption with leakage-resilient CCA security. Theoretical Computer Science, 795(??):57–80, November 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930386X.

Dolev:2019:AAC

[2308] Shlomi Dolev, Niv Gilboa, and Ximing Li. Accumulating automata and cascaded equations automata for communicationless information theoretically secure multi-party computation. Theoretical Computer Science, 795(??): 81–99, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303883.

Gurski:2019:HPB

[2309] Frank Gurski, Carolin Rehs, and Jochen Rethmann. On the hardness of palletizing bins using FIFO queues. Theoretical Computer Science, 795(??): 100–107, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303895.

Bang-Jensen:2019:PCL

[2310] J. Bang-Jensen, Kristine V. K. Knudsen, Saket Saurabh, and Meirav Zehavi. The parameterized complexity landscape of finding 2partitions of digraphs. Theoretical Computer Science, 795(??):108– 114, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303901.

Kuo:2019:CFH

[2311] Che-Nan Kuo and Yu-Huei Cheng. Cycles in folded hypercubes with two adjacent faulty vertices. Theoretical Computer Science, 795(??):115– 118, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303913.

Pham:2019:IMP

[2312] Van Huy Pham and Kien Trung Nguyen. Inverse 1-median problem on trees under mixed rectilinear and Chebyshev norms. Theoretical Computer Science, 795(??):119– 127, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303925.

Chiarelli:2019:NAW

[2313] Nina Chiarelli, Tatiana Romina Hartinger, Valeria Alejandra Leoni, Maria Inés Lopez Pujato, and Martin Milanic. New algorithms for weighted k-domination and total k-domination problems in proper interval graphs. Theoretical Computer Science, 795(??): 128–141, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303937.

Brimkov:2019:PDT

[2314] Boris Brimkov, Joshua Carlson, Illya V. Hicks, Rutvik Patel, and Logan Smith. Power domination throttling. Theoretical Computer Science, 795(??): 142–153, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303949.

Hu:2019:AHD

[2315] Yining Hu and Guoniu Wei-Han.

On the automaticity of the Hankel determinants of a family of automatic sequences. Theoretical Computer Science, 795(??):154–164, November 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303950.

Ramya:2019:LPV

[2316] C. Ramya and B. V. Raghavendra Rao. Linear projections of the Vandermonde polynomial. Theoretical Computer Science, 795(??):165– 182, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303962.

Berge:2019:PCS

[2317] Pierre Bergé, Arpad Rimmel, and Joanna Tomasik. On the parameterized complexity of separating certain sources from the target. Theoretical Computer Science, 795(??):183– 193, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303974.

Terrier:2019:CCT

[2318] Véronique Terrier. Communication complexity tools on recognizable picture languages. Theoretical Computer Science, 795(??):194– 203, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303986.

Sdroievski:2019:HSP

[2319] Nicollas M. Sdroievski, Murilo V. G. da Silva, and André L. Vignatti. The Hidden Subgroup Problem and MKTP. Theoretical Computer Science, 795(??): 204–212, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519303998.

Miao:2019:SFT

[2320] Lu Miao, Shurong Zhang, Rong hua Li, and Weihua Yang. Structure fault tolerance of k-ary n-cube networks. Theoretical Computer Science, 795(??): 213–218, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304001.

Cheng:2019:VDP

[2321] Dongqin Cheng. Vertex-disjoint paths joining adjacent vertices in faulty hypercubes. Theoretical Computer Science, 795(??):219–224, November 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304025.

Khalili:2019:STS

[2322] Mojtaba Khalili, Mohammad Dakhilalian, and Carla Ràfols. Short tightly secure signatures for signing a vector of group elements: a new approach. Theoretical Computer Science, 795(??): 225–239, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930427X.

Latteux:2019:POR

[2323] Michel Latteux and Yves Roos. On prefixal one-rule string rewrite systems. Theoretical Computer Science, 795(??): 240–256, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930430X.

Otto:2019:DOR

[2324] Friedrich Otto. On deterministic ordered restart-delete automata. Theoretical Computer Science, 795(??):257– 274, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304347.

Cai:2019:TED

[2325] Leizhen Cai and Junjie Ye. Two edgedisjoint paths with length constraints. Theoretical Computer Science, 795(??): 275–284, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304359.

Calinescu:2019:IAA

[2326] Gruia Calinescu, Guy Kortsarz, and Zeev Nutov. Improved approximation algorithms for minimum power covering problems. Theoretical Computer Science, 795(??):285– 300, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304360.

Glen:2019:MPF

[2327] Amy Glen, Jamie Simpson, and W. F. Smyth. More properties of the Fibonacci word on an infinite alphabet. Theoretical Computer Science, 795(??): 301–311, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304372.

Andres:2019:OCG

[2328] Stephan Dominique Andres, Melissa Huggan, Fionn Mc Inerney, and Richard J. Nowakowski. The orthogonal colouring game. Theoretical Computer Science, 795(??):312– 325, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304499. See

corrigendum [2418].

Zhang:2019:TSS

[2329] Xiao Zhang, Shengli Liu, Jiaxin Pan, and Dawu Gu. Tightly secure signature schemes from the LWE and subset sum assumptions. Theoretical Computer Science, 795(??):326– 344, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304505.

Ferraioli:2019:SPO

[2330] Diodato Ferraioli and Carmine Ventre.

Social pressure in opinion dynamics. Theoretical Computer Science, 795(??): 345–361, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304529.

Czyzowicz:2019:EOB

[2331] Jurek Czyzowicz, Krzysztof Diks, Jean Moussi, and Wojciech Rytter. Energyoptimal broadcast and exploration in a tree using mobile agents. Theoretical Computer Science, 795(??):362– 374, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304530.

Angelinia:2019:GRD

[2332] Patrizio Angelinia, Michael A. Bekosa, Walter Didimo, Luca Grilli, Philipp Kindermann, Tamara Mchedlidze, Roman Prutkin, Antonios Symvonis, and Alessandra Tappini. Greedy rectilinear drawings. Theoretical Computer Science, 795(??):375–397, November 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304542.

Abam:2019:GSG

[2333] Mohammad Ali Abam and Mahnaz Sadat Qafari. Geometric spanner games. Theoretical Computer Science, 795(??): 398–407, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304554.

Bekosa:2019:PDF

[2334] Michael A. Bekosa, Felice De Luca, Walter Didimo, Tamara Mchedlidze, Martin Nöllenburg, Antonios Symvonis, and Ioannis Tollis. Planar drawings of fixed-mobile bigraphs. Theoretical Computer Science, 795(??):408– 419, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304608.

Khuller:2019:SPI

[2335] Samir Khuller, Jingling Li, Pascal Sturmfels, Kevin Sun, and Prayaag Venkat. Select and permute: an improved online framework for scheduling to minimize weighted completion time. Theoretical Computer Science, 795(??): 420–431, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930461X.

Peng:2019:GCS

[2336] Liqiang Peng and Atsushi Takayasu. Generalized cryptanalysis of small CRT-exponent RSA. Theoretical Computer Science, 795(??):432–458, November 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304669.

Amir:2019:PPG

[2337] Michael Amir and Alfred M. Bruckstein. Probabilistic pursuits on graphs. Theoretical Computer Science, 795(??): 459–477, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304906.

Crampton:2019:PRP

[2338] Jason Crampton, Gregory Gutin, Martin Koutecký, and Rémi Watrigant. Parameterized resiliency problems. Theoretical Computer Science, 795(??):478– 491, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304918.

Garncarek:2019:OPS

[2339] Paweł Garncarek, Tomasz Jurdziński, Dariusz R. Kowalski, and Krzysztof Loryś. Online packet scheduling under adversarial errors. Theoretical Computer Science, 795(??):492– 509, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930492X.

Kamiyama:2019:DNM

[2340] Naoyuki Kamiyama. Discrete Newton methods for the evacuation problem. Theoretical Computer Science, 795(??): 510–519, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519304931.

Jansen:2019:CCN

[2341] Bart M. P. Jansen and Jesper Nederlof. Computing the chromatic number using graph decompositions via matrix rank. Theoretical Computer Science, 795(??):520– 539, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304955.

Xu:2019:ETD

[2342] Min Xu and Yulong Wei. The h-edge tolerable diagnosability of balanced hypercubes. Theoretical Computer Science, 795(??):540–546, November 26, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304967.

Legersky:2019:CAP

[2343] Jan Legerský and Milena Svobodová. Construction of algorithms for parallel addition in expanding bases via Extending Window Method. Theoretical Computer Science, 795(??):547– 569, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304992.

Hara:2019:SBR

[2344] Keisuke Hara, Fuyuki Kitagawa, Takahiro Matsuda, Goichiro Hanaoka, and Keisuke Tanaka. Simulationbased receiver selective opening CCA secure PKE from standard computational assumptions. Theoretical Computer Science, 795(??):570– 597, November 26, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305055.

Anonymous:2019:PDa

[2345] Anonymous. Pages 1–308 (3 December 2019). Theoretical Computer Science, 796(??):1–308, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2019:EBbs

[2346] Anonymous. Editorial Board. Theoretical Computer Science, 796(??): ii, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306656.

Banerjee:2019:PRD

[2347] S. Banerjee, J. Mark Keil, and D. Pradhan. Perfect Roman domination in graphs. Theoretical Computer Science, 796(??):1–21, December 3, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305067.

DeLuca:2019:RDS

[2348] Felice De Luca, Md Iqbal Hossain, Stephen Kobourov, Anna Lubiw, and Debajyoti Mondal. Recognition and drawing of stick graphs. Theoretical Computer Science, 796(??):22– 33, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305079.

DeBonis:2019:NSL

[2349] Annalisa De Bonis. New selectors and locally thin families with applications to multi-access channels supporting simultaneous transmissions. Theoretical Computer Science, 796(??): 34–50, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305092.

Kotzing:2019:FHT

[2350] Timo Kötzing and Martin S. Krejca. First-hitting times under drift. Theoretical Computer Science, 796(??): 51–69, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305109.

Wang:2019:CPF

[2351] Yucheng Wang, Qi Bao, and Zhongzhi Zhang. Combinatorial properties of Farey graphs. Theoretical Computer Science, 796(??):70–89, December 3, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305110.

Yang:2019:TAS

[2352] Liya Yang and Xiwen Lu. Two-agent scheduling problems with the general position-dependent processing time. Theoretical Computer Science, 796(??): 90–98, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305122.

Hanneke:2019:OSN

[2353] Steve Hanneke and Aryeh Kontorovich. Optimality of SVM: Novel proofs and tighter bounds. Theoretical Computer Science, 796(??):99– 113, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305262.

Siddiqui:2019:RFT

[2354] Hafiz Muhammad Afzal Siddiqui, Sakander Hayat, Asad Khan, Muhammad Imran, Ayesha Razzaq, and Jia-Bao Liu. Resolvability and faulttolerant resolvability structures of convex polytopes. Theoretical Computer Science, 796(??):114–128, December 3, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305286.

Docker:2019:DTA

[2355] Janosch Döcker, Simone Linz, and Charles Semple. Displaying trees across two phylogenetic networks. Theoretical Computer Science, 796(??):129– 146, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930547X.

Lian:2019:PEH

[2356] Guanqin Lian, Shuming Zhou, Sun-Yuan Hsieh, Jiafei Liu, Gaolin Chen, and Yihong Wang. Performance evaluation on hybrid fault diagnosability of regular networks. Theoretical Computer Science, 796(??):147– 153, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305481.

Tellache:2019:TMO

[2357] Nour El Houda Tellache, Mourad Boudhar, and Farouk Yalaoui. Twomachine open shop problem with agreement graph. Theoretical Computer Science, 796(??):154–168, December 3, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305493.

Li:2019:SCS

[2358] Dong Li, Xiaolan Hu, and Huiqing Liu. Structure connectivity and substructure connectivity of twisted hypercubes. Theoretical Computer Science, 796(??):169–179, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/pii/S0304397519305511.

Rad:2019:CMB

[2359] Nader Jafari Rad. On the complexity of multiple bondage in graphs. Theoretical Computer Science, 796(??): 180–186, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305523.

Janczewski:2019:CNR

[2360] Robert Janczewski, Paweł Obszarski, and Krzysztof Turowski. 2-coloring number revisited. Theoretical Computer Science, 796(??):187–195, December 3, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305535.

Schweser:2019:DDC

[2361] Thomas Schweser. DP-degree colorable hypergraphs. Theoretical Computer Science, 796(??):196–206, December 3, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305547.

Majumdar:2019:TKE

[2362] Diptapriyo Majumdar, Rian Neogi, Venkatesh Raman, and S. Vaishali. Tractability of König edge deletion problems. Theoretical Computer Science, 796(??):207–215, December 3, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305559.

Jaffke:2019:MWI

[2363] Lars Jaffke, O joung Kwon, Torstein J. F. Strømme, and Jan Arne Telle. Mim-width III. Graph powers and generalized distance domination problems. Theoretical Computer Science, 796(??):

216–236, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305560.

Moore:2019:FDC

[2364] Matthew Moore. Finite degree clones are undecidable. Theoretical Computer Science, 796(??):237– 271, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305584.

Castelli:2019:CIS

[2365] Mauro Castelli, Riccardo Dondi, Giancarlo Mauri, and Italo Zoppis. Comparing incomplete sequences via longest common subsequence. Theoretical Computer Science, 796(??): 272–285, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305663.

Filmus:2019:ALD

[2366] Yuval Filmus. Another look at degree lower bounds for polynomial calculus. Theoretical Computer Science, 796(??): 286–293, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305675.

Hanaka:2019:MWM

[2367] Tesshu Hanaka, Hans L. Bodlaender, Tom C. van der Zanden, and Hirotaka Ono. On the maximum weight minimal separator. Theoretical Computer Science, 796(??):294– 308, December 3, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519305766.

Anonymous:2019:EBbt

[2368] Anonymous. Editorial Board. Theoretical Computer Science, 797(??): ii, December 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306760.

Schewe:2019:SIT

[2369] Sven Schewe, Thomas Schneider, and Jef Wijsen. Special issue on Temporal Representation and Reasoning (TIME 2017). Theoretical Computer Science, 797(??):1, December 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306723.

Cairo:2019:DCS

[2370] Massimo Cairo and Romeo Rizzi. Dynamic controllability of simple temporal networks with uncertainty: Simple rules and fast real-time execution. Theoretical Computer Science, 797(??): 2–16, December 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518306765.

Sioutis:2019:CSB

[2371] Michael Sioutis, Anastasia Paparrizou, and Jean-François Condotta.

Collective singleton-based consistency for qualitative constraint networks: Theory and practice. Theoretical Computer Science, 797(??):17– 41, December 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519301343.

Amarilli:2019:CPC

[2372] Antoine Amarilli, Mouhamadou Lamine Ba, Daniel Deutch, and Pierre Senellart. Computing possible and certain answers over order-incomplete data. Theoretical Computer Science, 797(??): 42–76, December 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303214.

Zavatteri:2019:CST

[2373] Matteo Zavatteri and Luca Viganò. Conditional simple temporal networks with uncertainty and decisions. Theoretical Computer Science, 797(??):77– 101, December 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397518305942.

Walega:2019:HFH

[2374] Przemysław Andrzej Wałega. Hybrid fragments of Halpern–Shoham logic and their expressive power. Theoretical Computer Science, 797(??):102– 128, December 10, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519300313.

Goranko:2019:ATT

[2375] Valentin Goranko, Antti Kuusisto, and Raine Rönnholm. Alternatingtime temporal logic ATL with finitely bounded semantics. Theoretical Computer Science, 797(??):129–155, December 10, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303780.

Anonymous:2019:EBbu

[2376] Anonymous. Editorial Board. Theoretical Computer Science, 798(??): ii, December 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306875.

Konstantinidis:2019:SSD

[2377] Stavros Konstantinidis and Giovanni Pighizzini. Special section on Descriptional Complexity of Formal Systems. Theoretical Computer Science, 798(??):1, December 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306796.

Yamakami:2019:SCC

[2378] Tomoyuki Yamakami. State complexity characterizations of parameterized degree-bounded graph connectivity, sub-linear space computation, and the linear space hypothesis. Theoretical Computer Science, 798(??): 2–22, December 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930550X.

Ibarra:2019:SGS

[2379] Oscar H. Ibarra and Ian McQuillan. State grammars with stores. Theoretical Computer Science, 798(??): 23–39, December 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519304165.

Cho:2019:SDI

[2380] Da-Jung Cho, Yo-Sub Han, Kai Salomaa, and Taylor J. Smith. Sitedirected insertion: Language equations and decision problems. Theoretical Computer Science, 798(??): 40–51, December 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930307X.

Jiraskova:2019:SCU

[2381] Galina Jirásková and Alexander Okhotin. State complexity of unambiguous operations on finite automata. Theoretical Computer Science, 798(??): 52–64, December 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519302890.

Okhotin:2019:FCP

[2382] Alexander Okhotin and Kai Salomaa. Further closure properties of inputdriven pushdown automata. Theoretical Computer Science, 798(??):65– 77, December 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930283X.

Beier:2019:PRO

[2383] Simon Beier and Markus Holzer. Properties of right one-way jumping finite automata. Theoretical Computer Science, 798(??):78–94, December 17, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302178.

Guillon:2019:LTL

[2384] Bruno Guillon and Luca Prigioniero.

Linear-time limited automata. Theoretical Computer Science, 798(??):95– 108, December 17, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519302105.

Hetzl:2019:CCF

[2385] Stefan Hetzl and Simon Wolfsteiner.

On the cover complexity of finite languages. Theoretical Computer Science, 798(??):109–125, December 17, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519303020.

Anonymous:2019:PDb

[2386] Anonymous. Pages 1–160 (24 December 2019). Theoretical Computer Science, 799(??):1–160, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic).

Anonymous:2019:EBbv

[2387] Anonymous. Editorial Board. Theoretical Computer Science, 799(??): ii, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307108.

DaLozzo:2019:PSG

[2388] Giordano Da Lozzo and Ignaz Rutter. Planarity of streamed graphs. Theoretical Computer Science, 799(??): 1–21, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305808.

Sabir:2019:ECE

[2389] Eminjan Sabir, Aygul Mamut, and Elkin Vumar. The extra connectivity of the enhanced hypercubes. Theoretical Computer Science, 799(??): 22–31, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519305869.

Rahbani:2019:NCL

[2390] Hadi Rahbani, Nader Jafari Rad, and Mohammad-Reza Sadeghi. A note on the complexity of locatingtotal domination in graphs. Theoretical Computer Science, 799(??):32– 39, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306036.

Boros:2019:SGF

[2391] Endre Boros, Vladimir Gurvich, Nhan Bao Ho, Kazuhisa Makino, and Peter Mursic. Sprague–Grundy function of matroids and related hypergraphs. Theoretical Computer Science, 799(??):40–58, December 24, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930605X.

Zhang:2019:HFD

[2392] Shunzhe Zhang, Huiqing Liu, and Xiaolan Hu. Hybrid fault diagnosis capability analysis of triangle-free graphs. Theoretical Computer Science, 799(??): 59–70, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306061.

Ibarra:2019:FFT

[2393] Oscar H. Ibarra and Ian McQuillan.

On families of full trios containing counter machine languages. Theoretical Computer Science, 799(??):71– 93, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306152.

Gu:2019:SMC

[2394] Mei-Mei Gu, Shengjie He, Rong-Xia Hao, and Eddie Cheng. Strongly Menger connectedness of data center network and (n, k)-star graph. Theoretical Computer Science, 799(??):94– 103, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306176.

Miao:2019:PTE

[2395] Runjie Miao, Jinjiang Yuan, and Fei Huang. Paths and trails in edgecolored weighted graphs. Theoretical Computer Science, 799(??):104– 114, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306188.

Murtaza:2019:LDS

[2396] Muhammad Murtaza, Muhammad Fazil, and Imran Javaid. Locatingdominating sets of functigraphs. Theoretical Computer Science, 799(??):115– 123, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306218.

Dose:2019:BPI

[2397] Titus Dose. Balance problems for integer circuits. Theoretical Computer Science, 799(??):124–139, December 24, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S030439751930622X.

Currie:2019:SFR

[2398] James Currie, Tero Harju, Pascal Ochem, and Narad Rampersad. Some further results on squarefree arithmetic progressions in infinite words. Theoretical Computer Science, 799(??):140– 148, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306292.

Pferschy:2019:SKG

[2399] Ulrich Pferschy, Gaia Nicosia, and Andrea Pacifici. A Stackelberg knapsack game with weight control. Theoretical Computer Science, 799(??):149– 159, December 24, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306309.

Anonymous:2019:EBbw

[2400] Anonymous. Editorial Board. Theoretical Computer Science, 800(??): ii, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519307169.

Droste:2019:P

[2401] Manfred Droste, Ilias Kotsireas, and Robert Rolland. Preface. Theoretical Computer Science, 800(??):1– 2, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306383.

Bard:2019:DWG

[2402] Gregory V. Bard. Determining whether a given cryptographic function is a permutation of another given cryptographic function — a problem in intellectual property. Theoretical Computer Science, 800(??):3– 14, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306395.

Caron:2019:SCC

[2403] Pascal Caron, Jean-Gabriel Luque, and Bruno Patrou. State complexity of combined operations involving catenation and binary Boolean operations: Beyond the Brzozowski conjectures. Theoretical Computer Science, 800(??): 15–30, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306425.

Colbourn:2019:DHF

[2404] Charles J. Colbourn, Ryan E. Dougherty, and Daniel Horsley. Distributing hash families with few rows. Theoretical Computer Science, 800(??): 31–41, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306437.

Doukas:2019:HFD

[2405] Nikolaos Doukas, Oleksandr P. Markovskyi, and Nikolaos G. Bardis. Hash function design for cloud storage data auditing. Theoretical Computer Science, 800(??): 42–51, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306449.

Duchamp:2019:KSP

[2406] G. H. E. Duchamp, V. Hoang Ngoc Minh, and Ngo Quoc Hoan. Kleene stars of the plane, polylogarithms and symmetries. Theoretical Computer Science, 800(??):52–72, December 31, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306450.

Fotiadis:2019:TRF

[2407] Georgios Fotiadis and Elisavet Konstantinou. TNFS resistant families of pairing-friendly elliptic curves. Theoretical Computer Science, 800(??): 73–89, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306462.

Kampel:2019:PAC

[2408] Ludwig Kampel, Manuel Leithner, Bernhard Garn, and Dimitris E. Simos. Problems and algorithms for covering arrays via set covers. Theoretical Computer Science, 800(??):90– 106, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306474.

Kampel:2019:SSA

[2409] Ludwig Kampel and Dimitris E. Simos.

A survey on the state of the art of complexity problems for covering arrays. Theoretical Computer Science, 800(??): 107–124, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306486.

Klima:2019:SSR

[2410] Ondrej Klíma and Libor Polák. Syntactic structures of regular languages. Theoretical Computer Science, 800(??): 125–141, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306498.

Lisonek:2019:KSS

[2411] Petr Lisonek. Kochen–Specker sets and Hadamard matrices. Theoretical Computer Science, 800(??):142– 145, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306504.

Moura:2019:UBS

[2412] Lucia Moura, Sebastian Raaphorst, and Brett Stevens. Upper bounds on the sizes of variable strength covering arrays using the Lovász local lemma. Theoretical Computer Science, 800(??): 146–154, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519306516.

Pin:2019:LFG

[2413] Jean-Éric Pin and Xaro Soler-Escrivà. Languages and formations generated by D_{4} and Q_{8}. Theoretical Computer Science, 800(??):155–172, December 31, 2019. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306528.

Sakkalis:2019:RQP

[2414] Takis Sakkalis, Kwanghee Ko, and Galam Song. Roots of quaternion polynomials: Theory and computation. Theoretical Computer Science, 800(??): 173–178, December 31, 2019. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S030439751930653X.

Sadhu:2020:CLT

[2415] Sanjib Sadhu, Xiaozhou He, Sasanka Roy, Subhas C. Nandy, and Suchismita Roy. Corrigendum to: “Linear time algorithm to cover and hit a set of line segments optimally by two axis-parallel squares” [Theor. Comput. Sci. 769 (2019) 63–74]. Theoretical Computer Science, 806(??): 632–640, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0304397519306140. See

[2013].

Lazard:2020:CEL

[2416] Sylvain Lazard, William J. Lenhart, and Giuseppe Liotta. Corrigendum to “On the edge-length ratio of outerplanar graphs” [Theoret. Comput. Sci. 770 (2019) 88–94]. Theoretical Computer Science, 806(??): 689, February 2, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

//www.sciencedirect.com/science/

                                        article/pii/S0304397519307443. See

[2022].

Dang:2020:ECC

[2417] Chuangyin Dang and Yinyu Ye. Erratum/correction to “On the complexity of an expanded Tarski’s fixed point problem under the componentwise ordering” [Theor. Comput. Sci. 732 (2018) 26–45]. Theoretical Computer Science, 817(??): 80, May 12, 2020. CODEN TC-SCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519301677. See

[1698].

Andres:2020:COC

[2418] Stephan Dominique Andres, Melissa Huggan, Fionn Mc Inerney, and Richard J. Nowakowski. Corrigendum to “The orthogonal colouring game” [Theor. Comput. Sci. 795 (2019) 312–325]. Theoretical Computer Science, 842(??):133– 135, November 24, 2020. CODEN TCSCDI. ISSN 0304-3975 (print), 1879-2294 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0304397519307698. See

[2328].